A Complete Bibliography of Publications in Technometrics for the decade 1980–1989

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

03 June 2024 Version 1.03

Title word cross-reference

#9312 [Hsi80].

(≥ 2) [Sin92]. 2^{k−p} [FH80]. A [HM84]. X̄ [RAAN88, San89, BR89, BR88]. β [Mee84a, OCO89]. C_{p} [Wei81]. D [CN80, DA88, GK80a, Hil80, JN83, Khu84, VDY81, Wel82]. F [LM80, WD81]. K [All84, Gen80, GL84a, GL84b, HLG85, Hoc84, Wei84, Hsi86, Sax81, Sin92]. k(≥ 2) [Sin83]. l [CO86]. Λ [RL80]. M [MRR80, DM87, DM88, Sax81]. Pr(Y < X) [RG86]. p [DM87, DM88]. p^{n−m} [Fra84, Fra85]. R [San89, DM87, DM88]. R^{2} [Agr86a, RP80]. S_{B} [Mag80]. t [BJ87, Koe83, Wil83]. |X′X| [Bat83].

-Clustering [All84, GL84b, Hoc84, Wei84, GL84a, HLG85]. -Content [Mee84a]. -Control [BR88, BR89]. -Estimators [MRR80]. -Expectation [Mee84a, OCO89]. -Optimal [CN80, HM84, Hil80, JN83, Khu84, VDY81].

-optimality [Wel82]. -Optimum [DA88, GK80a]. [?]out-of-Saxena:1981:DFT. -to-Enter [WD81]. -Type [Agr86a].

1 [Har83]. 105D [JT83]. 12 [Pri81]. 12th [Fed81j]. 15 [Hsi80]. 17th [Zie89c]. 1964 [Boa85b]. 1979 [Ber83, Fed80c]. 1980 [Fed81d, Fed84d, Par83]. 1981 [Aro84a, Han84a]. 1982 [Fed84e].

21 [Kit81]. 22 [FHL81]. 23 [EL83, HL84]. 24 [LM85, Fed82b]. 25 [Sin92]. 28 [BM87b, Cro89]. 29 [DM88, Ste90].

3 [Aro84c, Fed81c, Liu89]. 30 [BR89, NP89, WB88a]. 31 [BT98]. 39 [IC89].

4 [Aro84d, Zie88-28].

5 [Aro86a, Sch86b]. 51 [Hsi80].

6 [Aro86b].

79 [Hea82].

80f [Kit81]. 81i [FHL81]. 82f [EL83]. 84e [LM85]. 84f [HL84]. 84g [Sin92]. 88e [Ste90].

90e [BR89]. 90f [BT98].

Aberration [Fra84, FH80]. Absolute [NW79, NW85, Par88a]. Abstract [Roh82]. Accelerated [Kan87a, Mee84b, Sch86d]. Acceptance [Ber82a, Car87a, NA88, SY87, Per84]. Accountancy [Hen86]. Accumulation [HW86]. Accuracy [Aro85c, CP81]. Accurate [FM80a]. ACE [dVS89]. Acquisitions [Sch84e]. Acreage [Chh86]. Action [Zie89w]. Ad [Ing81]. Adaptive [HMW80, Lou84, MRR80, Win87]. Added [JM87]. Added-Variable [JM87]. Adding [Wan85]. Additive [FS89a, FS89b, Has89, Haw89]. Adequacy [Lee80b]. Adequate [HJF80, SD80]. Adjustment [Boa87, BF87, Deh87b, DA88, Eas87, LSK87b, PR87, Wu87, LSK87a]. Administration [PS87]. Advanced [Zie89h, vT88, Jac85a]. Advances [FTK89, Web83, Yfa87, Roh80]. Aerial [Hai87b]. Afifi [Cru82, Jac85b]. Age [J.80e]. Agostino [Jac85c]. Agresti [McC85]. Agricultural [Koc81]. Agriculture [Fed81i]. AIC [Kit81, Kit79]. Aickin [Gar85]. Aided [CN89, Wel84, Jac85b]. Aikin [Zie87e]. Air [Cle80, Hor80a, Hor80b]. Aitchison [Pie88a]. Aitkin [Ano81e]. Akaike [Bro89]. Alan [Gbu83, Gos86, McC85, Zie89h]. Alfréd [Jon89a]. Algebra [Cas80, Jac80c, Nel81, Sch85e]. Algorithm [AW80, BS81, DA88, Hai87a, NGG83]. Algorithms [CN80, Lak87, Lin81a, Gra87b, Mee86, Zie88w]. Ali [Maz89]. Allan [Iye87, Jac80d, Tuc82]. Allen [MM83b, War87b, Zie89c]. Allied [Scl81a]. Allocating [Wei81]. Allocations [BT83]. Aloke [Nac87b, Pro89c]. Alternative [DG87]. Alternatives [Ski80, Low83b]. Alvey [Dri89]. Aman [Web83]. Amendola [Rig89]. American [Fed84a]. among [Far85, QK82]. Amount [PC85, Pie88b]. Amsden [Lee89a, Lee89a]. Analyses [GM87, HI87, MS89, St.89, Yfa87, Zie88b]. Analysis [AW89, Ait81, Aro84c, Aro84d, Aro86b, Aro86a, Atk83, BNC87, BLSF88, Ber89b, Bin89, Bis83, BM86a, Caf87, Cal88, CR85, CCMC86, Chh86, Cle80, Cre87, Cru82, Dic82, DGH85, Eas81, EK82, Fed80a, Fed81k, Fed81e, Fed84e, Gar85, Gho88, Goo87, Gun88a, Gun85b, HMS81, HW86, Han85b, Han86, Hei80b, Hil81, Hin80, Hoe87, HI87, Hor80a, Hor80b, HL81b, HL81a, Ize89, Jac85c, Jac85e, Jac86b, Jac86a, Kad84, Kle82, Kot88, Lac82, LC81, Law81, Len89, Lin81a, Lom88, Low83b, Luc81, Lwi81, LM82, Mac81, MWF88, Maz89, Mea89, Mey88, Mie82, Mil80, Mil81, Mul89, Nac87a, NW85, Nel85, Nel88, Orb81, Pag85, Par83, Pie88a, Pre83b, Pro87d, Ree88, RL80, RS81, Sal86, Sch87b, Sch86d, Shu87, Sid84, Ste89b, TW87, Tup89, Wal89, War87b, Wei83, WG79]. Analysis [Win87, Zem86, Zie86c, Zie86b, Zie87b, Zie88m, Zie88-27, Zie88c, Zie89s, Zie89x, dLNB^{+}81, HL84, LM85, Zie88k, Agr88, Bea83, Ber89a, Bow87, Col88a, Cro80, Deh87a, Dun89, Fed80a, Fed81g, Gib83, Goo88, Gri89, Gri88, Gup82, Gup83, Hah82, Ima86, Jac84, Jac85b, Jac89, Jon88, Lac81, Lak87, Lee81, Lee82, Lee83, Mar88b, Moe83, Par88c, Pry89, Rab89, Sch83a, Sch84f, Sch86a, Sch89c, She89, Sin81, SG80, Tuc82, Zie89d, Zie89r, Zie89l, Abr81, Gok83, Kat82, McC85, Roc87a, Sch85a]. Analysis-Plackett [Zie87b]. Analysis-V [Fed80a]. Analytes [Ven82]. Analytical [Min87, Zie89m, Yfa87]. Analyze [Kan87b, Pro87b]. Analyzing [NP88b, NP89, Oma84, Sin87, MM83b]. Anatomy [Eas80a]. Andersen [Aro85c]. Anderson [Aro84a, Aro86b, Par83, Pro88a, Ree88, Zah89, Aro84c, Aro84d, Aro86a, Fed80c, Fed81d, Fed84e, J.80b, Sch86a, Zie85a, Zie89m]. Andersson [Gro89]. André [St.89]. Andrew [She89]. Andrews [Sch87a, Zie87a, Lit85]. Angewandte [Sch84b]. Angular [Lom86]. Ann [Hin87, Min89]. Annals [Fed81a]. Annealing [BJS86, Hai87a]. Annick [Pie89]. Annino [Zie89q]. Anniversary [Lam84b]. Annotated [Sch85g, Zie89x, Zie89x]. Annual [Fed81j]. ANOVA [Ams87, BT98, BT89, Mee84a, SM84, WTA80, Zie89x]. Anscombe [Gun85a, GK80b]. Anthony [Zie86a]. Any [GZ79]. Aoki [Wel89]. APL [Gun85a, Sch82, Sch82]. APL-STAT [Sch82]. Applicable [Che83]. Application [BW85a, BW85b, BJ87, Cle80, DPM80, GH80, Hai87a, Hai87b, HMW80, Hen86, Hoe87, Hor80a, Hor80b, Hub89, JT83, Lee80c, Mac81, McL85, Mee87b, PR82, San89, TW87, Tor89, Yan85, Zem86, Zie85b, dLNB^{+}81, Vin89]. Applications [Ano84e, CJJ81, CJ86, Cro82, HR88, Hod87, Ize87, Mee85, Min86, Sch89a, Sch85d, Sid82, Tuc82, Zie88w, Zie89q, Zie89j, Jac86b, Sto80, Aro85a, Bor86, Bre89, Cor87, Fed81a, Gok83, Gun88a, Ize88, Job89, Kad84, Ker85, Lee89c, McC87, Mee87a, Pen81, Pro88a, Roh80, Han85a, Lew87]. Applied [Ams87, Ano80g, Jac85c, McD79, Rig88, Sch84a, Sin80, St.87, Zie88u, Agr88, Aro84a, Bea83, Bec87, Boe83, Car88, Dun89, Fed81b, Hah82, Lam84a, Mee86, Nac87a, Rus88, Sch81b, Sch85b, Sid84, Wal89, Zie85a, Zie88b, Zie88c]. Apply [Pro87c]. Applying [Agr86a]. Approach [BME85, Boy80, Cha87, LRS82, Lom86, LV86, Rup87, WMS85, Woo84, Yan83, Yfa87, Sin80, Agr89, Cru82, Jac85c, Mac81, She87, St.87, Wat84]. Approximate [BE81, BA89, BR87, DiC87, EM82, MB88, RHM87]. Approximating [Cor80]. Approximation [BN81, FM80a, GZ79, Koe83, MTL82, Ran82]. Aproach [Gru88b]. Araujo [Ald81]. Archetti [Fed81h]. Architecture [Han84b]. ARIMA [MHW84]. ARMA [DP81, Gla82, Jon80, PD80a]. Arnold [Dal87, Gup82, Nac87a]. Aroian [Spr83]. Art [Har89b, Zie87d]. Arthur [Sch89a]. Artificial [Zie89b]. Ascher [Eas85]. Asit [Zie88f]. Aspects [Che85, Jac80d, Tit80a, Gle84]. ASQC [She81a, Fed84a]. Assessing [BH84, SW86]. Assessment [DGH85, GJM89]. Assigning [PBT85]. Assignment [Ber89a]. assignments [Wu81]. Associated [YSOE84]. Association [J.80d]. Assumptions [Col88b]. Assurance [Zie89o]. Asymmetric [MS89]. Asymmetrical [AD83, Che89, GND82]. Asymptotic [Ano90b, Fed80b, Gok81, HZ80, Koz80, SS89, Sen80]. Atkinson [Zie86a, Wei87]. Attribute [Sch84c]. Attributes [SY87]. Audit [Lor80]. Auditing [Min87]. Augmented [AM80, Mal86]. August [Aro84a, Par83]. Autocorrelations [Gla82]. Autoregressions [New82]. Autoregressive [Vec85]. Available [SMV82]. Average [Cro87, VoR85, Vec85]. Awardee [LD80]. Azen [Cru82]. Azencott [Gra87a].

B [Aro83, Bin89, Che83, Coh82, Fed81l, Gup83, Hah81, Han85b, Han86, Hea82, Hub89, Jac85c, Jac85a, Kat82, Liu89, MM83b, Par82, Per82, Pre83b, Sch81a, Sch82, Sch84c, Ste89b, Zie88e, Zie88a, Zie89w]. Back [Ano80a, Ano80b, Ano80c, Ano80d, Ano81a, Ano81b, Ano81c, Ano81d, Ano82a, Ano82b, Ano82c, Ano82d, Ano83a, Ano83b, Ano83c, Ano83d, Ano84a, Ano84b, Ano84c, Ano84d, Ano85a, Ano85b, Ano85c, Ano85d, Ano86a, Ano86b, Ano86c, Ano86d, Ano87a, Ano87b, Ano87c, Ano87d, Ano88a, Ano88b, Ano88c, Ano88d, Ano89a, Ano89b, Ano89c, Ano89d]. Backwards [Ano90a, Sim88a]. Backwards-Stepping [Ano90a, Sim88a]. Bain [Zie88s]. Balakrishnan [Roc87b]. Balanced [BT89, Mee84a, SM81, BT98, Wu81]. Baldessori [Joh82]. Banach [Ald81]. Bancroft [Sin82]. Bands [CI83, Che87, CI88, Dal83, Nai84, Pie86]. Banks [Mor84]. Barker [Law86c]. Barndorff [Mat80]. Barndorff-Nielsen [Mat80]. Barnett [Sch86c, You80]. Barrie [Gun88a]. Bártfai [Fed82b]. Bartholomew [Jac89]. Based [Bai80b, BE81, BLSF88, EBW81, Har89a, HGB89, Hor86, Hsi86, KE80, LM80, MRR80, New80, Pie87, Rup87, SH83, ST86, Sin83, Tit80b, Sin92, Wel82, Lee89b]. Bases [Rig89]. BASIC [Smi88, Gro89, Har80, Low82, She81a, Sto80, Sin80, Maz87, McL89, Gru88b, Hei80b, Mye80]. Basics [Ams87, Was89]. Basu [Zie88f]. Bates [BSBW86]. Bathtub [Hjo80]. Bathtub-Shaped [Hjo80]. Batschelet [Wat82]. Bayes [Bru86, GM87, Mil89, Han89]. Bayesian [Ber89b, Car87a, Eas81, Hil81, HL81a, HL84, Law81, Lwi81, Orb81, RS81, Ber87, Cha87, CB80, EN80, HI87, HL81b, MWF88, Mil80, RPL81, SS81, WMS85, Ber83, Sch83a]. Bayne [Zie88e]. Be [HW86]. Beattie [WW89]. Bechhofer [Ano82e]. Beckman [Ano90c]. Behavioral [Bow83, Hei80b, Mul89]. Behrens [BR87]. Bendat [Iye87, Tuc82]. Bensoussan [Fed81b]. Berbee [Sys82]. Berger [Hsi87, Ber82c]. Bernard [Cre87, Deh87a, J.80a, Obr82a]. Bernardo [Ber83]. Bernhard [Sch84b]. Bernoulli [KN84]. Berry [Obr82a]. Bertrand [Vin89]. Best [Dud80, RP80]. Betteley [Scl81a]. between [Atk81, EJG84, MHW84]. Beyer [Zie89f]. Beyond [Ams87]. Bias [DS88, McL80, RL80, YB87]. Biased [HSH86]. Bibby [Cop83, Lee81]. Bibliographical [Far85]. Bibliography [Sch85g, J.80a]. Bidding [Ano84e, CJJ81]. Biles [Bis81]. Binary [BA89, GC85, Mea89]. Binomial [Atw86, JT83, Kre87, MWF88]. Bio [Syl81]. Bio-Sciences [Syl81]. Bioenergetics [YSOE84]. Biology [Wat82]. Biomedical [Hea82, Mye80, Zie89s]. Biostatistics [Hei80b, Klo82, Zie88m]. Birnbaum [Eas80b, EBW81]. Birth [Hun83]. Bivar [You82]. Bivariate [LM80]. Blanton [Syl87]. Blending [Sne81, Zem86]. Blinchikoff [Zie88p]. Block [BT81, Dea81, HM84, Pro89c, SM81]. Blocked [Lee80c]. Blocking [CN89]. BMDP [Zie88b, Cru84, Hea82]. BMDP-79 [Hea82]. Board [EJG84]. Bocklebank [Dow88]. Body [WP88]. Boen [Gun84]. Bolton [Ker85]. Bolyai [Fed82b]. Bonferroni [CP81]. Book [Abr81, Agr89, Agr88, Ald81, Ams87, And86, Ang88, Ang89, Aro83, Aro84a, Aro84c, Aro84d, Aro84b, Aro85a, Aro85d, Aro85b, Aro85c, Aro86b, Aro86a, Asc89, Bai80a, Bas86, Bas88, Bat88, Bea83, Bec87, Bem89, Ber89a, Ber82b, Ber89b, Ber82c, Ber83, Bil84, Bil88a, Bil88b, Bin89, Bis89, Bis81, Bis83, Boa85b, Boa85a, Boe83, Bon81, Boo88, Bor86, Bor87, Bor89, Bow87, Bow83, Boy80, Bre89, Bro89, Bry89, Bul89, Caf87, Cal88, Car88, Car87a, Car87b, Cas80, Cas82, Cha89, Che83, Chr89, Coh81, Coh82, Col87, Col88b, Col88a, Cor81, Cor87, Cre87, Cro80, Cro88a, Cru82, Cru84, Czi89, Dal87, DeG81, DeP88b, DeP88a, Dea81, Dea82, Dea89, Deh87a, Der88, Der89, Dic89, Dic82, Dow88, Dri89, Dud82, Dun89, Eas80b, Eas80a, Eas85, Edd87]. Book [Ell82, Ens80, Far88a, Far88b, Fed80c, Fed80d, Fed80a, Fed80b, Fed81h, Fed81j, Fed81g, Fed81f, Fed81d, Fed81k, Fed81c, Fed81i, Fed81b, Fed81l, Fed81e, Fed81a, Fed82b, Fed82a, Fed82c, Fed84e, Fed84a, Fed84d, Fed84b, Fed84c, Fre88, Ful89, Gar85, Gau88, Gbu81, Gbu83, Gho88, Gib83, Gib88, Gle84, Gok81, Gok83, Gol88, Goo87, Goo88, Gos86, Gra87a, Gra87b, Gri89, Gri88, Gro89, Gru88b, Gru88a, Gun87, Gun88a, Gun83, Gun84, Gun85a, Gun85b, Gun85c, Gup82, Gup83, Gup89, Hac80, Hah81, Hah82, Han89, Han84a, Han85b, Han85a, Han84b, Han84c, Han86, Han88, Har89a, Har89b, Har86, Har87, Har83, Har80, Has88, Hau87, Hau89, Hea89, Hea82, Hei80b, Hei80a, Hel87, Her89, Hid86, Hin87, Hix87, Hod87, Hod89, Hor86, Hor82, Hsi87, Hub89, Igl88, Ima86, Ing81, Iye87]. Book [Ize87, Ize88, Ize89, J.80d, J.80a, J.80b, J.80e, J.80c, Jac80b, Jac80a, Jac80d, Jac80c, Jac81, Jac84, Jac85f, Jac85d, Jac85c, Jac85e, Jac85b, Jac85a, Jac86b, Jac86a, Jac89, Job89, Joh82, Joh83, Jol88, Jon88, Jon89a, Jon89b, Kad84, Kal83, Kan87b, Kan87a, Kat82, Ker85, Khu87, Kim82, Kin87, Kle82, Klo82, Koc81, Kot88, Lac81, Lac82, Lak87, Lak88, Lam84c, Lam84b, Lam84a, Lan87, Lan88, Lar89, Law86c, Lee81, Lee82, Lee83, Lee89a, Lee89b, Lee89c, Lew87, Lin81b, Lin81a, Lin84, Lit88, Lit89, Liu89, Loc87, Low82, Low83a, Low83b, Luc81, Mac81, Mac83, Mar88a, Mar88b, Mar83, Mas88, Mat80, Maz87, Maz88, Maz89, McC87, McC85, MM83b, McL88, McL89, Mee87a, Mee82, Mee85, Mee86, MM82b, Mey88, Mie82, Mil81, Min86, Min87, Min88, Min89]. Book [Moe83, Mor84, Mor88a, Mor88b, Mul83, Mul89, Mye80, Nac89, Nac87a, Nac87b, Nai88a, Nel81, Nyc89, Obr81, Obr82b, Obr82a, Odo80, Owe88, Owe89, Par89, Par88c, Par82, Par83, Pen81, Per89, Per82, Per84, Per88, Pie88a, Pie89, Pie90, Pig88, Pil83b, Pil83a, Pla88, Por82a, Por82b, Pre83b, Pri81, Pri85, Pro87a, Pro87b, Pro87c, Pro87d, Pro88b, Pro88a, Pro89a, Pro89b, Pro89c, Pry89, Que88, Rab89, Ran82, Ran86, Ran87, Rao80, Rea89, Ree81, Ree88, Rel82, Rey88, Rig88, Rig89, Roc87a, Roc87b, Roh80, Roh81, Roh82, Roh83a, Roh83b, Rup87, Rus88, Sal86, Sal89, Sar89, Sav83, Sch88, Sch81a, Sch83b, Sch83a, Sch82, Sch89a, Sch81b, Sch84g, Sch84e, Sch84b, Sch84a, Sch84f, Sch84d, Sch84c, Sch85d, Sch85e, Sch85b, Sch85f, Sch85g, Sch85c, Sch85a]. Book [Sch86c, Sch86b, Sch86a, Sch87a, Sch87b, Sch89d, Sch89c, Scl81a, Scl81b, Sei82, Sen80, Sen89, Set84, Sha82, Sha84, She88, She81a, She81b, She87, She89, Shu87, Sid80, Sid82, Sid84, Sin80, Sin81, Sin82, Sin87, Ski83, SG80, Smi87, Smi88, Spr83, St.87, St.89, Sta88, Ste89a, Ste89b, Sto80, Swa81a, Swa80, Syl81, Syl87, Syl88, Sys82, Tay83, Ter89, Thi88, Tho86, Tho88, Tho89, Tuc82, Tuc84, Tup89, Tur89, Vin89, Voe89, Wal89, War87a, War87b, War88, Was89, Wat82, Wat84, Web83, Wei87, Wei90, Wel89, Win87, Woo86, Yan89, Yfa87, You80, Zah89, Zai80, Zai87, Zie85a, Zie86c, Zie86a, Zie86b, Zie87c, Zie87b, Zie87e, Zie87h, Zie87g, Zie87f, Zie87d, Zie87a, Zie88e, Zie88b, Zie88i, Zie88k, Zie88o, Zie88l, Zie88y, Zie88m, Zie88p]. Book [Zie88t, Zie88v, Zie88n, Zie88s, Zie88u, Zie88x, Zie88r, Zie88q, Zie88z, Zie88w, Zie88-27, Zie88f, Zie88d, Zie88a, Zie88j, Zie88c, Zie88g, Zie88h, Zie89d, Zie89h, Zie89i, Zie89b, Zie89c, Zie89q, Zie89u, Zie89p, Zie89r, Zie89s, Zie89e, Zie89a, Zie89f, Zie89w, Zie89m, Zie89k, Zie89l, Zie89x, Zie89v, Zie89n, Zie89o, Zie89t, Zie89g, Zie89j, vT88]. Bootstrap [BB89, Rob83]. Borg [Jac81]. Bos [Zie88z]. Both [GC82]. bound [Wel82]. Bounded [BS81, Oma83]. Bounds [JSOS85, MB88, Wal86, Sax81]. Bowen [Sin81]. Bowerman [Hau89, St.87]. Box [Fre88, Boa85a, Boa85a]. Bozdogan [Zie89l]. Bradley [Klo82, Low83b]. Branch [Wel82]. Branch-and-bound [Wel82]. Branching [Sys82]. Break [EJG84]. Breakdown [Ham85]. Brelsford [Ell82]. Brian [Boo88, Lac81, Zie87d]. Brillinger [Aro85b, Boa85b]. Brockwell [Dic89]. Broechaert [Zai87]. Brook [Dal87, Nac87a]. Brookes [Scl81a]. Brown [Klo82, Gun87, Hea82]. Bruce [Hau89, Pig88, Sin81, St.87]. Brunner [Jac80d]. Bruno [Joh82, Sch85c, Swa81a]. Brushing [BC87]. Building [Fre88, HJ86, Gra87a]. Bulk [Bil88a, Bil88b]. Bunke [Zie88g, Zie88g]. Bunn [Har86]. Burman [Zie87b]. Burr [CJ86]. Bury [Rig88]. Business [Sal89, Sch84d]. Butler [Lee89a]. Byrkit [Zie89u]. Byron [Klo82].

C [Ber89b, Boa85a, Bre89, Cas80, Dal87, Dow88, Fed81b, Gbu81, Gib83, Gos86, Gun85b, Han84a, Has88, Jac81, Jon88, Koc81, Lee83, Mas88, McL89, Mey88, Mil81, Moe83, Nac89, Nac87a, Nyc89, Per89, Pri81, Pro87d, Que88, Roh83b, Sal89, Sar89, Sch81b, Scl81a, Sen80, Shu87, Syl81, Sys82, Tho86, Tho88, Wei87, Zah89, Zie86a, Zie88e, Zie88y]. Cady [MM83b]. Calculating [Pie83a]. Calculus [Zie89j]. Calibration [Ano90b, Buo86, CSS88, Eas81, GS82, Hil81, HL81b, HL81a, KSSY84, Law81, LRS82, Lwi81, LM82, MD80, Næs85, Oma84, Orb81, RS81, SS89, SB89, VIC89, HL84, LM85]. Calibration-Curve [CSS88]. Call [PS87]. Call-Record [PS87]. Calvin [Car87a]. Campbell [Aro83, Bow83]. Canada [Fed84d]. Capability [GJM89]. Carbone [Aro85c]. Carl [Hel87, Rao80]. Carl-Erik [Rao80]. Carlo [Ang88, Sei82, Was89, BM87a, SM84]. Carlsson [Han84a]. Carroll [Wei90, Ano90c]. Carson [Mor84]. Carter [Sch84a]. Carvalho [Abr81]. Case [RG86]. Casebook [Klo82]. Casella [Zie88y]. Cases [CW80, CW83, HW88, Wei81]. Cassel [Rao80]. Castelle [Gra87a]. Casti [Czi89]. Catalytic [HMW80]. Categorical [Agr86a, Agr86b, Agr88, Bis83, BJ86, BR85, Gok83, McC85, McC86, Nai86a, Nai86b, Tit80b, Wu81]. Categories [Jac86a]. Caulcott [Zie88i]. Causal [Nyc89]. Cause [Atw86]. Celebration [Zie86a]. Cell [Bai80b]. Cells [Ano90a, Sim88a]. Censored [Ait81, BS83, HMS81, Hsi86, Koz80, Mor88a, Nai84, OS88, San85, SGN85, Sch89b, SH83, Sin83, Sin92]. Censoring [Mee84b, Pie87, Rob83, Kim82]. Censues [Por82a]. Center [Dra82]. Central [Ald81]. Centroids [Pie83a]. Chain [BA89, Car87a, Kat81, Woo84]. Chalmer [Pig88]. Chambers [Woo86]. Chance [Sha82]. Change [GM86, Hen86, Lom86, Lom88]. Change-Point [Hen86, Lom86]. Changing [CB80, VIC89]. Characteristic [She81b]. Characteristics [Lor80]. Characterizations [CW80]. Charles [Fed81i, Joh82, Kad84]. Chart [BLSF88, San89]. Charts [Ano85e, BR88, CW87a, Cro87, LV86, RAAN88, Roc89, San89, WLP85, WLV86, BR89]. Chase [Gun87]. Chatfield [Lee83, Mil81]. Chatterjee [Maz89]. Check [SKC83]. Chemical [Bac80, MW80b, MW80a, TM85, Yfa87, Zai87, Zie89o]. Chemistry [Scl81a]. Chemists [Zie88e, Zie89m]. Chemometric [Agr89]. Cheremisinoff [Cro88a, Pro89a]. Chiang [McC87]. Chien [Sin82]. Chien-Pai [Sin82]. Chin [Maz88, McC87]. Chin-Kuei [Maz88]. Cho [Maz88]. Choice [EK82, GZ79]. Cholesky [HE82]. Chomet [Roh83a]. Chosen [CR80]. Chris [Boy80]. Christensen [Bor87, Smi87]. Chronological [Sch85g]. Circuit [Mee87b]. Circular [Wat82]. Civil [Lee89b]. Claes [Rao80]. Claes-Magnus [Rao80]. Claire [Jon89b]. Clark [Sch89d, Edd87, Jac85b, Zie88c]. Clarke [Sto80]. Class [LLS80, NA88, Par87]. Classical [Bru86, Hur86, Cor87]. Classification [BG84, CW87b, Fed84b, JG80, Lin81a, SK84, SK85, Gra87b, Lac82]. Classifications [J.80d]. Classified [J.80a, Jac86a]. Classifiers [GS82]. Cleveland [Col87, Zie89v]. Clinical [Ker85]. Cluster [Lin81b, Zem86, Lac81, Lin81a]. Clustering [All84, GL84a, GL84b, HLG85, Hoc84, Wei84, Lan87]. Clyde [Asc89]. Cochran [Cor81, Sav83]. Code [Mul83]. Coefficient [New80, dVS89]. Cohen [Mar83, Mul89]. Coles [Syl81]. Colin [Zie89d]. Collected [Boa85b, Boa85a, Zie88-27, Zie88-28, Zie89v]. Collection [Han84a, Sch87a, Zie87a]. Collinearities [MG85]. Collinearity [Par81]. Collins [Lee83]. Colloquia [Fed82b]. Colombo [Lee89c]. Combinations [BG84]. Combinatorial [Ber89a, Fed81a]. Combined [Ham85, Miy82]. Combining [HR88]. Comment [Atk84, Bec84, Cop83, Cor84, Cox84, Hah84b, Joi84, Pik84, SH84a, SH84b, Zie84]. Common [Atw86, SB83, SS81, YSOE84]. Compact [Cas80]. Comparative [Ano84e, CJJ81, Mee82, Nai84, Tit80b, Har86]. Compare [SC82]. Comparing [BT81, BT83, Hsi86, Lee80a, Tor89]. Comparison [Atk81, Bjø80, Bru86, Cam83, CN80, DG87, Khu86, Lit89, Mee84b, SM84, WTA80]. Comparisons [BD82a, HM84, Chr89]. Competing [Eas80b]. Complete [BS83, MTL82]. Complex [Hah84a]. Component [EK82, Goo88, HGB89, Maz80, Miy82, Pie82]. Components [Bre89, BG84, Cha87, EK82, Fel86, GC82, HJ86, JG80, MWF88, Swa81b, SM84, Sax81]. Composite [Dra85, ETM80, Rub80]. Compositional [Pie88a]. Compound [LM82, MLG82, LM85]. Comprehensive [Zie89u]. Compstat [Zie88n]. Computation [MG84, Ste89a]. Computational [DS87, Sch82, Zie88n]. Computer [CN89, Cru82, Der89, Fed81j, GK80a, Hau89, Hea82, Hod89, J.80e, Luc81, Mul83, SSW89, Wel84, Zie88v, Zie89x, DeP88a, Fed82a, Jac85b, Zie89c]. Computer-Aided [CN89, Wel84, Jac85b]. Computers [Cas80, Han84b, Zie89q]. Computing [Edd87, Ell82, Har89a, Kao88, Ste89a, Rel82, Zie87d, Gun85a]. Concentration [SMV82]. Concentrations [SAJ89, Ven82]. Concepts [Ber82c, Gro89, Hau89, Pro87a, Pro88a]. Conceptual [She87]. Conditioning [BH84]. Conference [Aro84a, Fed80c, Fed81d, Fed81k, Fed84e, Par83, Zie89n, Fed81i]. Confidence [Bai80b, BE81, BA89, BG84, CCMC86, CI83, Che87, CI88, Cro82, Dal83, DS87, DL87, GJBR88, JT83, JG80, MD80, McD81, Nai84, Pie86, Rob83, San85, Sch84c, SGN85, Sch87b, ST86, Spu81]. Conflict [DeG81]. Confounded [Fra85]. Confounding [Col89]. Conolly [Che83]. Conover [Obr81]. Consequences [MK84]. Considerations [Buo86]. Consistent [Pie83b]. Consonance [Sal80]. Constance [Aro84b]. Constant [Hjo80]. Constantin [Har89a]. Constituents [SMV82]. Constrained [Cro86a, Pie82, Pie83a]. Constraint [Pie83b]. Constructing [CN80, FMM80, Fra84, JN83, PG86]. Construction [DA88, Hai87a, NG85, Wu81]. Consultant [Zie89w]. Consulting [Gun84]. Contagious [Kat82]. Contemporaneously [DP81]. Contemporary [Sch88]. Content [Mee84a]. Continental [Ano84e, CJJ81]. Contingency [Ano90a, Odo80, Sim88a]. Continuous [Car87a, CLS86, CI83, Haw80, She81a, Sys82, Woo83, Yan83]. Contrasts [BD82a, Fra85]. Contributions [Sav83]. Control [AW89, Ano85e, Bac80, BR88, BH80, BT81, BT83, BLSF88, Ber87, Bry89, CW87a, Cro86b, Cro88b, DPM80, Fed81b, Gun83, HMW80, HM84, Hsi87, Koc81, Liu89, Loc87, LV86, LC82, MW80b, MW80a, MHW84, Pie90, Roc89, San89, Syl87, Vin89, Wal86, WLP85, WN85, WLV86, Zie88x, BR89, Cro89, Fed84a, Maz88, Per82, Per84, Per88, Pro87d, Que88, Sta88, Tuc84, Voe89, Woo86, Zie88f]. Control-Chart [San89]. Control-Test [HM84]. Controllers [HMW80]. Controversies [Pro87a]. Controversy [LM82, LM85]. Convergence [BW81]. Convex [JN83]. Cook [Lin84]. Cooke [Edd87, Sto80]. Cooley [Zie88k]. Coomans [Zai87]. Coombs [Asc89]. Coordinate [Han89]. Coordinate-Free [Han89]. Cornell [Dic82, Pro87c, St.89, Zie87h]. Cornwell [Spr83]. Corrected [Gan81]. Correction [Ano86e, Ano90a, Ano90b, BR89, BT98, BM87b, Cro89, IC89, LM85, McL80, NP89, Sin92, Ste90, YB87]. Corrections [Ano85e, Ano90c, DM88, WB88a]. Correlated [DP81, Hen86, Khu86, OL81, Tan87, WMS85]. Correlation [EJG84, HE82, IC87, IC89, Tuc82]. Corrigenda [Ano80e, Ano80f, Ano81e, Ano82e, Ano82f, FHL81, Kit81]. Corrigendum [Ano80g, Ano83e, Ano84e, Ano84f, EL83, HL84]. Cost [Gau88]. Counted [Luc85]. Counterexamples [EL81, EL83, Zie88o]. Counting [Nel83, Far88a]. Course [Far88b, Goo87, MM82b, Roh83a, Sto80, Zie88k, Hix87]. Covariance [HE82, Low83b, Næs85, War87b]. covariates [Wu81]. Cox [Boe83, BJ81, Col88a, Has88]. Coxon [Jac85f, Jac85d]. Craven [Edd87]. Credal [Sha82]. Crew [Fed82a]. Critchlow [Sin87]. Criteria [Atk81, Box88a, Box88b, CR88, Gun88b, Han84a, Kat81, NP88a, Par88b, STL88, Ull88, Wel82]. Criterion [BW81, Bro89, HH81, HW85, NW79, You82]. Criticisms [EL81, EL83]. Crocker [Pro87d]. Crop [Chh86]. Cross [EK82, J.80d, Jac86a, JG80]. Cross-Classification [JG80]. Cross-Classified [Jac86a]. Cross-Validatory [EK82]. Crossings [Ked87]. Crow [Job89]. Cryer [Bow87]. Csörgö [Fed84d]. CSP [Car87a, She81a]. Cubic [Dra83a]. Cugiani [Fed81h]. Cumulative [CI83, CI88, Cro86b, Cro88b, HH81, Yas89, Cro89]. Cureton [Jac85c]. Current [BE80]. Curtailed [Ano85e, WLP85]. Curve [CSS88, DV86, MD80, She81b]. Curves [CLS86, Nel85, VIC89]. CUSUM [Hea87, LC82, Luc85, VoR85, Wal86, Woo83, Woo84, WN85]. Cut [DV85]. Cuthbert [Luc81]. Cyclic [Mor88b].

D [Abr81, Aro84a, Aro84c, Aro84d, Aro85b, Aro86b, Aro86a, Ber83, Boe83, Boo88, Bow87, Bul89, Car88, Coh82, Col88a, Dea82, Eas80a, Edd87, Fed80c, Fed80d, Fed81d, Fed81l, Fed84e, Fed84d, Gho88, Gup89, Hac80, Hah81, Har86, Har87, Hei80b, Hsi80, J.80b, J.80c, Jac80c, Jac85c, Kim82, Kle82, Lac82, Lee89a, Mee87a, Mee86, Nel81, Odo80, Par83, Roh83a, Sch81b, Sch84c, Sch87a, Sin80, Sin81, Sto80, Ter89, Wal89, Web83, Wei90, Woo86, Zai80, Zai87, Zie87e, Zie87a, Zie89p, Zie89e, Zie89a, Zie89w]. Dacunha [Gra87a]. Dacunha-Castelle [Gra87a]. Dagnelie [Pil83b, Pil83a]. D’Agostino [Kin87]. Dallas [Sch85a]. Dan [Fed81i, Har89a]. Daniel [Agr88, Ell82, Kad84, Koc81, Luc81, She81b, Win87, Hei80b, Pen81, Zie88m]. Data [Agr86a, Agr86b, Agr88, Ait81, Ano84e, BJ87, Ber84, Ber89a, BJ86, BR85, Chh86, CJJ81, CJ86, Cre87, Deh87a, Fed81k, GM86, Gok83, GC85, Gri89, GH80, HMS81, Hai87b, HBK84, Hoe87, Hsi86, Hur86, Jac80d, Jac85e, Jac85a, Jac86a, Kad84, Koz80, Lak87, Lan87, Lee80c, Len81b, Lin81a, Lom86, Luc81, Luc85, Mac81, McC86, MM83b, Mea89, Miy82, Moe83, Mor87, Nai84, Nai86a, Nai86b, Par88c, Pro87b, Pry89, Rab89, Rig89, San85, Sch85a, Sch89c, SH83, She89, SW86, Shu87, SAJ89, Sid82, Sin81, SG80, SS81, Suz85, TM85, Tho86, Tit80b, Tup89, ZG80, Zie88k, Zie88-27, Zie89d, Zie89r, Zie89l, Zie89x, dLNB^{+}81, Bin89, Bis83, Col87, Cre87, Dic82, Gar85, Gho88, Hac80, Hin87, Ing81, Iye87, Jac81, Kan87b, Kle82, Kot88, Lac82, Luc81]. Data [McC85, Pie88a, Ree88, Roc87a, Sch83b, Sch86c, Sin87, Ste89b, Tur89, You80, Zie89s, Bor87, Sch87a, Zie87a]. Data-Dependent [Hur86]. Data-Oriented [Mac81]. Davenport [Hea89, Khu87]. David [Boa85b, Dow88, Jac89, MM83b, Pro87a, Pro88a, Sch84g, Sch87b, Sid84, St.87, Tho86, Wat84, Zie89c, Chr89, Mac83]. Davies [Jac85d]. Davis [Ano80e, Dic89, Shu87, Tuc84]. Dawson [Fed84d]. Dean [Bea83, Dun89, Fed81l]. DeAntoni [Zie88n]. Decade [Lam84b]. Decision [Ber82c, Cal88, Dud82, Han84a, Han85b, Zai87]. Decisions [Har87]. Deconvolution [SD85]. Decreasing [DV86, FG88, Hjo80]. deFinetti [Swa81a]. Defining [Fra85, Pie83b]. DeGroot [Ber83, Zie87f]. Dehnad [Pie90]. DeMaio [Low82]. Deming [Agr89]. Dennis [Gri88, Lam84c, Lin84, Mee87a, Pro88a]. Density [DV86, Tit80b, Deh87a]. Departures [MK84]. Dependent [Hur86, SC82]. Depends [PC85]. DePriest [Sch84e, Mie82, Thi88]. Derivative [Bat83]. Dertougos [J.80e]. Design [Agr89, Ano85e, Atk81, Atk84, BR88, BT83, Bec84, Bis89, Buo86, Che87, Cor84, Cox84, Dea82, FTK89, Gup89, Hah84a, Hah84b, HW85, Han86, JN83, Joi84, Lee89b, LV86, Mor87, Pie90, Pik84, Sen89, SH84a, SH84b, Tan87, Tit80a, Wel84, WLP85, WLV86, Zem86, Zie84, Zie88y, BR89, Law86c, Nac87a, Caf87, Gun85b, Mey88, Zie89d]. Designed [Sch85a]. Designing [Par89]. Designs [AD83, BT81, BD82b, BM86b, CN80, CN82, CN89, Dic82, DA88, Dra82, Dra83a, Dra85, DS88, DJ88, Fed81a, Fra84, FH80, GK80a, GJM89, Hai87a, HM84, Hil80, JN83, Khu84, Khu88, MM83c, NGG83, Par87, SSW89, San89, SM81, SS82, St.89, VDY81, Zie88e, BM87b, Wel82, Dea81, Ens80, Mor88b, Nac87b, Pro89c, Zie85a, Zie87b]. Destructive [EJG84, SY87]. Detect [Sim88b]. Detecting [Ano90a, Boo87, CW80, CP81, Lom88, Mar85, MM83c, Sim88a, Wan87]. Detection [AC89, All84, Dav79, GL84a, GL84b, HLG85, Hoc84, Jai81, Kit79, SAJ89, TM85, TB80, Wei84, Kit81, Pie89]. Deterioration [MM82a, McD81]. determination [OCO89]. Determining [Lor80]. Deterministic [MHW84]. DETMAX [GK80a]. Developing [Maz88, Sne81]. Development [Box84, Han84c, Zie88i]. Developments [Fed81k, Fri83, Hoc83a, Hoc83b, Mit83, Sne83, Wel83, Fed81c]. Deviations [Cor80, Gok81]. Devore [Kot88, Gol88]. d’Expérimentation [Pil83a]. Dey [Nac87b, Pro89c]. Diagnostic [Atk83, Atk86, YY87]. Diagnostics [BNC87, BH84, CR87, CWC^{+}89, HGB89, Sim88b, Wei83]. Diary [Jon89a]. Dickey [Dow88, St.87]. Dickinson [Ima86, Ran86]. Dictionary [Gru88a, Zie89a]. Didier [Gra87a]. Dietrich [Gru88b, Far88b, McL88, Zie89t]. Differences [Spu81]. Different [BM87a]. Differential [BW85a, BW85b, Kal83, McL85, Yan85, Zie85b, Zie89j]. Differing [Gal88]. Difficulty [BME85]. Diffusion [Kal83]. Diffusivity [dVS89]. Digital [DeP88b, Sch89a, Sid82, Per89]. Dillon [Jac86b]. Dimensionality [Fed84b]. Directional [Len81b]. Directory [Gos86, Zie89e]. Discontinuous [Haw80]. Discovering [Nyc89]. Discovery [NW89]. Discrete [Fed81a, Gar85, Kad84, Jac84, Mor84]. Discrete-Event [Mor84]. Discriminant [Jac84, LC81, Pag85, RL80]. Discriminating [Atk81]. Discrimination [McL80, Zie87e]. Discussion [AL82, Agr86b, All84, Asc83, Atk84, Bac80, Bar83, BSBW86, Bec84, Bir83, Boa87, BJ86, BF87, BM88, CR88, Cle80, Cor84, Cox84, Deh87b, Dra83b, Eas81, EDGH86, Eas87, Eng83, Eva83, Fer82, Fri83, FM87, GGP87, Gun88b, Hah84b, HW86, Has89, Haw83, Haw89, Hil81, Hoc84, Hog83, Jae82, Joi84, Law81, LD82, Lwi81, McC86, MM83a, McL85, Mee83, Mit83, Mog82, Moo82, Mor82, Nai88b, NP88a, Orb81, Par88b, PR87, Pik84, Pre83a, Rob88, RS81, Sch83c, STL88, Sne83, Tho83, Ull88, Wei83, Wei84, Wel83, Wu87, Yan85, Zie84, Zie85b]. Discussions [CWC^{+}89]. Dispersion [BM86b, BM87b, Khu86, NP88b, NP89]. Display [Car87b, Gal88]. Distinguishing [FG88]. Distributed [Miy82, Spr83, VoR85]. Distribution [BE81, BJ87, BN81, CA81, CI83, CI88, CG82, CO86, DiC87, EBW81, EB82, EN80, FM80a, FMM80, GM86, HH81, Hjo80, HWW85, HW87, Koe83, LNN61, Lik82, LM80, Lin81b, Mee89a, MH80, Mil80, MTL82, New80, ÖD85, San85, Sax81, SGN85, Sch86d, SH83, SB83, SBE88, Ski80, Ven82, Wal86, Woo83, Zie87c, Zie89h, OCO89, Aro83, J.80a, Ski83]. Distribution-Free [Sch86d, Sax81, J.80a, Ski83]. Distribution-Regression [Zie87c]. Distributional [BE81, Col88b]. Distributions [BM87a, CJ86, Cro87, Hsi81, Hsi86, Job89, Joh82, Koz80, KE80, Law80, MS89, MLG82, Mee84b, QK82, Sch89b, Hsi80, Bor87, Coh82, Kat82, Zie88a]. Disturbance [Ano90b, SS89]. Disturbances [MHW84]. Divisible [Cas82]. Division [Fed84a]. Dixon [Hea82, Cru84]. Do [Sch82]. Do-It-Yourself [Sch82]. Dodge [Roc87a, Zie89r]. Does [JO86]. Doktor [Fed80d]. Domain [Aro85b, Hau87]. Domains [Zie88p]. Donald [Ize88, Obr82a, Pro89b, Smi88, Zie89u]. Douglas [Gbu81, Gib83, Gun84, Gun85b, Mie82, Pro87d, Que88, Sch84e, Sin87, Thi88, Kat82]. Dowdy [Maz87]. Down [IC87, IC89]. Downing [EDGH86]. Draper [Hah82, Fre88]. Driver [Zie89q]. Dual [Gok83]. Duality [MHW84]. Dublin [Fed84e]. Dudewicz [Mul83, Zie89k]. Dugger [Hsi80]. Duncan [Zie88x]. Dunn [Zie88c, Jac85a, Mye80]. Dynamic [CWC^{+}89, Tit80a, TB80]. Dynamical [Czi89]. Dynamics [Ang89].

each [DM87, DM88]. Eason [Syl81]. Easy [CSS88, Len89]. Eckschlager [Yfa87]. Econometric [Kad84]. Econometrics [Fed81b, Jac80c, Zie89g, Zie89j]. Economic [Abr81, Ano85e, BR88, BR89, BHP84, LV86, San89, Tan87, WLP85, WLV86]. Economical [AB85]. Economics [Sal89, Sch84d]. Ed [Cru84, Lam84b]. Edgington [Her89]. Edition [Gun83]. editor [Hsi80, Eas83a, Eas83b, JN81, JRT^{+}80, KECH81, Law85, Law86a, Law86b, Mee89b, SGO82, SB80]. Edosomwan [Zie88r]. Edward [Car87b, Ing81, Mie82, Mul83, Per84, Thi88, Wat82, Zie88a]. Edwards [War87b]. Edwin [Bow83, Job89]. Edzard [Nai88a]. Effect [Che89, GND82, NG85, PM89]. Effects [BT89, BM86b, Fra85, NP88b, Pie82, WTA80, BT98, BM87b, NP89]. Efficiencies [Gok81]. Efficiency [GM86]. Efficient [BSA81]. Efron [Klo82]. Egerton [Cop83]. Eigenvector [BH84]. Elandt [Moe83]. Elandt-Johnson [Moe83]. Electrical [Har86]. Elemental [HBK84, Hin80]. Elementary [Smi88, Sto80, Obr82a, War87a]. Elements [Col87, Ste89a]. Elizabeth [Gib83]. Elliot [Bor89]. Elliott [Gos86, Zie89e]. Embleton [Bin89, Ste89b]. Emission [MM82a, McD81]. Emissions [Lor80]. Empirical [CW80, GM87, Mil89, Fre88, Hod87]. Encyclopedia [Ano86e, Har83, Sch85c, Sch86b, Zie87c, Zie87b]. Enders [Sid82]. Eng [Sin80]. Engelhardt [Zie88s]. Engineering [Asc89, Der88, Der89, Gbu81, Gol88, Hel87, Par89, Par82, Zie88v, Zie89q, Bem89, Lee89b, Han88, Tuc82]. Engineers [Cro88a, Pro89a, Rey88]. English [Gun83]. Ennis [Koc81, She81b]. Enrick [Loc87]. Enter [WD81]. Enterprise [Sha82]. Entropy [Hur86, Ber89b]. Envelopes [FF89]. Environmental [SAJ89, War88]. Equal [Car80, SB83, SBE88]. Equality [Hsi81, Sin83, Hsi80, Sin92]. Equation [HJF80, SD80]. Equations [BW85a, BW85b, GC85, Kal83, Luc81, McL85, Yan85, Zie85b]. Erickson [Ber89b]. Erik [Rao80]. Erkki [Ber82b]. Error [BS81, BD84, Chh86, Dea89, MS89, Næs85, Pag85, RPL81, SK84, SK85, TM85]. Error-Bounded [BS81]. Error-in-Variables [RPL81]. Error-Rate [Pag85, SK85]. Errors [Boo87, Gal88, Ket83, NW77, NW79, NW85, Par88a, ST85, WP88, Sal86]. Errors-in-Variables [Ket83]. ESD [Ros83]. Essay [Sha82]. Essays [Roh83b]. Essentials [Lan88]. Estimate [Boo84]. Estimated [HZ80]. Estimates [BR85, Gal88, Hur86, PG86, Pie87, Tit80b]. Estimating [EJG84, Kat81, Ket83, Miy82, Mor87, Par81, Par87, ST85, SAJ89, YSOE84]. Estimation [BW85a, BW85b, BM87a, Ber82b, Ber89b, CR87, Cha87, CR80, CTC88, CA81, Chh86, DeP88b, Deh87a, DV85, DS88, EBW81, Fel86, GCR86, Gup89, HW85, HGB89, HSH86, HWW85, HW87, JO86, KL88a, KL88b, KN84, KE80, LM82, McL80, McL85, McL89, Nai88b, NW89, New82, Oma83, ÖD85, Pag85, Rob88, Sch89a, SS80, SMV82, Sta88, SS81, Suz85, TW87, Tuc84, Vec85, Weg81, Wel84, Yan85, Zie85b, dVS89, LM85, Dea82, Hid86, Sch81a, Bre89]. Estimator [BSA81, Kre87]. Estimators [Ano90b, AM80, BE81, BD84, DSH80, Far85, GK80b, MRR80, New80, PR82, RDR82, SH83, ST86, SK84, SK85, SS89, Swa81b, SM84]. Eubank [Yan89]. Eugene [Her89]. Eutectics [GC85]. Evaluated [SK84]. Evaluating [EM82]. Evaluation [FF89, SK85]. Evalution [Nac89]. Event [GM87, Mor84]. Everitt [Coh82, Jac85a, Hub89, Lac81, Zie89w]. Exact [Car80, CW87a, CN80, Dal83, DA88, Hai87a, Hsi86, JN83, Khu86, Sch86d]. Examination [DGH85]. Example [Cre87, Ful89]. Examples [BD82a, Boe83]. Exercises [Ree81, Sch84g, Pil83b]. Exhibiting [GC85]. Expectation [Mee84a, OCO89]. Experience [DS87]. Experiences [Swa80]. Experiment [BJ87, HMS81]. Experimental [Atk84, Bec84, Bis89, Cor84, Cox84, DA88, FTK89, Hac80, Hah84a, Hah84b, Joi84, Law86c, MM83b, Mor87, Nac87a, Pik84, Sen89, SH84a, SH84b, WP88, Zem86, Zie84, Zie88e, Wel82, Agr89]. Experimentation [Bis81, Pro89b]. Experiments [Agr86b, Atk81, BT83, BJ86, Cro84, Cro86a, Fra85, GC82, HW86, MS82, Mau84, Maz87, McC86, Nai86a, Nai86b, NP88b, Pie82, Pie83a, Pie83b, PC85, Pie88b, Roc87a, SSW89, Ste88, VDY81, Wel84, Zem86, Zie87h, NP89, Caf87, Gun85b, Gup89, Mey88, Mie82, Sch85a, Dic82, Han86]. Explain [Pro88b]. Explained [Gro89]. explanation [LSK87a]. Explanatory [BT82, GJBR88]. Explicit [Mag80]. Exploration [BT82, CJ86, Jac85a, Kot88]. Exploratory [Gri89, Lac82, Mea89, SG80]. Exploring [Hin87, Jon89b, Min89, Ter89, Tho86, Tup89]. Exponential [EN80, Fai88, Hsi81, Hsi86, Lik82, Mat80, MDH80, MH80, MK84, Pie87, Sen80, Sin83, Hsi80, Sin92]. Exponentiality [BS83, LLS80, LM80, SS87]. Exponentially [Cro87, Miy82, VoR85]. Extended [Car80, Gan81]. Extension [BH80, LSK87a]. Extreme [Boo84, Cle80, EB82, FM80a, FMM80, Hor80a, Hor80b, HWW85, NGG83, Rus88, SH83]. Extreme-Value [EB82, HWW85, SH83].

F [Asc89, Ber83, Che83, Cro80, Fed81j, Gup82, Han85a, Igl88, Jac86b, Kad84, Mac81, McL88, Ree81, Rup87, Sch83b, Sch83a, Sch85d, Sch87a, Sch89c, Sid84, Sin80, Sin81, Zie87a, Zie88n, Zie89s, Zie89t]. Faà [Sch85c]. Faces [Sha84]. Factor [Jac89, JG80, MS82, dLNB^{+}81, Jac85c]. Factorial [CN89, Fra85, HMS81, Nac87b, NP88b, Ste88, Zie85a, NP89]. Factorials [AD83, BM86a, Che89, GND82, Len89]. Factorization [HE82]. Factors [MM82a, McD81]. Failure [ASVR85, Atw86, Hjo80, Kle82, MDH80, Mee87b, Miy82, Sch89b]. Failure-Censored [Sch89b]. Failure-Free [ASVR85]. Failures [MB88].

Families [Mat80, Sen80]. Family [Zie87c, Zie88h]. Farmer [Har86]. Fascination [Dal87]. Fast [LC82]. Fatality [Ano80g, GH80, Lee80c, McD79]. Fatigue [EBW81]. Feedback [TB80]. Feingold [Eas85]. Fewer [DJ80]. Field [KL88a, KL88b, Nai88b, Rob88, Suz85]. Field-Performance [KL88a, KL88b, Nai88b, Rob88]. Fields [Sch87a, Zie87a]. Fienberg [Zie87f, Zie86a]. Fifth [Fed80a]. Fildes [Aro85c]. Filter [DPM80, Weg81]. Filtering [Ize88, Lou84, Win87, Zie88p]. Finding [GK80a, Gen80, Len81a]. Findley [Sid84]. Fingleton [Cre87]. Finite [DSH80, Low82, Coh82, Lak87]. First [Ber83, Cor80, Far88b, Hix87, PD80a, SS82]. First-Order [SS82]. Fisher [Bin89, Ste89b, BR87, Car80]. Fit [BD82b, Dra83a, DG87, Khu85, OS88, SKC83, YY87, Kin87]. Fits [MO86]. Fitting [BJ87, GC85, Jon80, Luc81]. Five [Eas83c]. Fixed [Pie86]. Flannery [Zie87d]. Flexible [FS89a, FS89b, Has89, Haw89]. Flury [Sch84b]. Folded [LNN61]. Folio [Zie88a]. Following [JT83]. Food [Koc81]. Forecasting [Aro85c, Dow88, Gra87a, Hau89, LA81, Swa80, Har86, Fed81d, J.80b, Pri81]. Forecasts [RHM87]. Form [Lit85]. Formation [Lam84c]. Formula [Sch85c]. Formula-Hypothesis [Sch85c]. Forward [WD81]. Foster [MM83b]. Foundation [Hei80b, Zie88m]. Foundations [Aro85a, Ber82c, Gup83, Tho81, Bul89, Rao80]. Four [Mag80]. Fourier [Lom88]. Fowlkes [Zie88a]. Fox [Min86]. Fractional [BM86a, BM86b, HMS81, Nac87b, Zie85a, BM87b]. Framework [Hur86]. Francesco [Fed81h]. Francis [DeP88a, Gun85a, Mee82]. Franck [Hsi80]. Frank [Coh81, Far88b, Gru88b, Por82a, Sch89a, Sta88, LD80]. Fred [Gun87, Luc81]. Frederick [Tho86]. Free [ASVR85, Sch86d, Sax81, Han89, J.80a, Ski83]. Freeman [Agr88]. Freezing [GC85]. Freezing-Point [GC85]. Frequencies [Bai80b]. Frequency [Aro85b, Zie88p]. Fresh [Sin80]. Freund [Hix87, Mar88a, Mas88, Scl81b]. Friedman [SM81]. Friedman-Type [SM81]. Front [Ano80h, Ano80i, Ano80j, Ano80k, Ano81f, Ano81g, Ano81h, Ano81i, Ano82g, Ano82h, Ano82i, Ano82j, Ano83f, Ano83g, Ano83h, Ano83i, Ano84g, Ano84h, Ano84i, Ano84j, Ano85f, Ano85g, Ano85h, Ano85i, Ano86f, Ano86g, Ano86h, Ano86i, Ano87e, Ano87f, Ano87g, Ano87h, Ano88e, Ano88f, Ano88g, Ano88h, Ano89e, Ano89f, Ano89g, Ano89h]. Frontiers [Liu89, Per82]. Frugal [Bon81]. Fujikoshi [Goo87]. Fuller [Dea89]. Function [BSMB82, BJS86, Cas80, CW80, San85]. Functions [CI83, CI88, Cor80, HJ86, Kao88, KC81, LRS82, Nai84, Weg81, Rup87]. Fundamental [Sal80]. Fundamentals [Scl81a]. Further [SD80]. Future [CO86, DM87, DM88]. Fuzzy [Har89a, Roh80, Han85b].

G [Ams87, Ang89, Box87, Bro89, Bul89, Caf87, Cor81, Cro80, Edd87, Fre88, Gro89, Gun88a, Han85a, Har80, Hsi80, Jac80c, Jac85a, Jac86b, Joh82, Klo82, Lee89c, Liu89, Mul83, Per82, Per84, Roh83b, Sav83, Scl81a, Sto80, Syl81, Tuc82, Zie88y, Zie89g]. Gabriel [Ano80e]. Gaines [Han85b]. Gale [Zie89b].

Galen [Hod87]. Gallant [Bat88]. Galwey [Dri89]. Game [DeG81, J.80c]. Gamma [DiC87, GM86, JSOS85, Law80, Mil80, PG86, SB83, SBE88, YB87]. Gamma-Distribution [SBE88]. Gani [Aro85d]. Gardner [EDGH86, Ize87]. Gareth [Gri88]. Gary [Ber89b, Gru88a, Pla88]. Gasoline [Sne81, Zem86]. Gauss [Han89]. Gaussian [CA81, CG82]. Gelsema [Nai88a]. Gene [Hei80b]. General [Bas86, BT81, Lak88, Gun87, Sin80]. Generalization [Gla82]. Generalizations [Cro88b]. Generalized [BJS86, Che85, CJ86, DiC87, HWW85, HW87, KN84, Law80, ÖD85, Ros83, Wan85, Wan87, YB87]. Generalizing [Hur86]. Generated [VDY81]. Generating [MG84]. Generation [Mul83]. Generator [Zem86]. Genichi [Bis89, Par89, Sen89]. Genstat [Dri89]. Gentle [Rel82]. Gentleman [Fed81j]. Geographers [Sch89d]. Geographical [Gri88]. Geology [Shu87]. Geometric [Jac81, Sid80]. Geometry [Cro84, Cro86a]. George [Boa85a, Sch85a]. Gerald [Kan87a, Low82, Sch82]. Gettinby [Syl81]. Ghare [Vin89]. Ghiselli [Bow83]. Ghosh [Roh83b]. Gibbons [Ima86, Ran86, Zie88d]. Gilbert [Bow83, War88]. Gilchrist [And86]. Gill [Kim82]. Gilmore [Nac89]. Giné [Ald81]. Giordano [Sch89a]. Giri [Mey88]. Give [LC82]. Gladen [GGP87]. Glass [Hei80b]. GLIM [Tho89]. GLM [Zie89x]. Glossary [Fed84a]. Glymour [Nyc89]. Gnanadesikan [Har89b]. Goal [Hei80a]. Godfrey [Syl87]. Gokhale [Odo80]. Goldberg [Zie88t]. Goldstein [Jac84]. Golhar [Fed81l]. Gonzalez [Per89]. Good [Aro85a, Aro85a]. Goodman [J.80d, Jac86a]. Goodness [YY87, Kin87]. Goodness-of-Fit [Kin87]. Gordon [Lac82]. Gothenberg [Fed82c]. Graduate [Goo87]. Graham [Cre87, Jac80b]. Grant [Asc89]. Graphical [Gal88, GJM89, Gri89]. Graphics [CWC^{+}89, Zie89v, Thi88]. Graphing [Col87]. Graupe [Win87]. Gray [Zie89e, Gos86]. Graybill [Sch85d]. Green [Hac80]. Gregory [Dal87, Nac87a]. Grenander [Roh82]. Grether [Abr81]. Griffith [Pla88]. Griffiths [Mee86, Zie89g]. Groeneboom [Gok81]. Gross [Gbu83, TM85]. Group [MS82, Mau84]. Grouped [BJ87, TW87]. Growth [Cro82, YSOE84]. GT2 [Spu81]. Guide [Hsi87, Jac85f, Par88c, Sch82, Zie88l, Low82, Gib88, vT88, Gun83, Voe89]. Guided [BT82, dVS89]. Guided-Transformation [dVS89]. Guidelines [JN83, Kan87a]. Gunst [Mac81]. Gupta [Dud82, Roh80, Fed81g, Zie89l]. Gwilym [Swa80]. Gy [Bil84].

H [Agr88, Ber83, Chr89, Cor87, Dal87, Edd87, Far88b, Gau88, Gos86, Han85b, Hsi87, Ize89, J.80d, Jac80d, Lam84a, Lee89a, Lit88, Liu89, Loc87, Mac83, McL88, Mee85, MM82b, Mor88a, Nac89, Owe89, Par82, Per82, Sal89, Sch85f, Sch85g, Set84, Sin80, Smi88, Sys82, Tuc84, Was89, Zie87e, Zie88k, Zie88p, Zie88v, Zie88g, Zie89e, Zie89a, Zie89f, Zie89l, Zie89g, Zie89j]. H.-J [Han85b, Liu89]. Hadamard [NG85]. Hadi [Maz89]. Hahn [Kan87a]. Hakan [Rao80]. Halmos [Han89]. Hamburg [Sin80, Cal88]. Hampel [Rup87]. Han [Sin82]. Hand [Zie89w, Coh82]. Handbook [Has88, Mul83, Sch85b, Zie88b, Goo87, Pla88, She88, Aro83, Aro85b, Asc89, Fed81e, Fed84b, Hau87, Ran87, Zie89f]. Hannon [Hau87]. Hans [Sch84b]. Hansen [Vin89]. Hare [Han86]. Harold [Eas85]. Harr [Lee89b]. Harrington [Gau88]. Harrison [Syl87]. Harry [Eas85, Hah82, Loc87, Sch83a]. Hart [Hsi87]. Harter [Sch85g]. Harvey [Low82]. Haseeb [Owe88]. Hassard [Dal87]. Hassett [Low83a, Min88]. Having [Gal88, Jac86a]. Hayakawa [Goo87]. Hazard [FG88]. Head [LC82]. Health [Hei80b, Hor82, Zie88m]. Healy [Tho89, Zie89i]. Heck [Sch89c]. Heinz [Sch84f]. Held [Aro84a, Ber83, Fed80c, Fed81d, Fed84e, Fed84d, Han84a, Par83]. Helmuth [Lin81a]. Help [JO86]. Henk [Jon88]. Henry [Lan88]. Herbert [Igl88, Sid80, Sin81]. Herzberg [Sch87a, Zie87a]. Heterogeneity [FG88]. Heteroscedastic [BS88]. Hettmansperger [Hor86]. Hewett [Hsi80]. Hibon [Aro85c]. Hidden [Gun85c]. High [Han84b]. Higher [Ked87]. Higher-Order [Ked87]. Highly [Tor89]. Hilary [Hei80a]. Hill [Zie89g, Mee86]. Hines [Gbu81]. History [Zie88j]. Hitoshi [Rea89]. Hoaglin [Tho86]. Hoc [Ing81]. Hochberg [Lit89]. Hocking [Zie86c]. Hodges [GK80b]. Hoffman [EDGH86]. Hogg [Bor89, Han88]. Höhne [Ano81e]. Homogeneity [Ano84e, BB89, CJJ81, MLG82]. Homoscedasticity [Haw81]. Honor [Roh83b]. Hooke [Pri85]. Hopkins [Hei80b]. Horton [Lak88]. Hosking [Sch89d]. Houston [Aro84a, Par83]. Hradesky [Lar89]. Hrishikesh [Web83]. Hsu [Sch89a]. Huber [Por82b]. Hubert [Ber89a]. Huitema [Low83b]. Human [Gun84, Zie88q]. Hunter [Box87]. Husková [Fed80b]. Hypercube [Ste87, Ste90]. Hypotheses [Zie87g]. Hypothesis [Ber82a, Lik82, Sch85c].

Ibbett [Han84b]. IBM [Zie88h]. Identification [Bac80, Ked87, MW80b, MW80a, PD80a, Win87, Set84]. Identifying [Gla82]. II [Sin92, Hsi86, HJ86, Sin83, Nai88a, Sid84]. Ikeda [Kal83]. Illustrated [Han89, LRS82]. Image [Per89, Thi88]. Implementation [Fed80d, LRS82, Hau89, Hsi87]. Importance [Box84, LA81]. Improved [CCMC86, Kre87, Spu81]. Improvement [Lar89, Rea89, Syl87]. Imputation [TW87]. Incidence [BJ81]. Include [Hur86]. Incomplete [BT81, HM84, SW86, Suz85, Dea81]. Increasing [Hjo80]. Increments [Sys82]. Independence [MK84, MTL82]. Independent [Boa87, BF87, Deh87b, Eas87, Lee80a, LSK87b, PR87, Wil83, Wu87, LSK87a]. Index [DG88]. Indexes [MG84]. Individual [HJ86, Wei81]. Induced [MG85]. Industrial [Agr86b, Bis81, BJ86, HR88, HW86, Jac85e, Lam84b, McC86, Nai86a, Nai86b, Pro89b, Zie88x]. Industries [Lew87]. Industry [Han86]. Inference [DiC87, Fed81k, Hor86, Ing81, Law80, Rao80, RG86, Sin82, Zie88g, Bor89, Har80, Joh83, Ran86, Roc87b, Roh82, Tup89, Zie88d]. Inferences [BE80, CB80, EBW81]. Inferential [Gru88b]. Infinite [She81b]. Infinitely [Cas82]. Inflation [RP80]. Influence [CW80, Lin84, Lit85, Rup87, SW86, WB88a, WB88b, Wei83]. Influential [All84, CW80, CW83, DJ81, GL84a, GL84b, HLG85, HW88, Hoc84, Par88a, Wei84]. Information [Ano80p, Ano81m, Ano82o, Ano83m, Ano84n, Ano85l, Ano86j, Ano87i, Ano88i, Ano89i, BME85, Bro89, HR88, Jon89a, Min89, Odo80, PG86, Yfa87, Car87b, Mat80]. Informative [Set84]. Ingwer [Jac81]. Initial [LC82]. Innovation [Dav79, Fed81l]. Innovations [DP81]. Inspections [Tro80]. Institutions [Jac80d]. Integrals [Kim82]. Integrated [McL89, Mee87b, Nac89]. Integrating [Zie88r]. Integration [Roh81, Sch85f]. Intelligence [Zie89b]. Interaction [MJ81]. Interactions [MM83c, Nel88]. Interactive [Cha89]. Intercept [Dal83]. Intercepts [Pie86]. Interface [Fed81j, Zie89c]. Interior [NW85]. Intermational [Fed80c]. International [Aro84a, Ber83, Fed80a, Fed81d, Fed81k, Fed81i, Fed84e, Fed84d, Par83, Zie89n]. Interpretation [GM84, Han86, PD80a]. Interruptible [SS82]. Interval [Cro82, DL87, MD80, PG86, Pie87]. Intervals [Bai80b, BA89, BG84, CCMC86, CO86, DS87, FMM80, JT83, JG80, McD81, MH80, Mil89, OL81, RHM87, RAAN88, Rob83, San85, Spu81, DM87, DM88, OCO89, Sch87b]. Introduction [Dea82, Ize88, Pry89, Sch84a, Sch86a, Sin81, Sta88, Tup89, Dri89, Mil81, Tho89, Zie89u, Boy80, Gib83, Hid86, Hub89, Ize87, Lee83, Par89, Que88, Rey88, Sch88, Syl88, Zie88s, Zie89g]. Introductory [Gib88, Has88, Low83a, Min88, Pen81, Sch84d]. Invariance [Joh83]. Inventory [DPM80]. Inverse [CA81, CG82, HE82]. Involving [SY87]. Ireland [Fed84e]. Ireson [Asc89]. Irregular [Gra87a]. Isaac [Sha82]. Ishiguro [Bro89]. Ishikawa [Gun83, Voe89]. ISP [Cha89]. Issue [Lam84b]. Items [Nel83]. Iterative [HMS81, PD80b, Wu81]. IV [AD83]. Ivor [Mee82].

J [Abr81, Aro85a, Aro85d, Aro85c, Aro86b, Bai80a, Bem89, Ber89a, Ber89b, Ber83, Bil88a, Bin89, Bis81, Boe83, Bor86, Cas80, Coh82, Cru84, Dal87, DeG81, Fed81f, Fed81l, Fed81a, Fed82b, Fed84d, Gbu83, Gib88, Gle84, Gru88b, Hac80, Han85b, Han86, Han88, Har80, Hau87, Hea89, Hea82, Hsi80, Igl88, Jac85e, Jac86b, Kan87a, Kat82, Khu87, Kle82, Lee81, Lee83, Lin81b, Liu89, Mar83, Mas88, McL88, Mee85, MM82b, Mie82, Mor88b, Mul83, Nac87a, Obr81, Par88c, Par82, Per82, Pie88a, Pie89, Pig88, Por82b, Pro88a, Pro89b, Ran82, Rel82, Roh81, Roh83b, Rup87, Sal86, Sch83b, Sch84e, Scl81a, She88, Sin80, Ski83, Ste89b, Sys82, Tay83, Thi88, Tho89, Tup89, Wei90, Woo86, Zie87f, Zie88b, Zie88k, Zie88o, Zie88t, Zie88s, Zie88u, Zie88x, Zie88r, Zie88q, Zie88c]. J [Zie88h, Zie89h, Zie89i, Zie89w, Zie89t, Zie89j, vT88]. Jack [Tho88]. Jackknife [ST86]. Jackknife-Based [ST86]. Jackknifing [FHL80, FHL81]. Jackson [Ano80e]. Jacob [Mul89]. Jaech [Sal86]. Jaffe [Igl88]. Jagdish [Aro83, Jac80a]. Jaiswal [Fed84c]. James [Ber82c, Bis81, Bon81, Far88b, Gau88, Gun84, Hin87, Jac81, Min89, Rel82, Sal89, Sch82, Gra87b]. Jan [Rao80]. Jane [Fed81j]. János [Fed82b]. Japanese [Gun85c]. Jay [Gol88, Kot88]. Jean [Ima86, Mye80, Ran86, Zie88d]. Jenkins [Swa80]. Jerrard [Zie89a]. Jerry [DeP88b, Mor84]. Jessen [Bai80a]. Jim [Har89b, Min89]. Joel [J.80e]. Johannes [Cas82]. John [Boa85b, Bow83, Czi89, Dic82, Dow88, Hix87, Lam84a, Lar89, McL89, Min86, Mor84, Per88, Pro87c, Rab89, Ree81, Sal86, Sar89, Scl81b, Shu87, SG80, St.89, Tho86, Zie87h, Zie88v, Zie88-27, Zie88-28, Zie89v, Zie89o, Dea81, Mor88b]. Johnson [Bea83, Dun89, Moe83, Har83, Jol88, Moe83, Sch85c, Sch85a, Sch86b, SS80, Zie87c, Zie87b]. Joint [WP88]. Jolliffe [Goo88]. Jon [Hod87]. Jonathan [Bow87]. Jonckheere [Ski80]. Jones [Zie88-27, DeG81, Zie88q]. Joseph [Cas82, Ize89]. Joshi [Gup89]. Journal [Hun83]. Jr [Klo82]. Judge [Zie89g]. Jügen [Bre89]. Julius [Iye87, Tuc82]. June [Ber83]. Jure [Lan87].

Kachigan [Mar88b]. Kadane [Zie87f]. Kalbfleisch [Kle82]. Kale [Sch81a]. Kallenberg [Sen80, Sys82]. Kallianpur [Roh83b]. Kalman [Ize88, DPM80]. Kalos [Was89]. Kanal [Fed84b, Nai88a]. Kaoru [Gun83]. Kapadia [Gro89]. Karl [Bil88b, Jac80d, Rig88]. Karson [Jac86b]. Kaye [Zie87e]. Kearns [Gru88b]. Keith [Zie89h]. Keller [Lee89c, Rig89]. Kelly [Coh81, Nyc89]. Kendall [Lee82, Zie89h]. Kennedy [Lin81b, Rel82]. Kenneth [Car87a, Hei80b, Ize88, Koc81, Min87, She81a, She81b, Syl87]. Kent [Lee81, Ano82e]. Kern [Zie88q]. Kernel [BR85, Tit80b]. Kernel-Based [Tit80b]. Kerstan [Cas82]. Key [Jac85d]. Khazanie [Sto80]. Khosrow [Pie90]. Khuri [St.89]. Kimber [She88]. Kinematics [WP88]. Kinetic [ZG80]. King [Bon81]. Kinnison [Rus88]. Kirk [Hid86]. Kitagawa [Bro89]. Kitchens [Tup89]. Klagge [Zie88h]. Klaus [Cas82]. Kleffe [Bre89]. Kleijnen [Tho88]. Kleinbaum [Wal89]. Kleindorfer [Fed81b]. Klonecki [Fed81f]. Klugh [Lan88]. Knowledge [Har89a, Jac80d, Sha82]. Knowledge-Based [Har89a]. Kochetkov [Han84a]. Kolmogorov [Sal80]. Koopmans [Sch88]. Kotz [Har83, Sch85c, Sch86b, Zie87c, Zie87b]. Kozek [Fed81f]. Kraft [Ens80]. Kres [Sch84f]. Krishnaiah [Fed84b, Hau87, Ran87, Roh83b, Aro85b, Fed80a, Fed81c, Fed81e, Zie86b]. Kruskal [J.80d]. Kuei [Maz88]. Kuiper [Koz80]. Kullback [Odo80]. Kumar [Joh83]. Kume [Rea89]. Kunio [Job89]. Kupper [Wal89]. Kushner [Low82]. Kutner [Lam84a].

L [Abr81, Agr89, Bas86, Bis83, Bor86, Cha89, Czi89, Fed80d, Fed84b, Gol88, Gos86, Gru88a, Han85b, Han86, Har89b, Har83, Hau89, Hei80a, J.80e, Jac80b, Job89, Jon89b, Kle82, Lar89, Loc87, Mac81, Mee87a, Moe83, Nac89, Ree88, Roc87b, Sch82, Sch81b, Sch84e, Sch84d, Sch85c, Sch86b, Sch89d, St.87, Sta88, Sys82, Vin89, Wal89, War87a, Yan89, Zie85a, Zie87g, Zie88q, Zie89d, Zie89m, Zie89g]. Labor [Lam84c]. Laboratory [Ize89]. Lack [BD82b, Dra83a, DG87, Khu85, SKC83]. Lambda [ÖD85]. Lambert [Sch88]. Landwehr [Hin87, Min89]. Lane [Dri89]. Large [Bai80b, Bin80, Boo84, Lan87, Nel83, Ste87, Ste90, WG79, Wil80, Gok81]. Larry [Spr83, Tup89]. Larsen [Bor86]. Larson [Par82]. Latent [WG79, Jac89, Sar89]. Latin [Ste90, Ste87]. Launer [Jac80b, Sch84e]. Laurence [She87]. Lauro [Zie88n]. Laveen [Nai88a]. Law [Zie87f]. Lawless [Sch83b]. Lawrence [Ber89a]. Laycock [Cop83]. Layout [GH80]. Layouts [GH81]. Leach [Boy80]. Leamer [Ing81]. Learning [HI87]. Least [BW81, DS87, GZ79, HMS81, ÖD85, Rub80, SMV82, DM87, DM88, Sch89a].

Ledolter [Han88]. Lee [Zie88s, Zie89g]. Left [EN80]. Lehmann [GK80b, Zie87g]. Lemeshow [Hor82]. Length [Cro87, Wal86, Woo83]. Lengths [VoR85]. Lennart [Fed82c]. Lenz [Liu89, Per82]. Leo [J.80d, Jac86a, Spr83]. Leon [Sch85g]. Leone [Ano80f]. Leroy [Pie89]. Leskiw [Ize88]. Lessons [Gun85c, DeP88b]. Lester [Loc87]. Letter [Hsi80, SB80]. Letters [JN81, JRT^{+}80, KECH81, SGO82]. Level [Dra83a, MM83c]. Levels [CP81, Sch86d]. Levi [Sha82]. Levin [Obr82b]. Levy [Hor82]. Lewandowski [Aro85c]. Lewi [Jac85e]. Lewis [Bin89, Sch81b, Ste89b, Sch86c, Sta88, You80]. Liars [Pri85]. Library [Ell82]. Life [ASVR85, Bas88, EBW81, Fai88, FM80b, Kan87a, Lik82, McC87, Mee84b, Mee87b, Mor87, Sch86d, Aro84b, Sch81a]. Life-Test [FM80b]. Lifetime [Sch83b, Suz85, TW87, Sax81]. Lifetimes [Miy82]. Likelihood [BE81, BJ81, CA81, EBW81, Jon80, McL80, MLG82, MTL82, NW89, Pie87, SGN85, Sen80, Sin83, Vec85, YY87, Sin92]. Likelihood-Based [Pie87]. likely [Gen80]. Limit [Ald81]. Limited [Mee87b]. Limits [BE81, BT89, CG82, EB82, GJBR88, Hah81, Mee84a, Mee89a, Sch84c, SGN85, SAJ89, BT98]. Lincoln [Klo82, Pro88b]. Linda [Zie88h]. Lindgren [Obr82a]. Lindley [Ber83, Har87]. Line [AW89, Lin81b]. Linear [Ano90b, BW85a, BW85b, Bec87, Ber82b, BD84, BG84, Cas80, Che85, CB80, CP81, CN82, Eas81, Fri83, Gib83, Gup82, Hai87a, Hil81, Hoc83a, Hoc83b, HL81b, HL81a, JO86, JM87, Ket83, Khu85, Lak88, Lam84a, Law81, Lee80b, Lwi81, LM82, Mar85, Mar88a, Maz89, MO86, MS89, McL85, Mit83, NW77, NW79, NGG83, Oma82, Orb81, PR82, RS81, SW86, Sim88b, Sne83, SS89, TB80, Wan85, Wan87, Wel83, YY87, Yan85, Zie85b, Zie86c, Zie88g, HL84, LM85, Cro80, Czi89, Ens80, Gar85, Gup89, Min86, St.87, Tuc84, Tur89]. Linear-Calibration [LM82, SS89, LM85]. Linear-Optimal [CN82]. Linear-Regression [Hai87a]. Linearized [GK80b]. Lineral [Bas86]. Lines [SC82]. Lingoes [Jac81]. Linhart [Lit88]. Lipow [Zai80]. Liski [Ber82b]. Literature [Zie88l]. Litigation [Zie87e]. Littell [Mar88a, Mas88]. Little [Gho88]. Lloyd [Zai80]. Load [Har86]. Local [Hai87b]. Locally [Swa81b]. Location [BH82, CR80, DL87, HBK84, Hsi86, KE80, Len81b, Rob83, Sin83, Sin92]. Location-Scale [KE80, Rob83]. Loehlin [Sar89]. Log [JSOS85, Law80, Lee80b, YB87]. Log-Gamma [YB87]. Logistic [CJ86, McL80]. Lognormal [Lik80, Mee84b, Sch89b, Job89]. Lohnes [Zie88k]. Long [McC87]. Loren [She88]. Lot [Car87a, She81b]. Lothar [Sch85b, Zie88l]. Lower [CO86, GZ79]. Loxston [Scl81a]. Loynes [Zah89]. Lütkepohl [Zie89g]. Lyle [Zie88-27]. Lyman [Pry89, War87a].

M [Abr81, Ang89, Aro84a, Aro85c, Ber83, Bil84, Bro89, Dal87, DeP88b, Dea81, Edd87, Ell82, Fed84d, Gup83, Hau87, Hea82, Hid86, Hin87, Ize88, Jac85f, Jac85d, Jon89b, Khu87, Kin87, Lak87, Lee81, Lee82, MM83b, Mey88, Nel81, Owe88, Owe89, Par83, Pie89, Pre83b, Rey88, Roc87b, Roh80, Rup87, Sch89a, Sch84a, Sch87a, Scl81a, Sen80, She81b, Sin80, Sto80, Swa80, Syl87, Sys82, Tho89, Tuc84, Vin89, Zah89, Zai80, Zie87f, Zie87a, Zie88t, Zie88u, Zie89i, Zie89p]. Machines [GS82]. Madan [Bas86]. Madsen [Sch84d]. Magazine [Lam84b]. Magnus [Rao80, Zie89j]. Main [Che89, GND82, NG85, PM89]. Main-Effect [Che89, GND82, NG85, PM89]. Major [Aro85c]. Making [Aro85d, Han84a, Cal88, Zai87, Har87]. Makridakis [Aro85c, Cha89, Pri81]. Mallows [BM88, Zie89d]. Malvin [Was89]. Management [Fed80d, Fed81b, Fed81l, Fed84c, Gbu81, Han85b, Mee85, Pri81, She88, Zai80, Asc89, Obr82b, Zie88r, Zie88q]. Managers [Bry89]. Mandel [Cor80]. Mandl [Fed80b]. Manning [Hix87]. Manski [Kad84]. Manual [She81b]. Manufactured [BHP84]. Manufacturing [Bry89, Gun85c]. Many [BT82, Jai81, Ros83, Sch87a, Zie87a]. Many-Outlier [Jai81, Ros83]. Marasinghe [FM87]. March [Fed80c, Fed81d, Fed84e]. Marco [Fed81h]. Mardia [Lee81]. Margerison [Hac80]. Marginal [YY87]. Marie [Fed80b]. Marija [Gib88, Par88c, vT88]. Maritz [Ski83]. Mark [Jol88]. Market [Lam84c]. Markov [BA89, Kat81, Woo84]. Martz [Sch83a]. Marvin [Jac86b]. Marx [Bor86]. Mary [McL89]. Masanao [Wel89]. Mason [Mac81]. Material [Hen86]. Materials [AL82, Bil84, Fer82, GPS82a, GPS82b, Jae82, LD82, Mog82, Moo82, Mor82, Ven82, Bil88b]. Mathematica [Fed82b]. Mathematical [DeG81, Hea89, Khu87, Lin81b, Ran82, Spr83, Zie88s, Zie89f, Zie89k, Fed81f, Mee87a, Rab89, Scl81b, Zie88t]. Mathematics [Che83, Eas80b, Fed81a, Low82, Scl81a, Zai80, Zie88v, Syl81]. Mathews [Zie88v]. Matlock [Zie89e]. Matloff [Hod89]. Matrices [GZ79, NG85, Sch85d, Zie89i]. Matrix [Cor80, HE82, Khu86, Næs85, Sch85e, Zie89j]. Matter [Ano80a, Ano80b, Ano80c, Ano80d, Ano80h, Ano80i, Ano80j, Ano80k, Ano81a, Ano81b, Ano81c, Ano81d, Ano81f, Ano81g, Ano81h, Ano81i, Ano82a, Ano82b, Ano82c, Ano82d, Ano82g, Ano82h, Ano82i, Ano82j, Ano83a, Ano83b, Ano83c, Ano83d, Ano83f, Ano83g, Ano83h, Ano83i, Ano84a, Ano84b, Ano84c, Ano84d, Ano84g, Ano84h, Ano84i, Ano84j, Ano85a, Ano85b, Ano85c, Ano85d, Ano85f, Ano85g, Ano85h, Ano85i, Ano86a, Ano86b, Ano86c, Ano86d, Ano86f, Ano86g, Ano86h, Ano86i, Ano87a, Ano87b, Ano87c, Ano87d, Ano87e, Ano87f, Ano87g, Ano87h, Ano88a, Ano88b, Ano88c, Ano88d, Ano88e, Ano88f, Ano88g, Ano88h, Ano89a, Ano89b, Ano89c, Ano89d, Ano89e, Ano89f, Ano89g, Ano89h]. Matthes [Cas82]. Matthew [Jac84, Low83a, Min88]. Matusita [Fed81k]. Max [Zie88s]. Maximum [BE81, CA81, EBW81, Hur86, Jon80, Lin81b, McL80, NW89, SGN85, Vec85, Ber89b]. Maximum-Entropy [Ber89b]. May [Ber83, Fed84d, Han84a]. McClave [Far88b, McL88, Zie89t]. McCulloch [Zie88y]. McFadden [Kad84]. McLean [Zie85a]. McNeill [Zie89a]. Md [Fed84d]. Mean [BSA81, BD84, Ham85, Kre87, PG86, SAJ89, SK84]. Means [Cor80, MH80, Nel85, Nel88, SBE88]. Measure [IC87, IC89, Khu88]. Measured [OL81]. Measurement [Sal86, Yfa87, Bow83, Dea89]. Measurements [Zie89o]. Measures [Agr86a, Boa87, BD82b, BF87, Deh87b, Eas87, Len81b, LSK87b, PR87, WB88a, WB88b, Wu87, LSK87a, J.80d].

Measuring [Pie82]. Mecke [Cas82]. Medhi [Tay83]. Medical [JT83, Ree81, Zai87]. Meeker [Kan87a, Spr83]. Meeting [Ber83]. Meltzer [Jac80d]. Melvin [Sch84d]. Memoriam [Box87]. Mendel [DeP88b]. Mendenhall [Der89, Mee87a, War87a, Syl88]. Merks [Bil88a]. Mervyn [Han89]. Messina [Bry89]. Messy [Sch85a]. Method [Chr89, Cro87, DZ88, DV85, EM82, HWW85, HJ86, PBT85, Spu81, dVS89, Pie90, Sei82]. Méthodes [Pil83b]. Methodologies [DGH85]. Methodology [Dud82, Fri83, Hoc83a, Hoc83b, Mit83, MKC89, Sne83, Wel83, Pro87c]. Methods [AL82, Ang89, Asc83, Bas86, Ber89a, Bir83, BB89, Cas80, Dav79, Eng83, Eva83, Fer82, GK80a, GCR86, GPS82a, GPS82b, HW86, Hel87, Hub89, Ima86, J.80a, Jac80a, Jac85a, Jac86b, Jae82, Lac82, Law83a, Law83b, LD82, Mee83, Min86, Mog82, Moo82, Mor82, Por82a, Pry89, Rea89, Sal89, Sch83b, Sch83c, Sch89d, ST85, Sin87, Syl87, Tho83, War88, Was89, Zai80, Zie87e, Zie88e, Zie88l, Zie88v, Zie89p, Zie89s, Ber82c, Cor81, Dic89, Gro89, Has88, Jac86b, Ran87, Sch88, Ski83, Wal89, Web83, Zie88t]. Metivier [Roh81]. Metric [Sin87]. Michael [J.80e, Lam84a]. Michel [Roh81]. Microbial [YSOE84]. Microcomputer [Gos86, Han84c, Zie89e]. Microprocessors [Han84c]. Mijnheer [Sys82]. Mike [Gra87b]. Mikel [Gar85, Zie87e]. MIL [JT83]. MIL-STD [JT83]. Miller [Ize88, Klo82, Ams87]. Milliken [Sch85a]. Milton [Lee89b]. Minimax [Bjø80, PR82]. Minimization [Cas80]. Minimum [DS88, Fra84, FH80, Lik82, NW77, NW79, NW85, Par88a, Swa81b]. Minoru [Goo87]. MINQE [CW87b]. Mishra [Zie89k]. Missing [Gho88, Jon80, Roc87a, SS81, dLNB^{+}81]. Missingness [Sim88b]. Misused [Igl88]. Mitchell [GK80a]. MIVQUE [SM84]. MIVQUE’s [Swa81b]. Mixed [BNC87, EM82, MDH80]. Mixed-Model [BNC87]. Mixing [MLG82]. Mixture [AW80, Ber87, Coh82, Cro84, Cro86a, Dic82, GC82, GC85, Hoe87, NGG83, Pie82, Pie83a, Pie83b, PC85, Pie88b, VDY81, Zie87h]. Mixture-Amount [Pie88b]. Mixtures [Dic82, Sne81, SMV82]. ML [SM84]. Model [AW89, Atw86, Bac80, BNC87, BSA81, BJ81, CR87, Cha87, CB80, CN82, Fre88, GC82, Gra87a, HGB89, HJ86, Ked87, Ket83, Khu85, MW80b, MW80a, MJ81, Mee84a, NW89, RPL81, Ven82, WTA80, YY87, YB87, Lak88, Lit88]. Model-Based [HGB89]. Model-Building [HJ86]. Modeling [AC89, BT82, DP81, FS89a, FS89b, Gbu83, Has89, Haw89, Hod89, PD80b, Swa80, Tor89, Wat84, Wel89, Smi87]. Modelling [Jon88, Lee89c, ZG80, Zie89l, And86, DeP88a, Jac85a]. Models [AW80, Atk81, BR88, BT89, Ber87, Che85, CP81, DeG81, Dic82, DGH85, Ens80, Gla82, Gup82, GJBR88, Hai87a, Hai87b, HW85, Har86, Hil80, Jac89, JG80, Jon80, JSOS85, Khu84, Khu86, Lee80b, MS89, Min86, Moe83, NGG83, Par82, PD80a, PC85, Pie88b, Rig88, Rob83, Sch83b, ST85, SKC83, Ski80, Sne81, St.87, Tsa88, Tur89, Vec85, Wan85, Wan87, YSOE84, BR89, BT98, Bas86, Bat88, Ber82b, Dea89, Lam84a, Mar88a, Sar89, Zie86c, Zie88z, Zie88g]. Modern [Cor87, Tup89, Goo87, Owe88, Owe89, Smi88, Syl87, Zie89k].

Modes [CLS86]. Modification [DZ88, Mie82]. Modified [WW89]. Moeschberger [Sch84d]. Moeschlin [J.80c]. Mohamed [Han84c]. Moment [New80]. Moments [HWW85]. Monitoring [WW89, War88]. Monotone [BJ81, FT84]. Monte [Sei82, BM87a, SM84, Ang88, Was89]. Montgomery [Gib83, Gbu81, Gun85b, Que88]. Moore [Pro87a]. Morgan [Agr89]. Morris [Bor86, Cal88]. Mortimer [Lam84b]. Moscow [Han84a]. Moses [J.80e, Klo82, Pro88b]. Most [AB85, Gen80]. Mosteller [Tho86]. MOSUM [BH80]. Motor [GH80, Lee80c]. Mottley [Loc87]. Moving [Cro87, Vec85]. MR [EL83, FHL81, Hsi80, Kit81]. MR0667653 [LM85]. MR0697694 [HL84]. MR0702171 [Sin92]. MR0887702 [Ste90]. MR0970999 [BR89]. MR1007294 [BT98]. Mrudulla [Har89b]. Much [JO86]. Muirhead [Gle84]. Mukherjee [Gup83]. Mullen [Koc81, She81b]. Muller [Wal89]. Multicriteria [Han84a]. Multidimensional [Jac85f, Jac85d]. Multifactor [Luc81, MM83c, Par87]. Multinomial [Bai80b]. Multiparameter [Ber82a]. Multiple [BD82a, BH82, CSS88, Han84a, HE82, HBK84, JO86, KC81, MRR80, New82, Oma82, RHM87, SW86, SD85, SY87, Wor83, Dud82, Lit89, War87b]. Multiple-Regression [HBK84]. Multiple-Use [CSS88]. Multiplicative [MJ81]. Multiresponse [BW85a, BW85b, Khu85, McL85, SS81, YSOE84, Yan85, ZG80, Zie85b]. Multistage [JT83, MF83, Mar85]. Multivariable [Wal89]. Multivariate [Bac80, Bea83, Cro88b, Dun89, Fed80a, Gle84, Goo87, Gup82, Gup83, Han89, Haw81, Hea87, Hin80, Jac85b, Lac82, Lee83, MW80b, MW80a, Næs85, Sch84b, Sch84a, Sch84f, Sch86a, SB89, WN85, Zie87b, Jac86b, Zie88k, Fed80a, Fed81g, Jac85e, Jac86b, Jol88, Lee81, Lee82, Sch89c, Smi87, Zie86b, Zie89l]. Murph [Sin80]. Murtagh [Sch89c]. Musgrave [Sch82]. MVUE [Lik80]. Myers [Cor87].

N [Agr89, Bin89, Cro88a, Fed84d, Fed84b, Fed84c, Fre88, Gup83, Han84b, Hea89, Jac80b, Mey88, Nai88a, Roc87b, Ste89b, Zie87c, Zie87b, Zie88n]. Nalimov [Sha84]. Narula [Ano80f]. Nash [McL89, Cas80]. nearly [Wu81]. Necessary [RDR82]. Neelamkavil [DeP88a]. Negoita [Har89a]. Neil [Col88a, Low83a, Min88]. Nelson [Kan87b, Pro87b]. Nerlove [Abr81]. Nested [BSA81]. Neter [Lam84a]. Network [PS87]. Networks [TM85, Ang88, Coh81]. Neudecker [Zie89j]. Neumann [BN81]. Never [WTA80]. Newbold [Zie88z]. Newby [Ano82f]. Newell [Han85a]. Newman [Jon89b]. Newton [Aro85c, Ize89]. Neyman [Aro84b]. Nicholas [Pro89a]. Nick [Dri89]. Nielsen [Mat80]. Nine [Gun85c]. Nishisato [Gok83]. no [BR89, BT98, BM87b, Cro89, DM88, EL83, FHL81, HL84, IC89, Kit81, LM85, NP89, Sin92, Ste90, WB88a]. Nobuyuki [Kal83]. Noise [Box88a, Box88b, CR88, Gun88b, NP88a, Par88b, STL88, Ull88, LSK87a]. Non [Che83, J.80a]. Non-Parametric [J.80a]. Non-Physical [Che83]. Nonexperimental [Ing81]. Nonlinear [BW81, BS88, DS87, FTK89, FHL80, HW85, Hil80, Khu84, ST85, ST86, Tsa88, FHL81, Bat88, McL89, Wat84].

Nonlinearity [Wan87]. Nonparametric [Ano80g, Boy80, GM86, KSSY84, Lee80c, Lom86, McD79, Obr81, Ran87, San85, Yan89, Zie89p, Bas86, Ima86, Ran86, Zie88d]. Nonparametrics [Joh83]. Nonrandom [Sim88b]. Nonstationary [Cle80, Hor80a, Hor80b]. Norbert [Loc87]. Norma [Bow83]. Normal [Aro83, BS81, CO86, FF89, Hah81, HZ80, LNN61, Mee89a, MH80, Mor88a, RG86, SGN85, DM87, DM88, OCO89]. Normality [Col88b, Haw81]. Normally [Spr83]. Norman [Dri89, Hah82, Har83, Hod89, Moe83, Sch85c, Sch86b]. Norusis [Gib88, Par88c, vT88]. Note [Ait81, Far85, GC82, GC85, Haw80, Hea87, Khu84, Lik82, McL80, PG86, Pie88b, PM89, SD80]. Nottingham [Fed80c, Fed81d]. np [Ano85e, WLP85]. np-Control [Ano85e, WLP85]. Nuclear [AL82, Fer82, GPS82a, GPS82b, Jae82, LD82, Mog82, Moo82, Mor82]. Null [Ski80]. Number [EK82, MB88, Mul83, Nel83]. Numbers [Igl88]. Numerical [Cas80, Sch85f, ST85, Ste89a, Fed81h, Hel87, Zie87d, Zie88v].

O [Aro84a, Aro84c, Aro84d, Aro86b, Aro86a, Ber82c, Fed80c, Fed81d, Fed84e, J.80b, J.80c, Par82, Par83, War88, Zie88c, Zie88g]. Oakland [Per88]. Objects [Lin81a]. Obremski [Jon89b]. Observation [WP88]. Observations [BT83, DJ81, Gra87a, GK80b, Hen86, Jon80, Par88a, SS81, SB89, VoR85, DM87, DM88, Jac86b]. occasions [DM87, DM88]. Occurring [MHW84]. O’Connell [St.87, Hau89]. Odeh [Hah81, Hea89, Khu87, Sch84c, Lin81b]. Offset [BW81]. O’Keeffe [Zie88h]. Olaf [Ens80]. Ole [Mat80]. Olive [Mye80]. Oliveira [Che83]. Oliveira-Pinto [Che83]. On-Line [AW89]. One [BG84, Cha87, CI88, CO86, CW87b, DM87, DM88, Koz80, Mee84a, MB88, San85, Tan87, Wal86, Woo83]. One-Sided [CI88, CO86, MB88, San85, Tan87, Wal86, Woo83, DM87, DM88]. One-Way [BG84, Cha87, CW87b, Mee84a]. Operating [She81b]. Operations [Mee85, Fed81i, Zie88w]. Optimal [BT83, CN80, CN82, Ens80, Fed81a, Hai87a, HM84, Hil80, JN83, Khu84, PS87, Par81, Tit80a, VDY81, Dea82, Sta88]. Optimality [Sen80, Wel82]. Optimally [CR80]. Optimization [Ang88, BJS86, Hub89, KC81, Lak87, Zie88e, Bis81, Der88, Mee85]. Optimizing [Jac80a]. Optimum [DA88, GK80a, Maz80]. ORAGWA [Fed81i]. Ord [Aro86b, Zie89h]. Order [Dra82, Dra83a, Kat81, Ked87, KE80, PD80a, Sch85g, SS82, VDY81, Mac83]. Ordered [Agr86a, Agr86b, BJ86, Jac86a, McC86, Nai86a, Nai86b, Ski80]. Ordering [Dud80]. Ordinal [McC85]. Oriented [Cru82, Mac81]. Origin [CA81, SS87]. Orthogonal [AD83, BD82a, Che89, GND82, NG85, PM89, Nac87b]. Orthogonality [BW81]. Osborn [Ree81]. Osborne [Lak87]. Other [BH84, Col88b, GS82, HR88, Sne81, Wal89, Wel82]. Ott [Pry89]. Ott. [War87a]. Ottawa [Fed84d]. Outer [Ano84e, CJJ81]. Outgoing [HH81].

Outlier [AC89, GH81, Jai81, MG85, Pie89, Ros83]. Outlier-Induced [MG85]. Outliers [AW80, BH82, CTC88, CP81, DJ80, DJ81, GCR86, HBK84, Kit79, Mar85, Gen80, Kit81, Sch86c, You80]. Outlying [Ano90a, Sim88a]. Output [Zie89x]. Overview [AL82, Dud80, Fer82, GPS82a, GPS82b, Jae82, LD82, Mog82, Moo82, Mor82]. Owen [Hah81, Sch84c].

P [Aro85b, Boa85a, Bow83, Car88, Coh81, Cro88a, Cru82, Dea81, Dic89, Fed80d, Fed80a, Fed81g, Fed81c, Fed81b, Fed81e, Fed82b, Fed84b, Fre88, Gok81, Hau87, Hea82, Hor86, Jac85f, Jac85d, Joh82, Liu89, MF83, Mar88a, Mee86, Nyc89, Per82, Pro89a, Ran87, Roh83b, Rup87, Sch89d, Sys82, Tho88, Zie86b, Zie88k, Zie88z, Zie89x, Zie89g]. P.-Th [Liu89, Per82]. Packages [Zie88h]. Pai [Sin82]. Paired [Chr89, LC81]. Pairwise [Spu81]. Pallaschke [J.80c]. Panchapakesan [Dud82]. Paper [Cop83]. Papers [Han84a]. Parameter [Dea82, Fai88, FM80b, HW87, JT83, McL89, Mil80, Mor87, Pie87, SB83, SS80, Zie88z]. Parameters [BE80, BA89, CTC88, CA81, EBW81, Hsi86, Ket83, Mag80, ÖD85, Par81, SGN85, SH83, ST85, SB83, SBE88, Sin83, SS87, SS81, Suz85, YSOE84, Sin92]. Parametric [J.80a, Mil89]. Pareto [CJ86, HW87]. Pareto-Logistic [CJ86]. Parsimonious [FS89a, FS89b, Has89, Haw89]. Parsimony [LA81]. Part [Owe88]. Partial [Gla82, Mal86]. Partially [Che85, Hil80, Khu84, PM89, Sin87]. Particle [BJ87]. Particulate [Bil84, Ven82]. Parzen [Aro85c]. Pascal [Edd87]. Passenger [Fed82a]. Patel [Aro83]. Path [Sys82]. Patil [Joh82]. Pattern [Cre87, Fed84b, Zai87, Nai88a]. Patterson [Sal89]. Paul [Hor82, Jac85e, Per89]. Paula [Was89]. PC [Zie88h]. Pearson [Hea89, Hel87]. Peck [Gib83, Kot88]. Pellaumail [Roh81]. Percentage [Jai81, Koz80, Ros83, Wil83]. Percentile [Mag80]. Percentiles [Boo84, Koe83]. Perform [Car87a, Col88a, She81a]. Performance [Boa87, BF87, Box88a, Box88b, CR88, Deh87b, Eas87, Gun88b, Han84b, KL88a, KL88b, LSK87a, LSK87b, MS82, Mau84, Nai88b, NP88a, Par88b, Pie87, PR87, Rob88, STL88, Ull88, Wu87]. Period [ASVR85]. Periodic [New82, Vec85]. Periodogram [BLSF88]. Perryman [Aro84a, Par83]. Personal [Zie89q]. Personalities [Far88a]. Perspective [LM82, LM85]. Perspectives [Zie88u]. Pertinent [Zie88l]. Peter [Dri89, Pie89, Por82b]. Peters [Far88a]. Petersen [Caf87]. Petr [Fed80b]. Pfaffenberger [Sal89]. Pharmaceutical [Ker85]. Phase [Kao88, Wor83]. Phase-Type [Kao88]. Phenomena [Che83]. Philip [Has88]. Phillip [Bem89]. Philosophy [Zie88-27]. Physical [Che83, Sid82]. Piegorsch [GGP87]. Piepel [Ano90c]. Pierre [Bil84]. Piersol [Iye87, Tuc82]. Pike [Der88]. Pinto [Che83]. Pitman [Har80]. Plackett [Zie87c, Zie87b, Bis83]. Plan [AB85, Kan87a, KN84, Kre87]. Planning [Lam84c, PS87]. Plans [Ano85e, Che89, EM82, FM80b, GND82, Hah81, HH81, Mee84b, NG85, PM89, Sch84c, Sch89b, SY87, WLP85, Yan83, She81b]. Plot [Mea89]. Plots [FF89, JM87, Wan87, Pro87b, Wei87]. Point [Cas82, Cre87, GC85, Hen86, Lom86, Fed82b]. Points [Dra82, Ham85, Jai81, Koz80, Lin81b, Lom88, Mag80, Ros83, Rub80, SKC83, Wil83]. Poisson [Lee80b, MB88]. Policy [Jac80d, Lam84c]. Policymaking [Jac80d]. Pollock [Jac80c]. Pollution [War88]. Polynomial [KC81, Pie86]. Pool [WTA80]. Poor [Gau88]. Poor-Quality [Gau88]. Population [HZ80, Mee87b, DM87, DM88]. Populations [FM80b, Sin83, Sin92, Dud82, Mor88a]. Possible [GCR86]. Potential [Zai87]. Potter [Zie88t]. Powders [Bil88b]. Power [Atk83, MG84, Mul89, Nel85, Tsa88]. Power-Generating [MG84]. Prabhakar [Vin89]. Practical [Kan87a, Ker85, Lee89a, Wat84, Zah89, Cro88a, Obr81, Pro89a, Swa80, Zie88e, Zie89m]. Practice [Aro84c, Aro84d, Aro86b, Aro86a, Box84, Fed84e, Han84a, Nai88a, Zie89g, Bil84, Jac85e, Zie88y]. Practitioners [Tho88]. Prague [Fed80b]. Pranab [Bas86, Joh83]. Pre [Sch85g]. Pre-1950 [Sch85g]. Precise [HW85]. Precision [GS82]. Prediction [BT82, BD84, CG82, CO86, EB82, EN80, FMM80, GJM89, HR88, MH80, NW77, OL81, RHM87]. Predictor [BD82b]. Pregibon [Lit85]. Prentice [Kle82]. Presence [CTC88, MM83c]. Presented [Han84a]. Press [Zie87d]. Pressure [LRS82]. Pressure-Volume [LRS82]. Priestley [Pre83b]. Primer [Mye80]. Principal [CLS86, EK82, McC84, Goo88]. Principes [Pil83a]. Principles [Boe83, Far88a, Joh83, Wei83, Zie88-27]. Pringle [Dal87]. Prior [PG86]. Priorities [PBT85]. Prizes [LD80]. Probabilistic [Par82]. Probabilities [Bai80b, BS81, Hei80a]. Probability [Aro85a, Fed81f, FF89, HWW85, QK82, Rey88, Roh83b, Sha82, Syl88, Zie88o, Zie88s, Jon89b, Sid80, Swa81a, Bor89, Gbu81, Gol88, Hod89]. Probability-Weighted [HWW85]. Probable [Bin80, Wil80]. Problem [BR87, Eas81, HR88, Hen86, Hil81, HL81b, HL81a, Law81, Lom86, Lwi81, Orb81, RS81, TW87, HL84]. Problems [Che87, Fed82b, J.80a, Oma84, Sch87a, Zie87a, Ber89b]. Procedure [AW89, CSS88, LC81, Lor80, Mad82, Mar85, Maz80, MRR80, NA88, PD80b, Ros83, Tan87, WW89, Woo84]. Procedures [Ano80g, Ber84, Bjø80, Cro82, Dud82, ETM80, GH80, Hea87, Jai81, LD80, McD79, WTA80, Woo83, WN85, Lit89]. Proceedings [Aro84a, Ber83, Fed80c, Fed80a, Fed81d, Fed81k, Fed81i, Fed84e, Fed84d, Par83, Zie88n, Zie89c, Fed80b, Fed81j, Fed82c, Zie89n]. Process [AW89, AB85, BE80, Cle80, Cro82, GC82, Hor80a, Hor80b, Hsi87, Lee80b, MDH80, Per88, TM85, VDY81, WW89, Woo86, Yan83, Sch84e]. Process-Control [AW89]. Processes [CLS86, DV85, Fed82b, Hod87, Ize87, Lee80a, MHW84, Sys82, Tor89, Cas82, Kal83, Par89, Roh83a, Sys82, Tay83]. Processing [Sid82, Thi88, Per89, Sch89a]. Producer [Bru86]. Product [BHP84, Nac89, Spr83, Zie87h]. Production [AB85]. Productivity [Zie88r, Lar89]. Products [Par89]. Professionals [Hor82]. Profit [Loc87]. Programming [Maz88, Smi88]. Programs [Hea82, Cha89]. Progression [Han89]. Progressive [Rob83]. Properties [DSH80, ETM80, MDH80, New80, Sal80, Ste87, Sys82, Ste90]. Proportion [HZ80]. Proportions [MLG82, Sch84c]. Pseudocomponents [Cro84]. Public [Fed81d, Fed82a]. Publications [Ano80l, Ano80m, Ano80n, Ano80o, Ano81j, Ano81k, Ano81l, Ano82k, Ano82l, Ano82m, Ano82n, Ano83j, Ano83k, Ano83l, Ano84k, Ano84l, Ano84m, Ano85j, Ano85k]. Pukkila [Zie89n]. Puntanen [Zie89n]. Puri [Bas86, Koc81, Zie88u].

Q [Kan87a, Spr83]. QC [Lew87]. QSAR [BME85]. Quadratic [HW85, Lit85, Swa81b]. Qualitative [DJ88]. Quality [BH80, Bem89, BLSF88, BHP84, Bry89, Cle80, Cro86b, Cro88b, Fed84a, Gun83, HH81, Hor80a, Hor80b, Koc81, Lar89, Liu89, LC82, Maz88, Par89, Per82, Per84, Pro87d, Que88, Rea89, Syl87, Voe89, Wal86, WN85, Zie88r, Zie88f, Cro89, Lee89a, Zie87h, Gau88, Law86c, Loc87, Maz88, Pie90, Pla88, She88, Vin89, Zie88x, Zie89o]. Quality-Control [Cro88b, LC82, WN85]. Quantal [Mor87]. Quantile [HW87]. Quantiles [CR80, KE80]. Quantitative [Car87b, Cre87, DJ88, Ima86]. Quasi [VDY81]. Queue [Mar83]. Queueing [Ang88, Han85a]. Queuing [Fed82b]. Quick [CSS88, Col89, Gan81, Len89].

R [Aro84a, Aro85b, Aro85c, Bis83, Boa85b, Boe83, Bon81, Car87b, Dic89, Fed80d, Fed80a, Fed81g, Fed81c, Fed81e, Fed84b, Fre88, Gho88, Gri89, Gro89, Gun84, Hac80, Han85b, Han86, Han88, Hau87, Hea89, Hod87, Jac86b, Khu87, Kim82, Kin87, Kle82, Lak87, Lee89a, Lin81b, Lin84, Mac81, Mar88a, Mas88, MM82b, Min87, Nyc89, Pro88a, Ran87, Roh80, Roh83b, Rup87, Rus88, Sch85e, Sin80, Tho89, Tur89, Wei90, Zah89, Zie86c, Zie86b, Zie88k, Zie89i, Zie89q, Zie89x]. Råde [Fed82c]. Radhakrishna [Bre89]. Rafael [Per89]. Rafiquzzaman [Han84c]. Ragade [Roh80]. Raghavarao [Ter89]. Ralescu [Har89a]. Ralley [Mul83]. Ralph [Der88, Jac85c, She88]. Ramakant [Sto80]. Ramsey [Sch82]. Randall [Yan89]. Random [AW89, Ald81, BT89, Bin80, Cha87, CI83, Dal83, Ize87, Mee84a, Mee89a, Mul83, Nel81, Spr83, Wil80, WTA80, Woo83, BT98, Sys82, Iye87]. Random-Effects [BT89, BT98]. Random-Walk [AW89]. Randomization [Her89]. Randomly [MHW84, VIC89]. Randomness [Sch84g]. Range [Wil83]. Rank [Ano84f, GZ79, IC79, MS89]. Ranked [Sin87]. Ranking [Dud80, Dud82]. Ranks [Hor86]. Rao [Bre89, Fed84d, Hau87, Roh83b]. Rasch [Zie89p]. Rate [Atw86, Lee80c, Lee80b, Pag85, SK84, SK85]. Rates [Ano80g, FG88, GM87, Hjo80, Lee80a, McD79, Miy82]. Ratio [BN81, MLG82, MTL82, Sen80, Sin83, Sin92]. Rational [GM84]. Ratios [Box88a, Box88b, CR88, Gun88b, MH80, NP88a, Par88b, STL88, Ull88, LSK87a]. Ratkowsky [Wat84]. Ray [Ber89b, Par83]. Raymond [Bai80a, Cor87, Lak88, She88]. Reactor [Bac80, HMW80, MW80b, MW80a]. Read [Aro83]. Real [Ald81]. Really [RDR82]. Received [Ano80l, Ano80m, Ano80n, Ano80o, Ano81j, Ano81k, Ano81l, Ano82k, Ano82l, Ano82m, Ano82n, Ano83j, Ano83k, Ano83l, Ano84k, Ano84l, Ano84m, Ano85j, Ano85k]. Recipes [Zie87d]. Recognition [Fed84b, Nai88a, Zai87]. Reconciliation [TM85]. Record [PS87]. Recorders [Cam83]. Rectangular [BS81]. Reduced [Ber82b]. Reduction [Fed84b, GC82, Lin81a]. Redundant [Maz80]. Rees [Bul89]. Reference [LD80, She81a, Sid82]. Regina [Moe83]. Region [Oma83]. Regional [Hai87b]. Regions [BS81, DS87, Pie83b, ST86]. Regression [Ait81, All84, Ano84f, Atk83, BT82, BS88, Ber84, Boo87, BS83, BD84, CR85, CR87, Che85, CW80, CW83, CWC^{+}89, Cor87, Cro80, Dal83, DJ81, EL81, FF89, FHL80, Fri83, Gib83, GCR86, GL84a, GL84b, Hah82, Hai87a, HW85, HE82, HBK84, Hil80, HJF80, HW88, Hoc83a, Hoc83b, Hoc84, HSH86, IC79, JO86, JM87, JSOS85, KC81, Khu84, Lam84a, Mar85, Mit83, MRR80, MB88, Nac87a, NW77, NW79, NW85, Oma82, Oma83, Par81, Par88a, PR82, Pie86, Pie89, Pro87d, RP80, SW86, ST86, Sim88b, SC82, Sne83, SD80, Tsa88, WB88a, WB88b, Wal89, War87b, Wat84, Web83, Wei83, Wei84, Wel83, WG79, WD81, Wor83, YY87, YB87, Zie87c, Zie88z, EL83, FHL81, Bec87, Lin84, Mas88, Maz89, MM83b, Wei87, Wei90, Yan89, Zie88c, Gun88a, Mac81]. Regression [MM82b, Zie88k]. Regressions [Haw80, Rub80]. Regressors [You82]. Reid [Aro84b]. Rejection [GH81, Ham85]. Related [Fed84d, GK80a, HW86, J.80c, Kao88, Min86]. Relational [Jac81]. Relations [Far85]. Relative [BW81, NW77]. Reliability [Asc83, Asc89, BE80, Bir83, Cro82, Eng83, Eva83, Hjo80, KL88a, KL88b, Kan87b, Law83a, Law83b, MWF88, MG84, Mee83, Mee87b, Nai88b, QK82, Rob88, Sch81a, Sch83a, Sch83c, Tho81, Tho83, Tor89, Car88, Eas85, Bas88, Lee89b, Lee89c, Rig89, Sch84e, Zai80, Zie88f]. Reliability-Based [Lee89b]. Reliable [Tor89]. Relles [Ell82]. REML [SM84]. Renewal [BSMB82, DV85, Kao88, Sys82, Tor89, Yan83]. Renewal-Process [Yan83]. Rényi [Jon89a]. Repairable [Eas85]. Reparameterizations [Tsa88]. Replicated [NP88b, PM89, NP89]. Reply [Bin80, HJF80]. Report [Eas83a, Law85, Law86a]. Representations [Jac81]. Represented [KC81]. Research [Fed81i, Fed81l, Sch87a, Zie87a, Zie88i, Zie88w, Lan88, Mee85, Min86, Ree81, Sin82]. Residual [Wan87]. Residuals [Boo87, FF89, Mal86, Oma84, Lin84]. Resolution [AD83]. Resolution-IV [AD83]. Resources [Fed81i, Zie88q]. Response [Ano83e, BW85b, BC83b, BD82b, Box88b, CR87, CCMC86, CN89, CWC^{+}89, DGH85, Dra82, Dra83a, DJ88, Fre88, FS89b, GJM89, GPS82b, GL84b, Hoc83b, Hor80b, HL81a, HJ86, KL88b, Khu86, Khu88, Law83b, LSK87b, LC82, MW80a, MKC89, Nai86b, Oma83, Par87, PC85, Pro87c, SKC83, SH84b, Wel84, St.89]. Response-Surface [CN89, DGH85, DJ88, Khu88]. Responses [KC81, Par81]. Restricted [SMV82]. Restrictions [Par81]. Resultant [Bin80, Wil80]. Results [BE81, CW87a, Han86]. Resurfacing [PBT85]. Reuven [Ang88, Sei82]. Reversibility [Coh81]. Review [Abr81, Agr89, Agr88, Ald81, Ams87, And86, Ang88, Ang89, Aro83, Aro84a, Aro84c, Aro84d, Aro84b, Aro85a, Aro85d, Aro85b, Aro85c, Aro86b, Aro86a, Asc89, Atk84, Bai80a, Bas86, Bas88, Bat88, Bea83, Bec84, Bec87, Bem89, Ber89a, Ber82b, Ber89b, Ber82c, Ber83, Bil84, Bil88a, Bil88b, Bin89, Bis89, Bis81, Bis83, Boa85b, Boa85a, Boe83, Bon81, Boo88, Bor86, Bor87, Bor89, Bow87, Boy80, Bre89, Bro89, Bry89, Bul89, Caf87, Cal88, Car88, Car87b, Cas80, Cas82, Cha89, Che83, Chr89, Coh81, Coh82, Col87, Col88b, Col88a, Cor81, Cor84, Cor87, Cox84, Cre87, Cro80, Cro88a, Cru82, Cru84, Czi89, Dal87, DeG81, DeP88b, DeP88a, Dea81, Dea82, Dea89, Deh87a, Der88, Der89, Dic89, Dic82, Dow88, Dri89, Dud82, Dun89, Eas80b, Eas80a]. Review [Eas85, Edd87, Ell82, Ens80, Far88a, Far88b, Fed80c, Fed80d, Fed80a, Fed80b, Fed81h, Fed81j, Fed81g, Fed81f, Fed81d, Fed81k, Fed81c, Fed81i, Fed81b, Fed81l, Fed81e, Fed81a, Fed82b, Fed82a, Fed82c, Fed84e, Fed84a, Fed84d, Fed84b, Fed84c, Fre88, Ful89, Gar85, Gau88, Gbu81, Gbu83, Gho88, Gib83, Gib88, Gle84, Gok81, Gok83, Gol88, Goo87, Goo88, Gos86, Gra87a, Gra87b, Gri89, Gri88, Gro89, Gru88b, Gru88a, Gun87, Gun88a, Gun83, Gun84, Gun85a, Gun85b, Gun85c, Gup82, Gup83, Gup89, Hac80, Hah81, Hah82, Hah84b, Han89, Han84a, Han85b, Han85a, Han84b, Han84c, Han86, Han88, Har89a, Har89b, Har86, Har87, Har83, Har80, Has88, Hau87, Hau89, Hea89, Hea82, Hei80a, Hel87, Her89, Hid86, Hin87, Hix87, Hod87, Hod89, Hor86, Hor82, Hsi87, Hub89, Igl88]. Review [Ima86, Ing81, Iye87, Ize87, Ize88, Ize89, J.80d, J.80a, J.80b, J.80e, J.80c, Jac80b, Jac80a, Jac80d, Jac80c, Jac81, Jac84, Jac85f, Jac85d, Jac85c, Jac85e, Jac85b, Jac85a, Jac86a, Jac89, Job89, Joh82, Joh83, Joi84, Jol88, Jon88, Jon89a, Jon89b, Kad84, Kal83, Kan87b, Kan87a, Kat82, Ker85, Khu87, Kim82, Kin87, Kle82, Klo82, Koc81, Kot88, Lac81, Lac82, Lak87, Lak88, Lam84c, Lam84b, Lam84a, Lan87, Lan88, Lar89, Law86c, Lee81, Lee82, Lee83, Lee89a, Lee89b, Lee89c, Lew87, Lin81b, Lin81a, Lin84, Lit88, Lit89, Liu89, Loc87, Low83a, Low83b, Luc81, Mac81, Mac83, Mar88a, Mar88b, Mar83, Mas88, Mat80, Maz87, Maz88, Maz89, McC87, McC85, MM83b, McL88, McL89, Mee87a, Mee82, Mee85, Mee86, MM82b, Mey88, Mie82, Mil81, Min86, Min87]. Review [Min88, Min89, Moe83, Mor84, Mor88a, Mor88b, Mul83, Mul89, Mye80, Nac89, Nac87a, Nac87b, Nai88a, Nel81, Nyc89, Obr81, Obr82b, Obr82a, Odo80, Owe88, Owe89, Par89, Par88c, Par82, Par83, Pen81, Per89, Per82, Per84, Per88, Pie88a, Pie89, Pie90, Pig88, Pik84, Pil83b, Pil83a, Pla88, Por82a, Por82b, Pre83b, Pri81, Pri85, Pro87a, Pro87b, Pro87c, Pro87d, Pro88b, Pro88a, Pro89a, Pro89b, Pro89c, Pry89, Que88, Rab89, Ran82, Ran86, Ran87, Rao80, Rea89, Ree81, Ree88, Rel82, Rey88, Rig88, Rig89, Roc87a, Roc87b, Roh80, Roh81, Roh82, Roh83a, Roh83b, Rup87, Rus88, Sal86, Sal89, Sar89, Sav83, Sch88, Sch81a, Sch83b, Sch83a, Sch82, Sch89a, Sch81b, Sch84g, Sch84e, Sch84b, Sch84a, Sch84f, Sch84d, Sch84c, Sch85d, Sch85e, Sch85b, Sch85f]. Review [Sch85g, Sch85c, Sch85a, Sch86c, Sch86b, Sch86a, Sch87a, Sch87b, Sch89d, Sch89c, Scl81a, Scl81b, Sei82, Sen80, Sen89, Set84, Sha82, Sha84, She88, She81a, She81b, She87, She89, Shu87, Sid80, Sid82, Sid84, Sin81, Sin82, Sin87, Ski83, SG80, Smi87, Smi88, St.87, St.89, Sta88, Ste89a, Ste89b, SH84a, SH84b, Swa81a, Swa80, Syl81, Syl87, Syl88, Tay83, Ter89, Thi88, Tho86, Tho88, Tho89, Tuc82, Tuc84, Tup89, Tur89, Vin89, Voe89, Wal89, War87a, War87b, War88, Was89, Wat82, Wat84, Web83, Wei87, Wei90, Wel89, Win87, Woo86, Yan89, Yfa87, You80, Zah89, Zai80, Zai87, Zie84, Zie85a, Zie86c, Zie86a, Zie86b, Zie87c, Zie87b, Zie87e, Zie87h, Zie87g, Zie87f, Zie87d, Zie87a, Zie88e, Zie88b, Zie88i, Zie88o, Zie88l, Zie88y, Zie88m]. Review [Zie88p, Zie88t, Zie88v, Zie88n, Zie88s, Zie88u, Zie88x, Zie88r, Zie88q, Zie88z, Zie88w, Zie88-27, Zie88f, Zie88d, Zie88a, Zie88j, Zie88c, Zie88g, Zie88h, Zie88-28, Zie89d, Zie89h, Zie89i, Zie89b, Zie89c, Zie89q, Zie89u, Zie89p, Zie89r, Zie89s, Zie89e, Zie89a, Zie89f, Zie89w, Zie89m, Zie89k, Zie89l, Zie89v, Zie89n, Zie89o, Zie89t, Zie89g, Zie89j, vT88, Mee82]. Reviews [Bow83, Car87a, Hei80b, Jac86b, Low82, Sin80, Spr83, Sto80, Sys82, Zie88k, Zie89x]. Revised [Gun83]. Rezzi [Zie88n]. Rice [Rab89]. Richard [Bea83, Bor86, Dal87, Dun89, Gun85c, Har89b, Hau89, Hix87, Jon89b, Mee87a, Nac87a, Obr82b, Sch84d, Smi88, St.87, War87a, War88]. Ridge [CCMC86, DSH80, EL81, GM84, Hoe87, PR82, WB88a, WB88b, EL83]. Riedwyl [Sch84b]. Riley [Zie88q]. Ripley [Boo88]. Risk [Ber82b, Bru86, Hei80a, HI87, Eas80a]. Risks [Eas80b]. Rizvi [Owe88, Owe89]. Road [PBT85]. Robb [Gle84]. Robert [Bor89, Cha89, Dal87, Gra87a, Hah81, Jac80b, Pri85, Ran82, Rus88, Sch84e, Sch84c, Zie85a, Zie89s, Zie89m]. Robinson [Aro86b, Sid82]. Robot [Lam84b]. Robotics [Lam84b]. Robust [AM80, CR85, CR87, CR80, CN82, Dav79, DL87, Fel86, GM86, GM87, Len81b, Pie90, RDR82, Roc89, Pie89, Por82b, Roc87b, Rup87]. Robustness [Che85, GK80b, Jac80b, Zie89p]. Roger [Caf87, Hsi87, Sal89]. Roland [Han84b, Zie88i]. Role [Lam84c]. Romano [Zie88o]. Ronald [Bat88, Bor87, Loc87, Low82, Scl81b, Smi87, Ste89a, Zie86c]. Ronchetti [Rup87]. Root [WG79]. Roots [Cor80]. Rosander [Lew87]. Rosiński [Fed81f]. Roskam [Jac81]. Ross [Bem89, Rey88]. Rotatability [DG88, Khu88]. Rothenberg [Low82]. Rousseeuw [Pie89, Rup87]. Rowe [Eas80a]. Roxy [Kot88]. Rubin [Gho88, Zie88e]. Rubinstein [Ang88, Sei82]. Rule [WD81]. Rules [CW87a, Ham85]. Run [Cro87, VoR85, Wal86, Woo83, Zie87h]. Run-Length [Cro87]. Runs [CW87a]. Rupert [Ams87, Klo82]. Ruppert [Wei90]. Rustagi [Jac80a, Mee85].

S [Aro85c, Bas88, Car87a, Coh82, Col88b, Cru82, Dea82, Dud82, Far88a, Fed81i, Fed81b, Gbu83, Gok83, Gri89, Hod89, Hor82, Hub89, Ize88, Jac80a, Jac80c, Jac85a, Luc81, Mee85, Mor84, Nai88a, Per88, Pri81, Pro87a, Ree88, Roh83a, Sch81a, Sch81b, Sch84g, Sch84a, Sch85e, Scl81a, She81a, Ski83, Syl87, Tuc82, Tur89, Woo86, Zie87c, Zie87b, Zie87f, Zie88y, Zie89w, Zie89k, Zie89x, Zie89v]. Sachs [Sch85b, Zie88l]. Safeguards [AL82, Fer82, GPS82a, GPS82b, Jae82, LD82, Mog82, Moo82, Mor82]. Sakamoto [Bro89]. Saleh [Fed84d]. Salsburg [Sch84g]. Sam [Mar88b]. Same [BM87a, EJG84]. Sample [Bai80b, CLS86, CR80, DSH80, Koz80, MH80, Nel83, New80, OCO89, Set84, SB83, Ste87, Tro80, WG79, Ste90, Sys82]. Sample-size [OCO89]. Samples [BS83, Car80, CO86, EB82, FMM80, Mee89a, Mor88a, OS88, Pie87, SGN85, Sin83, Ven82, Sin92, Min89]. Sampling [Ano85e, Ber82a, Bon81, Car87a, ETM80, EM82, FM80a, FM80b, Hah81, HH81, KN84, Kre87, NW89, NA88, PS87, Per84, RAAN88, Sch84c, Sch89b, She81a, She81b, SY87, Ste87, WLP85, Yan83, Ste90, Car87a, Rao80, War87a, Bil84, Bil88a, Bil88b, Hor82, Por82a]. Samprit [Maz89]. Samuel [Col88b, Gbu83, Har83, Sch85c, Sch86b]. Sanders [Sin80]. Sanford [Bec87, Ker85, Lin84]. Sarndal [Rao80]. SAS [Zie89x, Dow88, Mar88a, Mas88]. Satellite [Chh86]. Saul [Zie87d]. Saunders [EBW81]. Saving [GK80a]. Scale [GM86, KE80, Mor87, Rob83, SB83, SS87]. Scales [Hai87b, Hsi86]. Scaling [Gok83, Jac85f, Jac85d]. Scattering [BJ87]. Scatterplots [BC87, FT84]. Scheaffer [Har89b, Mee87a, War87a, Jon89b]. Scheduling [Fed82a, Fed82a]. Scheffé [LRS82]. Scheines [Nyc89]. Scheme [Cro86b, Cro89]. Schemes [Cro88b, LC82, VoR85, Wal86, Bon81]. Schilling [Per84]. Schneider [Mor88a]. Schonberger [Gun85c]. Schueler [Ang89]. Schultz [Fed80d]. Schwager [Zie88y]. Schwartz [Nac89]. Science [Fed81j, Gun85a, Mee85, Zie88v, Zie89c, Fed80d, Fed81b, Gbu81, Hel87, Rig88, Sha84]. Sciences [Ano86e, Fed80d, Fed81l, Har83, Pri81, Sch85c, Sch86b, Zie87c, Zie87b, Bow83, Hei80b, Boy80, Der89, Gol88, Han85b, Hei80b, Mul89, Mye80, Par82, Syl81, Zie88m, Zie89f]. Scientific [Joh82, Zie87d, Zie89a, Fed84c, Zie89q]. Scientists [Cro88a, Koc81, Pro89a, Rey88]. Sclove [Ree88]. Score [Che85]. Scores [MS89]. Screening [Mad82, MS82, Mau84, OL81, Tan87, WMS85, Hah81]. Seal [Hei80a]. search [Wel82]. Searches [Ing81]. Searle [Zie89x, Sch85e, Tur89, Zie88y]. Seber [Cro80, Jac86b]. Second [Dra82, Dra83a, Fed80b, Gun83, MM82b, VDY81, Zie88k, Zie89n]. Second-Order [Dra82, Dra83a]. Sections [PBT85]. Segmented [Haw80]. Seismic [Sid82]. Select [Dud80]. Selected [J.80a, Hea89, Khu87, Lin81b, Spr83]. Selecting [Dud82, Fra85, QK82, SKC83, You82]. Selection [Ano80g, AB85, BHP84, Bjø80, BME85, Dud80, GH80, Hin80, HSH86, LD80, Mad82, McD79, NW79, Owe88, SS80, WD81, Lit88, Owe89]. Selections [Lee80c]. Self [HMW80, Low82]. Self-Teaching [Low82]. Self-Tuning [HMW80]. Sen [Bas86, Joh83, Ran87]. Sensitivity [Ang88, Maz89]. Sequential [Kre87, MF83, Mor87, Joh83, Sch87b]. Sequentially [VDY81]. Serfling [Ran82]. Series [Abr81, AC89, Aro84a, Aro84c, Aro84d, Aro85b, Aro86b, Aro86a, Boa85b, Bow87, Cam83, CTC88, Dav79, Dic89, Fed80c, Fed84e, Hau87, Hau89, Ize89, Jon80, Ked87, LA81, MWF88, Maz80, Mil81, MK84, Par83, SD85, Sid82, Sid84, Win87, Aro85c, Dow88, Hea82, Pre83b, She81a, Swa80, Wel89, Gra87a]. Server [Mar83]. Service [Lew87]. Set [CJ86, Roh80]. Sets [Han85b, HBK84, Sal80, Lan87]. Several [HBK84, Lee80a, Mar85, Sch86d]. Shanti [Dud82]. Shape [SB83, SBE88]. Shaped [Hjo80]. Shapes [Tho86]. Shapiro [Gbu83, Col88b]. Shaw [Gri88]. Sheldon [Bow83, Rey88]. Shelf [Ano84e, CJJ81]. Shewhart [CW87a]. Shifted [Atk83]. Shifts [Dav79]. Shimizu [Job89]. Shinozuka [Ang89]. Shinzo [Kal83]. Shirley [Maz87]. Shock [BR88, BR89]. Shorack [Hod87]. Short [DV85]. Short-Cut [DV85]. Should [HW86]. Shrinkage [Cop83, EL81, EL83]. Shrinking [Oma82]. Shubert [Par82]. Side [Gun84]. Sided [BT89, CI88, CO86, Cro86b, MB88, San85, Tan87, Wal86, Woo83, Woo84, BT98, Cro89, DM87, DM88, OCO89]. Siegel [She89, Zie88o]. Siegmund [Sch87b]. Signal [Box88a, Box88b, CR88, Gun88b, Len81a, NP88a, Par88b, Sch89a, STL88, Ull88, LSK87a]. Signal-to-Noise [Box88a, Box88b, CR88, Gun88b, NP88a, Par88b, STL88, Ull88, LSK87a]. Signals [Ize87, Sid82]. Significance [CP81, WD81]. Silverman [Deh87a]. Silvey [Dea82]. Simo [Zie89n]. Simple [BE81, Cro87, Ket83, Koe83, Pro87b, SH83, SC82, YY87]. Simplicity [Gun85c]. Simplified [Lee89a]. Simulated [BJS86, FF89]. Simulation [Ang88, BM87a, DeP88a, HSH86, Tho88, Boo88, Har89b, Hod89, Jol88, Mor84, Har89a, Sei82]. Simulations [Ste87, Ste90]. Simultaneous [Bai80b, CO86, Dal83, KC81, RHM87, Spu81]. Sincich [Der89, Ful89]. Singer [J.80a]. Single [GH81, Mar83, Sch86d, She81b]. Singly [SGN85]. Sinha [Sch81a, Bas88]. Siotani [Goo87]. Size [Car80, Lin81b, She81b, OCO89, Set84]. Skewed [Boo87]. Skip [Car87a]. Skip-Lot [Car87a]. Slevin [Fed80d]. Slope [Par87]. Small [Ano90b, Dra85, Nel83, SS89, Tro80]. Small-Disturbance [Ano90b, SS89]. Smirnov [Sal80]. Smith [Ber89b, Ber83, Hah82, McL89, Sch85f]. Smoothed [SK85]. Smoothing [FT84, FS89a, FS89b, Has89, Haw89, MO86, Yan89]. Snedecor [Cor81]. Snee [Han86]. Snell [Boe83, Zie88b]. Social [Boy80, Min86]. Societatis [Fed82b]. Society [Fed84a]. Software [Maz88, Mee82, Cru84, Gos86, Zie89e]. Soil [dVS89]. Soil-Water [dVS89]. Solids [Bil88a]. Solomon [Odo80, Sid80]. Solution [BR87, Mag80]. Some [AD83, Che89, CB80, EL81, EL83, GCR86, Ham85, JN83, Kan87a, Kat81, Sal80, Wei83, YSOE84, Har80]. Something [Far88a]. Sometimes [WTA80]. Sommer [Bil88b]. Sorted [Mea89]. Sorting [GS82]. Source [Len81a]. Sources [HR88]. Space [DP81, GK80a, PD80a, PD80b, Wel89, Sys82]. Space-Saving [GK80a]. Space-Time [DP81, PD80a, PD80b]. Spaces [JN83]. Spain [Ber83]. Späth [Lin81a]. Spatial [Cre87]. SPC [Lee89a]. Special [BW85a, BW85b, Lam84b, LD80, McL85, Sid82, Yan85, Zie85b]. Specification [Ing81]. Spector [Mar88a]. Spectra [SMV82]. Spectral [BLSF88, Ber89b, Hur86, New82, Tuc82, Pre83b]. Spherical [Bin89, Ste89b]. Spirer [Igl88]. Spline [LRS82, Yan89]. Splines [Weg81]. Split [MO86]. Springer [Nel81]. Sprites [Nyc89]. SPSS [Par88c, Gib88, vT88]. SPSS-X [Gib88, vT88]. SPSSX [Zie89x]. Spurious [GK80b]. Spyros [Cha89]. Square [Sch89a]. Squared [BD84, SK84]. Squares [BW81, DS87, GZ79, HMS81, ÖD85, Rub80, SMV82]. Srivastava [Sch84a, Fed81a]. Stable [Sys82]. Stage [BSA81, Che87, Fai88, MS82, Mau84, PD80b]. Stahel [Rup87]. Stainton [Sch81b]. Standard [Cor80, Gal88, Kat82, MD80, VIC89]. Stanley [Agr89, Hor82]. Start [LC82]. STAT [Sch82]. State [Ano80g, McD79, Sys82, Wel89]. Stationary [BA89, Sid82, Sys82]. Statistic [FM80a, Lit85, MTL82, Ski80, SM81, Wei81]. Statistical [AL82, Ano86e, Asc83, Bat88, Bea83, Bir83, Bor89, Cha89, Col88a, Cru84, DZ88, Dun89, Ell82, Eng83, Ens80, Eva83, Far88a, Fed81g, Fed81k, Fed84a, Fer82, Gar85, Gle84, GPS82a, GPS82b, Goo87, Gos86, GS82, Gun84, Gun85a, Har83, Har80, Has88, Jac86b, Jae82, Joh83, Jol88, Kle82, Law83a, Law83b, LD82, Liu89, Lor80, Mat80, Maz88, Mee83, Min86, Mog82, Moo82, Mor82, Owe88, Owe89, Per82, PBT85, Pie88a, Pry89, Que88, Ran86, Ree88, RG86, San89, Sch88, Sch83c, Sch81b, Sch85c, Sch86c, Sch86b, Sch86a, Ski83, Smi87, St.87, Ste89a, Tho83, Woo86, You80, Zie87c, Zie87b, Zie87g, Zie88l, Zie88d, Zie89p, Zie89e, Zie89w, Zie89l, And86, Bai80a, Ber82c, Bin89, Bry89, Cal88, Cor81, Cru82, Edd87, Gbu83, Gho88, Gri88, Hac80, Hor86]. Statistical [Hsi87, Joh82, Koc81, Mar88b, Maz87, Mee82, Mie82, Min87, Mul89, Per88, Rea89, Ree81, Rel82, Rig88, Sal86, Sal89, Sch83b, Sch84f, Sch89d, Sin82, Ste89b, Thi88, Tho88, War88, Zie87e, Zie88y, Zie88g, Zie88h, Zie89r, Zie89s]. Statisticians [Aro85d, Pri85, Sch84g]. Statistics [Aro85b, Ber83, Boe83, Box84, Boy80, Cro88a, Deh87a, Fed81j, Fed81f, Fed81k, Fed81c, Fed81e, Fed84a, Fed84d, Fed84b, Gbu81, Gib88, Gol88, Gru88b, Han89, Hau87, Hea89, Hei80b, Igl88, Ker85, Khu87, Koz80, KE80, Lin81b, Mee87a, Mee85, Mee86, Mye80, Pen81, Pro89a, Rab89, Ran87, Rey88, Rup87, Sch82, Sch84d, Sch85b, Sch85g, Scl81a, Smi88, Spr83, Sto80, Syl81, Tup89, Wat82, Zie88b, Zie88c, Zie89h, Zie89c, Zie89m, Zie89j, vT88, Sin80, Ams87, Bro89, Bul89, Dal87, Far88b, Fed80b, Fed82c, Gru88a, Gun87, Han88, Har89a, Hod87, Hub89, Jac80b, Jac80a, Low82, Low83a, Mac83, MM82b, Min88, Obr81, Obr82a, Pig88, Por82b, Pro88b, Ran82, Rus88, Sav83, Sch84a, Sch85d, Sch85e, Scl81b, Sin80, Sto80, Syl88, Ter89, Zah89, Zie86a, Zie88k, Zie88o, Zie88n]. Statistics [Zie88s, Zie88u, Zie88x, Zie88j, Zie89i, Zie89b, Zie89p, Zie89k, Zie89n, Bor86, Bow83, Der89, Fed84d, Ful89, Gro89, Hix87, Kot88, Lam84c, Lan88, McL88, Obr82b, Pro87a, Pro88a, Roh83b, She87, She89, Shu87, Zie87f, Zie88i, Zie89u, Zie89t]. Statistik [Sch84b]. Statistiques [Pil83b]. Statlib [Ell82]. STD [JT83]. Stein [JO86]. Stepanek [Yfa87]. Stephen [Agr89, Zie86a, Zie88j]. Stephens [Car87a, Syl87, Kin87, She81a]. Stepping [Ano90a, Sim88a]. Steps [Lee89a]. Stepwise [RL80]. Steven [Gup82]. Stewart [Ano82f, Min87]. Steyn [Gri89]. Stigler [Zie88j]. Stochastic [Bac80, Cle80, Coh81, Cop83, EL81, Fed81h, Fed81b, Hor80a, Hor80b, Kim82, MW80b, MW80a, MHW84, Roh83a, Sta88, Tuc84, EL83, Ang89, Boo88, Jon88, Kal83, Roh81, Tay83, Zie88z]. Stone [Han89]. Stoodley [Sch81b].

Stopping [WD81]. Straight [Igl88]. Stratified [Mee89a]. Strength [GJBR88]. Stress [GJBR88, Sch86d]. Stress-Strength [GJBR88]. Stringer [Min87]. Structural [Ang89, Kad84]. Structure [BT82, Nyc89]. Structured [Næs85]. Stuart [Zie89h]. Student [Sch87a, Zie87a, BJ87]. Student- [BJ87]. Studentized [FM80a]. Students [Scl81a]. Studies [BME85, Fed80d, Fed81l, Han85b, KL88a, KL88b, Nai88b, Pri81, Rob88]. Study [Ano84e, CJJ81, Nai84, RPL81, Tit80b]. Studying [Cro87]. Stumpf [Gri89]. Subash [Koc81]. Subhypotheses [MJ81]. Subject [MHW84, Pie87]. Subjects [Scl81a]. Subset [Bjø80, Far85, GH80, HSH86, RP80]. Subsets [All84, GL84a, GL84b, HLG85, Hoc84, Wei84]. Subspaces [Oma82]. Substituent [BME85]. Subsystems [MWF88, Maz80]. Successive [NW89]. Suitable [MS89]. Sum [Cro86b, Cro88b, NW77, NW79, NW85, Par88a, Yas89, Cro89]. Supplement [Smi88]. Supplementary [CW87a]. Surface [BD82b, CN89, DGH85, Dra82, Dra83a, DJ88, GJM89, Hai87b, Khu88, MKC89, Pro87c, SKC83]. Surfaces [CCMC86, Par87, St.89, Fre88]. Survey [Bai80a, Hai87b, Rao80, War87a]. Surveys [Min89, Por82a]. Survival [DV86, Nai84, Hei80a, Moe83]. Suspension [Tro80]. Swain [Bis81]. Sweeney [Pro88a]. Swift [Min89, Har89b]. Switch [WTA80]. Symposium [Fed80a, Fed80b, Fed81j, Fed82c, Fed84d, Zie89c]. Synthesis [Abr81]. Synthesizing [HJ86]. System [BE80, Dow88, Mar88a, Mas88, Maz80, MG84, McL89, Mor84, SS80, Sax81, Bis89, Sen89]. Systems [BW85a, BW85b, Eas85, GC85, MWF88, McL85, Miy82, MK84, Tit80a, Tor89, Tro80, TB80, Yan85, Zie85b, Czi89, Der88, Fed81h, Fed84c, Han84c, Ize87].

T [Bin89, Bow83, Far88b, Goo88, Hau89, Lee81, McL88, MM82b, Ree88, Sch81b, Sch86a, Sin82, St.87, Ste89b, Zie88k, Zie88z, Zie89g]. Table [McC87, Wil83]. Tables [Ano90a, BD82a, BT83, BH82, Cor80, DJ80, Fed84a, Fra84, Gan81, Hea89, Khu87, Lin81b, Sch84f, Sch84c, Sim88a, Spr83, Tho86, Gen80, Odo80, Hah81, Sch85f]. Tabulation [BSMB82]. Taguchi [Pie90, AW89, Bis89, DZ88, LSK87a, Par89, Sen89, Bem89]. Taillie [Joh82]. Takes [Pie88b]. Takesi [Goo87]. Takeuchi [Gup83]. Talk [Igl88]. Tamhane [Lit89]. Tampere [Zie89n]. Tan [Roc87b]. Tanis [Bor89]. Tapiero [Fed81i, Fed81b]. Tarmo [Zie89n]. Taylor [Zie89o]. Teaching [Fed82c, Low82, Zah89]. Technician [Pla88]. Technique [BH80, Mea89, Yas89]. Techniques [Bem89, Fed81h, Gri88, Gun85c, Har89b, HSH86, Hoe87, Min87, Bai80a, Gbu83, Kin87, Sch81b, Sch85b]. Technometrics [BR89, BT98, BM87b, Cro89, DM88, EL83, FHL81, Hsi80, HL84, IC89, Kit81, LM85, NP89, Sin92, Ste90, WB88a, Eas83c]. Tell [Pri85]. Tenth [Lam84b]. Ternary [GC85]. Terry [Der89, Ful89]. Test [BS83, Car80, Col88b, Fai88, FM80b, Gan81, GM86, Haw81, HM84, Hsi86, JT83, Kan87a, Khu85, LM80, MLG82, Mee84b, SB83, SBE88, Sin83, YY87, Hsi80, Sin92]. Testing [Agr86b, Ber82a, BB89, BJ86, DJ80, EJG84, Hsi81, Lee80b, Lik82, MJ81, Maz88, Maz80, McC86, Mor87, MTL82, Mul83, Nac89, Nai86a, Nai86b, Nel88, OS88, Sch81a, Sch85c, Sch86d, SKC83, SY87, Wor83, YSOE84, Bas88, Zie87g]. TESTRAND [Mul83]. Tests [ASVR85, Ano84e, Atk86, Che85, CJJ81, DG87, Khu86, LLS80, MF83, Mee87b, Sch84c, Sch87b, Sen80, SC82, SS87, WD81, Her89]. Tetrads [BH82]. Teukolsky [Zie87d]. Texas [Aro84a, Par83]. Texts [Jac85d]. Th [Liu89, Per82]. Their [Fed81a, Hub89]. Theorem [Ald81, Cop83]. Theorems [Ran82]. Theoretic [Yfa87]. Theoretical [Smi88, Zie88u]. Théorie [Pil83b]. Theory [Ano90b, Aro84c, Aro84d, Aro86b, Aro86a, BT81, Ber82c, Bil84, Boo84, Bow83, DeG81, Dea82, Dic89, Dud82, Fed84e, Gup82, Har80, J.80c, Job89, Roh80, Roh83a, Sin82, SS89, Zie88y, Zie89h, Zie89g, Sys82, DeP88b, Fed81f, Gle84, Han85a, Hei80a, Jon89a, Mat80, Zie89h, Han84a, Pro89c, Swa81a]. There [GCR86]. Think [Pro88b]. Thinking [Aro85a]. Thisted [Ste89a]. Thomas [Car87a, Dal87, Gru88b, Hor86, Jon89b, Law86c, Mul83, Pro88a]. Three [Bjø80, Cor80, DJ80, Dra83a, GH81, JG80, KE80, NA88, PD80b, Jac80d]. Three-Class [NA88]. Three-Factor [JG80]. Three-Level [Dra83a]. Three-Stage [PD80b]. Three-Way [GH81]. Tiao [Boa85a]. Tietjen [Gru88a]. Tigelaar [Set84]. Tijms [Jon88]. Tiku [Roc87b, Zie89p]. Time [Abr81, AC89, Aro84a, Aro85b, Aro85c, BE80, Boa85b, Cam83, CTC88, Dav79, DP81, Dow88, GK80a, Hau87, Ize89, Jon80, Ked87, Kle82, LA81, Mil81, PD80a, PD80b, Pre83b, SD85, Sid82, Sid84, Ste88, Swa80, Wel89, Zie88p, Aro84c, Aro84d, Aro86b, Aro86a, Bow87, Dic89, Fed80c, Fed84e, Hau89, Par83, Win87]. Time- [GK80a]. TIMESLAB [Ize89]. TIMS [Han85b, Pri81, Fed80d, Fed81l]. Tobias [Car88]. Toby [Sch86c, You80]. Today [Zie89u]. Toit [Gri89]. Tokyo [Fed81k]. Tolerance [BE81, BT89, Col88a, Hah81, JSOS85, Mee84a, Mee89a, Mil89, MB88, BT98, OCO89, Sax81]. Tolerancing [DZ88]. Tomkó [Fed82b]. Tool [All84, AB85, GL84a, GL84b, Hoc84, Wei84]. Tool-Wear [AB85]. Tools [Tho88]. Top [IC87, IC89]. Top-Down [IC87, IC89]. Topical [Gru88a]. Topics [BH84, Fed84d, J.80c]. Total [PC85, Pie88b]. Toutenburg [Cop83]. Trace [GM84, Ven82]. Traffic [Ano80g, McD79]. Transform [Ano84f, IC79]. Transformation [SAJ89, dVS89]. Transformations [Atk83, Atk86, Atk88, Bai80b, BD82b, Box88a, Box88b, CR85, CR88, CW83, Gun88b, HW88, NP88a, Par88a, Par88b, STL88, Tsa88, Ull88, Wei87, Wei90]. Transforming [CR87, OS88]. Transport [Fed82a, Fed84c]. Treatment [Hac80, HM84]. Treatments [BT81, BT83]. Trend [Hai87b]. Trend-Surface [Hai87b]. Trends [Ste88, Tho86]. Trevor [Min87]. Trewin [Lam84c]. Trials [JT83, KN84]. Trindade [Car88]. Trout [Han86]. Truncated [BE80, EN80, HZ80, Koz80, Mor88a]. Tufte [Car87b]. Tukey [Boa85b, Zie88-27, Zie88-28, Zie89v, Boa85b, Gan81, SG80, Tho86, Zie88-27]. Tuning [HMW80]. Twenty [Eas83c]. Twenty-Five [Eas83c]. Twice [EJG84]. Twisted [Igl88]. Two [Ano90a, Atk81, BT89, BT98, BSA81, BH82, Car80, Che87, Cro86b, DJ80, Fai88, FM80b, GH81, GH80, Hsi81, Koz80, KE80, MS82, Mau84, Mil80, MM83c, Pie87, SB83, Sim88a, Ski80, SC82, SMV82, Spr83, Wal86, WTA80, Woo84, Wor83, Cro89, Gen80, Hsi80, OCO89]. Two- [GH81]. Two-Level [MM83c]. Two-Parameter [FM80b, Mil80, Pie87]. Two-Phase [Wor83]. Two-Sample [SB83]. Two-Sided [BT89, Cro86b, Wal86, Woo84, BT98, Cro89, OCO89]. Two-Stage [BSA81, Che87, Fai88, MS82, Mau84]. Two-Way [Ano90a, BH82, DJ80, GH80, Sim88a, Ski80, WTA80, Gen80]. Type [Agr86a, Kao88, Sin92, SM81, Hsi86, Mee84b, Pie87, Sin83].

Ulf [Roh82]. Ullah [Web83]. Unbalanced [BG84, Cha87, SM81, Tur89, Zie89x]. Unbiased [Swa81b]. Uncertainty [DGH85]. Unconditional [SK84]. Understanding [Low82, Pig88, Pro89b, Sch84g, Woo86]. Unequal [Hsi86, SBE88]. Unified [Hau89, LV86, Wat84]. Units [Zie89a]. University [Fed80c, Fed81d]. Unknown [CA81, Khu86, San85, SB83, SBE88, SS87, SMV82]. Unmasked [Atk88]. Unreplicated [BM86a, Len89]. Unsuspected [TB80]. Upton [Cre87]. Uranium [CJ86]. Urban [Fed82a]. Use [Ano84f, Ano85e, CSS88, GS82, Hin80, IC79, Kit79, PG86, Pro87d, SM81, WLP85, YY87, Kit81]. Used [HW86, QK82, Ski80]. Useful [Gla82, HJ86, Sch85e]. User [Jac85f]. Uses [Bat83]. Using [BM87a, Boo84, BH82, CR80, Chh86, DZ88, FF89, HMS81, HBK84, LRS82, Mad82, Mag80, Maz87, MB88, NW79, Nel83, Nel88, New82, NG85, OL81, San85, Sch82, SGN85, Ste87, Tan87, Ste90]. Utilities [Fed81d].

V [Ber83, Bor89, Fed80a, Fed81l, Han88, Har87, Hei80b, Lee81, Odo80, Sch85f, Sch89d, Sha84, Yfa87, Zie88c]. Valencia [Ber83]. Validating [Ber84]. Validatory [EK82]. Value [Boo84, EB82, FM80a, FMM80, HWW85, Pie88b, Rus88, SH83]. Valued [Ald81]. Values [Cle80, Hor80a, Hor80b, MF83]. Variable [BD82b, Hin80, Jac89, JM87, RAAN88, Sar89, Tan87, Wan85]. Variables [BT82, CI83, DJ88, EJG84, GC82, GJBR88, Ket83, LC81, McC84, NW79, NA88, RPL81, Sch89b, ST85, SB89, VDY81, Wil83, WMS85, Woo83, Spr83, Ald81, Nel81]. Variables-Sampling [Sch89b]. Variance [BNC87, Bre89, BG84, Cha87, Dav79, Fel86, HZ80, Hid86, HGB89, JG80, Lik80, Swa81b, SM84, War87b, Zie88c, Zie89x, Fed81e]. Variance-Component [HGB89]. Variances [Ano84e, BB89, CJJ81, MH80, Swa81b]. Variate [Mad82, OL81]. Variation [CLS86, Hai87b, New80]. Various [HR88]. Vector [Weg81]. Vehicle [Fed82a, GH80, Lee80c, Lor80, MM82a, McD81]. Vertices [NGG83]. Vetterling [Zie87d]. VI [Lin81b, Zie86b]. Via [Ano90a, Sim88a]. Vic [Sch86c, You80]. View [Mee85]. VII [Spr83]. Vilaplana [Zie88u]. Vinod [Web83]. Virgil [Har89a, Zie85a]. Virginia [Jac85b]. Visual [Car87b]. Vol [Bis89, Fed80d, Fed81l, Fed81a, Fed82b, Han85b, Hea89, Khu87, Pri81, Sch85g, Sch85c, Sen89, She81a, Spr83, Zie88-27, Zie89h, Zie89v]. Vols [Ano86e]. Volume [Ano80p, Ano81m, Ano82o, Ano83m, Ano84n, Ano85l, Ano86j, Ano87i, Ano88i, Ano89i, Boa85b, Cre87, Fed81c, Fed81e, Fed84b, Har83, LRS82, Lin81b, Ran87, Sch85a, Sch86b, Was89, Zie87c, Zie87b, Car87a, Car87a, Col88b, Col88a, Kan87b, Kan87a, Pro87b, Pro87c, Pro87d, Zie87h]. Volumes [Zie88-28].

W [Asc89, Bea83, Bil88a, Boa85b, Car87a, Che83, Cor81, Cru84, Dea81, Der88, Dun89, Eas80b, Fed81f, Gbu81, Har86, Hea82, Hei80b, Hsi87, Hsi80, Lin81b, Lit88, Low82, Mar83, Min87, Obr81, Obr82a, Pen81, Ree88, Roc87b, Rup87, Sch84d, Sch86a, Sch89d, Sen80, SG80, Smi88, Spr83, Syl81, Tho86, Zah89, Zie88k, Zie88u, Zie88-27, Zie88-28, Zie89e, Zie89v, Zie89g]. Wackerly [Mee87a]. Wadsworth [Syl87]. Walk [AW89, Bin80, Wil80]. Walker [McL89]. Walker-Smith [McL89]. Walks [Sys82]. Waller [Sch83a]. Walpole [Scl81b]. Walsh [She88]. Warren [And86, Gun87]. Wasserman [Lam84a]. Watanabe [Kal83]. Water [Fed81i, dVS89]. Watkins [Hin87, Min89]. Watts [BSBW86]. Way [Ano90a, BH82, BG84, Cha87, CW87b, DJ80, GH81, GH80, Mee84a, Sim88a, Ski80, WTA80, Gen80]. Wayne [Dea89, Gos86, Hei80b, Kan87b, Pro87b, Zie88m, Zie88w]. Weaknesses [BH84, WLV86]. Wear [AB85]. Weather [Mie82]. Wegman [Mie82, Thi88]. Weibull [BR89, BE80, BE81, BR88, Cro82, EB82, FMM80, FM80b, Lee80a, Lee80b, Mee84b, New80, Sch89b]. Weighing [Bil88a]. Weight [Nel83]. Weighted [Cro87, GCR86, HWW85, NW79, Rub80, Yas89]. Weighting [Wei90]. Weights [GZ79]. Weinman [Hsi80]. Weisberg [Bec87, Lin84, Sin81]. Weiss [Low83a, Min88]. Welch [BR87]. Weldon [She87]. Wellner [Hod87]. Wentzell [Roh83a]. Wertz [Zie88u]. Wetherill [Liu89, Per82, Gun88a]. Wheeler [Woo86, Gri88, Pro89b]. Wheelwright [Pri81]. White [Ano80f]. Whitlock [Was89]. Wichern [Bea83, Dun89]. Wilcoxon [LD80]. Wilkinson [Jac80b]. Wilks [RL80]. William [Bis81, Bry89, Col87, Der89, Eas80a, Ell82, Far88a, Gbu81, Ize87, J.80d, Jac86b, Kan87a, Lam84a, Mee87a, Rel82, Spr83, Syl88, Zie87d, Zie89b, Zie89f, Box87]. Williams [Pro88a, Smi88]. Willian [Sav83]. Wilrich [Liu89, Per82]. Windows [Hur86]. Winkler [Aro85c, Cha89]. Winston [Zie88w]. Wintz [Per89]. Wishart [Cor80]. Wm [Klo82]. Wolter [Hid86]. Wonnacott [MM82b, Zie88k, MM82b]. Wood [Luc81]. Woodward [Gos86, Zie89e]. Woolson [Zie89s]. Work [Joh82]. Worker [Sch87a, Zie87a]. Works [Boa85b, Boa85a, Zie88-27, Zie88-28, Zie89v]. Workshop [Han84a]. World [Hah84a, Sto80]. Wren [Fed82a]. Wretman [Rao80]. Wrong [Zie87c]. Wurster [She88].

X [Gib88, vT88].

Yadolah [Roc87a]. Yager [Roh80]. Yanai [Gup83]. Yaron [Fed81i].

Yasunori [Goo87]. Yates [Por82a]. Years [Eas83c]. Yerex [Zie89x, Zie89x]. Youden [LD80]. Yourself [Sch82].

Z [Eas80b, Lee89c]. Zadeh [Han85b]. Zahn [Gun84]. Zanakis [Mee85]. Zedeck [Bow83]. Zero [Pie88b]. Zimmerman [Smi88]. Zimmermann [Han85b]. Zucchini [Lit88]. Zupan [Lan87]. Zuwaylif [Sin80]. Zverev [Zie88p].

References

Arcelus:1985:SME

[AB85] F. J. Arcelus and P. K. Banerjee. Selection of the most economical production plan in a tool-wear process. Technometrics, 27 (4):433–437, November 1985. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1270210.

Abraham:1981:BRB

[Abr81] Bovas Abraham. Book review: Analysis of Economic Time Series: A Synthesis by M. Nerlove; D. M. Grether; J. L. Carvalho. Technometrics, 23(1):114, February 1981. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267989.

Abraham:1989:ODT

[AC89] Bovas Abraham and Alice Chuang. Outlier detection and time series modeling. Technometrics, 31(2):241–248, May 1989. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268821.

Agrawal:1983:ORI

[AD83] Veena Agrawal and A. Dey. Orthogonal Resolution-IV designs for some asymmetrical factorials. Technometrics, 25(2):197–199, May 1983. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1268554.

Agresti:1986:ATM

[Agr86a] Alan Agresti. Applying R^{2}-type measures to ordered categorical data. Technometrics, 28(2):133–138, May 1986. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1270449.

Agresti:1986:TIE

[Agr86b] Alan Agresti. [Testing in Industrial Experiments with Ordered Categorical Data]: Discussion. Technometrics, 28(4):292–294, November 1986. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1268975. See

[Nai86a].

Agresti:1988:BRB

[Agr88] Alan Agresti. Book review: Applied Categorical Data Analysis by Daniel H. Freeman. Technometrics, 30(4):456, November 1988. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1269816.

Agreda:1989:BRB

[Agr89] Carla Agreda. Book review: Experimental Design: A Chemometric Approach by Stanley N. Deming; Stephen L. Morgan. Technometrics, 31(2):262–263, May 1989. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268831.

Aitkin:1981:NRA

[Ait81] Murray Aitkin. A note on the regression analysis of censored data. Technometrics, 23(2):161–163, May 1981. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1268032. See corrigenda [Ano81e].

Abramson:1982:SMN

[AL82] Lee R. Abramson and Dan Lurie. [Statistical Methods for Nuclear Materials Safeguards: An Overview]: Discussion. Technometrics, 24(4):280–281, November 1982. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267817. See [GPS82a].

Aldous:1981:BRB

[Ald81] David Aldous. Book review: The Central Limit Theorem for Real and Banach Valued Random Variables by A. Araujo; E. Giné. Technometrics, 23(3):307, August 1981. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267796.

Allen:1984:CDT

[All84] David M. Allen. [K-Clustering as a Detection Tool for Influential Subsets in Regression]: Discussion. Technometrics, 26(4):319–320, November 1984. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1269491.

See [GL84a].

Askin:1980:ARE

[AM80] Ronald G. Askin and Douglas C. Montgomery. Augmented robust estimators. Technometrics, 22(3):333–341, August 1980. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268317.

Amster:1987:BRB

[Ams87] S. J. Amster. Book review: Beyond ANOVA, Basics of Applied Statistics by Rupert G. Miller. Technometrics, 29(3):387, August 1987. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1269361.

Anderson:1986:BRB

[And86] Oliver D. Anderson. Book review: Statistical Modelling by Warren Gilchrist. Technometrics, 28(4):403, November 1986. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268994.

Angus:1988:BRB

[Ang88] John E. Angus. Book review: Monte Carlo Optimization, Simulation and Sensitivity of Queueing Networks by Reuven Y. Rubinstein. Technometrics, 30(4):465–467, November 1988. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1269828.

Angus:1989:BRB

[Ang89] John Angus. Book review: Stochastic Methods in Structural Dynamics by G. Schueler; M. Shinozuka. Technometrics, 31(3):395– 397, August 1989. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
3556170.

Anonymous:1980:BMa

[Ano80a] Anonymous. Back matter. Technometrics, 22(1):137–138, February 1980. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1268408.

Anonymous:1980:BMb

[Ano80b] Anonymous. Back matter. Technometrics, 22(2):??, May 1980. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1268483.

Anonymous:1980:BMc

[Ano80c] Anonymous. Back matter. Technometrics, 22(3):??, August 1980. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1268343.

Anonymous:1980:BMd

[Ano80d] Anonymous. Back matter. Technometrics, 22(4):??, November 1980. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1268214.

Anonymous:1980:CGJ

[Ano80e] Anonymous. Corrigenda to Gabriel, Jackson and Davis. Technometrics, 22(1):136, February 1980. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/ stable/1268407. See [GZ79, Jac80d, Dav79].

Anonymous:1980:CLW

[Ano80f] Anonymous. Corrigenda to Leone, White, Narula and Narula. Technometrics, 22(3):452, August 1980. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/ stable/1268342. See [LNN61, WG79, NW77, NW79].

Anonymous:1980:CNS

[Ano80g] Anonymous. Corrigendum: “Nonparametric Selection Procedures Applied to State Traffic Fatality Rates”. Technometrics, 22(2):285, May 1980. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1268482. See

[McD79].

Anonymous:1980:FMa

[Ano80h] Anonymous. Front matter. Technometrics, 22(1):??, February 1980. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1268375.

Anonymous:1980:FMb

[Ano80i] Anonymous. Front matter. Technometrics, 22(2):??, May 1980. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1268452.

Anonymous:1980:FMc

[Ano80j] Anonymous. Front matter. Technometrics, 22(3):451, August 1980. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1268312.

Anonymous:1980:FMd

[Ano80k] Anonymous. Front matter. Technometrics, 22(4):i–iv, November 1980. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1268179.

Anonymous:1980:PRa

[Ano80l] Anonymous. Publications received. Technometrics, 22(1):134, February 1980. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1268405.

Anonymous:1980:PRb

[Ano80m] Anonymous. Publications received. Technometrics, 22(2):282–283, May 1980. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1268480.

Anonymous:1980:PRc

[Ano80n] Anonymous. Publications received. Technometrics, 22(3):449, August 1980. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1268340.

Anonymous:1980:PRd

[Ano80o] Anonymous. Publications received. Technometrics, 22(4):635–636, November 1980. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1268213.

Anonymous:1980:VI

[Ano80p] Anonymous. Volume information. Technometrics, 22(4):637–640, November 1980. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1268178.

Anonymous:1981:BMa

[Ano81a] Anonymous. Back matter. Technometrics, 23(1):116–118, February 1981. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1267991.

Anonymous:1981:BMb

[Ano81b] Anonymous. Back matter. Technometrics, 23(2):??, May 1981. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1268060.

Anonymous:1981:BMc

[Ano81c] Anonymous. Back matter. Technometrics, 23(3):??, August 1981. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1267811.

Anonymous:1981:BMd

[Ano81d] Anonymous. Back matter. Technometrics, 23(4):432, November 1981. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1268246.

Anonymous:1981:CAH

[Ano81e] Anonymous. Corrigenda to Aitkin and Höhne. Technometrics, 23(4):427, November 1981. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268245. See [Ait81].

Anonymous:1981:FMa

[Ano81f] Anonymous. Front matter. Technometrics, 23(1):??, February 1981. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1267968.

Anonymous:1981:FMb

[Ano81g] Anonymous. Front matter. Technometrics, 23(2):??, May 1981. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1268026.

Anonymous:1981:FMc

[Ano81h] Anonymous. Front matter. Technometrics, 23(3):318–319, August 1981. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1267783.

Anonymous:1981:FMd

[Ano81i] Anonymous. Front matter. Technometrics, 23(4):??, November 1981. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1268216.

Anonymous:1981:PRa

[Ano81j] Anonymous. Publications received. Technometrics, 23(2):211–214, May 1981. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1268057.

Anonymous:1981:PRb

[Ano81k] Anonymous. Publications received. Technometrics, 23(3):315–316, August 1981. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1267808.

Anonymous:1981:PRc

[Ano81l] Anonymous. Publications received. Technometrics, 23(4):423–426, November 1981. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1268244.

Anonymous:1981:VI

[Ano81m] Anonymous. Volume information. Technometrics, 23(4):428–431, November 1981. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1268215.

Anonymous:1982:BMa

[Ano82a] Anonymous. Back matter. Technometrics, 24(1):92–94, February 1982. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1267596.

Anonymous:1982:BMb

[Ano82b] Anonymous. Back matter. Technometrics, 24(2):??, May 1982. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1268509.

Anonymous:1982:BMc

[Ano82c] Anonymous. Back matter. Technometrics, 24(3):263–266, August 1982. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1268699.

Anonymous:1982:BMd

[Ano82d] Anonymous. Back matter. Technometrics, 24(4):351, November 1982. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1267843.

Anonymous:1982:CBK

[Ano82e] Anonymous. Corrigenda to Bechhofer and Kent. Technometrics, 24(2):171, May 1982. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/ stable/1268508. See [BT81, QK82].

Anonymous:1982:CNS

[Ano82f] Anonymous. Corrigenda to Newby and Stewart. Technometrics, 24(1):90–91, February 1982. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/ stable/1267595. See [New80, SS81].

Anonymous:1982:FMa

[Ano82g] Anonymous. Front matter. Technometrics, 24(1):??, February 1982. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1267571.

Anonymous:1982:FMb

[Ano82h] Anonymous. Front matter. Technometrics, 24(2):??, May 1982. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1268484.

Anonymous:1982:FMc

[Ano82i] Anonymous. Front matter. Technometrics, 24(3):??, August 1982. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1268675.

Anonymous:1982:FMd

[Ano82j] Anonymous. Front matter. Technometrics, 24(4):??, November 1982. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1267813.

Anonymous:1982:PRa

[Ano82k] Anonymous. Publications received. Technometrics, 24(1):87–89, February 1982. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1267594.

Anonymous:1982:PRb

[Ano82l] Anonymous. Publications received. Technometrics, 24(2):169–170, May 1982. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1268507.

Anonymous:1982:PRc

[Ano82m] Anonymous. Publications received. Technometrics, 24(3):255–257, August 1982. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1268697.

Anonymous:1982:PRd

[Ano82n] Anonymous. Publications received. Technometrics, 24(4):343–344, November 1982. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1267841.

Anonymous:1982:VI

[Ano82o] Anonymous. Volume information. Technometrics, 24(4):347–350, November 1982. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1267812.

Anonymous:1983:BMa

[Ano83a] Anonymous. Back matter. Technometrics, 25(1):118, February 1983. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1267746.

Anonymous:1983:BMb

[Ano83b] Anonymous. Back matter. Technometrics, 25(2):??, May 1983. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1268573.

Anonymous:1983:BMc

[Ano83c] Anonymous. Back matter. Technometrics, 25(3):??, August 1983. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1268618.

Anonymous:1983:BMd

[Ano83d] Anonymous. Back matter. Technometrics, 25(4):391–398, November 1983. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1267873.

Anonymous:1983:COR

[Ano83e] Anonymous. Corrigendum: [Outlier. . . . . . . . . s]: Response. Technometrics, 25(4):390, November 1983. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267872. See [BC83b].

Anonymous:1983:FMa

[Ano83f] Anonymous. Front matter. Technometrics, 25(1):??, February 1983. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1267719.

Anonymous:1983:FMb

[Ano83g] Anonymous. Front matter. Technometrics, 25(2):i–ii, May 1983. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1268539.

Anonymous:1983:FMc

[Ano83h] Anonymous. Front matter. Technometrics, 25(3):302, August 1983. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1268602.

Anonymous:1983:FMd

[Ano83i] Anonymous. Front matter. Technometrics, 25(4):??, November 1983. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1267845.

Anonymous:1983:PRa

[Ano83j] Anonymous. Publications received. Technometrics, 25(1):116, February 1983. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1267745.

Anonymous:1983:PRb

[Ano83k] Anonymous. Publications received. Technometrics, 25(2):216, May 1983. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1268571.

Anonymous:1983:PRc

[Ano83l] Anonymous. Publications received. Technometrics, 25(4):389, November 1983. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1267871.

Anonymous:1983:VI

[Ano83m] Anonymous. Volume information. Technometrics, 25(4):393–396, November 1983. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1267844.

Anonymous:1984:BMa

[Ano84a] Anonymous. Back matter. Technometrics, 26(1):70, February 1984. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1268423.

Anonymous:1984:BMb

[Ano84b] Anonymous. Back matter. Technometrics, 26(2):196, May 1984. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1268126.

Anonymous:1984:BMc

[Ano84c] Anonymous. Back matter. Technometrics, 26(3):303, August 1984. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1267568.

Anonymous:1984:BMd

[Ano84d] Anonymous. Back matter. Technometrics, 26(4):421–428, November 1984. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1269515.

Anonymous:1984:CCS

[Ano84e] Anonymous. Corrigendum: “A Comparative Study of Tests for Homogeneity of Variances, with Applications to the Outer Continental Shelf Bidding Data”. Technometrics, 26(3):302, August 1984. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1267567. See [CJJ81].

Anonymous:1984:CUR

[Ano84f] Anonymous. Corrigendum: “The Use of the Rank Transform in Regression”. Technometrics, 26(4):419, November 1984. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1269514. See [IC79].

Anonymous:1984:FMa

[Ano84g] Anonymous. Front matter. Technometrics, 26(1):??, February 1984. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1268409.

Anonymous:1984:FMb

[Ano84h] Anonymous. Front matter. Technometrics, 26(2):??, May 1984. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1268096.

Anonymous:1984:FMc

[Ano84i] Anonymous. Front matter. Technometrics, 26(3):??, August 1984. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1267544.

Anonymous:1984:FMd

[Ano84j] Anonymous. Front matter. Technometrics, 26(4):??, November 1984. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1269489.

Anonymous:1984:PRa

[Ano84k] Anonymous. Publications received. Technometrics, 26(1):67, February 1984. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1268421.

Anonymous:1984:PRb

[Ano84l] Anonymous. Publications received. Technometrics, 26(2):195, May 1984. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1268125.

Anonymous:1984:PRc

[Ano84m] Anonymous. Publications received. Technometrics, 26(4):418, November 1984. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1269513.

Anonymous:1984:VI

[Ano84n] Anonymous. Volume information. Technometrics, 26(4):423–426, November 1984. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1269488.

Anonymous:1985:BMa

[Ano85a] Anonymous. Back matter. Technometrics, 27(1):??, February 1985. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1270483.

Anonymous:1985:BMb

[Ano85b] Anonymous. Back matter. Technometrics, 27(2):??, May 1985. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1268777.

Anonymous:1985:BMc

[Ano85c] Anonymous. Back matter. Technometrics, 27(3):328, August 1985. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1269726.

Anonymous:1985:BMd

[Ano85d] Anonymous. Back matter. Technometrics, 27(4):447–454, November 1985. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1270218.

Anonymous:1985:CUC

[Ano85e] Anonymous. Corrections: “Use of Curtailed Sampling Plans in the Economic Design of np-Control Charts”. Technometrics, 27(3):327, August 1985. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1269725.

See [WLP85].

Anonymous:1985:FMa

[Ano85f] Anonymous. Front matter. Technometrics, 27(1):??, February 1985. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1270462.

Anonymous:1985:FMb

[Ano85g] Anonymous. Front matter. Technometrics, 27(2):??, May 1985. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1268756.

Anonymous:1985:FMc

[Ano85h] Anonymous. Front matter. Technometrics, 27(3):??, August 1985. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1269701.

Anonymous:1985:FMd

[Ano85i] Anonymous. Front matter. Technometrics, 27(4):??, November 1985. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1270197.

Anonymous:1985:PRa

[Ano85j] Anonymous. Publications received. Technometrics, 27(1):90–91, February 1985. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1270482.

Anonymous:1985:PRb

[Ano85k] Anonymous. Publications received. Technometrics, 27(2):211, May 1985. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1268776.

Anonymous:1985:VI

[Ano85l] Anonymous. Volume information. Technometrics, 27(4):449–451, November 1985. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1270196.

Anonymous:1986:BMa

[Ano86a] Anonymous. Back matter. Technometrics, 28(1):94, February 1986. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1269613.

Anonymous:1986:BMb

[Ano86b] Anonymous. Back matter. Technometrics, 28(2):??, May 1986. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1270461.

Anonymous:1986:BMc

[Ano86c] Anonymous. Back matter. Technometrics, 28(3):??, August 1986. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1269090.

Anonymous:1986:BMd

[Ano86d] Anonymous. Back matter. Technometrics, 28(4):411–418, November 1986. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1269001.

Anonymous:1986:CES

[Ano86e] Anonymous. Correction: “Encyclopedia of Statistical Sciences (Vols. 3 and 5)”. Technometrics, 28(3):280, August 1986. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1269089. See [Sch85c, Sch86b].

Anonymous:1986:FMa

[Ano86f] Anonymous. Front matter. Technometrics, 28(1):??, February 1986. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1269596.

Anonymous:1986:FMb

[Ano86g] Anonymous. Front matter. Technometrics, 28(2):??, May 1986. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1270444.

Anonymous:1986:FMc

[Ano86h] Anonymous. Front matter. Technometrics, 28(3):??, August 1986. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1269073.

Anonymous:1986:FMd

[Ano86i] Anonymous. Front matter. Technometrics, 28(4):??, November 1986. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1268972.

Anonymous:1986:VI

[Ano86j] Anonymous. Volume information. Technometrics, 28(4):413–415, November 1986. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1268971.

Anonymous:1987:BMa

[Ano87a] Anonymous. Back matter. Technometrics, 29(1):126, February 1987. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1269903.

Anonymous:1987:BMb

[Ano87b] Anonymous. Back matter. Technometrics, 29(2):??, May 1987. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1269794.

Anonymous:1987:BMc

[Ano87c] Anonymous. Back matter. Technometrics, 29(3):392, August 1987. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1269366.

Anonymous:1987:BMd

[Ano87d] Anonymous. Back matter. Technometrics, 29(4):507–515, November 1987. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1269487.

Anonymous:1987:FMa

[Ano87e] Anonymous. Front matter. Technometrics, 29(1):??, February 1987. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1269877.

Anonymous:1987:FMb

[Ano87f] Anonymous. Front matter. Technometrics, 29(2):??, May 1987. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1269767.

Anonymous:1987:FMc

[Ano87g] Anonymous. Front matter. Technometrics, 29(3):??, August 1987. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1269329.

Anonymous:1987:FMd

[Ano87h] Anonymous. Front matter. Technometrics, 29(4):487, November 1987. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1269448.

Anonymous:1987:VI

[Ano87i] Anonymous. Volume information. Technometrics, 29(4):510–513, November 1987. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1269447.

Anonymous:1988:BMa

[Ano88a] Anonymous. Back matter. Technometrics, 30(1):??, February 1988. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1270356.

Anonymous:1988:BMb

[Ano88b] Anonymous. Back matter. Technometrics, 30(2):??, May 1988. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1270195.

Anonymous:1988:BMc

[Ano88c] Anonymous. Back matter. Technometrics, 30(3):364, August 1988. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1270110.

Anonymous:1988:BMd

[Ano88d] Anonymous. Back matter. Technometrics, 30(4):471–472, November 1988. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1269833.

Anonymous:1988:FMa

[Ano88e] Anonymous. Front matter. Technometrics, 30(1):113, February 1988. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1270310.

Anonymous:1988:FMb

[Ano88f] Anonymous. Front matter. Technometrics, 30(2):??, May 1988. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1270157.

Anonymous:1988:FMc

[Ano88g] Anonymous. Front matter. Technometrics, 30(3):??, August 1988. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1270078.

Anonymous:1988:FMd

[Ano88h] Anonymous. Front matter. Technometrics, 30(4):??, November 1988. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1269796.

Anonymous:1988:VI

[Ano88i] Anonymous. Volume information. Technometrics, 30(4):473–477, November 1988. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1269795.

Anonymous:1989:BMa

[Ano89a] Anonymous. Back matter. Technometrics, 31(1):135–136, February 1989. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1270397.

Anonymous:1989:BMb

[Ano89b] Anonymous. Back matter. Technometrics, 31(2):??, May 1989. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1268851.

Anonymous:1989:BMc

[Ano89c] Anonymous. Back matter. Technometrics, 31(3):??, August 1989. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/3556176.

Anonymous:1989:BMd

[Ano89d] Anonymous. Back matter. Technometrics, 31(4):504–510, November 1989. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1270026.

Anonymous:1989:FMa

[Ano89e] Anonymous. Front matter. Technometrics, 31(1):111, February 1989. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1270357.

Anonymous:1989:FMb

[Ano89f] Anonymous. Front matter. Technometrics, 31(2):253, May 1989. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1268812.

Anonymous:1989:FMc

[Ano89g] Anonymous. Front matter. Technometrics, 31(3):379, August 1989. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/3556139.

Anonymous:1989:FMd

[Ano89h] Anonymous. Front matter. Technometrics, 31(4):479, November 1989. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1269990.

Anonymous:1989:VI

[Ano89i] Anonymous. Volume information. Technometrics, 31(4):506–510, November 1989. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1269989.

Anonymous:1990:CDO

[Ano90a] Anonymous. Correction: “Detecting Outlying Cells in Two-Way Contingency Tables Via Backwards-Stepping”. Technometrics, 32 (1):115, February 1990. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1269875. See [Sim88a].

Anonymous:1990:CSD

[Ano90b] Anonymous. Correction: “Small-Disturbance Asymptotic Theory for Linear Calibration Estimators”. Technometrics, 32(4):468, November 1990. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1270154.

See [SS89].

Anonymous:1990:CPB

[Ano90c] Anonymous. Corrections to Piepel, Beckman and Carroll. Technometrics, 32(2):241, May 1990. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/ stable/1268895. See [Pie90, BNC87, Wei90].

Aroian:1983:BRB

[Aro83] Leo A. Aroian. Book review: Handbook of the Normal Distribution by Jagdish K. Patel; Campbell B. Read. Technometrics, 25(1):112–115, February 1983. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267742.

Aroian:1984:BRBa

[Aro84a] Leo A. Aroian. Book review: Applied Time Series: Proceedings of the International Conference Held at Houston, Texas, August 1981 by O. D. Anderson; M. R. Perryman. Technometrics, 26(2):187– 189, May 1984. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1268112.

Aroian:1984:BRBd

[Aro84b] Leo A. Aroian. Book review: Neyman: From Life by Constance Reid. Technometrics, 26(4):417–418, November 1984. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1269511.

Aroian:1984:BRBb

[Aro84c] Leo A. Aroian. Book review: Time Series Analysis: Theory and Practice 3 by O. D. Anderson. Technometrics, 26(4):411–412, November 1984. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1269504.

Aroian:1984:BRBc

[Aro84d] Leo A. Aroian. Book review: Time Series Analysis: Theory and Practice 4 by O. D. Anderson. Technometrics, 26(4):412–413, November 1984. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1269505.

Aroian:1985:BRBa

[Aro85a] Leo A. Aroian. Book review: Good Thinking: The Foundations of Probability and Its Applications by J. J. Good. Technometrics, 27(1):89–90, February 1985. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1270479.

Aroian:1985:BRBc

[Aro85b] Leo A. Aroian. Book review: Handbook of Statistics 3: Time Series in the Frequency Domain by D. R. Brillinger; P. R. Krishnaiah. Technometrics, 27(3):321, August 1985. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1269723.

Aroian:1985:BRBd

[Aro85c] Leo A. Aroian. Book review: The Forecasting Accuracy of Major Time Series by S. Makridakis; A. Andersen; R. Carbone; R. Fildes; M. Hibon; R. Lewandowski; J. Newton; E. Parzen; R. Winkler. Technometrics, 27(4):442–443, November 1985. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1270215.

Aroian:1985:BRBb

[Aro85d] Leo A. Aroian. Book review: The Making of Statisticians by J. Gani. Technometrics, 27(2):210–211, May 1985. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268775.

Aroian:1986:BRBb

[Aro86a] Leo A. Aroian. Book review: Time Series Analysis: Theory and Practice 5 by O. D. Anderson. Technometrics, 28(2):183–186, May 1986. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1270460.

Aroian:1986:BRBa

[Aro86b] Leo A. Aroian. Book review: Time Series Analysis: Theory and Practice 6 by O. D. Anderson; J. K. Ord; E. A. Robinson. Technometrics, 28(2):181–183, May 1986. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1270459.

Ascher:1983:SMR

[Asc83] Harold Ascher. [Statistical Methods in Reliability]: Discussion. Technometrics, 25(4):320–326, November 1983. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1267848. See [Law83a].

Ascher:1989:BRB

[Asc89] Harold Ascher. Book review: Handbook of Reliability Engineering and Management by W. Grant Ireson; Clyde F. Coombs. Technometrics, 31(4):492–494, November 1989. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1270012.

Angus:1985:FFP

[ASVR85] John E. Angus, Ray E. Schafer, Sally Van Den Berg, and Herbert C. Rutemiller. Failure-free period life tests. Technometrics, 27(1):49–56, February 1985. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1270469.

Atkinson:1981:CTC

[Atk81] A. C. Atkinson. A comparison of two criteria for the design of experiments for discriminating between models. Technometrics, 23(3):301–305, August 1981. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267795.

Atkinson:1983:DRA

[Atk83] A. C. Atkinson. Diagnostic regression analysis and shifted power transformations. Technometrics, 25(1):23–33, February 1983. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1267723.

Atkinson:1984:EDR

[Atk84] A. C. Atkinson. [Experimental Design: Review and Comment]: Discussion. Technometrics, 26(2):117–118, May 1984. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268102. See [Hah84a].

Atkinson:1986:DTT

[Atk86] A. C. Atkinson. Diagnostic tests for transformations. Technometrics, 28(1):29–37, February 1986. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1269601.

Atkinson:1988:TU

[Atk88] A. C. Atkinson. Transformations unmasked. Technometrics, 30 (3):311–318, August 1988. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1270085.

Atwood:1986:BFR

[Atw86] Corwin L. Atwood. The binomial failure rate common cause model. Technometrics, 28(2):139–148, May 1986. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1270450.

Aitkin:1980:MMO

[AW80] Murray Aitkin and Granville Tunnicliffe Wilson. Mixture models, outliers, and the EM algorithm. Technometrics, 22(3):325–331, August 1980. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1268316.

Adams:1989:ATL

[AW89] Benjamin M. Adams and William H. Woodall. An analysis of Taguchi’s on-line process-control procedure under a random-walk model. Technometrics, 31(4):401–413, November 1989. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1269991.

Bedrick:1989:ACI

[BA89] Edward J. Bedrick and Jorge Aragón. Approximate confidence intervals for the parameters of a stationary binary Markov chain. Technometrics, 31(4):437–448, November 1989. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1269994.

Bacon:1980:DMM

[Bac80] David W. Bacon. Discussion of “Multivariate Model Identification and Stochastic Control of a Chemical Reactor”. Technometrics, 22(4):465–467, November 1980. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268181. See [MW80b].

Bailar:1980:BRB

[Bai80a] Barbara A. Bailar. Book review: Statistical Survey Techniques by Raymond J. Jessen. Technometrics, 22(1):132, February 1980. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1268400.

Bailey:1980:LSS

[Bai80b] B. J. R. Bailey. Large sample simultaneous confidence intervals for the multinomial probabilities based on transformations of the cell frequencies. Technometrics, 22(4):583–589, November 1980. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1268196.

Barnett:1983:OD

[Bar83] Vic Barnett. [Outlier. . . . . . . . . s]: Discussion. Technometrics, 25 (2):150–152, May 1983. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268542. See [BC83a].

Basu:1986:BRB

[Bas86] Asit P. Basu. Book review: Nonparametric Methods in General Lineral Models by Madan L. Puri; Pranab K. Sen. Technometrics, 28(4):400, November 1986. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268990.

Basu:1988:BRB

[Bas88] Asit Basu. Book review: Reliability and Life Testing by S. K. Sinha. Technometrics, 30(1):131, February 1988. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1270351.

Bates:1983:DU

[Bat83] Douglas Bates. The derivative of |X′X| and its uses. Technometrics, 25(4):373–376, November 1983. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267859.

Bates:1988:BRB

[Bat88] Douglas Bates. Book review: Nonlinear Statistical Models by A. Ronald Gallant. Technometrics, 30(4):453–454, November 1988. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1269811.

Boos:1989:BMT

[BB89] Dennis D. Boos and Cavell Brownie. Bootstrap methods for testing homogeneity of variances. Technometrics, 31(1):69–82, February 1989. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1270366.

Beckman:1983:O

[BC83a] R. J. Beckman and R. D. Cook. Outlier. . . . . . . . . s. Technometrics, 25(2):119–149, May 1983. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/ stable/1268541. See discussion [Bar83, MM83a, Haw83, Pre83a, Hog83, Dra83b] and response [BC83b].

Beckman:1983:OR

[BC83b] R. J. Beckman and R. D. Cook. [Outlier. . . . . . . . . s]: Response. Technometrics, 25(2):161–163, May 1983. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://

www.jstor.org/stable/1268548. See [BC83a] and corrigendum

[Ano83e].

Becker:1987:BS

[BC87] Richard A. Becker and William S. Cleveland. Brushing scatterplots. Technometrics, 29(2):127–142, May 1987. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www. jstor.org/stable/1269768. With discussion and a reply by the authors.

Bechhofer:1982:MCO

[BD82a] Robert E. Bechhofer and Charles W. Dunnett. Multiple comparisons for orthogonal contrasts: Examples and tables. Technometrics, 24(3):213–222, August 1982. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268681.

Box:1982:MLF

[BD82b] G. E. P. Box and N. R. Draper. Measures of lack of fit for response surface designs and predictor variable transformations. Technometrics, 24(1):1–8, February 1982. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267572.

Bunke:1984:EMS

[BD84] O. Bunke and B. Droge. Estimators of the mean squared error of prediction in linear regression. Technometrics, 26(2):145–155, May 1984. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1268109.

Bain:1980:IPC

[BE80] Lee J. Bain and Max Engelhardt. Inferences on the parameters and current system reliability for a time truncated Weibull process. Technometrics, 22(3):421–426, August 1980. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1268327.

Bain:1981:SAD

[BE81] Lee J. Bain and Max Engelhardt. Simple approximate distributional results for confidence and tolerance limits for the Weibull distribution based on maximum likelihood estimators. Technometrics, 23(1):15–20, February 1981. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267970.

Beauchamp:1983:BRB

[Bea83] John J. Beauchamp. Book review: Applied Multivariate Statistical Analysis by Richard A. Johnson; Dean W. Wichern. Technometrics, 25(4):385–386, November 1983. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267865.

Bechhofer:1984:EDR

[Bec84] Robert E. Bechhofer. [Experimental Design: Review and Comment]: Discussion. Technometrics, 26(2):125–127, May 1984. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1268105. See [Hah84a].

Beckman:1987:BRB

[Bec87] Richard J. Beckman. Book review: Applied Linear Regression by Sanford Weisberg. Technometrics, 29(1):119–120, February 1987. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1269895.

Bement:1989:BRB

[Bem89] Thomas R. Bement. Book review: Taguchi Techniques for Quality Engineering by Phillip J. Ross. Technometrics, 31(2):254–255, May 1989. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1268823.

Berger:1982:MHT

[Ber82a] Roger L. Berger. Multiparameter hypothesis testing and acceptance sampling. Technometrics, 24(4):295–300, November 1982. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267823.

Berliner:1982:BRB

[Ber82b] Mark Berliner. Book review: On Reduced Risk Estimation in Linear Models by Erkki Liski. Technometrics, 24(3):253–254, August 1982. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1268694.

Berry:1982:BRB

[Ber82c] Donald A. Berry. Book review: Statistical Decision Theory, Foundations, Concepts, and Methods by James O. Berger. Technometrics, 24(1):80–82, February 1982. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267583.

Berry:1983:BRB

[Ber83] Donald A. Berry. Book review: Bayesian Statistics, Proceedings of the First International Meeting Held in Valencia (Spain), May 28 to June 2, 1979 by J. M. Bernardo; M. H. DeGroot; D. V. Lindley; A. F. M. Smith. Technometrics, 25(2):206–207, May 1983. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268557.

Berk:1984:VRP

[Ber84] Kenneth N. Berk. Validating regression procedures with new data. Technometrics, 26(4):331–338, November 1984. CODEN TCMTA2.

ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1269495.

Berliner:1987:BCM

[Ber87] L. Mark Berliner. Bayesian control in mixture models. Technometrics, 29(4):455–460, November 1987. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1269457.

Berk:1989:BRB

[Ber89a] Robert H. Berk. Book review: Assignment Methods in Combinatorial Data Analysis by Lawrence J. Hubert. Technometrics, 31(2):272–273, May 1989. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268845.

Berliner:1989:BRB

[Ber89b] L. Mark Berliner. Book review: Maximum-Entropy and Bayesian Spectral Analysis and Estimation Problems by C. Ray Smith; Gary J. Erickson. Technometrics, 31(2):271–272, May 1989. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268843.

Box:1987:PMI

[BF87] George Box and Conrad A. Fung. [Performance Measures Independent of Adjustment]: Discussion. Technometrics, 29(3):270, August 1987. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1269334. See

[LSK87a].

Burdick:1984:CIL

[BG84] Richard K. Burdick and Franklin A. Graybill. Confidence intervals on linear combinations of variance components in the unbalanced one-way classification. Technometrics, 26(2):131–136, May 1984. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1268107.

Bauer:1980:EMT

[BH80] Peter Bauer and Peter Hackl. An extension of the MOSUM technique for quality control. Technometrics, 22(1):1–7, February 1980. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1268376.

Bradu:1982:LMO

[BH82] Dan Bradu and Douglas M. Hawkins. Location of multiple outliers in two-way tables, using tetrads. Technometrics, 24(2):103–108, May 1982. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1268486.

Belsley:1984:EWO

[BH84] David A. Belsley and R. R. Hocking. Eigenvector weaknesses and other topics for assessing conditioning diagnostics. Technometrics, 26(3):297–301, August 1984. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267566.

Bisgaard:1984:ESQ

[BHP84] Søren Bisgaard, William G. Hunter, and Lars Pallesen. Economic selection of quality of manufactured product. Technometrics, 26(1):9–18, February 1984. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268411.

Bilonick:1984:BRB

[Bil84] R. A. Bilonick. Book review: Sampling of Particulate Materials: Theory and Practice by Pierre M. Gy. Technometrics, 26 (3):293–294, August 1984. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267561.

Bilonick:1988:BRBa

[Bil88a] R. A. Bilonick. Book review: Sampling and Weighing of Bulk Solids by J. W. Merks. Technometrics, 30(1):132–134, February 1988. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1270354.

Bilonick:1988:BRBb

[Bil88b] R. A. Bilonick. Book review: Sampling of Powders and Bulk Materials by Karl Sommer. Technometrics, 30(3):349–351, August 1988. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1270092.

Bingham:1980:HPL

[Bin80] N. H. Bingham. [How Probable Is a Large Resultant in a Random Walk?]: Reply. Technometrics, 22(4):630, November 1980. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268204. See [Wil80].

Bingham:1989:BRB

[Bin89] Christopher Bingham. Book review: Statistical Analysis of Spherical Data by N. I. Fisher; T. Lewis; B. J. J. Embleton. Technometrics, 31(4):494–495, November 1989. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1270013.

Birnbaum:1983:SMR

[Bir83] Z. W. Birnbaum. [Statistical Methods in Reliability]: Discussion. Technometrics, 25(4):330–331, November 1983. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1267851. See [Law83a].

Bishop:1981:BRB

[Bis81] Thomas A. Bishop. Book review: Optimization and Industrial Experimentation by William E. Biles; James J. Swain. Technometrics, 23(2):203, May 1981. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268045.

Bishop:1983:BRB

[Bis83] Yvonne M. Bishop. Book review: The Analysis of Categorical Data by R. L. Plackett. Technometrics, 25(2):212–213, May 1983. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1268566.

Bisgaard:1989:BRB

[Bis89] Soren Bisgaard. Book review: System of Experimental Design (Vol. 2) by Genichi Taguchi. Technometrics, 31(2):257–260, May 1989. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1268826.

Bryson:1981:IML

[BJ81] Maurice C. Bryson and Mark E. Johnson. The incidence of monotone likelihood in the Cox model. Technometrics, 23(4):381–383, November 1981. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1268228.

Box:1986:TIE

[BJ86] George Box and Stephen Jones. [Testing in Industrial Experiments with Ordered Categorical Data]: Discussion. Technometrics, 28 (4):295–301, November 1986. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268976. See [BW85a].

Beckman:1987:FSD

[BJ87] Richard J. Beckman and Mark E. Johnson. Fitting the Student-t distribution to grouped data, with application to a particle scattering experiment. Technometrics, 29(1):17–22, February 1987. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1269879.

Bjornstad:1980:CTM

[Bjø80] Jan F. Bjørnstad. Comparison of three minimax subset selection procedures. Technometrics, 22(4):617–620, November 1980. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1268201.

Bohachevsky:1986:GSA

[BJS86] Ihor O. Bohachevsky, Mark E. Johnson, and Myron L. Stein. Generalized simulated annealing for function optimization. Technometrics, 28(3):209–217, August 1986. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1269076.

Beneke:1988:SAQ

[BLSF88] Margarita Beneke, Lawrence M. Leemis, Robert E. Schlegel, and Bobbie L. Foote. Spectral analysis in quality control: A control chart based on the periodogram. Technometrics, 30(1):63–70, February 1988. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1270321.

Box:1986:AUF

[BM86a] George E. P. Box and R. Daniel Meyer. An analysis for unreplicated fractional factorials. Technometrics, 28(1):11–18, February 1986. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1269599.

Box:1986:DEF

[BM86b] George E. P. Box and R. Daniel Meyer. Dispersion effects from fractional designs. Technometrics, 28(1):19–27, February 1986. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1269600. See correction

[BM87b].

Beckman:1987:MCE

[BM87a] Richard J. Beckman and Michael D. McKay. Monte Carlo estimation under different distributions using the same simulation. Technometrics, 29(2):153–160, May 1987. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1269770.

Box:1987:CDE

[BM87b] George E. P. Box and R. Daniel Meyer. Correction: “Dispersion effects from fractional designs” [Technometrics 28 (1986), no. 1, 19–27]. Technometrics, 29(2):250, May 1987. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1269793. See [BM86b].

Brien:1988:DM

[BM88] Carl M. O’ Brien and C. L. Mallows. Discussion of Mallows (1986). Technometrics, 30(1):135–136, February 1988. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1270355. See [Mal86].

Borth:1985:DIA

[BME85] David M. Borth, Robert J. McKay, and J. Richard Elliott. A difficulty information approach to substituent selection in QSAR studies. Technometrics, 27(1):25–35, February 1985. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1270466.

Bingham:1981:ADN

[BN81] Christopher Bingham and Lloyd S. Nelson. An approximation for the distribution of the von Neumann ratio. Technometrics, 23(3):285–288, August 1981. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267792.

Beckman:1987:DMM

[BNC87] Richard J. Beckman, Christopher J. Nachtsheim, and R. Dennis Cook. Diagnostics for mixed-model analysis of variance. Technometrics, 29(4):413–426, November 1987. CODEN TCMTA2.

ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1269452. See corrections[Ano90c].

Boardman:1985:BRBb

[Boa85a] Thomas J. Boardman. Book review: The Collected Works of George E. P. Box by George C. Tiao; George E. P. Box. Technometrics, 27(3):319, August 1985. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1269719.

Boardman:1985:BRBa

[Boa85b] Thomas J. Boardman. Book review: The Collected Works of John W. Tukey, Volume I, Time Series: 1949–1964 by David R. Brillinger; John W. Tukey. Technometrics, 27(3):319, August 1985. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1269718.

Boardman:1987:PMI

[Boa87] Thomas J. Boardman. [Performance Measures Independent of Adjustment]: Discussion. Technometrics, 29(3):266, August 1987. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1269332. See [LSK87a].

Boen:1983:BRB

[Boe83] James R. Boen. Book review: Applied Statistics: Principles and Examples by D. R. Cox; E. J. Snell. Technometrics, 25(1):110, February 1983. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1267735.

Bonis:1981:BRB

[Bon81] Austin J. Bonis. Book review: Frugal Sampling Schemes by James R. King. Technometrics, 23(3):309, August 1981. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267801.

Boos:1984:UEV

[Boo84] Dennis D. Boos. Using extreme value theory to estimate large percentiles. Technometrics, 26(1):33–39, February 1984. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268413.

Boos:1987:DSE

[Boo87] Dennis D. Boos. Detecting skewed errors from regression residuals. Technometrics, 29(1):83–90, February 1987. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1269885.

Booker:1988:BRB

[Boo88] Jane M. Booker. Book review: Stochastic Simulation by Brian D. Ripley. Technometrics, 30(2):231–232, May 1988. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1270171.

Borowiak:1986:BRB

[Bor86] Dale S. Borowiak. Book review: Statistics and Its Applications by Richard J. Larsen; Morris L. Marx. Technometrics, 28(4):403, November 1986. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1268995.

Borowiak:1987:BRB

[Bor87] Dale Borowiak. Book review: Data Distributions by Ronald Christensen. Technometrics, 29(2):243, May 1987. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1269784.

Borowiak:1989:BRB

[Bor89] Dale Borowiak. Book review: Probability and Statistical Inference by Robert V. Hogg; Elliot A. Tanis. Technometrics, 31(3):391–392, August 1989. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/3556162.

Bowker:1983:BRB

[Bow83] Rosedith Sitgreaves Bowker. Book reviews: Statistics by Norma Gilbert; Measurement Theory for the Behavioral Sciences by Edwin T. Ghiselli; John P. Campbell; Sheldon Zedeck. Technometrics, 25(1):111, February 1983. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267738.

Bowerman:1987:BRB

[Bow87] Bruce Bowerman. Book review: Time Series Analysis by Jonathan D. Cryer. Technometrics, 29(2):240, May 1987. CODEN TCMTA2.

ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1269781.

Box:1984:IPD

[Box84] George E. P. Box. The importance of practice in the development of statistics. Technometrics, 26(1):1–8, February 1984. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268410.

Box:1987:MWG

[Box87] George Box. In memoriam: William G. Hunter, 1937–1986. Technometrics, 29(3):251–252, August 1987. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1269330.

Box:1988:SNRa

[Box88a] George Box. Signal-to-noise ratios, performance criteria, and transformations. Technometrics, 30(1):1–17, February 1988. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1270311. See discussion [STL88,

Par88b, NP88a, CR88, Gun88b, Ull88] and response [Box88b].

Box:1988:SNRb

[Box88b] George Box. [Signal-to-Noise Ratios, Performance Criteria, and Transformations]: Response. Technometrics, 30(1):38–40, February 1988. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1270318. See

[Box88a].

Boyer:1980:BRB

[Boy80] John E. Boyer, Jr. Book review: Introduction to Statistics: A Nonparametric Approach for the Social Sciences by Chris Leach. Technometrics, 22(3):445, August 1980. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268333.

Brown:1985:KEC

[BR85] Philip J. Brown and Peter W. K. Rundell. Kernel estimates for categorical data. Technometrics, 27(3):293–299, August 1985. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1269711.

Best:1987:WAS

[BR87] D. J. Best and J. C. W. Rayner. Welch’s approximate solution for the Behrens–Fisher problem. Technometrics, 29(2):205–210, May 1987. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1269775.

Banerjee:1988:EDC

[BR88] P. K. Banerjee and M. A. Rahim. Economic design of x̄-control charts under Weibull shock models. Technometrics, 30(4):407–414, November 1988. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1269803.

See correction [BR89].

Banerjee:1989:CED

[BR89] P. K. Banerjee and M. A. Rahim. Correction: “Economic design of x̄-control charts under Weibull shock models” [Technometrics 30 (1988), no. 4, 407–414; MR0970999 (90e:62135)]. Technometrics, 31(3):399, August 1989. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/3556175. See [BR88].

Bremer:1989:BRB

[Bre89] Ronald Bremer. Book review: Estimation of Variance Components and Applications by C. Radhakrishna Rao; Jügen Kleffe. Technometrics, 31(4):496, November 1989. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1270015.

Brooks:1989:BRB

[Bro89] Daniel G. Brooks. Book review: Akaike Information Criterion Statistics by Y. Sakamoto; M. Ishiguro; G. Kitagawa. Technometrics, 31(2):270–271, May 1989. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268842.

Brush:1986:CCB

[Bru86] Gary G. Brush. A comparison of classical and Bayes Producer’s risk. Technometrics, 28(1):69–72, February 1986. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1269605.

Bryce:1989:BRB

[Bry89] G. Rex Bryce. Book review: Statistical Quality Control for Manufacturing Managers by William Messina. Technometrics, 31(1): 119–120, February 1989. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1270378.

Bohrer:1981:EBA

[BS81] Robert Bohrer and Mark J. Schervish. An error-bounded algorithm for normal probabilities of rectangular regions. Technometrics, 23(3):297–300, August 1981. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267794.

Brain:1983:RTE

[BS83] Carlos W. Brain and Samuel S. Shapiro. A regression test for exponentiality: Censored and complete samples. Technometrics, 25(1):69–76, February 1983. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267728.

Beal:1988:HNR

[BS88] S. L. Beal and L. B. Sheiner. Heteroscedastic nonlinear regression. Technometrics, 30(3):327–338, August 1988. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1270087.

Birkes:1981:EEM

[BSA81] David Birkes, Justus Seely, and Abdul-Mordy Azzam. An efficient estimator of the mean in a two-stage nested model. Technometrics, 23(2):143–148, May 1981. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268029.

Beal:1986:DBW

[BSBW86] Stuart Beal, Lewis Sheiner, Douglas M. Bates, and Donald G. Watts. Discussion of Bates and Watts (1985). Technometrics, 28(4):407–408, November 1986. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268999. See [BW85a].

Baxter:1982:TRF

[BSMB82] Laurence A. Baxter, Ernest M. Scheuer, Denis J. McConalogue, and Wallace R. Blischke. On the tabulation of the renewal function. Technometrics, 24(2):151–156, May 1982. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1268495.

Bechhofer:1981:IBD

[BT81] Robert E. Bechhofer and Ajit C. Tamhane. Incomplete block designs for comparing treatments with a control: General theory. Technometrics, 23(1):45–57, February 1981. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1267975. See corrigenda [Ano82e].

Baskerville:1982:GRM

[BT82] J. C. Baskerville and J. H. Toogood. Guided regression modeling for prediction and exploration of structure with many explanatory variables. Technometrics, 24(1):9–17, February 1982. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267573.

Bechhofer:1983:DEC

[BT83] Robert E. Bechhofer and Ajit C. Tamhane. Design of experiments for comparing treatments with a control: Tables of optimal allocations of observations. Technometrics, 25(1):87–95, February 1983. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1267730.

Beckman:1989:TST

[BT89] Richard J. Beckman and Gary L. Tietjen. Two-sided tolerance limits for balanced random-effects ANOVA models. Technometrics, 31(2):185–197, May 1989. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/ stable/1268816. See correction [BT98].

Beckman:1998:CTS

[BT98] Richard J. Beckman and Gary L. Tietjen. Correction: “Twosided tolerance limits for balanced random-effects ANOVA models” [Technometrics 31 (1989), no. 2, 185–197; MR1007294 (90f:62230)]. Technometrics, 40(3):269, August 1998. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1271208. See [BT89].

Bulmahn:1989:BRB

[Bul89] Barbara J. Bulmahn. Book review: Foundations of Statistics by D. G. Rees. Technometrics, 31(3):388–389, August 1989. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/3556158.

Buonaccorsi:1986:DCC

[Buo86] J. P. Buonaccorsi. Design considerations for calibration. Technometrics, 28(2):149–155, May 1986. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1270451.

Bates:1981:ROO

[BW81] Douglas M. Bates and Donald G. Watts. A relative offset orthogonality convergence criterion for nonlinear least squares. Technometrics, 23(2):179–183, May 1981. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268035.

Bates:1985:MESa

[BW85a] Douglas M. Bates and Donald G. Watts. Multiresponse estimation with special application to linear systems of differential equations. Technometrics, 27(4):329–339, November 1985. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1270198. See discussion

[McL85, Yan85, Zie85b, BW85b, BSBW86].

Bates:1985:MESb

[BW85b] Douglas M. Bates and Donald G. Watts. [Multiresponse Estimation with Special Application to Linear Systems of Differential Equations]: Response. Technometrics, 27(4):358–360, November 1985. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1270202. See [BW85a].

Cheng:1981:MLE

[CA81] R. C. H. Cheng and N. A. K. Amin. Maximum likelihood estimation of parameters in the inverse Gaussian distribution, with unknown origin. Technometrics, 23(3):257–263, August 1981. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1267789.

Caffrey:1987:BRB

[Caf87] Stephen J. Caffrey. Book review: Design and Analysis of Experiments by Roger G. Petersen. Technometrics, 29(3):379–380, August 1987. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1269349.

Calvin:1988:BRB

[Cal88] Thomas W. Calvin. Book review: Statistical Analysis for Decision Making by Morris Hamburg. Technometrics, 30(2):235–236, May 1988. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1270178.

Cameron:1983:CTS

[Cam83] M. A. Cameron. The comparison of time series recorders. Technometrics, 25(1):9–22, February 1983. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267722.

Carr:1980:FET

[Car80] Wendell E. Carr. Fisher’s exact test extended to more than two samples of equal size. Technometrics, 22(2):269–270, May 1980. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1268467.

Carlson:1987:BRB

[Car87a] Richard R. Carlson. Book reviews: Volume 2: How to Perform Continuous Sampling (CSP) by Kenneth S. Stephens; Volume 4: How to Perform Skip-Lot and Chain Sampling by Kenneth S. Stephens; Volume 7: How and When to Perform Bayesian Acceptance Sampling by Thomas W. Calvin. Technometrics, 29(4):488, November 1987. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1269462.

Carr:1987:BRB

[Car87b] Daniel Carr. Book review: The Visual Display of Quantitative Information by Edward R. Tufte. Technometrics, 29(1):118–119, February 1987. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1269894.

Carey:1988:BRB

[Car88] Michele Carey. Book review: Applied Reliability by P. A. Tobias; D. Trindade. Technometrics, 30(2):231, May 1988. CODEN TCMTA2.

ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1270170.

Castro:1980:BRB

[Cas80] Peter Castro. Book review: Compact Numerical Methods for Computers: Linear Algebra and Function Minimization by J. C. Nash. Technometrics, 22(1):132–133, February 1980. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1268401.

Castro:1982:BRB

[Cas82] Peter E. Castro. Book review: Infinitely Divisible Point Processes by Klaus Matthes; Johannes Kerstan; Joseph Mecke. Technometrics, 24(1):83, February 1982. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267587.

Choy:1980:SBI

[CB80] J. H. Chin Choy and L. D. Broemeling. Some Bayesian inferences for a changing linear model. Technometrics, 22(1):71–78, February 1980. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1268385.

Carter:1986:CII

[CCMC86] Walter H. Carter, Jr., Vernon M. Chinchilli, Raymond H. Myers, and Eleanor D. Campbell. Confidence intervals and an improved ridge analysis of response surfaces. Technometrics, 28(4):339–346, November 1986. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1268983.

Chhikara:1982:PLI

[CG82] Raj S. Chhikara and Irwin Guttman. Prediction limits for the inverse Gaussian distribution. Technometrics, 24(4):319–324, November 1982. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1267827.

Chaloner:1987:BAE

[Cha87] Kathryn Chaloner. A Bayesian approach to the estimation of variance components for the unbalanced one-way random model. Technometrics, 29(3):323–337, August 1987. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1269342.

Champ:1989:BRB

[Cha89] Charles W. Champ. Book review: ISP: Interactive Statistical Programs by Spyros Makridakis; Robert L. Winkler. Technometrics, 31(1):128–129, February 1989. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1270391.

Chesson:1983:BRB

[Che83] Jean Chesson. Book review: Applicable Mathematics of Non-Physical Phenomena by F. Oliveira-Pinto; B. W. Conolly. Technometrics, 25(1):112, February 1983. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267741.

Chen:1985:RAS

[Che85] Chan Fu Chen. Robustness aspects of score tests for generalized linear and partially linear regression models. Technometrics, 27(3):277–283, August 1985. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1269709.

Cheng:1987:CBT

[Che87] R. C. H. Cheng. Confidence bands for two-stage design problems. Technometrics, 29(3):301–309, August 1987. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1269340.

Cheng:1989:SOM

[Che89] Ching-Shui Cheng. Some orthogonal main-effect plans for asymmetrical factorials. Technometrics, 31(4):475–477, November 1989. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1269998.

Chhikara:1986:EAC

[Chh86] Raj S. Chhikara. Error analysis of crop acreage estimation using satellite data. Technometrics, 28(1):73–80, February 1986. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1269606.

Christensen:1989:BRB

[Chr89] Ronald Christensen. Book review: The Method of Paired Comparisons by H. A. David. Technometrics, 31(4):495–496, November 1989. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1270014.

Cheng:1983:CBC

[CI83] R. C. H. Cheng and T. C. Iles. Confidence bands for cumulative distribution functions of continuous random variables. Technometrics, 25(1):77–86, February 1983. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267729.

Cheng:1988:OSC

[CI88] R. C. H. Cheng and T. C. Iles. One-sided confidence bands for cumulative distribution functions. Technometrics, 30(2):155–159, May 1988. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1270160.

Cook:1986:GBP

[CJ86] R. Dennis Cook and Mark E. Johnson. Generalized Burr–Paretologistic distributions with applications to a uranium exploration data set. Technometrics, 28(2):123–131, May 1986. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1270448.

Conover:1981:CST

[CJJ81] W. J. Conover, Mark E. Johnson, and Myrle M. Johnson. A comparative study of tests for homogeneity of variances, with applications to the outer continental shelf bidding data. Technometrics, 23(4):351–361, November 1981. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/ stable/1268225. See corrigendum [Ano84e].

Cleveland:1980:DEV

[Cle80] William S. Cleveland. Discussion of “Extreme Values from a Nonstationary Stochastic Process: Application to Air Quality Analysis”. Technometrics, 22(4):479–481, November 1980. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268184. See [Hor80a].

Castro:1986:PMV

[CLS86] P. E. Castro, W. H. Lawton, and E. A. Sylvestre. Principal modes of variation for processes with continuous sample curves. Technometrics, 28(4):329–337, November 1986. CODEN TCMTA2.

ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1268982.

Cook:1980:CAC

[CN80] R. Dennis Cook and Christopher J. Nachtsheim. A comparison of algorithms for constructing exact D-optimal designs. Technometrics, 22(3):315–324, August 1980. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268315.

Cook:1982:MRL

[CN82] R. Dennis Cook and Christopher J. Nachtsheim. Model robust, linear-optimal designs. Technometrics, 24(1):49–54, February 1982. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1267577.

Cook:1989:CAB

[CN89] R. Dennis Cook and Christopher J. Nachtsheim. Computer-aided blocking of factorial and response-surface designs. Technometrics, 31(3):339–346, August 1989. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/3556143.

Chou:1986:OSS

[CO86] Youn-Min Chou and D. B. Owen. One-sided simultaneous lower prediction intervals for l future samples from a normal distribution. Technometrics, 28(3):247–251, August 1986. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1269080.

Cohen:1981:BRB

[Coh81] Joel E. Cohen. Book review: Reversibility and Stochastic Networks by Frank P. Kelly. Technometrics, 23(2):206–207, May 1981. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1268051.

Cohen:1982:BRB

[Coh82] A. Clifford Cohen. Book review: Finite Mixture Distributions by B. S. Everitt; D. J. Hand. Technometrics, 24(4):339, November 1982. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1267834.

Coleman:1987:BRB

[Col87] David E. Coleman. Book review: Elements of Graphing Data by William Cleveland. Technometrics, 29(1):116–118, February 1987. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1269893.

Coleman:1988:BRBb

[Col88a] David Coleman. Book review: Volume 11: How to Perform Statistical Tolerance Analysis by Neil D. Cox. Technometrics, 30 (1):114–115, February 1988. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1270328.

Coleman:1988:BRBa

[Col88b] David Coleman. Book review: Volume 3: How to Test Normality and Other Distributional Assumptions by Samuel S. Shapiro. Technometrics, 30(1):114, February 1988. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1270327.

Collings:1989:QC

[Col89] Bruce Jay Collings. Quick confounding. Technometrics, 31(1): 107–110, February 1989. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1270370.

Copas:1983:SS

[Cop83] J. B. Copas. Stochastic shrinkage: A comment on the paper by Egerton and Laycock and a theorem of Bibby and Toutenburg. Technometrics, 25(3):303–304, August 1983. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1268616.

Cornelius:1980:FAM

[Cor80] P. L. Cornelius. Functions approximating Mandel’s tables for the means and standard deviations of the first three roots of a Wishart matrix. Technometrics, 22(4):613–616, November 1980. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268200.

Cornell:1981:BRB

[Cor81] John A. Cornell. Book review: Statistical Methods by G. W. Snedecor; W. G. Cochran. Technometrics, 23(3):312–313, August 1981. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1267805.

Cornell:1984:EDR

[Cor84] John A. Cornell. [Experimental Design: Review and Comment]: Discussion. Technometrics, 26(2):119–122, May 1984. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268103. See [Hah84a].

Cornell:1987:BRB

[Cor87] John A. Cornell. Book review: Classical and Modern Regression with Applications by Raymond H. Myers. Technometrics, 29(3):377–378, August 1987. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1269347.

Cox:1984:EDR

[Cox84] D. R. Cox. [Experimental Design: Review and Comment]: Discussion. Technometrics, 26(2):116, May 1984. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1268101. See [Hah84a].

Cook:1981:ABS

[CP81] R. Dennis Cook and P. Prescott. On the accuracy of Bonferroni significance levels for detecting outliers in linear models. Technometrics, 23(1):59–63, February 1981. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267976.

Chan:1980:REL

[CR80] Lai K. Chan and Lennart S. Rhodin. Robust estimation of location using optimally chosen sample quantiles. Technometrics, 22(2):225– 237, May 1980. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1268462.

Carroll:1985:TRR

[CR85] R. J. Carroll and David Ruppert. Transformations in regression: A robust analysis. Technometrics, 27(1):1–12, February 1985. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1270463.

Carroll:1987:DRE

[CR87] R. J. Carroll and David Ruppert. Diagnostics and robust estimation when transforming the regression model and the response. Technometrics, 29(3):287–299, August 1987. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1269339.

Carroll:1988:SNR

[CR88] Raymond J. Carroll and David Ruppert. [Signal-to-Noise Ratios, Performance Criteria, and Transformations]: Discussion. Technometrics, 30(1):30–31, February 1988. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1270315. See [Box88a].

Cressie:1987:BRB

[Cre87] Noel Cressie. Book review: Spatial Data Analysis by Example: Volume 1: Point Pattern and Quantitative Data by Graham Upton; Bernard Fingleton. Technometrics, 29(1):114–116, February 1987. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1269891.

Crocker:1980:BRB

[Cro80] Douglas C. Crocker. Book review: Linear Regression Analysis by G. A. F. Seber. Technometrics, 22(1):130, February 1980. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268395.

Crow:1982:CIP

[Cro82] Larry H. Crow. Confidence interval procedures for the Weibull process with applications to reliability growth. Technometrics, 24(1):67–72, February 1982. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267580.

Crosier:1984:MEG

[Cro84] Ronald B. Crosier. Mixture experiments: Geometry and pseudocomponents. Technometrics, 26(3):209–216, August 1984. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267546.

Crosier:1986:GCM

[Cro86a] Ronald B. Crosier. The geometry of constrained mixture experiments. Technometrics, 28(2):95–102, May 1986. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1270445.

Crosier:1986:NTS

[Cro86b] Ronald B. Crosier. A new two-sided cumulative sum quality control scheme. Technometrics, 28(3):187–194, August 1986. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1269074. See correction [Cro89].

Crowder:1987:SMS

[Cro87] Stephen V. Crowder. A simple method for studying run-length distributions of exponentially weighted moving average charts. Technometrics, 29(4):401–407, November 1987. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1269450.

Crosby:1988:BRB

[Cro88a] David Crosby. Book review: Practical Statistics for Engineers and Scientists by N. P. Cheremisinoff. Technometrics, 30(2):234–235, May 1988. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1270175.

Crosier:1988:MGC

[Cro88b] Ronald B. Crosier. Multivariate generalizations of cumulative sum quality-control schemes. Technometrics, 30(3):291–303, August 1988. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1270083.

Crosier:1989:CNT

[Cro89] Ronald B. Crosier. Correction: “A new two-sided cumulative sum quality control scheme” [Technometrics 28 (1986), no. 3, 187–194]. Technometrics, 31(4):503, November 1989. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1270025. See [Cro86b].

Crump:1982:BRB

[Cru82] Phelps P. Crump. Book review: Statistical Analysis: A Computer Oriented Approach by A. A. Afifi; S. P. Azen. Technometrics, 24(3):249–250, August 1982. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268686.

Crump:1984:BRB

[Cru84] Phelps P. Crump. Book review: BMDP Statistical Software (1981 Ed.) by W. J. Dixon. Technometrics, 26(4):416–417, November 1984. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1269510.

Carroll:1988:QEM

[CSS88] R. J. Carroll, C. H. Spiegelman, and J. Sacks. A quick and easy multiple-use calibration-curve procedure. Technometrics, 30 (2):137–141, May 1988. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1270158.

Chang:1988:ETS

[CTC88] Ih Chang, George C. Tiao, and Chung Chen. Estimation of time series parameters in the presence of outliers. Technometrics, 30(2):193–204, May 1988. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1270165.

Cook:1980:CEI

[CW80] R. Dennis Cook and Sanford Weisberg. Characterizations of an empirical influence function for detecting influential cases in regression. Technometrics, 22(4):495–508, November 1980. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268187.

Cook:1983:TIC

[CW83] R. D. Cook and P. C. Wang. Transformations and influential cases in regression. Technometrics, 25(4):337–343, November 1983. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1267855.

Champ:1987:ERS

[CW87a] Charles W. Champ and William H. Woodall. Exact results for Shewhart control charts with supplementary runs rules. Technometrics, 29(4):393–399, November 1987. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1269449.

Conerly:1987:MOW

[CW87b] M. D. Conerly and J. T. Webster. MINQE for the one-way classification. Technometrics, 29(2):229–236, May 1987. CODEN TCMTA2.

ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1269778.

Cook:1989:RDD

[CWC^{+}89] R. Dennis Cook, Sanford Weisberg, Daniel B. Carr, Daryl Pregibon, A. C. Atkinson, Luke Tierney, and Roy E. Welsch. Regression diagnostics with dynamic graphics: [with discussions and response]. Technometrics, 31(3):277–291 + 293–301 + 303–311, August 1989. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/3556140. With discussion

and a reply by the authors.

Czitrom:1989:BRB

[Czi89] Veronica Czitrom. Book review: Linear Dynamical Systems by John L. Casti. Technometrics, 31(1):125–126, February 1989. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1270384.

Donev:1988:AAC

[DA88] A. N. Donev and A. C. Atkinson. An adjustment algorithm for the construction of exact D-optimum experimental designs. Technometrics, 30(4):429–433, November 1988. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1269806.

Dalal:1983:ESC

[Dal83] S. R. Dalal. Exact simultaneous confidence bands for random intercept regression. Technometrics, 25(3):263–269, August 1983. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1268611.

Dalal:1987:BRB

[Dal87] Jayesh Dalal. Book review: The Fascination of Statistics by Richard J. Brook; Gregory C. Arnold; Thomas H. Hassard; Robert M. Pringle. Technometrics, 29(2):246, May 1987. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1269787.

Davis:1979:RMD

[Dav79] William W. Davis. Robust methods for detection of shifts of the innovation variance of a time series. Technometrics, 21(3):313–320, August 1979. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1267753.

See [Ano80e].

Dean:1981:BRB

[Dea81] Angela Dean. Book review: Incomplete Block Designs by P. W. M. John. Technometrics, 23(4):416–417, November 1981. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268236.

Dean:1982:BRB

[Dea82] Angela M. Dean. Book review: Optimal Design: An Introduction to the Theory for Parameter Estimation by S. D. Silvey. Technometrics, 24(2):168, May 1982. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268505.

Deaton:1989:BRB

[Dea89] Michael Deaton. Book review: Measurement Error Models by Wayne A. Fuller. Technometrics, 31(1):112, February 1989. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1270371.

DeGroot:1981:BRB

[DeG81] Morris H. DeGroot. Book review: Game Theory: Mathematical Models of Conflict by A. J. Jones. Technometrics, 23(2):206, May 1981. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1268050.

Dehnad:1987:BRB

[Deh87a] Khosrow Dehnad. Book review: Density Estimation for Statistics and Data Analysis by Bernard Silverman. Technometrics, 29 (4):495, November 1987. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1269475.

Dehnad:1987:PMI

[Deh87b] Khosrow Dehnad. [Performance Measures Independent of Adjustment]: Discussion. Technometrics, 29(3):271–273, August 1987. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1269335. See [LSK87a].

DePriest:1988:BRBb

[DeP88a] Douglas J. DePriest. Book review: Computer Simulation and Modelling by Francis Neelamkavil. Technometrics, 30(4):463–464, November 1988. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1269826.

DePriest:1988:BRBa

[DeP88b] Douglas J. DePriest. Book review: Lessons in Digital Estimation Theory by Jerry M. Mendel. Technometrics, 30(1):128–129, February 1988. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1270347.

Derringer:1988:BRB

[Der88] George Derringer. Book review: Optimization for Engineering Systems by Ralph W. Pike. Technometrics, 30(1):131–132, February 1988. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1270352.

Derringer:1989:BRB

[Der89] George C. Derringer. Book review: Statistics for the Engineering and Computer Sciences by William Mendenhall; Terry Sincich. Technometrics, 31(3):387–388, August 1989. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/3556157.

Draper:1987:CAT

[DG87] Norman R. Draper and Irwin Guttman. A comparison of alternative tests for lack of fit. Technometrics, 29(1):91–93, February 1987. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1269886.

Draper:1988:IR

[DG88] Norman R. Draper and Irwin Guttman. An index of rotatability. Technometrics, 30(1):105–111, February 1988. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1270326.

Downing:1985:ERS

[DGH85] D. J. Downing, R. H. Gardner, and F. O. Hoffman. An examination of response-surface methodologies for uncertainty analysis in assessment models. Technometrics, 27(2):151–163, May 1985. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1268763. See dis-

cussion [EDGH86].

Dickinson:1982:BRB

[Dic82] Allan W. Dickinson. Book review: Experiments with Mixtures: Designs, Models, and the Analysis of Mixture Data by John A. Cornell. Technometrics, 24(2):161–162, May 1982. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1268497.

DiCiccio:1987:AIG

[DiC87] T. J. DiCiccio. Approximate inference for the generalized gamma distribution. Technometrics, 29(1):33–40, February 1987. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1269881.

Dickey:1989:BRB

[Dic89] David Dickey. Book review: Time Series Theory and Methods by P. Brockwell; R. Davis. Technometrics, 31(1):121, February 1989. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1270380.

Draper:1980:TTF

[DJ80] N. R. Draper and J. A. John. Testing for three or fewer outliers in two-way tables. Technometrics, 22(1):9–15, February 1980. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1268377.

Draper:1981:IOO

[DJ81] N. R. Draper and J. A. John. Influential observations and outliers in regression. Technometrics, 23(1):21–26, February 1981. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267971.

Draper:1988:RSD

[DJ88] Norman R. Draper and J. A. John. Response-surface designs for quantitative and qualitative variables. Technometrics, 30(4): 423–428, November 1988. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1269805.

DuMond:1987:RCI

[DL87] Charles E. Du Mond and Russell V. Lenth. A robust confidence interval for location. Technometrics, 29(2):211–219, May 1987. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1269776.

deLigny:1981:AFA

[dLNB^{+}81] C. L. de Ligny, G. H. E. Nieuwdorp, W. K. Brederode, W. E. Hammers, and J. C. van Houwelingen. An application of factor analysis with missing data. Technometrics, 23(1):91–95, February 1981. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1267981.

Davis:1987:OSI

[DM87] Charles B. Davis and Roger J. McNichols. One-sided intervals for at least p of m observations from a normal population on each of r future occasions. Technometrics, 29(3):359–370, August 1987. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1269345. See cor-

rection [DM88].

Davis:1988:COS

[DM88] Charles B. Davis and Roger J. McNichols. Corrections to: “Onesided intervals for at least p of m observations from a normal population on each of r future occasions” [Technometrics 29 (1987), no. 3, 359–370]. Technometrics, 30(4):469, November 1988. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1269832. See [DM87].

Downing:1988:BRB

[Dow88] Darryl J. Downing. Book review: SAS System for Forecasting Time Series by John C. Bocklebank; David A. Dickey. Technometrics, 30(1):118, February 1988. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1270332.

Deutsch:1981:STA

[DP81] Stuart Jay Deutsch and Phillip E. Pfeifer. Space-time ARMA modeling with contemporaneously correlated innovations. Technometrics, 23(4):401–409, November 1981. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1268231.

Downing:1980:AKF

[DPM80] D. J. Downing, D. H. Pike, and G. W. Morrison. Application of the Kalman filter to inventory control. Technometrics, 22(1):17–22, February 1980. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1268378.

Draper:1982:CPS

[Dra82] Norman R. Draper. Center points in second-order response surface designs. Technometrics, 24(2):127–133, May 1982. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268490.

Draper:1983:CLF

[Dra83a] N. R. Draper. Cubic lack of fit for three-level second-order response surface designs. Technometrics, 25(2):217, May 1983. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268572.

Draper:1983:OD

[Dra83b] Norman R. Draper. [Outlier. . . . . . . . . s]: Discussion. Technometrics, 25(2):159–160, May 1983. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268547. See [BC83a].

Draper:1985:SCD

[Dra85] Norman R. Draper. Small composite designs. Technometrics, 27(2):173–180, May 1985. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268765.

Driscoll:1989:BRB

[Dri89] Michael F. Driscoll. Book review: Genstat 5: An Introduction by Peter Lane; Nick Galwey; Norman Alvey. Technometrics, 31 (4):485–486, November 1989. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1270007.

Donaldson:1987:CEC

[DS87] Janet R. Donaldson and Robert B. Schnabel. Computational experience with confidence regions and confidence intervals for nonlinear least squares. Technometrics, 29(1):67–82, February 1987. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1269884.

Draper:1988:DMB

[DS88] Norman R. Draper and Elizabeth R. Sanders. Designs for minimum bias estimation. Technometrics, 30(3):319–325, August 1988. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1270086.

Dwivedi:1980:FSP

[DSH80] T. D. Dwivedi, V. K. Srivastava, and R. L. Hall. Finite sample properties of ridge estimators. Technometrics, 22(2):205–212, May 1980. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1268460.

Dudewicz:1980:ROS

[Dud80] Edward J. Dudewicz. Ranking (ordering) and selection: An overview of how to select the best. Technometrics, 22(1):113–119, February 1980. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1268390.

Dudewicz:1982:BRB

[Dud82] Edward J. Dudewicz. Book review: Multiple Decision Procedures: Theory and Methodology of Selecting and Ranking Populations by Shanti S. Gupta; S. Panchapakesan. Technometrics, 24(2):162–163, May 1982. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1268498.

Dunn:1989:BRB

[Dun89] Charles Dunn. Book review: Applied Multivariate Statistical Analysis by Richard A. Johnson; Dean W. Wichern. Technometrics, 31(2):265–266, May 1989. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268834.

Denby:1985:SCM

[DV85] L. Denby and Y. Vardi. A short-cut method for estimation in renewal processes. Technometrics, 27(4):361–373, November 1985. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1270203.

Denby:1986:SCD

[DV86] L. Denby and Y. Vardi. The survival curve with decreasing density. Technometrics, 28(4):359–367, November 1986. CODEN TCMTA2.

ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1268985.

deVeaux:1989:AGT

[dVS89] Richard D. de Veaux and J. Michael Steele. ACE guidedtransformation method for estimation of the coefficient of soil-water diffusivity. Technometrics, 31(1):91–98, February 1989. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1270368.

DErrico:1988:STU

[DZ88] John R. D’Errico and Nicholas A. Zaino, Jr. Statistical tolerancing using a modification of Taguchi’s method. Technometrics, 30(4):397–405, November 1988. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1269802.

Easterling:1980:BRBb

[Eas80a] Robert G. Easterling. Book review: An Anatomy of Risk by William D. Rowe. Technometrics, 22(2):278–279, May 1980. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1268471.

Easterling:1980:BRBa

[Eas80b] Robert G. Easterling. Book review: On the Mathematics of Competing Risks by Z. W. Birnbaum. Technometrics, 22(1):131–132, February 1980. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1268399.

Easterling:1981:BAL

[Eas81] Robert G. Easterling. [A Bayesian Analysis of the Linear Calibration Problem]: Discussion. Technometrics, 23(4):343–344, November 1981. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1268223. See

[HL81b].

Easterling:1983:ER

[Eas83a] Robert G. Easterling. Editor’s report. Technometrics, 25(1): 107–108, February 1983. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267733.

Easterling:1983:NE

[Eas83b] Robert G. Easterling. New Editor. Technometrics, 25(2):iii, May 1983. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1268540.

Easterling:1983:TFY

[Eas83c] Robert G. Easterling. Twenty-five years of “Technometrics”. Technometrics, 25(1):1, February 1983. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267720.

Easterling:1985:BRB

[Eas85] Robert G. Easterling. Book review: Repairable Systems Reliability by Harold Ascher; Harry Feingold. Technometrics, 27(4):439–440, November 1985. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1270211.

Easterling:1987:PMI

[Eas87] Robert G. Easterling. [Performance Measures Independent of Adjustment]: Discussion. Technometrics, 29(3):267–269, August 1987. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1269333. See [LSK87a].

Engelhardt:1982:PLS

[EB82] Max Engelhardt and Lee J. Bain. On prediction limits for samples from a Weibull or extreme-value distribution. Technometrics, 24(2):147–150, May 1982. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268494.

Engelhardt:1981:IPB

[EBW81] Max Engelhardt, Lee J. Bain, and F. T. Wright. Inferences on the parameters of the Birnbaum–Saunders fatigue life distribution based on maximum likelihood estimation. Technometrics, 23(3):251–256, August 1981. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267788.

Eddy:1987:BRB

[Edd87] William F. Eddy. Book review: Statistical Computing in Pascal by D. Cooke; A. H. Craven; G. M. Clark. Technometrics, 29(1):123–124, February 1987. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1269901.

Easterling:1986:DDG

[EDGH86] Robert G. Easterling, Darryl Downing, Robert Gardner, and Owen Hoffman. Discussion of Downing, Gardner, and Hoffman (1985). Technometrics, 28(1):91–93, February 1986. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1269612. See [DGH85].

Evans:1984:ECB

[EJG84] James W. Evans, Richard A. Johnson, and David W. Green. Estimating the correlation between variables under destructive testing, or how to break the same board twice. Technometrics, 26 (3):285–290, August 1984. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267555.

Eastment:1982:CVC

[EK82] H. T. Eastment and W. J. Krzanowski. Cross-validatory choice of the number of components from a principal component analysis. Technometrics, 24(1):73–77, February 1982. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1267581.

Egerton:1981:SCS

[EL81] M. F. Egerton and P. J. Laycock. Some criticisms of stochastic shrinkage and ridge regression, with counterexamples. Technometrics, 23(2):155–159, May 1981. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/ stable/1268031. See corrigendum [EL83].

Egerton:1983:CSC

[EL83] M. F. Egerton and P. J. Laycock. Corrigendum: “Some criticisms of stochastic shrinkage and ridge regression, with counterexamples” [Technometrics 23 (1981), no. 2, 155–159; MR 82f:62130]. Technometrics, 25(3):304, August 1983. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268617. See [EL81].

Ellickson:1982:BRB

[Ell82] Bryan Ellickson. Book review: Statlib: A Statistical Computing Library by William M. Brelsford; Daniel A. Relles. Technometrics, 24(4):336–337, November 1982. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267832.

Elder:1982:AME

[EM82] Robert S. Elder and H. David Muse. An approximate method for evaluating mixed sampling plans. Technometrics, 24(3):207–211, August 1982. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1268680.

Evans:1980:BPL

[EN80] I. G. Evans and A. H. M. Nigm. Bayesian prediction for the left truncated exponential distribution. Technometrics, 22(2):201–204, May 1980. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1268459.

Engelhardt:1983:SMR

[Eng83] Max Engelhardt. [Statistical Methods in Reliability]: Discussion. Technometrics, 25(4):332, November 1983. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1267852. See [Law83a].

Enslein:1980:BRB

[Ens80] Kurt Enslein. Book review: Linear Statistical Models and Optimal Designs by Olaf Kraft. Technometrics, 22(1):131, February 1980. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1268396.

Elder:1980:PCS

[ETM80] Robert S. Elder, William O. Thompson, and Raymond H. Myers. Properties of composite sampling procedures. Technometrics, 22(2):179–186, May 1980. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268456.

Evans:1983:SMR

[Eva83] Ralph A. Evans. [Statistical Methods in Reliability]: Discussion. Technometrics, 25(4):333, November 1983. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1267853. See [Law83a].

Fairbanks:1988:TSL

[Fai88] Kenneth Fairbanks. A two-stage life test for the exponential parameter. Technometrics, 30(2):175–180, May 1988. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1270163.

Farebrother:1985:RAS

[Far85] R. W. Farebrother. Relations among subset estimators: A bibliographical note. Technometrics, 27(1):85–86, February 1985. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1270474.

Farebrother:1988:BRB

[Far88a] R. W. Farebrother. Book review: Counting for Something: Statistical Principles and Personalities by William S. Peters. Technometrics, 30(3):362–363, August 1988. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1270108.

Farnsworth:1988:BRB

[Far88b] David L. Farnsworth. Book review: A First Course in Statistics by James T. McClave; Frank H. Dietrich. Technometrics, 30(1):124, February 1988. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1270340.

Feder:1980:BRBc

[Fed80a] Paul I. Feder. Book review: Multivariate Analysis-V: Proceedings of the Fifth International Symposium on Multivariate Analysis by P. R. Krishnaiah. Technometrics, 22(3):447–448, August 1980. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1268338.

Feder:1980:BRBd

[Fed80b] Paul I. Feder. Book review: Proceedings of the Second Prague Symposium on Asymptotic Statistics by Petr Mandl; Marie Hušková. Technometrics, 22(3):448–449, August 1980. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1268339.

Feder:1980:BRBa

[Fed80c] Paul I. Feder. Book review: Time Series: Proceedings of the Intermational Conference Held at Nottingham University, March 1979 by O. D. Anderson. Technometrics, 22(3):446–447, August 1980. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1268336.

Feder:1980:BRBb

[Fed80d] Paul I. Feder. Book review: TIMS Studies in the Management Sciences, Vol. 13: The Implementation of Management Science by R. Doktor; R. L. Schultz; D. P. Slevin. Technometrics, 22(3):447, August 1980. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1268337.

Feder:1981:BRBl

[Fed81a] Paul I. Feder. Book review: Annals of Discrete Mathematics, Vol. 6: Combinatorial Mathematics, Optimal Designs and Their Applications by J. Srivastava. Technometrics, 23(4):422–423, November 1981. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1268243.

Feder:1981:BRBi

[Fed81b] Paul I. Feder. Book review: Applied Stochastic Control in Econometrics and Management Science by A. Bensoussan; P. Kleindorfer; C. S. Tapiero. Technometrics, 23(4):419, November 1981. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268240.

Feder:1981:BRBg

[Fed81c] Paul I. Feder. Book review: Developments in Statistics, Volume 3 by P. R. Krishnaiah. Technometrics, 23(3):313–314, August 1981. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1267806.

Feder:1981:BRBe

[Fed81d] Paul I. Feder. Book review: Forecasting Public Utilities: Proceedings of the International Conference Held at Nottingham University, March 1980 by O. D. Anderson. Technometrics, 23(2):210–211, May 1981. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1268055.

Feder:1981:BRBk

[Fed81e] Paul I. Feder. Book review: Handbook of Statistics, Volume 1: Analysis of Variance by P. R. Krishnaiah. Technometrics, 23 (4):420–422, November 1981. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268242.

Feder:1981:BRBd

[Fed81f] Paul I. Feder. Book review: Mathematical Statistics and Probability Theory by W. Klonecki; A. Kozek; J. Rosiński. Technometrics, 23(2):202–203, May 1981. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268044.

Feder:1981:BRBc

[Fed81g] Paul I. Feder. Book review: Multivariate Statistical Analysis by R. P. Gupta. Technometrics, 23(2):201–202, May 1981. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268043.

Feder:1981:BRBa

[Fed81h] Paul I. Feder. Book review: Numerical Techniques for Stochastic Systems by Francesco Archetti; Marco Cugiani. Technometrics, 23(2):199, May 1981. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268041.

Feder:1981:BRBh

[Fed81i] Paul I. Feder. Book review: Operations Research in Agriculture and Water Resources: Proceedings of the ORAGWA International Conference by Dan Yaron; Charles S. Tapiero. Technometrics, 23(3):314–315, August 1981. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267807.

Feder:1981:BRBb

[Fed81j] Paul I. Feder. Book review: Proceedings of the Computer Science and Statistics: 12th Annual Symposium on the Interface by Jane F. Gentleman. Technometrics, 23(2):199–201, May 1981. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268042.

Feder:1981:BRBf

[Fed81k] Paul I. Feder. Book review: Recent Developments in Statistical Inference and Data Analysis: Proceedings of the International Conference in Statistics in Tokyo by K. Matusita. Technometrics, 23(2):

211, May 1981. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1268056.

Feder:1981:BRBj

[Fed81l] Paul I. Feder. Book review: TIMS Studies in the Management Sciences, Vol. 15: Management of Research and Innovation by B. V. Dean; J. D. Golhar. Technometrics, 23(4):419–420, November 1981. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1268241.

Feder:1982:BRBb

[Fed82a] Paul I. Feder. Book review: Computer Scheduling of Public Transport: Urban Passenger Vehicle and Crew Scheduling by A. Wren. Technometrics, 24(1):87, February 1982. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1267593.

Feder:1982:BRBa

[Fed82b] Paul I. Feder. Book review: Point Processes and Queuing Problems, Colloquia Mathematica Societatis János Bolyai, Vol. 24 by P. Bártfai; J. Tomkó. Technometrics, 24(1):86–87, February 1982. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1267592.

Feder:1982:BRBc

[Fed82c] Paul I. Feder. Book review: Proceedings of the Gothenberg Symposium on Teaching Statistics by Lennart Råde. Technometrics, 24(2): 168, May 1982. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1268506.

Feder:1984:BRBb

[Fed84a] Paul I. Feder. Book review: Glossary and Tables for Statistical Quality Control by ASQC Statistics Division, American Society for Quality Control. Technometrics, 26(1):64, February 1984. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268418.

Feder:1984:BRBd

[Fed84b] Paul I. Feder. Book review: Handbook of Statistics, Volume 2, Classification, Pattern Recognition, and Reduction of Dimensionality by P. R. Krishnaiah; L. N. Kanal. Technometrics, 26(1):65–67, February 1984. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1268420.

Feder:1984:BRBe

[Fed84c] Paul I. Feder. Book review: Scientific Management of Transport Systems by N. K. Jaiswal. Technometrics, 26(2):194–195, May 1984. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1268123.

Feder:1984:BRBc

[Fed84d] Paul I. Feder. Book review: Statistics and Related Topics, Proceedings of the International Symposium on Statistics and Related Topics Held in Ottawa, Canada, May 1980 by M. Csörgö; D. A. Dawson; J. N. K. Rao; A. K. Md. E. Saleh. Technometrics, 26(1):64–65, February 1984. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1268419.

Feder:1984:BRBa

[Fed84e] Paul I. Feder. Book review: Time Series Analysis: Theory and Practice 2, Proceedings of the International Conference Held in Dublin, Ireland, March, 1982 by O. D. Anderson. Technometrics, 26(1):63–64, February 1984. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268417.

Fellner:1986:REV

[Fel86] William H. Fellner. Robust estimation of variance components. Technometrics, 28(1):51–60, February 1986. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1269603.

Ferris:1982:SMN

[Fer82] Yvonne M. Ferris. [Statistical Methods for Nuclear Materials Safeguards: An Overview]: Discussion. Technometrics, 24(4):275–276, November 1982. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1267815.

See [GPS82a].

Flack:1989:USE

[FF89] Virginia F. Flack and Rafael A. Flores. Using simulated envelopes in the evaluation of normal probability plots of regression residuals. Technometrics, 31(2):219–225, May 1989. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1268819.

Follmann:1988:DHD

[FG88] Dean A. Follmann and Matthew S. Goldberg. Distinguishing heterogeneity from decreasing hazard rates. Technometrics, 30 (4):389–396, November 1988. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1269801.

Fries:1980:MAD

[FH80] Arthur Fries and William G. Hunter. Minimum aberration 2^{k−p} designs. Technometrics, 22(4):601–608, November 1980. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268198.

Fox:1980:JNR

[FHL80] Terry Fox, David Hinkley, and Kinley Larntz. Jackknifing in nonlinear regression. Technometrics, 22(1):29–33, February 1980. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1268380. See corrigenda

[FHL81].

Fox:1981:CJN

[FHL81] Terry Fox, David Hinkley, and Kinley Larntz. Corrigenda: “Jackknifing in nonlinear regression” [Technometrics 22 (1980), no. 1, 29–33; MR 81i:62114]. Technometrics, 23(2):220, May 1981. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1268059. See [FHL80].

Fertig:1980:AAS

[FM80a] K. W. Fertig and Nancy R. Mann. An accurate approximation to the sampling distribution of the Studentized extreme value statistic. Technometrics, 22(1):83–97, February 1980. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1268387.

Fertig:1980:LTS

[FM80b] K. W. Fertig and Nancy R. Mann. Life-test sampling plans for two-parameter Weibull populations. Technometrics, 22(2):165–177, May 1980. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1268455.

Fung:1987:DM

[FM87] W. K. Fung and Mervyn G. Marasinghe. Discussion of Marasinghe (1985). Technometrics, 29(1):125, February 1987. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1269902. See [Mar85].

Fertig:1980:CPI

[FMM80] K. W. Fertig, M. Eugene Meyer, and Nancy R. Mann. On constructing prediction intervals for samples from a Weibull or extreme value distribution. Technometrics, 22(4):567–573, November 1980. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1268194.

Franklin:1984:CTM

[Fra84] M. F. Franklin. Constructing tables of minimum aberration p^{n−m} designs. Technometrics, 26(3):225–232, August 1984. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267548.

Franklin:1985:SDC

[Fra85] M. F. Franklin. Selecting defining contrasts and confounded effects in p^{n−m} factorial experiments. Technometrics, 27(2):165–172, May 1985. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1268764.

Freeny:1988:BRB

[Fre88] Anne Freeny. Book review: Empirical Model Building and Response Surfaces by G. E. P. Box; N. R. Draper. Technometrics, 30(2):229– 231, May 1988. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1270169.

Friedman:1983:DLR

[Fri83] Jerome H. Friedman. [Developments in Linear Regression Methodology: 1959–1982]: Discussion. Technometrics, 25(3):246–247, August 1983. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1268608. See

[Hoc83a].

Friedman:1989:FPSa

[FS89a] Jerome H. Friedman and Bernard W. Silverman. Flexible parsimonious smoothing and additive modeling. Technometrics, 31(1):3–21, February 1989. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1270359.

See discussion [Has89, Haw89] and response [FS89b].

Friedman:1989:FPSb

[FS89b] Jerome H. Friedman and Bernard W. Silverman. [Flexible Parsimonious Smoothing and Additive Modeling]: Response. Technometrics, 31(1):35–39, February 1989. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1270362. See [FS89a].

Friedman:1984:MSS

[FT84] Jerome Friedman and Robert Tibshirani. The monotone smoothing of scatterplots. Technometrics, 26(3):243–250, August 1984. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1267550.

Ford:1989:RAN

[FTK89] Ian Ford, D. M. Titterington, and Christos P. Kitsos. Recent advances in nonlinear experimental design. Technometrics, 31 (1):49–60, February 1989. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1270364.

Fulkerson:1989:BRB

[Ful89] William F. Fulkerson. Book review: Statistics by Example by Terry Sincich. Technometrics, 31(4):496–497, November 1989. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1270016.

Galbraith:1988:GDE

[Gal88] R. F. Galbraith. Graphical display of estimates having differing standard errors. Technometrics, 30(3):271–281, August 1988. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1270081.

Gans:1981:CET

[Gan81] Daniel J. Gans. Corrected and extended tables for Tukey’s quick test. Technometrics, 23(2):193–195, May 1981. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1268038.

Gart:1985:BRB

[Gar85] John J. Gart. Book review: Linear Statistical Analysis of Discrete Data by Mikel Aickin. Technometrics, 27(3):316–317, August 1985.

CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1269715.

Gaudard:1988:BRB

[Gau88] Marie Gaudard. Book review: Poor-Quality Cost by H. James Harrington. Technometrics, 30(3):358, August 1988. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1270101.

Gbur:1981:BRB

[Gbu81] Edward Gbur. Book review: Probability and Statistics in Engineering and Management Science by William W. Hines; Douglas C. Montgomery. Technometrics, 23(2):205–206, May 1981. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268049.

Gbur:1983:BRB

[Gbu83] Edward E. Gbur. Book review: Statistical Modeling Techniques by Samuel S. Shapiro; Alan J. Gross. Technometrics, 25(1):115–116, February 1983. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1267743.

Gorman:1982:NMR

[GC82] John W. Gorman and John A. Cornell. A note on model reduction for experiments with both mixture components and process variables. Technometrics, 24(3):243–247, August 1982. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268685.

Gorman:1985:NFE

[GC85] John W. Gorman and John A. Cornell. A note on fitting equations to freezing-point data exhibiting eutectics for binary and ternary mixture systems. Technometrics, 27(3):229–239, August 1985. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1269704.

Giltinan:1986:SNE

[GCR86] David M. Giltinan, Raymond J. Carroll, and David Ruppert. Some new estimation methods for weighted regression when there are possible outliers. Technometrics, 28(3):219–230, August 1986. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1269077.

Gentleman:1980:FML

[Gen80] Jane F. Gentleman. Finding the K most likely outliers in two-way tables. Technometrics, 22(4):591–600, November 1980. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268197.

Grieve:1987:DPG

[GGP87] A. P. Grieve, Beth C. Gladen, and Walter W. Piegorsch. Discussion of Piegorsch and Gladen (1986). Technometrics, 29(4):504–506, November 1987. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1269486.

See [PG86].

Gupta:1980:SSP

[GH80] Shanti S. Gupta and Jason C. Hsu. Subset selection procedures with application to motor vehicle fatality data in a two-way layout. Technometrics, 22(4):543–546, November 1980. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1268191.

Galpin:1981:RSO

[GH81] Jacqueline S. Galpin and Douglas M. Hawkins. Rejection of a single outlier in two- or three-way layouts. Technometrics, 23(1):65–70, February 1981. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1267977.

Ghosh:1988:BRB

[Gho88] Subir Ghosh. Book review: Statistical Analysis with Missing Data by R. Little; D. Rubin. Technometrics, 30(4):455, November 1988. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1269814.

Gibbs:1983:BRB

[Gib83] Paul H. Gibbs. Book review: Introduction to Linear Regression Analysis by Douglas C. Montgomery; Elizabeth A. Peck. Technometrics, 25(4):388, November 1983. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267869.

Gibson:1988:BRB

[Gib88] Martin G. Gibson. Book review: SPSS-X Introductory Statistics Guide by Marija J. Norusis. Technometrics, 30(2):237, May 1988.

CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1270180.

Guttman:1988:CLS

[GJBR88] I. Guttman, R. A. Johnson, G. K. Bhattacharyya, and B. Reiser. Confidence limits for stress-strength models with explanatory variables. Technometrics, 30(2):161–168, May 1988. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1270161.

Giovannitti-Jensen:1989:GAP

[GJM89] Ann Giovannitti-Jensen and Raymond H. Myers. Graphical assessment of the prediction capability of response surface designs. Technometrics, 31(2):159–171, May 1989. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1268814.

Galil:1980:TSS

[GK80a] Z. Galil and J. Kiefer. Time- and space-saving computer methods, related to Mitchell’s DETMAX, for finding D-optimum designs. Technometrics, 22(3):301–313, August 1980. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1268314.

Guttman:1980:RSO

[GK80b] Irwin Guttman and Charles H. Kraft. Robustness to spurious observations of linearized Hodges–Lehmann estimators and Anscombe estimators. Technometrics, 22(1):55–63, February 1980. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268383.

Gray:1984:CDTa

[GL84a] J. Brian Gray and Robert F. Ling. K-clustering as a detection tool for influential subsets in regression. Technometrics, 26(4):305–318, November 1984. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1269490.

See discussion [All84, Hoc84, Wei84] and response [GL84b].

Gray:1984:CDTb

[GL84b] J. Brian Gray and Robert F. Ling. [K-Clustering as a Detection Tool for Influential Subsets in Regression]: Response. Technometrics, 26(4):326–330, November 1984. CODEN TCMTA2.

ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1269494. See [GL84a].

Glasbey:1982:GPA

[Gla82] C. A. Glasbey. A generalization of partial autocorrelations useful in identifying ARMA models. Technometrics, 24(3):223–228, August 1982. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1268682.

Gleser:1984:BRB

[Gle84] Leon Jay Gleser. Book review: Aspects of Multivariate Statistical Theory by Robb J. Muirhead. Technometrics, 26(2):191–192, May 1984. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1268117.

Gibbons:1984:RIR

[GM84] Diane I. Gibbons and Gary C. McDonald. A rational interpretation of the ridge trace. Technometrics, 26(4):339–346, November 1984. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1269496.

Gastwirth:1986:ERN

[GM86] Joseph L. Gastwirth and Hosam Mahmoud. An efficiency robust nonparametric test for scale change for data from a gamma distribution. Technometrics, 28(1):81–84, February 1986. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1269607.

Gaver:1987:REB

[GM87] Donald P. Gaver and I. G. O’ Muircheartaigh. Robust empirical Bayes analyses of event rates. Technometrics, 29(1):1–15, February 1987. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1269878.

Gupta:1982:OME

[GND82] V. K. Gupta, A. K. Nigam, and A. Dey. Orthogonal main-effect plans for asymmetrical factorials. Technometrics, 24(2):135–137, May 1982. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1268491.

Gokhale:1981:BRB

[Gok81] D. V. Gokhale. Book review: Large Deviations and Asymptotic Efficiencies by P. Groeneboom. Technometrics, 23(2):204–205, May 1981. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1268048.

Gokhale:1983:BRB

[Gok83] D. V. Gokhale. Book review: Analysis of Categorical Data: Dual Scaling and Its Applications by S. Nishisato. Technometrics, 25(1):112, February 1983. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267740.

Goldman:1988:BRB

[Gol88] A. Goldman. Book review: Probability and Statistics for Engineering and the Sciences by Jay L. Devore. Technometrics, 30(2):235, May 1988. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1270176.

Goodall:1987:BRB

[Goo87] Colin Goodall. Book review: Modern Multivariate Statistical Analysis: A Graduate Course and Handbook by Minoru Siotani; Takesi Hayakawa; Yasunori Fujikoshi. Technometrics, 29(2):242–243, May 1987. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1269783.

Goodall:1988:BRB

[Goo88] Colin Goodall. Book review: Principal Component Analysis by I. T. Jolliffe. Technometrics, 30(3):351–352, August 1988. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1270093.

Goss:1986:BRB

[Gos86] Thomas I. Goss. Book review: Directory of Statistical Microcomputer Software by Wayne A. Woodward; Alan C. Elliott; H. L. Gray. Technometrics, 28(2):178–179, May 1986. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1270455.

Goldman:1982:SMNa

[GPS82a] Aaron S. Goldman, Richard R. Picard, and James P. Shipley. Statistical methods for nuclear materials safeguards: An overview. Technometrics, 24(4):267–275, November 1982. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www. jstor.org/stable/1267814. See discussion [Fer82, LD82, AL82, Mog82, Mor82, Jae82, Moo82] and response [GPS82b].

Goldman:1982:SMNb

[GPS82b] Aaron S. Goldman, Richard R. Picard, and James P. Shipley. [Statistical Methods for Nuclear Materials Safeguards: An Overview]: Response. Technometrics, 24(4):291–294, November 1982. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267822. See [GPS82a].

Grambsch:1987:BRB

[Gra87a] Patricia Grambsch. Book review: Series of Irregular Observations: Forecasting and Model Building by Robert Azencott; Didier Dacunha-Castelle. Technometrics, 29(4):496–497, November 1987. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1269477.

Gray:1987:BRB

[Gra87b] J. Brian Gray. Book review: Classification Algorithms by Mike James. Technometrics, 29(4):499, November 1987. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1269479.

Griffith:1988:BRB

[Gri88] Daniel A. Griffith. Book review: Statistical Techniques in Geographical Analysis by Gareth Shaw; Dennis Wheeler. Technometrics, 30(3):348–349, August 1988. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1270091.

Grice:1989:BRB

[Gri89] John V. Grice. Book review: Graphical Exploratory Data Analysis by S. du Toit; S. Steyn; R. Stumpf. Technometrics, 31(1):116–117, February 1989. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1270374.

Groggel:1989:BRB

[Gro89] David J. Groggel. Book review: Statistics Explained: Basic Concepts and Methods by R. Kapadia; G. Andersson. Technometrics, 31(3):386–387, August 1989. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/3556155.

Gruber:1988:BRBb

[Gru88a] Marvin Gruber. Book review: A Topical Dictionary of Statistics by Gary L. Tietjen. Technometrics, 30(2):240–241, May 1988. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1270184.

Gruber:1988:BRBa

[Gru88b] Marvin H. J. Gruber. Book review: Basic Statistics: An Inferential Aproach by Frank Dietrich; Thomas J. Kearns. Technometrics, 30(1):124–125, February 1988. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1270341.

Grassia:1982:SPC

[GS82] A. Grassia and R. Sundberg. Statistical precision in the calibration and use of sorting machines and other classifiers. Technometrics, 24(2):117–121, May 1982. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268488.

Gunter:1983:BRB

[Gun83] B. Gunter. Book review: Guide to Quality Control, Second Revised English Edition by Kaoru Ishikawa. Technometrics, 25(4):384–385, November 1983. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1267864.

Gunter:1984:BRB

[Gun84] Bert Gunter. Book review: The Human Side of Statistical Consulting by James R. Boen; Douglas A. Zahn. Technometrics, 26(2):190– 191, May 1984. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1268116.

Gunter:1985:BRBa

[Gun85a] Bert Gunter. Book review: Computing in Statistical Science Through APL by Francis Anscombe. Technometrics, 27(2):209–210, May 1985. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1268772.

Gunter:1985:BRBb

[Gun85b] Bert Gunter. Book review: Design and Analysis of Experiments by Douglas C. Montgomery. Technometrics, 27(3):313–315, August 1985. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1269713.

Gunter:1985:BRBc

[Gun85c] Bert Gunter. Book review: Japanese Manufacturing Techniques: Nine Hidden Lessons in Simplicity by Richard Schonberger. Technometrics, 27(4):440–442, November 1985. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1270213.

Gunst:1987:BRB

[Gun87] Richard F. Gunst. Book review: General Statistics by Warren Chase; Fred Brown. Technometrics, 29(3):381, August 1987. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1269351.

Gunst:1988:BRB

[Gun88a] Richard F. Gunst. Book review: Regression Analysis with Applications by G. Barrie Wetherill. Technometrics, 30(4):451–452, November 1988. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1269810.

Gunter:1988:SNR

[Gun88b] Berton Gunter. [Signal-to-Noise Ratios, Performance Criteria, and Transformations]: Discussion. Technometrics, 30(1):32–35, February 1988. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1270316. See

[Box88a].

Gupta:1982:BRB

[Gup82] A. K. Gupta. Book review: The Theory of Linear Models and Multivariate Analysis by Steven F. Arnold. Technometrics, 24(3):250, August 1982. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1268687.

Gupta:1983:BRB

[Gup83] A. K. Gupta. Book review: The Foundations of Multivariate Analysis by K. Takeuchi; M. Yanai; B. N. Mukherjee. Technometrics, 25(4):388–389, November 1983. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267870.

Gupta:1989:BRB

[Gup89] Sudhir C. Gupta. Book review: Linear Estimation and Design of Experiments by D. D. Joshi. Technometrics, 31(3):392–393, August 1989. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/3556164.

Gabriel:1979:LRA

[GZ79] K. Ruben Gabriel and S. Zamir. Lower rank approximation of matrices by least squares with any choice of weights. Technometrics, 21(4):489–498, November 1979. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/ stable/1268288. See corrigenda [Ano80e].

Hackney:1980:BRB

[Hac80] Olga P. Hackney. Book review: Statistical Treatment of Experimental Data by J. R. Green; D. Margerison. Technometrics, 22(1):131, February 1980. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1268397.

Hahn:1981:BRB

[Hah81] Gerald J. Hahn. Book review: Tables for Normal Tolerance Limits, Sampling Plans and Screening by Robert E. Odeh; D. B. Owen. Technometrics, 23(1):114–115, February 1981. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1267990.

Hahn:1982:BRB

[Hah82] Gerald J. Hahn. Book review: Applied Regression Analysis by Norman Draper; Harry Smith. Technometrics, 24(4):337–339, November 1982. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1267833.

Hahn:1984:EDC

[Hah84a] Gerald J. Hahn. Experimental design in the complex world. Technometrics, 26(1):19–31, February 1984. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://

www.jstor.org/stable/1268412. See discussion [Pik84, Hah84b,

Cox84, Atk84, Cor84, Joi84, Bec84, Zie84] and response [SH84b].

Hahn:1984:EDR

[Hah84b] Gerald J. Hahn. [Experimental Design: Review and Comment]: Discussion. Technometrics, 26(2):110–115, May 1984. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268100. See [Hah84a].

Haines:1987:AAA

[Hai87a] Linda M. Haines. The application of the annealing algorithm to the construction of exact optimal designs for linear-regression models. Technometrics, 29(4):439–447, November 1987. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1269455.

Haining:1987:TSM

[Hai87b] Robert Haining. Trend-surface models with regional and local scales of variation with an application to aerial survey data. Technometrics, 29(4):461–469, November 1987. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1269458.

Hampel:1985:BPM

[Ham85] Frank R. Hampel. The breakdown points of the mean combined with some rejection rules. Technometrics, 27(2):95–107, May 1985. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1268758.

Hannan:1984:BRB

[Han84a] Ed Hannan. Book review: Theory and Practice of Multiple Criteria Decision Making, Collection of Papers Presented at a Workshop on Multicriteria Decision Making, Held in Moscow, May 1981 by C. Carlsson; Y. Kochetkov. Technometrics, 26(2):193–194, May 1984. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1268120.

Hannay:1984:BRBa

[Han84b] Dave Hannay. Book review: The Architecture of High Performance Computers by Roland N. Ibbett. Technometrics, 26(2):194, May 1984. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1268121.

Hannay:1984:BRBb

[Han84c] David Hannay. Book review: Microprocessors and Microcomputer Development Systems by Mohamed Rafiquzzaman. Technometrics, 26(3):291–292, August 1984. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267557.

Hannan:1985:BRBb

[Han85a] Edward L. Hannan. Book review: Applications of Queueing Theory by G. F. Newell. Technometrics, 27(3):320, August 1985. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1269721.

Hannan:1985:BRBa

[Han85b] Edward L. Hannan. Book review: Fuzzy Sets and Decision Analysis, TIMS Studies in the Management Sciences (Vol. 20) by H.- J. Zimmermann; L. A. Zadeh; B. R. Gaines. Technometrics, 27(3):319–320, August 1985. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1269720.

Hansen:1986:BRB

[Han86] J. L. Hansen. Book review: Experiments in Industry, Design, Analysis, and Interpretation of Results by R. E. Snee; L. B. Hare; J. R. Trout. Technometrics, 28(4):402, November 1986. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268992.

Hansen:1988:BRB

[Han88] J. L. Hansen. Book review: Engineering Statistics by R. V. Hogg; J. Ledolter. Technometrics, 30(3):347–348, August 1988. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1270089.

Hand:1989:BRB

[Han89] D. J. Hand. Book review: Coordinate-Free Multivariate Statistics: An Illustrated Progression from Halmos to Gauss and Bayes by Mervyn Stone. Technometrics, 31(2):266, May 1989. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268835.

Hartigan:1980:BRB

[Har80] John A. Hartigan. Book review: Some Basic Theory for Statistical Inference by E. J. G. Pitman. Technometrics, 22(4):633–634, November 1980. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1268210.

Harter:1983:BRB

[Har83] H. Leon Harter. Book review: Encyclopedia of Statistical Sciences, Volume 1 by Samuel Kotz; Norman L. Johnson. Technometrics, 25(4):387–388, November 1983. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267867.

Harris:1986:BRB

[Har86] Thomas Harris. Book review: Comparative Models for Electrical Load Forecasting by D. W. Bunn; E. D. Farmer. Technometrics, 28(4):403–404, November 1986. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268996.

Harris:1987:BRB

[Har87] Bernard Harris. Book review: Making Decisions by D. V. Lindley. Technometrics, 29(2):247–248, May 1987. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1269790.

Hardin:1989:BRBa

[Har89a] J. Michael Hardin. Book review: Simulation, Knowledge-Based Computing and Fuzzy Statistics by Constantin Virgil Negoita; Dan Ralescu. Technometrics, 31(1):129–130, February 1989. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1270392.

Hardin:1989:BRBb

[Har89b] J. Michael Hardin. Book review: The Art and Techniques of Simulation by Mrudulla Gnanadesikan; Richard L. Scheaffer; Jim Swift. Technometrics, 31(4):483, November 1989. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1270003.

Hassett:1988:BRB

[Has88] Nancy C. Hassett. Book review: A Handbook of Introductory Statistical Methods by C. Philip Cox. Technometrics, 30(2):233–234, May 1988. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1270174.

Hastie:1989:FPS

[Has89] Trevor Hastie. [Flexible Parsimonious Smoothing and Additive Modeling]: Discussion. Technometrics, 31(1):23–29, February 1989. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1270360. See [FS89a].

Haugh:1987:BRB

[Hau87] Larry D. Haugh. Book review: Handbook of Statistics 5: Time Series in the Time Domain by E. J. Hannon; P. R. Krishnaiah; M. M. Rao. Technometrics, 29(3):388–389, August 1987. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1269363.

Haugh:1989:BRB

[Hau89] Larry D. Haugh. Book review: Time Series Forecasting: Unified Concepts and Computer Implementation by Bruce L. Bowerman; Richard T. O’Connell. Technometrics, 31(2):264–265, May 1989. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1268833.

Hawkins:1980:NCD

[Haw80] Douglas M. Hawkins. A note on continuous and discontinuous segmented regressions. Technometrics, 22(3):443–444, August 1980. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1268331.

Hawkins:1981:NTM

[Haw81] Douglas M. Hawkins. A new test for multivariate normality and homoscedasticity. Technometrics, 23(1):105–110, February 1981. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1267983.

Hawkins:1983:OD

[Haw83] Douglas M. Hawkins. [Outlier. . . . . . . . . s]: Discussion. Technometrics, 25(2):155–156, May 1983. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268544. See [BC83a].

Hawkins:1989:FPS

[Haw89] Douglas M. Hawkins. [Flexible Parsimonious Smoothing and Additive Modeling]: Discussion. Technometrics, 31(1):31–34, February 1989. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1270361. See

[FS89a].

Hawkins:1984:LSO

[HBK84] Douglas M. Hawkins, Dan Bradu, and Gordon V. Kass. Location of several outliers in multiple-regression data using elemental sets.

Technometrics, 26(3):197–208, August 1984. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1267545.

Hawkins:1982:CFI

[HE82] Douglas M. Hawkins and W. J. R. Eplett. The Cholesky factorization of the inverse correlation or covariance matrix in multiple regression. Technometrics, 24(3):191–198, August 1982. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268678.

Hearne:1982:BRB

[Hea82] F. Terry Hearne. Book review: BMDP-79 Biomedical Computer Programs, P Series by W. J. Dixon; M. B. Brown. Technometrics, 24(1):82, February 1982. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267584.

Healy:1987:NMC

[Hea87] John D. Healy. A note on multivariate CUSUM procedures. Technometrics, 29(4):409–412, November 1987. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1269451.

Healy:1989:BRB

[Hea89] John Healy. Book review: Selected Tables in Mathematical Statistics (Vol. 11) by R. Odeh; J. Davenport; N. Pearson. Technometrics, 31(3):397, August 1989. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/3556171.

Heil:1980:BRBb

[Hei80a] Peter F. Heil. Book review: Survival Probabilities: The Goal of Risk Theory by Hilary L. Seal. Technometrics, 22(3):445–446, August 1980. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1268334.

Heil:1980:BRBa

[Hei80b] Peter F. Heil. Book reviews: Basic Statistics for the Behavioral Sciences by Kenneth D. Hopkins; Gene V. Glass; Biostatistics: A Foundation for Analysis in the Health Sciences by Wayne W. Daniel. Technometrics, 22(2):280, May 1980. CODEN TCMTA2.

ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1268475.

Helton:1987:BRB

[Hel87] Jon C. Helton. Book review: Numerical Methods in Engineering and Science by Carl E. Pearson. Technometrics, 29(2):248–249, May 1987. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1269792.

Henderson:1986:CPP

[Hen86] Robin Henderson. Change-point problem with correlated observations, with an application in material accountancy. Technometrics, 28(4):381–389, November 1986. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268987.

Hertsgaard:1989:BRB

[Her89] Doris Hertsgaard. Book review: Randomization Tests by Eugene Edgington. Technometrics, 31(1):127, February 1989. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1270387.

Hocking:1989:VCE

[HGB89] R. R. Hocking, J. W. Green, and R. H. Bremer. Variancecomponent estimation with model-based diagnostics. Technometrics, 31(2):227–239, May 1989. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268820.

Hall:1981:CDO

[HH81] James E. Hall and M. Zia Hassan. On the cumulative distribution of outgoing quality: A new criterion for sampling plans. Technometrics, 23(4):395–400, November 1981. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1268230.

Hora:1987:BAL

[HI87] Stephen C. Hora and Ronald L. Iman. Bayesian analysis of learning in risk analyses. Technometrics, 29(2):221–228, May 1987. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1269777.

Hidiroglou:1986:BRB

[Hid86] Michael A. Hidiroglou. Book review: Introduction to Variance Estimation by Kirk M. Wolter. Technometrics, 28(3):277–278, August 1986. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1269086.

Hill:1980:ODP

[Hil80] Peter D. H. Hill. D-optimal designs for partially nonlinear regression models. Technometrics, 22(2):275–276, May 1980. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268469.

Hill:1981:BAL

[Hil81] Bruce M. Hill. [A Bayesian Analysis of the Linear Calibration Problem]: Discussion. Technometrics, 23(4):335–338, November 1981. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1268220. See [HL81b].

Hintze:1980:UEA

[Hin80] J. L. Hintze. On the use of “elemental analysis” in multivariate variable selection. Technometrics, 22(4):609–612, November 1980. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1268199.

Hinkelmann:1987:BRB

[Hin87] Klaus Hinkelmann. Book review: Exploring Data by James M. Landwehr; Ann E. Watkins. Technometrics, 29(3):380–381, August 1987. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1269350.

Hixson:1987:BRB

[Hix87] Marilyn M. Hixson. Book review: Statistics: A First Course by John E. Freund; Richard Manning. Technometrics, 29(4):493–494, November 1987. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1269473.

Hunter:1986:UMM

[HJ86] William G. Hunter and Andrzej P. Jaworski. A useful method for model-building II: Synthesizing response functions from individual components. Technometrics, 28(4):321–327, November 1986. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1268981.

Hill:1980:REA

[HJF80] R. Carter Hill, George G. Judge, and Thomas B. Fomby. Is the regression equation adequate?: A reply. Technometrics, 22 (1):127–128, February 1980. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268393. See [SD80].

Hjorth:1980:RDI

[Hjo80] Urban Hjorth. A reliability distribution with increasing, decreasing, constant and bathtub-shaped failure rates. Technometrics, 22(1):99–107, February 1980. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268388.

Hunter:1981:BALb

[HL81a] William G. Hunter and Warren F. Lamboy. [A Bayesian Analysis of the Linear Calibration Problem]: Response. Technometrics, 23(4):344–350, November 1981. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268224. See [HL81b].

Hunter:1981:BALa

[HL81b] William G. Hunter and Warren F. Lamboy. A Bayesian analysis of the linear calibration problem. Technometrics, 23(4):323–328, November 1981. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1268217.

See discussion [RS81, Law81, Hil81, Lwi81, Orb81, Eas81], response [HL81a], and corrigendum [HL84].

Hunter:1984:CBA

[HL84] W. G. Hunter and W. F. Lamboy. Corrigendum: “A Bayesian analysis of the linear calibration problem” [Technometrics 23 (1981), no. 4, 323–350; MR0697694 (84f:62045)]. Technometrics, 26(1):69, February 1984. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1268422.

See [HL81b].

Hadi:1985:CDI

[HLG85] Ali S. Hadi, Robert F. Ling, and J. Brian Gray. K-clustering and the detection of influential subsets. Technometrics, 27(3):323–325, August 1985. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1269724.

Hedayat:1984:OIB

[HM84] A. S. Hedayat and Dibyen Majumdar. A-optimal incomplete block designs for control-test treatment comparisons. Technometrics, 26 (4):363–370, November 1984. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1269499.

Hahn:1981:AFF

[HMS81] Gerald J. Hahn, Carolyn B. Morgan, and Josef Schmee. The analysis of a fractional factorial experiment with censored data using iterative least squares. Technometrics, 23(1):33–36, February 1981. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1267973.

Harris:1980:STA

[HMW80] T. J. Harris, J. F. MacGregor, and J. D. Wright. Self-tuning and adaptive controllers: An application to catalytic reactor control. Technometrics, 22(2):153–164, May 1980. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1268454.

Hocking:1983:DLRa

[Hoc83a] R. R. Hocking. Developments in linear regression methodology: 1959–1982. Technometrics, 25(3):219–230, August 1983. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268603. See discussion [Sne83,

Mit83, Mit83, Fri83] and response [Hoc83b].

Hocking:1983:DLRb

[Hoc83b] R. R. Hocking. [Developments in Linear Regression Methodology: 1959–1982]: Response. Technometrics, 25(3):248–249, August 1983. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1268609. See [Hoc83a].

Hocking:1984:CDT

[Hoc84] R. R. Hocking. [K-Clustering as a Detection Tool for Influential Subsets in Regression]: Discussion. Technometrics, 26(4):321–323, November 1984. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1269492.

See [GL84a].

Hodgson:1987:BRB

[Hod87] Vincent Hodgson. Book review: Empirical Processes with Applications to Statistics by Galen R. Shorack; Jon A. Wellner. Technometrics, 29(4):499, November 1987. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1269480.

Hodgson:1989:BRB

[Hod89] Vincent Hodgson. Book review: Probability Modeling and Computer Simulation by Norman S. Matloff. Technometrics, 31(3):394–395, August 1989. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/3556168.

Hoerl:1987:ART

[Hoe87] Roger W. Hoerl. The application of ridge techniques to mixture data: Ridge analysis. Technometrics, 29(2):161–172, May 1987. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1269771.

Hogg:1983:OD

[Hog83] Robert V. Hogg. [Outlier. . . . . . . . . s]: Discussion. Technometrics, 25(2):158–159, May 1983. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268546. See [BC83a].

Horowitz:1980:EVNa

[Hor80a] Joel Horowitz. Extreme values from a nonstationary stochastic process: An application to air quality analysis. Technometrics, 22(4):469–478, November 1980. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/ stable/1268183. See discussion [Cle80].

Horowitz:1980:EVNb

[Hor80b] Joel Horowitz. [Extreme Values from a Nonstationary Stochastic Process: An Application to Air Quality Analysis]: Response. Technometrics, 22(4):482, November 1980. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1268185.

Horn:1982:BRB

[Hor82] Susan D. Horn. Book review: Sampling for Health Professionals by Paul S. Levy; Stanley Lemeshow. Technometrics, 24(3):253, August 1982. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1268693.

Hora:1986:BRB

[Hor86] Stephen C. Hora. Book review: Statistical Inference Based on Ranks by Thomas P. Hettmansperger. Technometrics, 28(2):179–180, May 1986. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1270456.

Hahn:1988:CIV

[HR88] G. J. Hahn and T. E. Raghunathan. Combining information from various sources: A prediction problem and other industrial applications. Technometrics, 30(1):41–52, February 1988. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1270319.

Hoerl:1986:SBE

[HSH86] Roger W. Hoerl, John H. Schuenemeyer, and Arthur E. Hoerl. A simulation of biased estimation and subset selection regression techniques. Technometrics, 28(4):369–380, November 1986. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268986.

Hsieh:1980:LET

[Hsi80] Hui Kuang Hsieh. Letter to the editor: “On a test for the equality of two exponential distributions” [Technometrics 15 (1973), 177– 182; MR 51 #9312] by D. G. Weinman, G. Dugger, W. E. Franck and J. E. Hewett. Technometrics, 22(2):284, May 1980. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268481.

Hsieh:1981:TET

[Hsi81] Hui Kuang Hsieh. On testing the equality of two exponential distributions. Technometrics, 23(3):265–269, August 1981. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267790.

Hsieh:1986:ETC

[Hsi86] H. K. Hsieh. An exact test for comparing location parameters of k exponential distributions with unequal scales based on Type II censored data. Technometrics, 28(2):157–164, May 1986. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1270452.

Hsiang:1987:BRB

[Hsi87] Thomas C. Hsiang. Book review: Statistical Process Control: A Guide for Implementation by Roger W. Berger; Thomas H. Hart. Technometrics, 29(4):491–492, November 1987. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1269468.

Hubele:1989:BRB

[Hub89] Norma Faris Hubele. Book review: Introduction to Optimization Methods and Their Application in Statistics by B. S. Everitt. Technometrics, 31(4):484, November 1989. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1270004.

Hunter:1983:BJ

[Hun83] J. Stuart Hunter. The birth of a journal. Technometrics, 25(1):3–7, February 1983. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1267721.

Hurvich:1986:DDS

[Hur86] Clifford M. Hurvich. Data-dependent spectral Windows: Generalizing the classical framework to include maximum entropy estimates. Technometrics, 28(3):259–268, August 1986. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1269082.

Hamilton:1985:QDC

[HW85] David C. Hamilton and Donald G. Watts. A quadratic design criterion for precise estimation in nonlinear regression models. Technometrics, 27(3):241–250, August 1985. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1269705.

Hamada:1986:DSA

[HW86] M. Hamada and C. F. J. Wu. Discussion: Should accumulation analysis and related methods be used for industrial experiments? Technometrics, 28(4):302–306, November 1986. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1268977.

Hosking:1987:PQE

[HW87] J. R. M. Hosking and J. R. Wallis. Parameter and quantile estimation for the generalized Pareto distribution. Technometrics, 29(3):339–349, August 1987. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1269343.

Hinkley:1988:MAT

[HW88] D. V. Hinkley and S. Wang. More about transformations and influential cases in regression. Technometrics, 30(4):435–440, November 1988. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1269807.

Hosking:1985:EGE

[HWW85] J. R. M. Hosking, J. R. Wallis, and E. F. Wood. Estimation of the generalized extreme-value distribution by the method of probability-weighted moments. Technometrics, 27(3):251–261, August 1985. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1269706.

Hansen:1980:AVE

[HZ80] James N. Hansen and Scott Zeger. The asymptotic variance of the estimated proportion truncated from a normal population. Technometrics, 22(2):271–274, May 1980. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268468.

Iman:1979:URT

[IC79] Ronald L. Iman and W. J. Conover. The use of the rank transform in regression. Technometrics, 21(4):499–509, November 1979. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1268289. See corrigendum

[Ano84f].

Iman:1987:MTC

[IC87] Ronald L. Iman and W. J. Conover. A measure of top-down correlation. Technometrics, 29(3):351–357, August 1987. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1269344. See correction [IC89].

Iman:1989:CMT

[IC89] Ronald L. Iman and W. J. Conover. Correction: “A Measure of Top-Down Correlation” [Technometrics 39 (1988), no. 3, 351– 357]. Technometrics, 31(1):133, February 1989. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1270396. See [IC87].

Iglewicz:1988:BRB

[Igl88] Boris Iglewicz. Book review: Misused Statistics: Straight Talk for Twisted Numbers by A. J. Jaffe; Herbert F. Spirer. Technometrics, 30(1):121, February 1988. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1270336.

Iman:1986:BRB

[Ima86] Ronald L. Iman. Book review: Nonparametric Methods for Quantitative Analysis by Jean Dickinson Gibbons. Technometrics, 28 (3):275–277, August 1986. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1269085.

Inglis:1981:BRB

[Ing81] James Inglis. Book review: Specification Searches: Ad Hoc Inference with Nonexperimental Data by Edward E. Leamer. Technometrics, 23(1):112–113, February 1981. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267987.

Iyer:1987:BRB

[Iye87] Hari Iyer. Book review: Random Data by Julius Bendat; Allan Piersol. Technometrics, 29(3):381–382, August 1987. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1269352.

Izenman:1987:BRB

[Ize87] Alan Julian Izenman. Book review: Introduction to Random Processes, with Applications to Signals and Systems by William A. Gardner. Technometrics, 29(2):245–246, May 1987. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1269786.

Izenman:1988:BRB

[Ize88] Alan Julian Izenman. Book review: An Introduction to Kalman Filtering with Applications by Kenneth S. Miller; Donald M. Leskiw. Technometrics, 30(2):243–244, May 1988. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1270190.

Izenman:1989:BRB

[Ize89] Alan Julian Izenman. Book review: TIMESLAB: A Time Series Analysis Laboratory by H. Joseph Newton. Technometrics, 31(3):380–382, August 1989. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/3556149.

J:1980:BRBb

[J.80a] J. E. J. Book review: Distribution-Free Methods for Non-Parametric Problems: A Classified and Selected Bibliography by Bernard Singer. Technometrics, 22(2):280–281, May 1980. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1268476.

J:1980:BRBc

[J.80b] J. E. J. Book review: Forecasting by O. D. Anderson. Technometrics, 22(2):281, May 1980. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268477.

J:1980:BRBe

[J.80c] J. E. J. Book review: Game Theory and Related Topics by O. Moeschlin; D. Pallaschke. Technometrics, 22(2):282, May 1980. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1268479.

J:1980:BRBa

[J.80d] J. E. J. Book review: Measures of Association for Cross Classifications by Leo A. Goodman; William H. Kruskal. Technometrics, 22(2):280, May 1980. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268474.

J:1980:BRBd

[J.80e] J. E. J. Book review: The Computer Age by Michael L. Dertougos; Joel Moses. Technometrics, 22(2):281–282, May 1980. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268478.

Jackson:1980:BRBb

[Jac80a] J. Edward Jackson. Book review: Optimizing Methods in Statistics by Jagdish S. Rustagi. Technometrics, 22(1):133–134, February 1980. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1268403.

Jackson:1980:BRBa

[Jac80b] J. Edward Jackson. Book review: Robustness in Statistics by Robert L. Launer; Graham N. Wilkinson. Technometrics, 22(1):133, February 1980. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1268402.

Jackson:1980:BRBd

[Jac80c] J. Edward Jackson. Book review: The Algebra of Econometrics by D. G. S. Pollock. Technometrics, 22(4):634, November 1980. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1268211.

Jackson:1980:BRBc

[Jac80d] J. Edward Jackson. Book review: Three Aspects of Policy and Policymaking: Knowledge, Data and Institutions by Karl Brunner; Allan H. Meltzer. Technometrics, 22(1):134, February 1980. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1268404. See corrigenda

[Ano80e].

Jackson:1981:BRB

[Jac81] J. Edward Jackson. Book review: Geometric Representations of Relational Data by James C. Lingoes; Edward E. Roskam; Ingwer Borg. Technometrics, 23(2):207–209, May 1981. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268052.

Jackson:1984:BRB

[Jac84] J. Edward Jackson. Book review: Discrete Discriminant Analysis by Matthew Goldstein. Technometrics, 26(4):416, November 1984. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1269509.

Jackson:1985:BRBf

[Jac85a] J. Edward Jackson. Book review: Advanced Methods of Data Exploration and Modelling by B. S. Everitt; G. Dunn. Technometrics, 27(4):442, November 1985. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1270214.

Jackson:1985:BRBe

[Jac85b] J. Edward Jackson. Book review: Computer-Aided Multivariate Analysis by A. A. Afifi; Virginia Clark. Technometrics, 27 (3):318–319, August 1985. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1269717.

Jackson:1985:BRBc

[Jac85c] J. Edward Jackson. Book review: Factor Analysis: An Applied Approach by Edward E. Cureton; Ralph B. D’ Agostino. Technometrics, 27(2):208–209, May 1985. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268770.

Jackson:1985:BRBb

[Jac85d] J. Edward Jackson. Book review: Key Texts in Multidimensional Scaling by P. M. Davies; A. P. M. Coxon. Technometrics, 27(1):88, February 1985. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1270476.

Jackson:1985:BRBd

[Jac85e] J. Edward Jackson. Book review: Multivariate Data Analysis in Industrial Practice by Paul J. Lewi. Technometrics, 27(2):209, May 1985. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1268771.

Jackson:1985:BRBa

[Jac85f] J. Edward Jackson. Book review: The User’s Guide to Multidimensional Scaling by A. P. M. Coxon. Technometrics, 27(1):87–88, February 1985. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1270475.

Jackson:1986:BRBb

[Jac86a] J. Edward Jackson. Book review: The Analysis of Cross-Classified Data Having Ordered Categories by Leo A. Goodman. Technometrics, 28(3):279–280, August 1986. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1269088.

Jackson:1986:BRBa

[Jac86b] J. Edward Jackson. Book reviews: Multivariate Statistical Methods by Marvin J. Karson; Multivariate Analysis: Methods and Applications by William R. Dillon; Multivariate Observations by G. A.

F. Seber. Technometrics, 28(1):85–87, February 1986. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1269608.

Jackson:1989:BRB

[Jac89] J. Edward Jackson. Book review: Latent Variable Models and Factor Analysis by David Bartholomew. Technometrics, 31(2):266–267, May 1989. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1268836.

Jaech:1982:SMN

[Jae82] John L. Jaech. [Statistical Methods for Nuclear Materials Safeguards: An Overview]: Discussion. Technometrics, 24(4):284–287, November 1982. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1267820.

See [GPS82a].

Jain:1981:PPM

[Jai81] Ram B. Jain. Percentage points of many-outlier detection procedures. Technometrics, 23(1):71–75, February 1981. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267978.

Jeyaratnam:1980:CIV

[JG80] S. Jeyaratnam and Franklin A. Graybill. Confidence intervals on variance components in three-factor cross-classification models. Technometrics, 22(3):375–380, August 1980. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1268322.

Johnson:1987:AVP

[JM87] Bradford W. Johnson and Robert E. McCulloch. Added-variable plots in linear regression. Technometrics, 29(4):427–433, November 1987. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1269453.

Jani:1981:LE

[JN81] P. N. Jani and David L. Nelson. Letters to the Editor. Technometrics, 23(3):317, August 1981. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267809.

Johnson:1983:SGC

[JN83] Mark E. Johnson and Christopher J. Nachtsheim. Some guidelines for constructing exact D-optimal designs on convex design spaces. Technometrics, 25(3):271–277, August 1983. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1268612.

Jennrich:1986:HMD

[JO86] Robert I. Jennrich and Samuel D. Oman. How much does Stein estimation help in multiple linear regression? Technometrics, 28(2):113–121, May 1986. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1270447.

Jobe:1989:BRB

[Job89] J. Marcus Jobe. Book review: Lognormal Distributions: Theory and Applications by Edwin L. Crow; Kunio Shimizu. Technometrics, 31(3):392, August 1989. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/3556163.

Johnson:1982:BRB

[Joh82] Norman L. Johnson. Book review: Statistical Distributions in Scientific Work by Charles Taillie; G. P. Patil; Bruno Baldessori. Technometrics, 24(4):343, November 1982. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1267840.

Johnson:1983:BRB

[Joh83] Richard A. Johnson. Book review: Sequential Nonparametrics: Invariance Principles and Statistical Inference by Pranab Kumar Sen. Technometrics, 25(2):214–215, May 1983. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1268568.

Joiner:1984:EDR

[Joi84] Brian L. Joiner. [Experimental Design: Review and Comment]: Discussion. Technometrics, 26(2):123–124, May 1984. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268104. See [Hah84a].

Jolliffe:1988:BRB

[Jol88] Ian Jolliffe. Book review: Multivariate Statistical Simulation by Mark E. Johnson. Technometrics, 30(2):232–233, May 1988. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1270172.

Jones:1980:MLF

[Jon80] Richard H. Jones. Maximum likelihood fitting of ARMA models to time series with missing observations. Technometrics, 22 (3):389–395, August 1980. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268324.

Jones:1988:BRB

[Jon88] P. W. Jones. Book review: Stochastic Modelling and Analysis by Henk C. Tijms. Technometrics, 30(3):361, August 1988. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1270106.

Jones:1989:BRBa

[Jon89a] P. W. Jones. Book review: A Diary on Information Theory by Alfréd Rényi. Technometrics, 31(3):395, August 1989. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/3556169.

Jones:1989:BRBb

[Jon89b] Peter W. Jones. Book review: Exploring Probability by Claire M. Newman; Thomas E. Obremski; Richard L. Scheaffer. Technometrics, 31(4):500, November 1989. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1270021.

Johnson:1980:LE

[JRT^{+}80] Norman L. Johnson, John S. Ramberg, Pandu R. Tadikamalla, Edward J. Dudewicz, and Edward F. Mykytka. Letters to the Editor. Technometrics, 22(1):135, February 1980. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1268406.

Jones:1985:TBL

[JSOS85] Robert A. Jones, F. W. Scholz, Mina Ossiander, and Galen R. Shorack. Tolerance bounds for log gamma regression models.

Technometrics, 27(2):109–118, May 1985. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1268759.

Jennison:1983:CIB

[JT83] Christopher Jennison and Bruce W. Turnbull. Confidence intervals for a binomial parameter following a multistage test with application to MIL-STD 105D and medical trials. Technometrics, 25(1):49–58, February 1983. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267726.

Kadane:1984:BRB

[Kad84] Joseph B. Kadane. Book review: Structural Analysis of Discrete Data with Econometric Applications by Charles F. Manski; Daniel McFadden. Technometrics, 26(4):415, November 1984. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1269507.

Kallianpur:1983:BRB

[Kal83] G. B. Kallianpur. Book review: Stochastic Differential Equations and Diffusion Processes by Nobuyuki Ikeda; Shinzo Watanabe. Technometrics, 25(2):208, May 1983. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1268562.

Kanarek:1987:BRBb

[Kan87a] Paula Kanarek. Book review: Volume 10: How to Plan an Accelerated Life Test: Some Practical Guidelines by William Q. Meeker; Gerald J. Hahn. Technometrics, 29(4):491, November 1987. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1269467.

Kanarek:1987:BRBa

[Kan87b] Paula Kanarek. Book review: Volume 6: How to Analyze Reliability Data by Wayne Nelson. Technometrics, 29(4):489, November 1987. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1269464.

Kao:1988:CPT

[Kao88] Edward P. C. Kao. Computing the phase-type renewal and related functions. Technometrics, 30(1):87–93, February 1988. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1270324.

Katz:1981:SCE

[Kat81] Richard W. Katz. On some criteria for estimating the order of a Markov chain. Technometrics, 23(3):243–249, August 1981. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1267787.

Katti:1982:BRB

[Kat82] S. K. Katti. Book review: Analysis with Standard Contagious Distributions by J. B. Douglas. Technometrics, 24(1):82, February 1982. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1267585.

Khuri:1981:SOM

[KC81] A. I. Khuri and M. Conlon. Simultaneous optimization of multiple responses represented by polynomial regression functions. Technometrics, 23(4):363–375, November 1981. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1268226.

Kubat:1980:EQL

[KE80] Peter Kubat and Benjamin Epstein. Estimation of quantiles of location-scale distributions based on two or three order statistics. Technometrics, 22(4):575–581, November 1980. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1268195.

Kirby:1981:LE

[KECH81] W. Kirby, Robert G. Easterling, Wendell E. Carr, and Hans Dietrich Höhne. Letters to the Editor. Technometrics, 23(2):215–219, May 1981. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1268058. See

corrigenda [Ano81e].

Kedem:1987:HOC

[Ked87] Benjamin Kedem. Higher-order crossings in time series model identification. Technometrics, 29(2):193–204, May 1987. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1269774.

Kershner:1985:BRB

[Ker85] Ron Kershner. Book review: Pharmaceutical Statistics: Practical and Clinical Applications by Sanford Bolton. Technometrics, 27(4):443–444, November 1985. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1270216.

Ketellapper:1983:EPS

[Ket83] Ronald H. Ketellapper. On estimating parameters in a simple linear errors-in-variables model. Technometrics, 25(1):43–47, February 1983. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1267725.

Khuri:1984:NOD

[Khu84] A. I. Khuri. A note on D-optimal designs for partially nonlinear regression models. Technometrics, 26(1):59–61, February 1984. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1268416.

Khuri:1985:TLF

[Khu85] A. I. Khuri. A test for lack of fit of a linear multiresponse model. Technometrics, 27(3):213–218, August 1985. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1269702.

Khuri:1986:ETC

[Khu86] A. I. Khuri. Exact tests for the comparison of correlated response models with an unknown dispersion matrix. Technometrics, 28 (4):347–357, November 1986. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268984.

Khuri:1987:BRB

[Khu87] André I. Khuri. Book review: Selected Tables in Mathematical Statistics (Vol. 10) by R. E. Odeh; J. M. Davenport. Technometrics, 29(3):386, August 1987. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1269359.

Khuri:1988:MRR

[Khu88] A. I. Khuri. A measure of rotatability for response-surface designs. Technometrics, 30(1):95–104, February 1988. CODEN TCMTA2.

ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1270325.

Kim:1982:BRB

[Kim82] Jee Soo Kim. Book review: Censoring and Stochastic Integrals by R. D. Gill. Technometrics, 24(3):250–251, August 1982. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268688.

King:1987:BRB

[Kin87] Terry King. Book review: Goodness-of-Fit Techniques by R. D’Agostino; M. Stephens. Technometrics, 29(4):493, November 1987. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1269472.

Kitagawa:1979:UAD

[Kit79] Genshiro Kitagawa. On the use of AIC for the detection of outliers. Technometrics, 21(2):193–199, May 1979. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1268516. See corrigenda [Kit81].

Kitagawa:1981:CUA

[Kit81] Genshiro Kitagawa. Corrigenda: “On the use of AIC for the detection of outliers” [Technometrics 21 (1979), no. 2, 193–199; MR 80f:62026]. Technometrics, 23(3):320–321, August 1981. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267810. See [Kit79].

Kalbfleisch:1988:ERFa

[KL88a] J. D. Kalbfleisch and J. F. Lawless. Estimation of reliability in field-performance studies. Technometrics, 30(4):365–378, November 1988. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1269797. See

discussion [Nai88b, Rob88] and response [KL88b].

Kalbfleisch:1988:ERFb

[KL88b] J. D. Kalbfleisch and J. F. Lawless. [Estimation of Reliability in Field-Performance Studies]: Response. Technometrics, 30 (4):387–388, November 1988. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1269800. See [KL88a].

Klein:1982:BRB

[Kle82] John P. Klein. Book review: The Statistical Analysis of Failure Time Data by J. D. Kalbfleisch; R. L. Prentice. Technometrics, 24(3):251, August 1982. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268689.

Klotz:1982:BRB

[Klo82] Jerome Klotz. Book review: Biostatistics Casebook by Rupert G. Miller,; Bradley Efron; Byron Wm. Brown, Jr.; Lincoln E. Moses. Technometrics, 24(1):83–84, February 1982. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1267588.

Kim:1984:EBT

[KN84] Seong In Kim and Joel A. Nachlas. Estimation in Bernoulli trials under a generalized sampling plan. Technometrics, 26(4):379–387, November 1984. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1269501.

Koch:1981:BRB

[Koc81] George V. Koch. Book review: Statistical Quality Control for Food and Agricultural Scientists by Subash C. Puri; Daniel Ennis; Kenneth Mullen. Technometrics, 23(2):210, May 1981. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268054.

Koehler:1983:SAP

[Koe83] Kenneth J. Koehler. A simple approximation for the percentiles of the t distribution. Technometrics, 25(1):103–105, February 1983. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1267732.

Kotnour:1988:BRB

[Kot88] Kenneth D. Kotnour. Book review: Statistics: The Exploration and Analysis of Data by Jay Devore; Roxy Peck. Technometrics, 30(4):460–461, November 1988. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1269820.

Koziol:1980:PPA

[Koz80] James A. Koziol. Percentage points of the asymptotic distributions of one and two sample Kuiper statistics for truncated or censored data. Technometrics, 22(3):437–442, August 1980. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268330.

Kremers:1987:IEM

[Kre87] Walter K. Kremers. An improved estimator of the mean for a sequential binomial sampling plan. Technometrics, 29(1):109–112, February 1987. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1269889.

Knafl:1984:NC

[KSSY84] George Knafl, Clifford Spiegelman, Jerome Sacks, and Donald Ylvisaker. Nonparametric calibration. Technometrics, 26(3):233– 241, August 1984. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1267549.

Ledolter:1981:PIT

[LA81] Johannes Ledolter and Bovas Abraham. Parsimony and its importance in time series forecasting. Technometrics, 23(4):411–414, November 1981. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1268232.

Lachenbruch:1981:BRB

[Lac81] Peter A. Lachenbruch. Book review: Cluster Analysis by Brian Everitt. Technometrics, 23(3):308, August 1981. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267799.

Lachenbruch:1982:BRB

[Lac82] Peter A. Lachenbruch. Book review: Classification: Methods for the Exploratory Analysis of Multivariate Data by A. D. Gordon. Technometrics, 24(4):340–341, November 1982. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1267836.

Lake:1987:BRB

[Lak87] Robin B. Lake. Book review: Finite Algorithms in Optimization and Data Analysis by M. R. Osborne. Technometrics, 29(4):501, November 1987. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1269483.

Lakshminarayanan:1988:BRB

[Lak88] Mani Y. Lakshminarayanan. Book review: The General Linear Model by Raymond Horton. Technometrics, 30(1):130, February 1988. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1270349.

Lambrinos:1984:BRBc

[Lam84a] James Lambrinos. Book review: Applied Linear Regression Models by John Neter; William Wasserman; Michael H. Kutner. Technometrics, 26(4):415–416, November 1984. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1269508.

Lambrinos:1984:BRBb

[Lam84b] James Lambrinos. Book review: Decade of Robotics, Special Tenth Anniversary Issue of the Industrial Robot Magazine by Ed Mortimer. Technometrics, 26(2):194, May 1984. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1268122.

Lambrinos:1984:BRBa

[Lam84c] James Lambrinos. Book review: Statistics in the Labor Market: The Role in Planning and Policy Formation by Dennis Trewin. Technometrics, 26(2):189–190, May 1984. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268114.

Landwehr:1987:BRB

[Lan87] James M. Landwehr. Book review: Clustering of Large Data Sets by Jure Zupan. Technometrics, 29(4):497–498, November 1987. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1269478.

Langenberg:1988:BRB

[Lan88] Patricia Langenberg. Book review: Statistics: The Essentials for Research by Henry Klugh. Technometrics, 30(4):459–460, November 1988. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1269819.

Larsen:1989:BRB

[Lar89] Greg Larsen. Book review: Productivity and Quality Improvement by John L. Hradesky. Technometrics, 31(4):487–488, November 1989. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1270009.

Lawless:1980:IGG

[Law80] J. F. Lawless. Inference in the generalized gamma and log gamma distributions. Technometrics, 22(3):409–419, August 1980. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268326.

Lawless:1981:BAL

[Law81] J. F. Lawless. [A Bayesian Analysis of the Linear Calibration Problem]: Discussion. Technometrics, 23(4):334–335, November 1981. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1268219. See [HL81b].

Lawless:1983:SMRa

[Law83a] J. F. Lawless. Statistical methods in reliability. Technometrics, 25(4):305–316, November 1983. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/ stable/1267846. See discussion [Mee83, Asc83, Tho83, Sch83c, Bir83, Eng83, Eva83] and response [Law83b].

Lawless:1983:SMRb

[Law83b] J. F. Lawless. [Statistical Methods in Reliability]: Response. Technometrics, 25(4):334–335, November 1983. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1267854. See [Law83a].

Lawless:1985:ER

[Law85] J. F. Lawless. Editor’s report. Technometrics, 27(2):93–94, May 1985. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1268757.

Lawless:1986:ER

[Law86a] J. F. Lawless. Editor’s report. Technometrics, 28(4):281, November 1986. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1268973.

Lawless:1986:NE

[Law86b] J. F. Lawless. New Editor. Technometrics, 28(1):1, February 1986. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1269597.

Lawson:1986:BRB

[Law86c] John Lawson. Book review: Quality by Experimental Design by Thomas Barker. Technometrics, 28(4):400–402, November 1986. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1268991.

Lam:1981:DAP

[LC81] Chan F. Lam and Michael Cox. A discriminant analysis procedure for paired variables. Technometrics, 23(2):185–187, May 1981. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1268036.

Lucas:1982:FIR

[LC82] James M. Lucas and Ronald B. Crosier. Fast initial response for CUSUM quality-control schemes: Give your CUSUM a head start. Technometrics, 24(3):199–205, August 1982. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1268679.

Lee:1980:ASP

[LD80] Young Jack Lee and Edward J. Dudewicz. Awardee selection procedures, with special reference to the Frank Wilcoxon and Jack Youden Prizes. Technometrics, 22(1):121–124, February 1980. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1268391.

Littell:1982:SMN

[LD82] Ramon C. Littell and Darryl J. Downing. [Statistical Methods for Nuclear Materials Safeguards: An Overview]: Discussion. Technometrics, 24(4):277–279, November 1982. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1267816. See [GPS82a].

Lee:1980:CRS

[Lee80a] Larry Lee. Comparing rates of several independent Weibull processes. Technometrics, 22(3):427–430, August 1980. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268328.

Lee:1980:TAW

[Lee80b] Larry Lee. Testing adequacy of the Weibull and log linear rate models for a Poisson process. Technometrics, 22(2):195–199, May 1980.

CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1268458.

Lee:1980:NSB

[Lee80c] Young Jack Lee. Nonparametric selections in blocked data: Application to motor vehicle fatality rate data. Technometrics, 22 (4):535–542, November 1980. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268190.

Lee:1981:BRB

[Lee81] Jack C. Lee. Book review: Multivariate Analysis by K. V. Mardia; J. T. Kent; J. M. Bibby. Technometrics, 23(3):309, August 1981. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1267802.

Lee:1982:BRB

[Lee82] Jack C. Lee. Book review: Multivariate Analysis by M. Kendall. Technometrics, 24(3):251–252, August 1982. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1268690.

Lee:1983:BRB

[Lee83] Jack C. Lee. Book review: Introduction to Multivariate Analysis by C. Chatfield; A. J. Collins. Technometrics, 25(4):384, November 1983. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1267862.

Lee:1989:BRB

[Lee89a] Cliff J. Lee. Book review: SPC Simplified: Practical Steps to Quality by R. Amsden; H. Butler; D. Amsden. Technometrics, 31(1):120–121, February 1989. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1270379.

Leemis:1989:BRBa

[Lee89b] Lawrence Leemis. Book review: Reliability-Based Design in Civil Engineering by Milton E. Harr. Technometrics, 31(1):126, February 1989. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1270385.

Leemis:1989:BRBb

[Lee89c] Lawrence Leemis. Book review: Reliability Modelling and Applications by A. G. Colombo; A. Z. Keller. Technometrics, 31(2):268, May 1989. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1268838.

Lenth:1981:FSS

[Len81a] Russell V. Lenth. On finding the source of a signal. Technometrics, 23(2):149–154, May 1981. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268030.

Lenth:1981:RML

[Len81b] Russell V. Lenth. Robust measures of location for directional data. Technometrics, 23(1):77–81, February 1981. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1267979.

Lenth:1989:QEA

[Len89] Russell V. Lenth. Quick and easy analysis of unreplicated factorials. Technometrics, 31(4):469–473, November 1989. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1269997.

Lewis:1987:BRB

[Lew87] Donald Lewis. Book review: Applications of QC in the Service Industries by A. Rosander. Technometrics, 29(1):113–114, February 1987. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1269890.

Likes:1980:VML

[Lik80] Jiří Likeš. Variance of the MVUE for lognormal variance. Technometrics, 22(2):253–258, May 1980. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268465.

Likes:1982:NML

[Lik82] Jiří Likeš. A note on minimum life hypothesis testing for the exponential distribution. Technometrics, 24(2):145–146, May 1982. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1268493.

Ling:1981:BRBb

[Lin81a] Robert F. Ling. Book review: Cluster Analysis Algorithms for Data Reduction and Classification of Objects by Helmuth Späth. Technometrics, 23(4):417–418, November 1981. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1268238.

Ling:1981:BRBa

[Lin81b] Robert F. Ling. Book review: Selected Tables in Mathematical Statistics, Volume VI, The Distribution of the Size of the Maximum Cluster of Points on a Line by W. J. Kennedy; R. E. Odeh. Technometrics, 23(4):417, November 1981. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268237.

Ling:1984:BRB

[Lin84] Robert F. Ling. Book review: Residuals and Influence in Regression by R. Dennis Cook; Sanford Weisberg. Technometrics, 26(4):413–415, November 1984. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1269506.

Little:1985:IQF

[Lit85] James K. Little. Influence and a quadratic form in the Andrews– Pregibon statistic. Technometrics, 27(1):13–15, February 1985. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1270464.

Littell:1988:BRB

[Lit88] Ramon C. Littell. Book review: Model Selection by H. Linhart; W. Zucchini. Technometrics, 30(1):115–116, February 1988. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1270329.

Littell:1989:BRB

[Lit89] Ramon C. Littell. Book review: Multiple Comparison Procedures by Y. Hochberg; A. Tamhane. Technometrics, 31(2):261–262, May 1989. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1268830.

Liu:1989:BRB

[Liu89] Regina Y. Liu. Book review: Frontiers in Statistical Quality Control 3 by H.-J. Lenz; G. B. Wetherill; P.-Th. Wilrich. Technometrics, 31(4):488–491, November 1989. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1270010.

Lee:1980:CTE

[LLS80] S. C. S. Lee, Charles Locke, and John D. Spurrier. On a class of tests of exponentiality. Technometrics, 22(4):547–554, November 1980. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1268192.

Lin:1980:TEB

[LM80] Ching-Chuong Lin and Govind S. Mudholkar. A test of exponentiality based on the bivariate F distribution. Technometrics, 22(1):79–82, February 1980. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268386.

Lwin:1982:ALC

[LM82] T. Lwin and J. S. Maritz. An analysis of the linear-calibration controversy from the perspective of compound estimation. Technometrics, 24(3):235–242, August 1982. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/ stable/1268684. See correction [LM85].

Lwin:1985:CAL

[LM85] T. Lwin and J. S. Maritz. Correction: “An analysis of the linearcalibration controversy from the perspective of compound estimation” [Technometrics 24 (1982), no. 3, 235–242; MR0667653 (84e:62017)]. Technometrics, 27(4):445, November 1985. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1270217. See [LM82].

Leone:1961:FND

[LNN61] F. C. Leone, L. S. Nelson, and R. B. Nottingham. The folded normal distribution. Technometrics, 3(4):543–550, November 1961. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1266560. See corrigenda

[Ano80f].

Lochner:1987:BRB

[Loc87] Robert H. Lochner. Book review: Quality Control for Profit by Ronald H. Lester; Norbert L. Enrick; Harry E. Mottley. Technometrics, 29(4):492, November 1987. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1269469.

Lombard:1986:CPP

[Lom86] F. Lombard. The change-point problem for angular data: A nonparametric approach. Technometrics, 28(4):391–397, November 1986. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1268988.

Lombard:1988:DCP

[Lom88] F. Lombard. Detecting change points by Fourier analysis. Technometrics, 30(3):305–310, August 1988. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1270084.

Lorenzen:1980:DSC

[Lor80] Thomas J. Lorenzen. Determining statistical characteristics of a vehicle emissions audit procedure. Technometrics, 22(4):483–493, November 1980. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1268186.

Louv:1984:AF

[Lou84] William C. Louv. Adaptive filtering. Technometrics, 26(4):399–409, November 1984. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1269503.

Low:1982:BRB

[Low82] Leone Y. Low. Book reviews: Understanding Basic Statistics by Harvey W. Kushner; Gerald DeMaio; Finite Mathematics, A Self-Teaching Guide by Ronald I. Rothenberg. Technometrics, 24(2): 164, May 1982. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1268500.

Low:1983:BRBa

[Low83a] Leone Y. Low. Book review: Introductory Statistics by Neil Weiss; Matthew Hassett. Technometrics, 25(2):215–216, May 1983. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1268570.

Low:1983:BRBb

[Low83b] Leone Y. Low. Book review: The Analysis of Covariance and Alternatives by Bradley E. Huitema. Technometrics, 25(4):384, November 1983. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1267863.

Lechner:1982:ISA

[LRS82] James A. Lechner, Charles P. Reeve, and Clifford H. Spiegelman. An implementation of the Scheffé approach to calibration using spline functions, illustrated by a pressure-volume calibration. Technometrics, 24(3):229–234, August 1982. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1268683.

Leon:1987:PMIa

[LSK87a] Ramón V. León, Anne C. Shoemaker, and Raghu N. Kacker. Performance measures independent of adjustment. An explanation and extension of Taguchi’s signal-to-noise ratios. Technometrics, 29(3):253–265, August 1987. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/ stable/1269331. See discussion [Boa87, Eas87, BF87, Deh87b, PR87, Wu87] and response [LSK87b].

Leon:1987:PMIb

[LSK87b] Ramón V. León, Anne C. Shoemaker, and Raghu N. Kacker. [Performance Measures Independent of Adjustment]: Response. Technometrics, 29(3):283–285, August 1987. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1269338.

Lucas:1981:BRB

[Luc81] James M. Lucas. Book review: Fitting Equations to Data: Computer Analysis of Multifactor Data by Cuthbert Daniel; Fred S. Wood. Technometrics, 23(2):203–204, May 1981. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268046.

Lucas:1985:CDC

[Luc85] James M. Lucas. Counted data CUSUM’s. Technometrics, 27 (2):129–144, May 1985. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268761.

Lorenzen:1986:EDC

[LV86] Thomas J. Lorenzen and Lonnie C. Vance. The economic design of control charts: A unified approach. Technometrics, 28(1):3–10, February 1986. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1269598.

Lwin:1981:BAL

[Lwi81] T. Lwin. [A Bayesian Analysis of the Linear Calibration Problem]: Discussion. Technometrics, 23(4):339–341, November 1981. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1268221. See [HL81b].

Mack:1981:BRB

[Mac81] Gregory A. Mack. Book review: Regression Analysis and Its Application: A Data-Oriented Approach by R. F. Gunst; R. L. Mason. Technometrics, 23(3):309, August 1981. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1267800.

Mack:1983:BRB

[Mac83] Gregory Mack. Book review: Order Statistics by H. A. David. Technometrics, 25(1):110, February 1983. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1267736.

Madsen:1982:SPU

[Mad82] Richard W. Madsen. A selection procedure using a screening variate. Technometrics, 24(4):301–306, November 1982. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267824.

Mage:1980:ESP

[Mag80] David T. Mage. An explicit solution for S_{B} parameters using four percentile points. Technometrics, 22(2):247–251, May 1980. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1268464.

Mallows:1986:APR

[Mal86] C. L. Mallows. Augmented partial residuals. Technometrics, 28 (4):313–319, November 1986. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/ stable/1268980. See discussion [BM88].

Marshall:1983:BRB

[Mar83] Clifford W. Marshall. Book review: The Single Server Queue by J. W. Cohen. Technometrics, 25(1):109–110, February 1983. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267734.

Marasinghe:1985:MPD

[Mar85] Mervyn G. Marasinghe. A multistage procedure for detecting several outliers in linear regression. Technometrics, 27(4):395–399, November 1985. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1270206.

See discussion [FM87].

Marasinghe:1988:BRB

[Mar88a] Mervyn G. Marasinghe. Book review: SAS System for Linear Models by R. Freund; R. Littell; P. Spector. Technometrics, 30 (1):116–117, February 1988. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1270330.

Marcucci:1988:BRB

[Mar88b] Mark Marcucci. Book review: Statistical Analysis by Sam Kachigan. Technometrics, 30(2):235, May 1988. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1270177.

Mason:1988:BRB

[Mas88] Robert L. Mason. Book review: SAS System for Regression by R. J. Freund; R. C. Littell. Technometrics, 30(1):117–118, February 1988. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1270331.

Mathai:1980:BRB

[Mat80] A. M. Mathai. Book review: Information and Exponential Families in Statistical Theory by Ole Barndorff-Nielsen. Technometrics, 22(2):280, May 1980. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268473.

Mauro:1984:PTS

[Mau84] Carl A. Mauro. On the performance of two-stage group screening experiments. Technometrics, 26(3):255–264, August 1984. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267552.

Mazumdar:1980:OCT

[Maz80] M. Mazumdar. An optimum component testing procedure for a series system with redundant subsystems. Technometrics, 22 (1):23–27, February 1980. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268379.

Mazu:1987:BRB

[Maz87] Michael J. Mazu. Book review: Statistical Experiments Using BA-SIC by Shirley Dowdy. Technometrics, 29(2):246–247, May 1987. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1269788.

Mazu:1988:BRB

[Maz88] Michael J. Mazu. Book review: Quality Programming: Developing and Testing Software with Statistical Quality Control by Chin-Kuei Cho. Technometrics, 30(4):467–468, November 1988. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1269829.

Mazu:1989:BRB

[Maz89] Michael J. Mazu. Book review: Sensitivity Analysis in Linear Regression by Samprit Chatterjee; Ali Hadi. Technometrics, 31 (4):481–482, November 1989. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1270001.

Moore:1988:AOS

[MB88] Leslie M. Moore and Richard J. Beckman. Approximate one-sided tolerance bounds on the number of failures using Poisson regression. Technometrics, 30(3):283–290, August 1988. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1270082.

McCabe:1984:PV

[McC84] George P. McCabe. Principal variables. Technometrics, 26(2):137– 144, May 1984. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1268108.

McCullagh:1985:BRB

[McC85] Peter McCullagh. Book review: Analysis of Ordinal Categorical Data by Alan Agresti. Technometrics, 27(3):317–318, August 1985.

CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1269716.

McCullagh:1986:TIE

[McC86] Peter McCullagh. [Testing in Industrial Experiments with Ordered Categorical Data]: Discussion. Technometrics, 28(4):307, November 1986. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1268978. See

[BW85a].

McCool:1987:BRB

[McC87] John McCool. Book review: The Life Table and Its Applications by Chin Long Chiang. Technometrics, 29(2):243–245, May 1987. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1269785.

McDonald:1979:NSP

[McD79] Gary C. McDonald. Nonparametric selection procedures applied to state traffic fatality rates. Technometrics, 21(4):515–523, November 1979. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1268291. See

corrigendum [Ano80g].

McDonald:1981:CIV

[McD81] Gary C. McDonald. Confidence intervals for vehicle emission deterioration factors. Technometrics, 23(3):239–242, August 1981. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1267786.

McLachlan:1980:NBC

[McL80] G. J. McLachlan. A note on bias correction in maximum likelihood estimation with logistic discrimination. Technometrics, 22 (4):621–627, November 1980. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268202.

McLean:1985:MES

[McL85] David D. McLean. [Multiresponse Estimation with Special Application to Linear Systems of Differential Equations]: Discussion. Technometrics, 27(4):340–348, November 1985. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1270199. See [BW85a].

McLaughlin:1988:BRB

[McL88] Gregory C. McLaughlin. Book review: Statistics by J. T. McClave; F. H. Dietrich. Technometrics, 30(1):125–126, February 1988. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1270342.

McLean:1989:BRB

[McL89] David D. McLean. Book review: Nonlinear Parameter Estimation: An Integrated System in BASIC by John C. Nash; Mary Walker-Smith. Technometrics, 31(1):114–115, February 1989. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1270373.

Makowski:1980:CIC

[MD80] Gary G. Makowski and Darryl Downing. Confidence interval calibration of a standard curve. Technometrics, 22(3):381–388, August 1980. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1268323.

McNolty:1980:PME

[MDH80] Frank McNolty, James Doyle, and Eldon Hansen. Properties of the mixed exponential failure process. Technometrics, 22(4):555–565, November 1980. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1268193.

Mead:1989:SBP

[Mea89] G. Alvin Mead. The sorted binary plot: A new technique for exploratory data analysis. Technometrics, 31(1):61–67, February 1989. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1270365.

Meeker:1982:BRB

[Mee82] William Q. Meeker, Jr. Book review: Statistical Software: A Comparative Review by Ivor Francis. Technometrics, 24(4):339–340, November 1982. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1267835.

Meeker:1983:SMR

[Mee83] William Q. Meeker, Jr. [Statistical Methods in Reliability]: Discussion. Technometrics, 25(4):316–320, November 1983. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267847. See [Law83a].

Mee:1984:ECT

[Mee84a] Robert W. Mee. β-expectation and β-content tolerance limits for balanced one-way ANOVA random model. Technometrics, 26(3):251–254, August 1984. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267551.

Meeker:1984:CAL

[Mee84b] William Q. Meeker, Jr. A comparison of accelerated life test plans for Weibull and lognormal distributions and Type I censoring. Technometrics, 26(2):157–171, May 1984. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1268110.

Meeker:1985:BRB

[Mee85] William Q. Meeker, Jr. Book review: Optimization in Statistics with a View toward Applications in Management Science and Operations Research by S. H. Zanakis; J. S. Rustagi. Technometrics, 27(2):207– 208, May 1985. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1268769.

Meeker:1986:BRB

[Mee86] William Q. Meeker, Jr. Book review: Applied Statistics Algorithms by P. Griffiths; I. D. Hill. Technometrics, 28(3):278–279, August 1986. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1269087.

Meek:1987:BRB

[Mee87a] Gary E. Meek. Book review: Mathematical Statistics with Applications by William Mendenhall; Richard L. Scheaffer; Dennis D. Wackerly. Technometrics, 29(3):384–385, August 1987. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1269356.

Meeker:1987:LFP

[Mee87b] William Q. Meeker, Jr. Limited failure population life tests: Application to integrated circuit reliability. Technometrics, 29(1):51–65, February 1987. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1269883.

Mee:1989:NDT

[Mee89a] Robert W. Mee. Normal distribution tolerance limits for stratified random samples. Technometrics, 31(1):99–105, February 1989.

CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1270369.

Meeker:1989:NE

[Mee89b] William Q. Meeker, Jr. New Editor. Technometrics, 31(1):1, February 1989. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1270358.

Meyer:1988:BRB

[Mey88] R. Daniel Meyer. Book review: Design and Analysis of Experiments by M. N. Das; N. C. Giri. Technometrics, 30(2):241–242, May 1988. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1270186.

Madsen:1983:PVM

[MF83] Richard W. Madsen and Kenneth B. Fairbanks. P values for multistage and sequential tests. Technometrics, 25(3):285–293, August 1983. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1268614.

Mazumdar:1984:CPG

[MG84] M. Mazumdar and D. P. Gaver. On the computation of powergenerating system reliability indexes. Technometrics, 26(2):173– 185, May 1984. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1268111.

Mason:1985:OIC

[MG85] Robert L. Mason and Richard F. Gunst. Outlier-induced collinearities. Technometrics, 27(4):401–407, November 1985. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1270207.

Meeker:1980:PIR

[MH80] William Q. Meeker, Jr. and Gerald J. Hahn. Prediction intervals for the ratios of normal distribution sample variances and exponential distribution sample means. Technometrics, 22(3):357–366, August 1980. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1268320.

MacGregor:1984:DBC

[MHW84] J. F. MacGregor, T. J. Harris, and J. D. Wright. Duality between the control of processes subject to randomly occurring deterministic disturbances and ARIMA stochastic disturbances. Technometrics, 26(4):389–397, November 1984. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1269502.

Mielke:1982:BRB

[Mie82] Paul W. Mielke, Jr. Book review: Statistical Analysis of Weather Modification Experiments by Edward J. Wegman; Douglas J. De-Priest. Technometrics, 24(1):84–85, February 1982. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267590.

Miller:1980:BAT

[Mil80] Robert B. Miller. Bayesian analysis of the two-parameter gamma distribution. Technometrics, 22(1):65–69, February 1980. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268384.

Miller:1981:BRB

[Mil81] Robert B. Miller. Book review: The Analysis of Time Series: An Introduction by C. Chatfield. Technometrics, 23(4):415, November 1981. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1268233.

Miller:1989:PEB

[Mil89] Robert W. Miller. Parametric empirical Bayes tolerance intervals. Technometrics, 31(4):449–459, November 1989. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1269995.

Minton:1986:BRB

[Min86] Paul Minton. Book review: Linear Statistical Models and Related Methods, with Applications to Social Research by John Fox. Technometrics, 28(2):180, May 1986. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1270457.

Minton:1987:BRB

[Min87] Paul D. Minton. Book review: Statistical Techniques for Analytical Review in Auditing by Kenneth W. Stringer; Trevor R. Stewart. Technometrics, 29(2):247, May 1987. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1269789.

Minton:1988:BRB

[Min88] Paul D. Minton. Book review: Introductory Statistics by Neil Weiss; Matthew Hassett. Technometrics, 30(3):354–356, August 1988. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1270096.

Minton:1989:BRB

[Min89] Paul D. Minton. Book review: Exploring Surveys and Information from Samples by James Landwehr; Jim Swift; Ann Watkins. Technometrics, 31(4):500–501, November 1989. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1270022.

Mitchell:1983:DLR

[Mit83] Toby J. Mitchell. [Developments in Linear Regression Methodology: 1959–1982]: Discussion. Technometrics, 25(3):237–239, August 1983. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1268605. See

[Hoc83a].

Miyamura:1982:ECF

[Miy82] Tetsuo Miyamura. Estimating component failure rates from combined component and systems data: Exponentially distributed component lifetimes. Technometrics, 24(4):313–318, November 1982. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1267826.

Marasinghe:1981:TSM

[MJ81] Mervyn G. Marasinghe and Dallas E. Johnson. Testing subhypotheses in the multiplicative interaction model. Technometrics, 23(4):385–393, November 1981. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268229.

Moeschberger:1984:CDI

[MK84] M. L. Moeschberger and John P. Klein. Consequences of departures from independence in exponential series systems. Technometrics, 26(3):277–284, August 1984. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267554.

Myers:1989:RSM

[MKC89] Raymond H. Myers, André I. Khuri, and Walter H. Carter, Jr. Response surface methodology: 1966–1988. Technometrics, 31(2):137– 157, May 1989. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1268813.

McLachlan:1982:LRT

[MLG82] G. J. McLachlan, C. R. O. Lawoko, and S. Ganesalingam. On the likelihood ratio test for compound distributions for homogeneity of mixing proportions. Technometrics, 24(4):331–334, November 1982. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1267829.

Mantel:1982:VED

[MM82a] Nathan Mantel and Gary C. McDonald. Vehicle emission deterioration factors. Technometrics, 24(4):345–346, November 1982. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1267842.

Meeter:1982:BRB

[MM82b] Duane A. Meeter and Charles E. McCulloch. Book review: Regression: A Second Course in Statistics by T. H. Wonnacott; R. J. Wonnacott. Technometrics, 24(1):85–86, February 1982. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267591.

McCulloch:1983:OD

[MM83a] Charles E. McCulloch and Duane Meeter. [Outlier. . . . . . . . . s]: Discussion. Technometrics, 25(2):152–155, May 1983. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268543. See [BC83a].

McCulloch:1983:BRB

[MM83b] Charles E. McCulloch and Duane A. Meeter. Book review: Analyzing Experimental Data by Regression by David M. Allen; Foster B. Cady. Technometrics, 25(2):205–206, May 1983. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268556.

Morris:1983:TLM

[MM83c] Max D. Morris and Toby J. Mitchell. Two-level multifactor designs for detecting the presence of interactions. Technometrics, 25(4):345–355, November 1983. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267856.

McDonald:1986:SSL

[MO86] John Alan McDonald and Art B. Owen. Smoothing with split linear fits. Technometrics, 28(3):195–208, August 1986. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1269075.

Moeschberger:1983:BRB

[Moe83] Melvin L. Moeschberger. Book review: Survival Models and Data Analysis by Regina C. Elandt-Johnson; Norman L. Johnson. Technometrics, 25(1):110–111, February 1983. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1267737.

Moglewer:1982:SMN

[Mog82] Sidney Moglewer. [Statistical Methods for Nuclear Materials Safeguards: An Overview]: Discussion. Technometrics, 24(4):282, November 1982. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1267818.

See [GPS82a].

Moore:1982:SMN

[Moo82] Roger H. Moore. [Statistical Methods for Nuclear Materials Safeguards: An Overview]: Discussion. Technometrics, 24(4):288–290, November 1982. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1267821.

See [GPS82a].

Morris:1982:SMN

[Mor82] Carl N. Morris. [Statistical Methods for Nuclear Materials Safeguards: An Overview]: Discussion. Technometrics, 24(4):283, November 1982. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1267819.

See [GPS82a].

Morgan:1984:BRB

[Mor84] Carolyn B. Morgan. Book review: Discrete-Event System Simulation by Jerry Banks; John S. Carson. Technometrics, 26(2):195, May 1984. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1268124.

Morris:1987:SED

[Mor87] Max D. Morris. A sequential experimental design for estimating a scale parameter from quantal life testing data. Technometrics, 29(2):173–181, May 1987. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1269772.

Morgan:1988:BRB

[Mor88a] Carolyn Morgan. Book review: Truncated and Censored Samples from Normal Populations by H. Schneider. Technometrics, 30(1):118–119, February 1988. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1270333.

Morris:1988:BRB

[Mor88b] Max D. Morris. Book review: Cyclic Designs by J. A. John. Technometrics, 30(4):455–456, November 1988. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1269815.

Moberg:1980:AMR

[MRR80] Thomas F. Moberg, John S. Ramberg, and Ronald H. Randles. An adaptive multiple regression procedure based on M-estimators. Technometrics, 22(2):213–224, May 1980. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1268461.

Mauro:1982:PTS

[MS82] Carl A. Mauro and Dennis E. Smith. The performance of two-stage group screening in factor screening experiments. Technometrics, 24(4):325–330, November 1982. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267828.

McKean:1989:RSS

[MS89] Joseph W. McKean and Gerald L. Sievers. Rank scores suitable for analyses of linear models under asymmetric error distributions. Technometrics, 31(2):207–218, May 1989. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268818.

Mudholkar:1982:ADL

[MTL82] Govind S. Mudholkar, Madhusudan C. Trivedi, and C. Thomas Lin. An approximation to the distribution of the likelihood ratio statistic for testing complete independence. Technometrics, 24(2):139–143, May 1982. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1268492.

Muller:1983:BRB

[Mul83] Mervin E. Muller. Book review: The Handbook of Random Number Generation and Testing with TESTRAND Computer Code by Edward J. Dudewicz; Thomas G. Ralley. Technometrics, 25(2):207– 208, May 1983. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1268561.

Muller:1989:BRB

[Mul89] Keith Muller. Book review: Statistical Power Analysis for the Behavioral Sciences by Jacob Cohen. Technometrics, 31(4):499–500, November 1989. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1270020.

MacGregor:1980:MMIb

[MW80a] J. F. MacGregor and A. K. L. Wong. [Multivariate Model Identification and Stochastic Control of a Chemical Reactor]: Response. Technometrics, 22(4):468, November 1980. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1268182.

MacGregor:1980:MMIa

[MW80b] John F. MacGregor and A. K. L. Wong. Multivariate model identification and stochastic control of a chemical reactor. Technometrics, 22(4):453–464, November 1980. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/ stable/1268180. See discussion [Bac80].

Martz:1988:BRA

[MWF88] H. F. Martz, R. A. Waller, and E. T. Fickas. Bayesian reliability analysis of series systems of binomial subsystems and components. Technometrics, 30(2):143–154, May 1988. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1270159.

Myers:1980:BRB

[Mye80] Max H. Myers. Book review: Basic Statistics: A Primer for the Biomedical Sciences by Olive Jean Dunn. Technometrics, 22(3):445, August 1980. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1268332.

Newcombe:1988:TCP

[NA88] P. A. Newcombe and O. B. Allen. A three-class procedure for acceptance sampling by variables. Technometrics, 30(4):415–421, November 1988. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1269804.

Nachtsheim:1987:BRBa

[Nac87a] Christopher J. Nachtsheim. Book review: Applied Regression Analysis and Experimental Design by Richard J. Brook; Gregory C. Arnold. Technometrics, 29(2):238–240, May 1987. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1269780.

Nachtsheim:1987:BRBb

[Nac87b] Christopher J. Nachtsheim. Book review: Orthogonal Fractional Factorial Designs by Aloke Dey. Technometrics, 29(3):387–388, August 1987. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1269362.

Nachlas:1989:BRB

[Nac89] Joel Nachlas. Book review: Integrated Product Testing and Evalution by H. L. Gilmore; H. C. Schwartz. Technometrics, 31 (3):384–385, August 1989. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/3556153.

Naes:1985:MCW

[Næs85] Tormod Næs. Multivariate calibration when the error covariance matrix is structured. Technometrics, 27(3):301–311, August 1985. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1269712.

Nair:1984:CBS

[Nai84] Vijayan N. Nair. Confidence bands for survival functions with censored data: A comparative study. Technometrics, 26(3):265–275, August 1984. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1267553.

Nair:1986:TIEa

[Nai86a] Vijayan N. Nair. Testing in industrial experiments with ordered categorical data. Technometrics, 28(4):283–291, November 1986. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1268974. See dis-

cussion [Agr86b, BJ86, McC86] and response [Nai86b].

Nair:1986:TIEb

[Nai86b] Vijayan N. Nair. [Testing in Industrial Experiments with Ordered Categorical Data]: Response. Technometrics, 28(4):308–311, November 1986. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1268979.

See [BW85a].

Naiman:1988:BRB

[Nai88a] Daniel Q. Naiman. Book review: Pattern Recognition in Practice II by Edzard S. Gelsema; Laveen N. Kanal. Technometrics, 30(1):128, February 1988. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1270346.

Nair:1988:ERF

[Nai88b] Vijayan N. Nair. [Estimation of Reliability in Field-Performance Studies]: Discussion. Technometrics, 30(4):379–383, November 1988. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1269798. See

[KL88a].

Nelson:1981:BRB

[Nel81] Peter R. Nelson. Book review: The Algebra of Random Variables by M. D. Springer. Technometrics, 23(2):197–198, May 1981. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268039.

Nelson:1983:CLN

[Nel83] Peter R. Nelson. Counting a large number of items using the weight of a small sample. Technometrics, 25(2):173–177, May 1983. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1268550.

Nelson:1985:PCA

[Nel85] Peter R. Nelson. Power curves for the analysis of means. Technometrics, 27(1):65–73, February 1985. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1270471.

Nelson:1988:TIU

[Nel88] Peter R. Nelson. Testing for interactions using the analysis of means. Technometrics, 30(1):53–61, February 1988. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1270320.

Newby:1980:PME

[New80] M. J. Newby. The properties of moment estimators for the Weibull distribution based on the sample coefficient of variation. Technometrics, 22(2):187–194, May 1980. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/ stable/1268457. See corrigenda [Ano82f].

Newton:1982:UPA

[New82] H. Joseph Newton. Using periodic autoregressions for multiple spectral estimation. Technometrics, 24(2):109–116, May 1982. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268487.

Nigam:1985:COM

[NG85] A. K. Nigam and V. K. Gupta. Construction of orthogonal maineffect plans using Hadamard matrices. Technometrics, 27(1):37–40, February 1985. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1270467.

Nigam:1983:NAE

[NGG83] A. K. Nigam, S. C. Gupta, and S. Gupta. A new algorithm for extreme vertices designs for linear mixture models. Technometrics, 25(4):367–371, November 1983. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267858.

Nair:1988:SNR

[NP88a] V. N. Nair and D. Pregibon. [Signal-to-Noise Ratios, Performance Criteria, and Transformations]: Discussion. Technometrics, 30(1):24–29, February 1988. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1270314. See [Box88a].

Nair:1988:ADE

[NP88b] Vijayan N. Nair and Daryl Pregibon. Analyzing dispersion effects from replicated factorial experiments. Technometrics, 30 (3):247–257, August 1988. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/ stable/1270079. See correction [NP89].

Nair:1989:CAD

[NP89] Vijayan N. Nair and Daryl Pregibon. Correction: “Analyzing dispersion effects from replicated factorial experiments” [Technometrics 30 (1988), no. 3, 247–257]. Technometrics, 31(1):133, February 1989. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1270396. See

[NP88b].

Narula:1977:PLR

[NW77] Subhash C. Narula and John F. Wellington. Prediction, linear regression and the minimum sum of relative errors. Technometrics, 19(2):185–190, May 1977. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/ stable/1268628. See corrigenda [Ano80f].

Narula:1979:SVL

[NW79] Subhash C. Narula and John F. Wellington. Selection of variables in linear regression using the minimum sum of weighted absolute errors criterion. Technometrics, 21(3):299–306, August 1979. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1267751. See corrigenda

[Ano80f].

Narula:1985:IAM

[NW85] Subhash C. Narula and John F. Wellington. Interior analysis for the minimum sum of absolute errors regression. Technometrics, 27(2):181–188, May 1985. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268766.

Nair:1989:MLE

[NW89] Vijayan N. Nair and Paul C. C. Wang. Maximum likelihood estimation under a successive sampling discovery model. Technometrics, 31(4):423–436, November 1989. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1269993.

Nychka:1989:BRB

[Nyc89] Douglas Nychka. Book review: Discovering Causal Structure by C. Glymour; R. Scheines; P. Sprites; K. Kelly. Technometrics, 31(2):273–274, May 1989. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268846.

Obremski:1981:BRB

[Obr81] Thomas E. Obremski. Book review: Practical Nonparametric Statistics by W. J. Conover. Technometrics, 23(4):415–416, November 1981. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1268234.

Obremski:1982:BRBb

[Obr82a] Thomas E. Obremski. Book review: Elementary Statistics by Bernard W. Lindgren; Donald A. Berry. Technometrics, 24(4): 342–343, November 1982. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267839.

Obremski:1982:BRBa

[Obr82b] Thomas E. Obremski. Book review: Statistics for Management by Richard I. Levin. Technometrics, 24(4):342, November 1982. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1267838.

Odeh:1989:SSD

[OCO89] Robert E. Odeh, Youn-Min Chou, and D. B. Owen. Sample-size determination for two-sided β-expectation tolerance intervals for a normal distribution. Technometrics, 31(4):461–468, November 1989. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1269996.

Ozturk:1985:LSE

[ÖD85] Aydin Öztürk and Robert F. Dale. Least squares estimation of the parameters of the generalized Lambda distribution. Technometrics, 27(1):81–84, February 1985. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1270473.

Odoroff:1980:BRB

[Odo80] Charles L. Odoroff. Book review: The Information in Contingency Tables by D. V. Gokhale; Solomon Kullback. Technometrics, 22(4):632, November 1980. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268207.

Owen:1981:PIS

[OL81] D. B. Owen and Loretta Li. Prediction intervals for screening using a measured correlated variate. Technometrics, 23(2):165–170, May 1981. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1268033.

Oman:1982:STS

[Oma82] Samuel D. Oman. Shrinking towards subspaces in multiple linear regression. Technometrics, 24(4):307–311, November 1982. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267825.

Oman:1983:REB

[Oma83] Samuel D. Oman. Regression estimation for a bounded response over a bounded region. Technometrics, 25(3):251–261, August 1983. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1268610.

Oman:1984:ARC

[Oma84] Samuel D. Oman. Analyzing residuals in calibration problems. Technometrics, 26(4):347–353, November 1984. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1269497.

Orban:1981:BAL

[Orb81] John E. Orban. [A Bayesian Analysis of the Linear Calibration Problem]: Discussion. Technometrics, 23(4):342–343, November 1981. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1268222. See

[HL81b].

OReilly:1988:TCS

[OS88] F. J. O’Reilly and M. A. Stephens. Transforming censored samples for testing fit. Technometrics, 30(1):79–86, February 1988. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1270323.

Owen:1988:BRB

[Owe88] Willis L. Owen. Book review: Modern Statistical Selection, Part I by M. Haseeb Rizvi. Technometrics, 30(1):127–128, February 1988. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1270345.

Owen:1989:BRB

[Owe89] Willis L. Owen. Book review: Modern Statistical Selection by M. H. Rizvi. Technometrics, 31(1):126–127, February 1989. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1270386.

Page:1985:ERE

[Pag85] Judith Tama Page. Error-rate estimation in discriminant analysis. Technometrics, 27(2):189–198, May 1985. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1268767.

Park:1981:COR

[Par81] Sung H. Park. Collinearity and optimal restrictions on regression parameters for estimating responses. Technometrics, 23(3):289–295, August 1981. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1267793.

Parzen:1982:BRB

[Par82] Emanuel Parzen. Book review: Probabilistic Models in Engineering Sciences by H. J. Larson; B. O. Shubert. Technometrics, 24(3):252, August 1982. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1268691.

Parzen:1983:BRB

[Par83] Emanuel Parzen. Book review: Time Series Analysis, Proceedings of the International Conference Held at Houston, Texas, August 1980 by O. D. Anderson; M. Ray Perryman. Technometrics, 25(4):383–384, November 1983. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267861.

Park:1987:CMD

[Par87] Sung H. Park. A class of multifactor designs for estimating the slope of response surfaces. Technometrics, 29(4):449–453, November 1987. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1269456.

Parker:1988:TIO

[Par88a] I. Parker. Transformations and influential observations in minimum sum of absolute errors regression. Technometrics, 30(2):215–220, May 1988. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1270167.

Parr:1988:SNR

[Par88b] William C. Parr. [Signal-to-Noise Ratios, Performance Criteria, and Transformations]: Discussion. Technometrics, 30(1):22–23, February 1988. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1270313. See

[Box88a].

Parunak:1988:BRB

[Par88c] Anita Parunak. Book review: The SPSS Guide to Data Analysis by Marija J. Norusis. Technometrics, 30(2):237–239, May 1988. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1270181.

Parr:1989:BRB

[Par89] William C. Parr. Book review: Introduction to Quality Engineering: Designing Quality into Products and Processes by Genichi Taguchi. Technometrics, 31(2):255–256, May 1989. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1268824.

Pfeffermann:1985:SMA

[PBT85] D. Pfeffermann and S. Ben-Tuvia. A statistical method for assigning priorities to road sections for resurfacing. Technometrics, 27(1):75–79, February 1985. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1270472.

Piepel:1985:MME

[PC85] Gregory F. Piepel and John A. Cornell. Models for mixture experiments when the response depends on the total amount. Technometrics, 27(3):219–227, August 1985. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1269703.

Pfeifer:1980:IIF

[PD80a] Phillip E. Pfeifer and Stuart Jay Deutsch. Identification and interpretation of first order space-time ARMA models. Technometrics, 22(3):397–408, August 1980. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268325.

Pfeifer:1980:TSI

[PD80b] Phillip E. Pfeifer and Stuart Jay Deutsch. A three-stage iterative procedure for space-time modeling. Technometrics, 22(1):35–47, February 1980. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1268381.

Pendleton:1981:BRB

[Pen81] Olga J. Pendleton. Book review: Introductory Statistics with Applications by W. W. Daniel. Technometrics, 23(1):111–112, February 1981. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1267985.

Perry:1982:BRB

[Per82] Robert L. Perry. Book review: Frontiers in Statistical Quality Control by H. J. Lenz; G. B. Wetherill; P.-Th. Wilrich. Technometrics, 24(3):252–253, August 1982. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268692.

Perry:1984:BRB

[Per84] Robert L. Perry. Book review: Acceptance Sampling in Quality Control by Edward G. Schilling. Technometrics, 26(3):293, August 1984. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1267560.

Perry:1988:BRB

[Per88] Robert L. Perry. Book review: Statistical Process Control by John S. Oakland. Technometrics, 30(4):454–455, November 1988. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1269812.

Percival:1989:BRB

[Per89] Donald B. Percival. Book review: Digital Image Processing by Rafael C. Gonzalez; Paul Wintz. Technometrics, 31(1):124–125, February 1989. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1270383.

Piegorsch:1986:NUP

[PG86] Walter W. Piegorsch and Beth C. Gladen. A note on the use of prior interval information in constructing interval estimates for a gamma mean. Technometrics, 28(3):269–273, August 1986. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1269083. See discussion

[GGP87].

Piepel:1982:MCE

[Pie82] Gregory F. Piepel. Measuring component effects in constrained mixture experiments. Technometrics, 24(1):29–39, February 1982. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1267575.

Piepel:1983:CCC

[Pie83a] Gregory F. Piepel. Calculating centroids in constrained mixture experiments. Technometrics, 25(3):279–283, August 1983. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268613.

Piepel:1983:DCC

[Pie83b] Gregory F. Piepel. Defining consistent constraint regions in mixture experiments. Technometrics, 25(1):97–101, February 1983. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267731.

Piegorsch:1986:CBP

[Pie86] Walter W. Piegorsch. Confidence bands for polynomial regression with fixed intercepts. Technometrics, 28(3):241–246, August 1986. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1269079.

Piegorsch:1987:PLB

[Pie87] Walter W. Piegorsch. Performance of likelihood-based interval estimates for two-parameter exponential samples subject to Type I censoring. Technometrics, 29(1):41–49, February 1987. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1269882.

Piepel:1988:BRB

[Pie88a] Gregory F. Piepel. Book review: The Statistical Analysis of Compositional Data by J. Aitchison. Technometrics, 30(1):120–121, February 1988. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1270335.

Piepel:1988:NMM

[Pie88b] Gregory F. Piepel. A note on models for mixture-amount experiments when the total amount takes a zero value. Technometrics, 30(4):449–450, November 1988. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1269809.

Piepel:1989:BRB

[Pie89] Gregory F. Piepel. Book review: Robust Regression and Outlier Detection by Peter J. Rousseeuw; Annick M. Leroy. Technometrics, 31(2):260–261, May 1989. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268828.

Piepel:1990:BRBa

[Pie90] Gregory F. Piepel. Book review: Quality Control, Robust Design, and the Taguchi Method by Khosrow Dehnad. Technometrics, 32(2): 221, May 1990. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1268868.

See corrections [Ano90c].

Pigeon:1988:BRB

[Pig88] Joseph G. Pigeon. Book review: Understanding Statistics by Bruce J. Chalmer. Technometrics, 30(1):121–122, February 1988. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1270337.

Pike:1984:EDR

[Pik84] Derek J. Pike. [Experimental Design: Review and Comment]: Discussion. Technometrics, 26(2):105–109, May 1984. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268099. See [Hah84a].

Pilon:1983:BRBb

[Pil83a] Pierre P. Pilon. Book review: Principes d’Expérimentation by Pierre Dagnelie. Technometrics, 25(2):207, May 1983. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268560.

Pilon:1983:BRBa

[Pil83b] Pierre P. Pilon. Book review: Théorie et Méthodes Statistiques Exercises by Pierre Dagnelie. Technometrics, 25(2):207, May 1983. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1268559.

Placzkowski:1988:BRB

[Pla88] Gene Placzkowski. Book review: Quality Technician’s Handbook by Gary Griffith. Technometrics, 30(1):132, February 1988. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1270353.

Pigeon:1989:NPR

[PM89] Joseph G. Pigeon and Paul R. McAllister. A note on partially replicated orthogonal main-effect plans. Technometrics, 31(2):249– 251, May 1989. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1268822.

Portier:1982:BRB

[Por82a] Kenneth M. Portier. Book review: Sampling Methods for Censues and Surveys by Frank Yates. Technometrics, 24(3):254–255, August 1982. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1268696.

Portnoy:1982:BRB

[Por82b] Stephen Portnoy. Book review: Robust Statistics by Peter J. Huber. Technometrics, 24(2):163–164, May 1982. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1268499.

Peele:1982:MLR

[PR82] Lawrence Peele and Thomas P. Ryan. Minimax linear regression estimators with application to ridge regression. Technometrics, 24(2):157–159, May 1982. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268496.

Pignatiello:1987:PMI

[PR87] Joseph J. Pignatiello, Jr. and John S. Ramberg. [Performance Measures Independent of Adjustment]: Discussion. Technometrics, 29(3):274–277, August 1987. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1269336. See [LSK87a].

Prescott:1983:OD

[Pre83a] Philip Prescott. [Outlier. . . . . . . . . s]: Discussion. Technometrics, 25(2):156–157, May 1983. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268545. See [BC83a].

Preston:1983:BRB

[Pre83b] David B. Preston. Book review: Spectral Analysis and Time Series by M. B. Priestley. Technometrics, 25(2):213–214, May 1983. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1268567.

Price:1981:BRB

[Pri81] Bertram Price. Book review: Forecasting: TIMS Studies in the Management Sciences. Vol. 12 by S. Makridakis; S. C. Wheelwright. Technometrics, 23(1):113–114, February 1981. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1267988.

Price:1985:BRB

[Pri85] Bertram Price. Book review: How to Tell the Liars from the Statisticians by Robert Hooke. Technometrics, 27(3):320–321, August 1985. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1269722.

Propst:1987:BRBa

[Pro87a] Annabeth L. Propst. Book review: Statistics: Concepts and Controversies by David S. Moore. Technometrics, 29(1):121–122, February 1987. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1269898.

Propst:1987:BRBb

[Pro87b] Annabeth L. Propst. Book review: Volume 1: How to Analyze Data with Simple Plots by Wayne Nelson. Technometrics, 29(4):488, November 1987. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1269461.

Propst:1987:BRBc

[Pro87c] Annabeth L. Propst. Book review: Volume 8: How to Apply Response Surface Methodology by John A. Cornell. Technometrics, 29(4):489–490, November 1987. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1269465.

Propst:1987:BRBd

[Pro87d] Annabeth L. Propst. Book review: Volume 9: How to Use Regression Analysis in Quality Control by Douglas C. Crocker. Technometrics, 29(4):490–491, November 1987. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1269466.

Propst:1988:BRBb

[Pro88a] Annabeth Propst. Book review: Statistics: Concepts and Applications by David R. Anderson; Dennis J. Sweeney; Thomas A. Williams. Technometrics, 30(4):461–462, November 1988. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1269821.

Propst:1988:BRBa

[Pro88b] Annabeth L. Propst. Book review: Think and Explain with Statistics by Lincoln E. Moses. Technometrics, 30(2):233, May 1988. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1270173.

Propst:1989:BRBa

[Pro89a] Annabeth L. Propst. Book review: Practical Statistics for Engineers and Scientists by Nicholas P. Cheremisinoff. Technometrics, 31(1):118, February 1989. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1270376.

Propst:1989:BRBb

[Pro89b] Annabeth L. Propst. Book review: Understanding Industrial Experimentation by Donald J. Wheeler. Technometrics, 31(4):480, November 1989. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1269999.

Proskin:1989:BRB

[Pro89c] Howard M. Proskin. Book review: Theory of Block Designs by Aloke Dey. Technometrics, 31(1):128, February 1989. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1270389.

Prybutok:1989:BRB

[Pry89] Victor R. Prybutok. Book review: An Introduction to Statistical Methods and Data Analysis by Lyman Ott. Technometrics, 31(3):389–390, August 1989. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/3556159.

Pack:1987:OCR

[PS87] C. D. Pack and S. Ben Saad. Optimal call-record sampling for network planning and administration. Technometrics, 29(3):311–321, August 1987. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1269341.

Quesenberry:1982:SAP

[QK82] C. P. Quesenberry and Jacqueline Kent. Selecting among probability distributions used in reliability. Technometrics, 24(1):59–65, February 1982. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1267579.

See [Ano82e].

Quesenberry:1988:BRB

[Que88] Charles Quesenberry. Book review: Introduction to Statistical Quality Control by Douglas C. Montgomery. Technometrics, 30 (3):356–357, August 1988. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1270099.

Reynolds:1988:CVS

[RAAN88] Marion R. Reynolds, Jr., Raid W. Amin, Jesse C. Arnold, and Joel A. Nachlas. X̄ charts with variable sampling intervals. Technometrics, 30(2):181–192, May 1988. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1270164.

Rabinowitz:1989:BRB

[Rab89] Larry Rabinowitz. Book review: Mathematical Statistics and Data Analysis by John A. Rice. Technometrics, 31(3):390–391, August 1989. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/3556160.

Randles:1982:BRB

[Ran82] Ronald H. Randles. Book review: Approximation Theorems of Mathematical Statistics by Robert J. Serfling. Technometrics, 24(3):254, August 1982. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268695.

Randles:1986:BRB

[Ran86] Ronald H. Randles. Book review: Nonparametric Statistical Inference by Jean Dickinson Gibbons. Technometrics, 28(3):275, August 1986. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1269084.

Randles:1987:BRB

[Ran87] Ronald H. Randles. Book review: Handbook of Statistics, Volume 4: Nonparametric Methods by P. R. Krishnaiah; P. K. Sen. Technometrics, 29(1):122–123, February 1987. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1269899.

Rao:1980:BRB

[Rao80] Poduri S. R. S. Rao. Book review: Foundations of Inference in Survey Sampling by Claes-Magnus Cassel; Carl-Erik Sarndal; Jan Hakan Wretman. Technometrics, 22(2):279–280, May 1980. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1268472.

Rocke:1982:RER

[RDR82] David M. Rocke, George W. Downs, and Alan J. Rocke. Are robust estimators really necessary? Technometrics, 24(2):95–101, May 1982. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1268485.

Read:1989:BRB

[Rea89] Sherry L. Read. Book review: Statistical Methods for Quality Improvement by Hitoshi Kume. Technometrics, 31(3):384, August 1989. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/3556152.

Reed:1981:BRB

[Ree81] Allen H. Reed. Book review: Statistical Exercises in Medical Research by John F. Osborn. Technometrics, 23(3):310, August 1981. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1267803.

Reed:1988:BRB

[Ree88] Allen H. Reed. Book review: The Statistical Analysis of Data by T. W. Anderson; S. L. Sclove. Technometrics, 30(2):236–237, May 1988. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1270179.

Relles:1982:BRB

[Rel82] Daniel A. Relles. Book review: Statistical Computing by William J. Kennedy,; James E. Gentle. Technometrics, 24(4):341–342, November 1982. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1267837.

Reynolds:1988:BRB

[Rey88] Marion R. Reynolds, Jr. Book review: Introduction to Probability and Statistics for Engineers and Scientists by Sheldon M. Ross. Technometrics, 30(4):457–459, November 1988. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1269818.

Reiser:1986:SIN

[RG86] Benjamin Reiser and Irwin Guttman. Statistical inference for Pr(Y < X): The normal case. Technometrics, 28(3):253–257, August 1986. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1269081.

Ravishanker:1987:ASP

[RHM87] Nalini Ravishanker, Yosef Hochberg, and Edward L. Melnick. Approximate simultaneous prediction intervals for multiple forecasts. Technometrics, 29(3):371–376, August 1987. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1269346.

Rigdon:1988:BRB

[Rig88] Steven E. Rigdon. Book review: Statistical Models in Applied Science by Karl Bury. Technometrics, 30(1):129–130, February 1988.

CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1270348.

Rigdon:1989:BRB

[Rig89] Steven E. Rigdon. Book review: Reliability Data Bases by A. Amendola; A. Keller. Technometrics, 31(2):268–269, May 1989. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268839.

Rencher:1980:BWS

[RL80] Alvin C. Rencher and Steven F. Larson. Bias in Wilks’ Λ in stepwise discriminant analysis. Technometrics, 22(3):349–356, August 1980. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1268319.

Robinson:1983:BCI

[Rob83] J. A. Robinson. Bootstrap confidence intervals in location-scale models with progressive censoring. Technometrics, 25(2):179–187, May 1983. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1268551.

Robinson:1988:ERF

[Rob88] Jeffrey A. Robinson. [Estimation of Reliability in Field-Performance Studies]: Discussion. Technometrics, 30(4):384–386, November 1988. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1269799. See

[KL88a].

Rocke:1987:BRBa

[Roc87a] David M. Rocke. Book review: Analysis of Experiments with Missing Data by Yadolah Dodge. Technometrics, 29(1):116, February 1987. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1269892.

Rocke:1987:BRBb

[Roc87b] David M. Rocke. Book review: Robust Inference by M. L. Tiku; W. Y. Tan; N. Balakrishnan. Technometrics, 29(4):495–496, November 1987. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1269476.

Rocke:1989:RCC

[Roc89] David M. Rocke. Robust control charts. Technometrics, 31(2):173– 184, May 1989. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1268815.

Rohatgi:1980:BRB

[Roh80] V. K. Rohatgi. Book review: Advances in Fuzzy Set Theory and Applications by M. M. Gupta; R. K. Ragade; R. R. Yager. Technometrics, 22(4):632–633, November 1980. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1268208.

Rohatgi:1981:BRB

[Roh81] V. K. Rohatgi. Book review: Stochastic Integration by Michel Metivier; J. Pellaumail. Technometrics, 23(2):209–210, May 1981. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1268053.

Rohatgi:1982:BRB

[Roh82] Vijay K. Rohatgi. Book review: Abstract Inference by Ulf Grenander. Technometrics, 24(1):82–83, February 1982. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267586.

Rohatgi:1983:BRBa

[Roh83a] Vijay K. Rohatgi. Book review: A Course in the Theory of Stochastic Processes by A. D. Wentzell; S. Chomet. Technometrics, 25(1):116, February 1983. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267744.

Rohatgi:1983:BRBb

[Roh83b] Vijay K. Rohatgi. Book review: Statistics and Probability: Essays in Honor of C. R. Rao by G. Kallianpur; P. R. Krishnaiah; J. K. Ghosh. Technometrics, 25(4):386–387, November 1983. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267866.

Rosner:1983:PPG

[Ros83] Bernard Rosner. Percentage points for a generalized ESD manyoutlier procedure. Technometrics, 25(2):165–172, May 1983. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1268549.

Rencher:1980:IRB

[RP80] Alvin C. Rencher and Fu Ceayong Pun. Inflation of R^{2} in best subset regression. Technometrics, 22(1):49–53, February 1980. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1268382.

Reilly:1981:BSE

[RPL81] Park M. Reilly and Hugo Patino-Leal. A Bayesian study of the error-in-variables model. Technometrics, 23(3):221–231, August 1981. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1267784.

Rosenblatt:1981:BAL

[RS81] Joan R. Rosenblatt and Clifford H. Spiegelman. [A Bayesian Analysis of the Linear Calibration Problem]: Discussion. Technometrics, 23(4):329–333, November 1981. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268218. See [HL81b].

Rubin:1980:CPW

[Rub80] Donald B. Rubin. Composite points in weighted least squares regressions. Technometrics, 22(3):343–348, August 1980. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268318.

Ruppert:1987:BRB

[Rup87] David Ruppert. Book review: Robust Statistics: The Approach Based on Influence Functions by F. R. Hampel; E. M. Ronchetti; P. J. Rousseeuw; W. A. Stahel. Technometrics, 29(2):240–241, May 1987. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1269782.

Rust:1988:BRB

[Rus88] Steven W. Rust. Book review: Applied Extreme Value Statistics by Robert R. Kinnison. Technometrics, 30(1):126, February 1988. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1270343.

Shumway:1989:EMC

[SAJ89] R. H. Shumway, A. S. Azari, and P. Johnson. Estimating mean concentrations under transformation for environmental data with detection limits. Technometrics, 31(3):347–356, August 1989. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/3556144.

Salvia:1980:SFP

[Sal80] Anthony A. Salvia. Some fundamental properties of Kolmogorov– Smirnov consonance sets. Technometrics, 22(1):109–111, February 1980. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1268389.

Salsbura:1986:BRB

[Sal86] D. Salsbura. Book review: Statistical Analysis of Measurement Errors by John J. Jaech. Technometrics, 28(2):177–178, May 1986. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1270454.

Salvia:1989:BRB

[Sal89] Anthony A. Salvia. Book review: Statistical Methods for Business and Economics by Roger C. Pfaffenberger; James H. Patterson. Technometrics, 31(2):261, May 1989. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1268829.

Sandford:1985:NOS

[San85] Martin D. Sandford. Nonparametric one-sided confidence intervals for an unknown distribution function using censored data. Technometrics, 27(1):41–48, February 1985. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1270468.

Saniga:1989:ESC

[San89] Erwin M. Saniga. Economic statistical control-chart designs with an application to X̄ and R charts. Technometrics, 31(3):313–320, August 1989. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/3556141.

Sarle:1989:BRB

[Sar89] Warren S. Sarle. Book review: Latent Variable Models by John C. Loehlin. Technometrics, 31(4):484–485, November 1989. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1270005.

Savage:1983:BRB

[Sav83] I. Richard Savage. Book review: Contributions to Statistics by Willian G. Cochran. Technometrics, 25(2):207, May 1983. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268558.

Saxena:1981:DFT

[Sax81] K. M. Lal Saxena. Distribution-free tolerance bounds for the lifetime of an m-out-of-k system of components. Technometrics, 23(1):97–103, February 1981. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267982.

Steiger:1980:LE

[SB80] W. L. Steiger and Peter Bloomfield. Letter to the Editor. Technometrics, 22(3):450, August 1980. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268341.

Shiue:1983:TST

[SB83] Wei Kei Shiue and Lee J. Bain. A two-sample test of equal gamma distribution scale parameters with unknown common shape parameter. Technometrics, 25(4):377–381, November 1983. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267860.

Sundberg:1989:MCM

[SB89] Rolf Sundberg and Philip J. Brown. Multivariate calibration with more variables than observations. Technometrics, 31(3):365–371, August 1989. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/3556146.

Shiue:1988:TEG

[SBE88] Wei Kei Shiue, Lee J. Bain, and Max Engelhardt. Test of equal gamma-distribution means with unknown and unequal shape parameters. Technometrics, 30(2):169–174, May 1988. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1270162.

Smith:1982:STC

[SC82] Philip J. Smith and Sung C. Choi. Simple tests to compare two dependent regression lines. Technometrics, 24(2):123–126, May 1982.

CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1268489.

Schafer:1981:BRB

[Sch81a] Ray E. Schafer. Book review: Life Testing and Reliability Estimation by S. K. Sinha; B. K. Kale. Technometrics, 23(3):310–311, August 1981. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1267804.

Schmee:1981:BRB

[Sch81b] Josef Schmee. Book review: Applied Statistical Techniques by K. D. C. Stoodley; T. Lewis; C. L. S. Stainton. Technometrics, 23 (4):418–419, November 1981. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268239.

Schatzoff:1982:BRB

[Sch82] Martin Schatzoff. Book review: APL-STAT, A Do-It-Yourself Guide to Computational Statistics Using APL by James B. Ramsey; Gerald L. Musgrave. Technometrics, 24(1):84, February 1982. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1267589.

Schafer:1983:BRBb

[Sch83a] Ray E. Schafer. Book review: Bayesian Reliability Analysis by Harry F. Martz; Ray A. Waller. Technometrics, 25(2):209–210, May 1983. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1268564.

Schafer:1983:BRBa

[Sch83b] Ray E. Schafer. Book review: Statistical Models and Methods for Lifetime Data by J. F. Lawless. Technometrics, 25(1):111–112, February 1983. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1267739.

Schafer:1983:SMR

[Sch83c] Ray E. Schafer. [Statistical Methods in Reliability]: Discussion. Technometrics, 25(4):329–330, November 1983. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1267850. See [Law83a].

Schmee:1984:BRBd

[Sch84a] Josef Schmee. Book review: An Introduction to Applied Multivariate Statistics by M. S. Srivastava; E. M. Carter. Technometrics, 26(3):295, August 1984. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267563.

Schmee:1984:BRBc

[Sch84b] Josef Schmee. Book review: Angewandte Multivariate Statistik by Bernhard Flury; Hans Riedwyl. Technometrics, 26(3):294–295, August 1984. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1267562.

Schmee:1984:BRBg

[Sch84c] Josef Schmee. Book review: Attribute Sampling Plans, Tables of Tests, and Confidence Limits for Proportions by Robert E. Odeh; D. B. Owen. Technometrics, 26(4):418, November 1984. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1269512.

Schmee:1984:BRBf

[Sch84d] Josef Schmee. Book review: Introductory Statistics for Business and Economics by Richard W. Madsen; Melvin L. Moeschberger. Technometrics, 26(3):296, August 1984. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267565.

Schmee:1984:BRBb

[Sch84e] Josef Schmee. Book review: Reliability in the Acquisitions Process by Douglas J. DePriest; Robert L. Launer. Technometrics, 26(2): 193, May 1984. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1268119.

Schmee:1984:BRBe

[Sch84f] Josef Schmee. Book review: Statistical Tables for Multivariate Analysis by Heinz Kres. Technometrics, 26(3):295–296, August 1984. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1267564.

Schmee:1984:BRBa

[Sch84g] Josef Schmee. Book review: Understanding Randomness: Exercises for Statisticians by David S. Salsburg. Technometrics, 26(2):189, May 1984. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1268113.

Schmee:1985:BRBg

[Sch85a] Josef Schmee. Book review: Analysis of Messy Data, Volume I: Designed Experiments by George A. Milliken; Dallas E. Johnson. Technometrics, 27(4):440, November 1985. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1270212.

Schmee:1985:BRBc

[Sch85b] Josef Schmee. Book review: Applied Statistics: A Handbook of Techniques by Lothar Sachs. Technometrics, 27(1):90, February 1985. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1270480.

Schmee:1985:BRBf

[Sch85c] Josef Schmee. Book review: Encyclopedia of Statistical Sciences (Vol. 3: Faà di Bruno’s Formula-Hypothesis Testing) by Samuel Kotz; Norman L. Johnson. Technometrics, 27(2):210, May 1985. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1268774. See cor-

rection [Ano86e].

Schmee:1985:BRBa

[Sch85d] Josef Schmee. Book review: Matrices with Applications in Statistics by F. A. Graybill. Technometrics, 27(1):88–89, February 1985. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1270477.

Schmee:1985:BRBb

[Sch85e] Josef Schmee. Book review: Matrix Algebra Useful for Statistics by S. R. Searle. Technometrics, 27(1):89, February 1985. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1270478.

Schmee:1985:BRBd

[Sch85f] Josef Schmee. Book review: Tables for Numerical Integration by H. V. Smith. Technometrics, 27(1):90, February 1985. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1270481.

Schmee:1985:BRBe

[Sch85g] Josef Schmee. Book review: The Chronological Annotated Bibliography of Order Statistics (Vol. 1: Pre-1950) by H. Leon Harter. Technometrics, 27(2):210, May 1985. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268773.

Schmee:1986:BRBc

[Sch86a] Josef Schmee. Book review: An Introduction to Multivariate Statistical Analysis by T. W. Anderson. Technometrics, 28(2):180–181, May 1986. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1270458.

Schmee:1986:BRBb

[Sch86b] Josef Schmee. Book review: Encyclopedia of Statistical Sciences: Volume 5 by Samuel Kotz; Norman L. Johnson. Technometrics, 28(1):90, February 1986. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/ stable/1269611. See correction [Ano86e].

Schmee:1986:BRBa

[Sch86c] Josef Schmee. Book review: Outliers in Statistical Data by Vic Barnett; Toby Lewis. Technometrics, 28(1):89–90, February 1986. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1269610.

Schmoyer:1986:EDF

[Sch86d] Richard L. Schmoyer. An exact distribution-free analysis for accelerated life testing at several levels of a single stress. Technometrics, 28(2):165–175, May 1986. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1270453.

Schmee:1987:BRBa

[Sch87a] Josef Schmee. Book review: Data: A Collection of Problems from Many Fields for the Student and Research Worker by D. F. Andrews; A. M. Herzberg. Technometrics, 29(1):120–121, February 1987. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1269897.

Schmee:1987:BRBb

[Sch87b] Josef Schmee. Book review: Sequential Analysis, Tests and Confidence Intervals by David Siegmund. Technometrics, 29(1):123, February 1987. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1269900.

Schaefer:1988:BRB

[Sch88] Robert Schaefer. Book review: Introduction to Contemporary Statistical Methods by Lambert Koopmans. Technometrics, 30 (3):353–354, August 1988. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1270095.

Schiess:1989:BRB

[Sch89a] James R. Schiess. Book review: Least Square Estimation with Applications to Digital Signal Processing by Arthur A. Giordano; Frank M. Hsu. Technometrics, 31(1):123–124, February 1989. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1270382.

Schneider:1989:FCV

[Sch89b] Helmut Schneider. Failure-censored variables-sampling plans for lognormal and Weibull distributions. Technometrics, 31(2):199– 206, May 1989. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1268817.

Schuenemeyer:1989:BRBb

[Sch89c] John H. Schuenemeyer. Book review: Multivariate Data Analysis by F. Murtagh; A. Heck. Technometrics, 31(3):393, August 1989. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/3556165.

Schuenemeyer:1989:BRBa

[Sch89d] John H. Schuenemeyer. Book review: Statistical Methods for Geographers by W. A. V. Clark; P. L. Hosking. Technometrics, 31(2): 269, May 1989. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1268840.

Sclove:1981:BRBa

[Scl81a] Stanley L. Sclove. Book review: Fundamentals of Mathematics and Statistics for Students of Chemistry and Allied Subjects by C. J. Brookes; I. G. Betteley; S. M. Loxston. Technometrics, 23(1):111, February 1981. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1267984.

Sclove:1981:BRBb

[Scl81b] Stanley L. Sclove. Book review: Mathematical Statistics by John E. Freund; Ronald E. Walpole. Technometrics, 23(1):112, February 1981. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1267986.

Suich:1980:REA

[SD80] Ronald Suich and George C. Derringer. Is the regression equation adequate?: A further note. Technometrics, 22(1):125–126, February 1980. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1268392. See

reply [HJF80].

Shumway:1985:DMT

[SD85] R. H. Shumway and Z. A. Der. Deconvolution of multiple time series. Technometrics, 27(4):385–393, November 1985. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1270205.

Seila:1982:BRB

[Sei82] Andrew F. Seila. Book review: Simulation and the Monte Carlo Method by Reuven Y. Rubinstein. Technometrics, 24(2):167–168, May 1982. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1268504.

Sen:1980:BRB

[Sen80] P. K. Sen. Book review: Asymptotic Optimality of Likelihood Ratio Tests in Exponential Families by W. C. M. Kallenberg. Technometrics, 22(4):633, November 1980. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268209.

Senturia:1989:BRB

[Sen89] Jerome Senturia. Book review: System of Experimental Design (Vol. 1) by Genichi Taguchi. Technometrics, 31(2):256–257, May 1989. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1268825.

Sethuraman:1984:BRB

[Set84] J. Sethuraman. Book review: Identification and Informative Sample Size by H. H. Tigelaar. Technometrics, 26(3):291, August 1984.

CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1267556.

Smith:1980:BRB

[SG80] J. T. Smith and M. P. Griffin. Book review: Exploratory Data Analysis, by John W. Tukey. Technometrics, 22(1):129–130, February 1980. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1268394.

Schmee:1985:CLP

[SGN85] Josef Schmee, David Gladstein, and Wayne Nelson. Confidence limits for parameters of a normal distribution from singly censored samples, using maximum likelihood. Technometrics, 27(2):119–128, May 1985. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1268760.

Sprott:1982:LE

[SGO82] D. A. Sprott, Berton H. Gunter, and John E. Orban. Letters to the Editor. Technometrics, 24(3):259–261, August 1982. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268698.

Schupbach:1983:SEP

[SH83] M. Schüpbach and J. Hüsler. Simple estimators for the parameters of the extreme-value distribution based on censored data. Technometrics, 25(2):189–192, May 1983. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268552.

Steinberg:1984:EDRa

[SH84a] David M. Steinberg and William G. Hunter. Experimental design: Review and comment. Technometrics, 26(2):71–97, May 1984. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1268097. With discussion

and a reply by the authors.

Steinberg:1984:EDRb

[SH84b] David M. Steinberg and William G. Hunter. [Experimental Design: Review and Comment]: Response. Technometrics, 26(2):128–130, May 1984. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1268106. See

[Hah84a].

Shafer:1982:BRB

[Sha82] Glenn Shafer. Book review: The Enterprise of Knowledge: An Essay on Knowledge, Credal Probability, and Chance by Isaac Levi. Technometrics, 24(2):164–165, May 1982. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1268501.

Shafer:1984:BRB

[Sha84] Glenn Shafer. Book review: Faces of Science by V. V. Nalimov. Technometrics, 26(3):292, August 1984. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1267558.

Sheesley:1981:BRBa

[She81a] J. H. Sheesley. Book review: How to Perform Continuous Sampling (CSP), Vol. 2: ASQC Basic Reference Series by Kenneth S. Stephens. Technometrics, 23(2):198–199, May 1981. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268040.

Sheesley:1981:BRBb

[She81b] John H. Sheesley. Book review: The Operating Characteristic Curve Manual: Infinite Lot Size Single Sampling Plans by Daniel M. Ennis; Kenneth Mullen. Technometrics, 23(3):308, August 1981. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1267798.

Sheldon:1987:BRB

[She87] Daniel D. Sheldon. Book review: Statistics: A Conceptual Approach by K. Laurence Weldon. Technometrics, 29(4):493, November 1987. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1269471.

Shea:1988:BRB

[She88] Gerald Shea. Book review: Quality Management Handbook by Loren Walsh; Ralph Wurster; Raymond J. Kimber. Technometrics, 30(2):243, May 1988. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1270188.

Shelton:1989:BRB

[She89] John T. Shelton. Book review: Statistics and Data Analysis by Andrew Siegel. Technometrics, 31(3):385–386, August 1989. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/3556154.

Shumway:1987:BRB

[Shu87] R. H. Shumway. Book review: Statistics and Data Analysis in Geology by John C. Davis. Technometrics, 29(4):492, November 1987. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1269470.

Siddiqui:1980:BRB

[Sid80] M. M. Siddiqui. Book review: Geometric Probability by Herbert Solomon. Technometrics, 22(3):446, August 1980. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268335.

Siddiqui:1982:BRB

[Sid82] M. M. Siddiqui. Book review: Physical Applications of Stationary Time Series (With Special Reference to Digital Data Processing of Seismic Signals) by Enders A. Robinson. Technometrics, 24(2):166– 167, May 1982. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1268503.

Siddiqui:1984:BRB

[Sid84] M. M. Siddiqui. Book review: Applied Time Series Analysis II by David F. Findley. Technometrics, 26(3):292–293, August 1984. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1267559.

Simonoff:1988:DOC

[Sim88a] Jeffrey S. Simonoff. Detecting outlying cells in two-way contingency tables via backwards-stepping. Technometrics, 30(3):339–345, August 1988. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1270088. See

correction [Ano90a].

Simonoff:1988:RDD

[Sim88b] Jeffrey S. Simonoff. Regression diagnostics to detect nonrandom missingness in linear regression. Technometrics, 30(2):205–214, May 1988. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1270166.

Singh:1980:BRB

[Sin80] J. Singh. Book reviews: General Applied Statistics by F. H. Zuwaylif; Basic Statistics by M. Hamburg; Statistics, A Fresh Approach by D. H. Sanders; R. F. Murph; R. J. Eng. Technometrics, 22(4):634–635, November 1980. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268212.

Singh:1981:BRB

[Sin81] Umed Singh. Book review: An Introduction to Data Analysis by Bruce D. Bowen; Herbert F. Weisberg. Technometrics, 23(3):307, August 1981. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1267797.

Singh:1982:BRB

[Sin82] Umed Singh. Book review: Statistical Theory and Inference in Research by T. A. Bancroft; Chien-Pai Han. Technometrics, 24 (4):335–336, November 1982. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267830.

Singh:1983:LRT

[Sin83] N. Singh. The likelihood ratio test for the equality of location parameters of k(≥ 2) exponential populations based on Type II censored samples. Technometrics, 25(2):193–195, May 1983. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268553. See correction [Sin92].

Singh:1987:BRB

[Sin87] Ashok K. Singh. Book review: Metric Methods for Analyzing Partially Ranked Data by Douglas E. Critchlow. Technometrics, 29(3):385, August 1987. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1269357.

Singh:1992:CLR

[Sin92] N. Singh. Correction to: “The likelihood ratio test for the equality of location parameters of k (≥ 2) exponential populations based on Type II censored samples” [Technometrics 25 (1983), no. 2, 193– 195; MR0702171 (84g:62031)]. Technometrics, 34(4):502, November 1992. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1268969. See

[Sin83].

Snapinn:1984:CER

[SK84] Steven M. Snapinn and James D. Knoke. Classification error rate estimators evaluated by unconditional mean squared error. Technometrics, 26(4):371–378, November 1984. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1269500.

Snapinn:1985:ESC

[SK85] Steven M. Snapinn and James D. Knoke. An evaluation of smoothed classification error-rate estimators. Technometrics, 27(2):199–206, May 1985. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1268768.

Shelton:1983:SCP

[SKC83] John T. Shelton, Andre I. Khuri, and John A. Cornell. Selecting check points for testing lack of fit in response surface models. Technometrics, 25(4):357–365, November 1983. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1267857.

Skillings:1980:NDJ

[Ski80] John H. Skillings. On the null distribution of Jonckheere’s statistic used in two-way models for ordered alternatives. Technometrics, 22(3):431–436, August 1980. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268329.

Skillings:1983:BRB

[Ski83] John H. Skillings. Book review: Distribution-Free Statistical Methods by J. S. Maritz. Technometrics, 25(2):208–209, May 1983. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1268563.

Skillings:1981:UFT

[SM81] John H. Skillings and Gregory A. Mack. On the use of a Friedmantype statistic in balanced and unbalanced block designs. Technometrics, 23(2):171–177, May 1981. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268034.

Swallow:1984:MCC

[SM84] William H. Swallow and John F. Monahan. Monte Carlo comparison of ANOVA, MIVQUE, REML, and ML estimators of variance components. Technometrics, 26(1):47–57, February 1984. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268415.

Smith:1987:BRB

[Smi87] Paul J. Smith. Book review: Multivariate Statistical Modeling by Ronald Christensen. Technometrics, 29(3):382–383, August 1987. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1269353.

Smith:1988:BRB

[Smi88] Paul J. Smith. Book review: Modern Elementary Statistics, with Theoretical Supplement and BASIC Programming by Donald W. Zimmerman; Richard H. Williams. Technometrics, 30(1):123, February 1988. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1270339.

Spjotvoll:1982:RLS

[SMV82] Emil Spjøtvoll, Harald Martens, and Rolf Volden. Restricted least squares estimation of the spectra and concentration of two unknown constituents available in mixtures. Technometrics, 24(3):173–180, August 1982. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1268676.

Snee:1981:DBM

[Sne81] Ronald D. Snee. Developing blending models for gasoline and other mixtures. Technometrics, 23(2):119–130, May 1981. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1268027.

Snee:1983:DLR

[Sne83] Ronald D. Snee. [Developments in Linear Regression Methodology: 1959–1982]: Discussion. Technometrics, 25(3):230–237, August 1983. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1268604. See

[Hoc83a].

Springer:1983:BRB

[Spr83] Melvin D. Springer. Book reviews: Selected Tables in Mathematical Statistics (Vol. VII), The Product of Two Normally Distributed Random Variables by William Q. Meeker,; Larry W. Cornwell; Leo A. Aroian. Technometrics, 25(2):211–212, May 1983. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268565.

Spurrier:1981:IGM

[Spu81] John D. Spurrier. An improved GT2 method for simultaneous confidence intervals on pairwise differences. Technometrics, 23(2):189– 192, May 1981. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1268037.

Slifker:1980:JSS

[SS80] James F. Slifker and Samuel S. Shapiro. The Johnson system: Selection and parameter estimation. Technometrics, 22(2):239–246, May 1980. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1268463.

Stewart:1981:BEC

[SS81] Warren E. Stewart and Jan P. Sørensen. Bayesian estimation of common parameters from multiresponse data with missing observations. Technometrics, 23(2):131–141, May 1981. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1268028. See corrigenda

[Ano82f].

Smith:1982:FOI

[SS82] Dennis E. Smith and Denise D. Schmoyer. First-order “interruptible” designs. Technometrics, 24(1):55–58, February 1982. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267578.

Spinelli:1987:TEW

[SS87] John J. Spinelli and Michael A. Stephens. Tests for exponentiality when origin and scale parameters are unknown. Technometrics, 29(4):471–476, November 1987. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1269459.

Srivastava:1989:SDA

[SS89] V. K. Srivastava and N. Singh. Small-disturbance asymptotic theory for linear-calibration estimators. Technometrics, 31(3):373–378, August 1989. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/3556147.

See correction [Ano90b].

Sacks:1989:DCE

[SSW89] Jerome Sacks, Susannah B. Schiller, and William J. Welch. Designs for computer experiments. Technometrics, 31(1):41–47, February 1989. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1270363.

Schwetlick:1985:NME

[ST85] Hubert Schwetlick and Volker Tiller. Numerical methods for estimating parameters in nonlinear models with errors in the variables. Technometrics, 27(1):17–24, February 1985. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1270465.

Simonoff:1986:JBE

[ST86] Jeffrey S. Simonoff and Chih-Ling Tsai. Jackknife-based estimators and confidence regions in nonlinear regression. Technometrics, 28(2):103–112, May 1986. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1270446.

StJohn:1987:BRL

[St.87] Ralph C. St. John. Book review: Linear Statistical Models: An Applied Approach by Bruce L. Bowerman; Richard T. O’Connell; David A. Dickey. Technometrics, 29(2):237–238, May 1987. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1269779.

StJohn:1989:BRR

[St.89] Ralph C. St. John. Book review: Response Surfaces: Designs and Analyses by André I. Khuri; John A. Cornell. Technometrics, 31(4):481, November 1989. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1270000.

Stanley:1988:BRB

[Sta88] Richard M. Stanley. Book review: Optimal Estimation with an Introduction to Stochastic Control by Frank L. Lewis. Technometrics, 30(3):359–361, August 1988. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1270104.

Stein:1987:LSP

[Ste87] Michael Stein. Large sample properties of simulations using Latin hypercube sampling. Technometrics, 29(2):143–151, May 1987. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1269769. See correction

[Ste90].

Steinberg:1988:FET

[Ste88] David M. Steinberg. Factorial experiments with time trends. Technometrics, 30(3):259–269, August 1988. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1270080.

Steele:1989:BRB

[Ste89a] Michael Steele. Book review: Elements of Statistical Computing: Numerical Computation by Ronald A. Thisted. Technometrics, 31(4):482–483, November 1989. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1270002.

Stein:1989:BRB

[Ste89b] Michael Stein. Book review: Statistical Analysis of Spherical Data by N. I. Fisher; T. Lewis; B. J. J. Embleton. Technometrics, 31(3):393–394, August 1989. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/3556167.

Stein:1990:CLS

[Ste90] Michael Stein. Correction: “Large sample properties of simulations using Latin hypercube sampling” [Technometrics 29 (1987), no. 2, 143–151; MR0887702 (88e:62055)]. Technometrics, 32(3):367, August 1990. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1269142.

See [Ste87].

Shoemaker:1988:SNR

[STL88] Anne C. Shoemaker, Kwok-Leung Tsui, and Ramon V. León. [Signal-to-Noise Ratios, Performance Criteria, and Transformations]: Discussion. Technometrics, 30(1):19–21, February 1988. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1270312. See [Box88a].

Stone:1980:BRB

[Sto80] R. R. Stone. Book reviews: A Basic Course in Statistics by G. M. Clarke; D. Cooke; Elementary Statistics in a World of Applications by Ramakant Khazanie. Technometrics, 22(4):632, November 1980. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1268206.

Suzuki:1985:ELP

[Suz85] Kazuyuki Suzuki. Estimation of lifetime parameters from incomplete field data. Technometrics, 27(3):263–271, August 1985. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1269707.

Shih:1986:AIM

[SW86] Weichung J. Shih and Sanford Weisberg. Assessing influence in multiple linear regression with incomplete data. Technometrics, 28(3):231–239, August 1986. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1269078.

Swanson:1980:BRB

[Swa80] Ronald E. Swanson. Book review: Practical Experiences with Modeling and Forecasting Time Series by Gwilym M. Jenkins. Technometrics, 22(2):277–278, May 1980. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268470.

Swaine:1981:BRB

[Swa81a] Ramona B. Swaine. Book review: Theory of Probability by Bruno deFinetti. Technometrics, 23(2):204, May 1981. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1268047.

Swallow:1981:VLM

[Swa81b] William H. Swallow. Variances of locally minimum variance quadratic unbiased estimators (“MIVQUE’s”) of variance components. Technometrics, 23(3):271–283, August 1981. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267791.

Speevak:1987:MAS

[SY87] T. Speevak and A. K. Yu. Multiple attributes sampling acceptance plans involving destructive testing. Technometrics, 29(1):103–107, February 1987. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1269888.

Sylwester:1981:BRB

[Syl81] David L. Sylwester. Book review: Mathematics and Statistics for the Bio-Sciences by G. Eason; C. W. Coles; G. Gettinby. Technometrics, 23(4):416, November 1981. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268235.

Sylwester:1987:BRB

[Syl87] David Sylwester. Book review: Modern Methods for Quality Control and Improvement by Harrison M. Wadsworth; Kenneth S. Stephens; A. Blanton Godfrey. Technometrics, 29(3):378–379, August 1987. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1269348.

Sylwester:1988:BRB

[Syl88] David Sylwester. Book review: Introduction to Probability and Statistics by William Mendenhall. Technometrics, 30(3):356, August 1988. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1270097.

Syski:1982:BRB

[Sys82] R. Syski. Book reviews: Sample Path Properties of Stable Processes by J. L. Mijnheer; Random Walks with Stationary Increments and Renewal Theory by H. C. P. Berbee; Branching Processes with Continuous State Space by P. J. M. Kallenberg. Technometrics, 24(2):165–166, May 1982. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268502.

Tang:1987:EDO

[Tan87] Kwei Tang. Economic design of a one-sided screening procedure using a correlated variable. Technometrics, 29(4):477–485, November 1987. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1269460.

Taylor:1983:BRB

[Tay83] Howard Taylor. Book review: Stochastic Processes by J. Medhi. Technometrics, 25(4):388, November 1983. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1267868.

Tsang:1980:DUF

[TB80] H. Y. Tsang and D. W. Bacon. Detection of unsuspected feedback in linear dynamic systems. Technometrics, 22(4):509–516, November 1980. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1268188.

Terpenning:1989:BRB

[Ter89] Irma Terpenning. Book review: Exploring Statistics by D. Raghavarao. Technometrics, 31(4):497–498, November 1989. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1270017.

Thisted:1988:BRB

[Thi88] Ronald Thisted. Book review: Statistical Image Processing and Graphics by Edward J. Wegman; Douglas J. DePriest. Technometrics, 30(1):126–127, February 1988. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1270344.

Thompson:1981:FR

[Tho81] W. A. Thompson, Jr. On the foundations of reliability. Technometrics, 23(1):1–13, February 1981. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267969.

Thompson:1983:SMR

[Tho83] W. A. Thompson, Jr. [Statistical Methods in Reliability]: Discussion. Technometrics, 25(4):326–329, November 1983. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267849. See [Law83a].

Thode:1986:BRB

[Tho86] Henry C. Thode, Jr. Book review: Exploring Data Tables, Trends, and Shapes by David C. Hoaglin; Frederick Mosteller; John W. Tukey. Technometrics, 28(4):399–400, November 1986. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268989.

Thode:1988:BRB

[Tho88] Henry C. Thode, Jr. Book review: Statistical Tools for Simulation Practitioners by Jack P. C. Kleijnen. Technometrics, 30(4):464, November 1988. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1269827.

Thode:1989:BRB

[Tho89] Henry C. Thode, Jr. Book review: GLIM: An Introduction by M. J. R. Healy. Technometrics, 31(4):485, November 1989. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1270006.

Titterington:1980:AOD

[Tit80a] D. M. Titterington. Aspects of optimal design in dynamic systems. Technometrics, 22(3):287–299, August 1980. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1268313.

Titterington:1980:CSK

[Tit80b] D. M. Titterington. A comparative study of kernel-based density estimates for categorical data. Technometrics, 22(2):259–268, May 1980. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1268466.

Tamhane:1985:DRG

[TM85] Ajit C. Tamhane and Richard S. H. Mah. Data reconciliation and Gross error detection in chemical process networks. Technometrics, 27(4):409–422, November 1985. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1270208.

Tortorella:1989:CRP

[Tor89] M. Tortorella. Comparing renewal processes, with application to reliability modeling of highly reliable systems. Technometrics, 31(3):357–364, August 1989. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/3556145.

Troxell:1980:SSS

[Tro80] Joseph R. Troxell. Suspension systems for small sample inspections. Technometrics, 22(4):517–533, November 1980. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1268189.

Tsai:1988:PTR

[Tsa88] Chih-Ling Tsai. Power transformations and reparameterizations in nonlinear regression models. Technometrics, 30(4):441–448, November 1988. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1269808.

Tucker:1982:BRB

[Tuc82] William T. Tucker. Book review: Engineering Applications of Correlation and Spectral Analysis by Julius S. Bendat; Allan G. Piersol. Technometrics, 24(1):79–80, February 1982. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1267582.

Tucker:1984:BRB

[Tuc84] William T. Tucker. Book review: Linear Estimation and Stochastic Control by M. H. A. Davis. Technometrics, 26(2):192–193, May 1984. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1268118.

Tuprah:1989:BRB

[Tup89] Kwami Tuprah. Book review: Exploring Statistics: A Modern Introduction to Data Analysis and Inference by Larry J. Kitchens. Technometrics, 31(4):498–499, November 1989. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1270018.

Turner:1989:BRB

[Tur89] David Turner. Book review: Linear Models for Unbalanced Data by S. R. Searle. Technometrics, 31(1):112–114, February 1989. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1270372.

Tanner:1987:AIE

[TW87] Martin A. Tanner and Wing Hung Wong. An application of imputation to an estimation problem in grouped lifetime analysis. Technometrics, 29(1):23–32, February 1987. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1269880.

Ullman:1988:SNR

[Ull88] Neil R. Ullman. [Signal-to-Noise Ratios, Performance Criteria, and Transformations]: Discussion. Technometrics, 30(1):36–37, February 1988. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1270317. See

[Box88a].

Vuchkov:1981:SGS

[VDY81] I. N. Vuchkov, D. L. Damgaliev, and Ch. A. Yontchev. Sequentially generated second order quasi D-optimal designs for experiments with mixture and process variables. Technometrics, 23(3):233–238, August 1981. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1267785.

Vecchia:1985:MLE

[Vec85] A. V. Vecchia. Maximum likelihood estimation for periodic autoregressive moving average models. Technometrics, 27(4):375–384, November 1985. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1270204.

Venter:1982:MDC

[Ven82] J. H. Venter. A model for the distribution of concentrations of trace analytes in samples from particulate materials. Technometrics, 24(1):19–27, February 1982. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267574.

Vecchia:1989:CRC

[VIC89] Dominic F. Vecchia, Hari K. Iyer, and Phillip L. Chapman. Calibration with randomly changing standard curves. Technometrics, 31(1):83–90, February 1989. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1270367.

Vining:1989:BRB

[Vin89] Geoff Vining. Book review: Quality Control and Application by Bertrand L. Hansen; Prabhakar M. Ghare. Technometrics, 31 (4):486–487, November 1989. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1270008.

Voelkel:1989:BRB

[Voe89] Joseph G. Voelkel. Book review: Guide to Quality Control by K. Ishikawa. Technometrics, 31(2):260, May 1989. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1268827.

Vardeman:1985:ARL

[VoR85] Stephen Vardeman and Di ou Ray. Average run lengths for CUSUM schemes when observations are exponentially distributed. Technometrics, 27(2):145–150, May 1985. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268762.

vonTress:1988:BRB

[vT88] Mark von Tress. Book review: SPSS-X Advanced Statistics Guide by Marija J. Norusis. Technometrics, 30(2):239–240, May 1988. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1270182.

Waldmann:1986:BDR

[Wal86] K.-H. Waldmann. Bounds for the distribution of the run length of one-sided and two-sided CUSUM quality control schemes. Technometrics, 28(1):61–67, February 1986. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1269604.

Walker:1989:BRB

[Wal89] Esteban Walker. Book review: Applied Regression Analysis and Other Multivariable Methods by D. Kleinbaum; L. Kupper; K. Muller. Technometrics, 31(1):117–118, February 1989. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1270375.

Wang:1985:AVG

[Wan85] P. C. Wang. Adding a variable in generalized linear models. Technometrics, 27(3):273–276, August 1985. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1269708.

Wang:1987:RPD

[Wan87] P. C. Wang. Residual plots for detecting nonlinearity in generalized linear models. Technometrics, 29(4):435–438, November 1987. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1269454.

Warde:1987:BRB

[War87a] William D. Warde. Book review: Elementary Survey Sampling by Richard L. Scheaffer; William Mendenhall; Lyman Ott. Technometrics, 29(2):248, May 1987. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1269791.

Wardrop:1987:BRB

[War87b] Daniel M. Wardrop. Book review: Multiple Regression and Analysis of Variance and Covariance by Allen Edwards. Technometrics, 29(1):120, February 1987. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1269896.

Warren:1988:BRB

[War88] John Warren. Book review: Statistical Methods for Environmental Pollution Monitoring by Richard O. Gilbert. Technometrics, 30(3):348, August 1988. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1270090.

Wasserstein:1989:BRB

[Was89] Ronald L. Wasserstein. Book review: Monte Carlo Methods, Volume 1: Basics by Malvin H. Kalos; Paula A. Whitlock. Technometrics, 31(2):269–270, May 1989. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268841.

Watson:1982:BRB

[Wat82] Geoffrey S. Watson. Book review: Circular Statistics in Biology by Edward Batschelet. Technometrics, 24(4):336, November 1982. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1267831.

Watts:1984:BRB

[Wat84] Donald G. Watts. Book review: Nonlinear Regression Modeling: A Unified Practical Approach by David A. Ratkowsky. Technometrics, 26(2):190, May 1984. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268115.

Walker:1988:CIM

[WB88a] Esteban Walker and Jeffrey B. Birch. Corrections to: “Influence Measures in Ridge Regression” [Technometrics 30 (1988), no. 2, 221–227]. Technometrics, 30(4):469–470, November 1988. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1269832. See [WB88b].

Walker:1988:IMR

[WB88b] Esteban Walker and Jeffrey B. Birch. Influence measures in ridge regression. Technometrics, 30(2):221–227, May 1988. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1270168. See corrections

[WB88a].

Wilkinson:1981:TSF

[WD81] Leland Wilkinson and Gerard E. Dallal. Tests of significance in forward selection regression with an F -to-enter stopping rule. Technometrics, 23(4):377–380, November 1981. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1268227.

Webster:1983:BRB

[Web83] John T. Webster. Book review: Recent Advances in Regression Methods by Hrishikesh D. Vinod; Aman Ullah. Technometrics, 25(2):215, May 1983. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268569.

Wegman:1981:VSE

[Weg81] Edward J. Wegman. Vector splines and the estimation of filter functions. Technometrics, 23(1):83–89, February 1981. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267980.

Weisberg:1981:SAI

[Wei81] Sanford Weisberg. A statistic for allocating C_{p} to individual cases. Technometrics, 23(1):27–31, February 1981. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1267972.

Weisberg:1983:DSP

[Wei83] Sanford Weisberg. Discussion: Some principles for regression diagnostics and influence analysis. Technometrics, 25(3):240–244, August 1983. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1268606.

Weisberg:1984:CDT

[Wei84] Sanford Weisberg. [K-Clustering as a Detection Tool for Influential Subsets in Regression]: Discussion. Technometrics, 26(4):324–325, November 1984. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1269493.

See [GL84a].

Weisberg:1987:BRB

[Wei87] Sanford Weisberg. Book review: Plots, Transformations, and Regression by A. C. Atkinson. Technometrics, 29(3):386–387, August 1987. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1269360.

Weisberg:1990:BRB

[Wei90] Sanford Weisberg. Book review: Transformations and Weighting in Regression by R. J. Carroll; D. Ruppert. Technometrics, 32(1):96–97, February 1990. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/ stable/1269852. See corrections [Ano90c].

Welch:1982:BBS

[Wel82] William J. Welch. Branch-and-bound search for experimental designs based on D-optimality and other criteria. Technometrics, 24(1):41–48, February 1982. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1267576.

Welsch:1983:DLR

[Wel83] Roy E. Welsch. [Developments in Linear Regression Methodology: 1959–1982]: Discussion. Technometrics, 25(3):245–246, August 1983. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1268607. See

[Hoc83a].

Welch:1984:CAD

[Wel84] William J. Welch. Computer-aided design of experiments for response estimation. Technometrics, 26(3):217–224, August 1984. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1267547.

Welch:1989:BRB

[Wel89] Michael E. Welch. Book review: State Space Modeling of Time Series by Masanao Aoki. Technometrics, 31(1):121–123, February 1989. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1270381.

White:1979:LRR

[WG79] J. W. White and R. F. Gunst. Latent root regression: Large sample analysis. Technometrics, 21(4):481–488, November 1979. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1268287. See corrigenda

[Ano80f].

Wilson:1980:HPL

[Wil80] A. J. C. Wilson. How probable is a large resultant in a random walk? Technometrics, 22(4):629–630, November 1980. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268203. See reply [Bin80].

Wilcox:1983:TPP

[Wil83] Rand R. Wilcox. A table of percentage points of the range of independent t variables. Technometrics, 25(2):201–204, May 1983. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1268555.

Wincek:1987:BRB

[Win87] Michael Wincek. Book review: Time Series Analysis, Identification, and Adaptive Filtering by Daniel Graupe. Technometrics, 29(3):389–391, August 1987. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1269364.

Williams:1985:UCS

[WLP85] William W. Williams, Stephen W. Looney, and Michael H. Peters. Use of curtailed sampling plans in the economic design of npcontrol charts. Technometrics, 27(1):57–63, February 1985. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1270470. See corrections

[Ano85e].

Woodall:1986:WED

[WLV86] William H. Woodall, Thomas J. Lorenzen, and Lonnie C. Vance. Weaknesses of the economic design of control charts. Technometrics, 28(4):408–410, November 1986. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1269000.

Wong:1985:SCV

[WMS85] Audrey Wong, Jeff B. Meeker, and Murray R. Selwyn. Screening on correlated variables: A Bayesian approach. Technometrics, 27(4):423–431, November 1985. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1270209.

Woodall:1985:MCQ

[WN85] William H. Woodall and Matoteng M. Ncube. Multivariate CUSUM quality-control procedures. Technometrics, 27(3):285–292, August 1985. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1269710.

Woodall:1983:DRL

[Woo83] William H. Woodall. The distribution of the run length of one-sided CUSUM procedures for continuous random variables. Technometrics, 25(3):295–301, August 1983. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268615.

Woodall:1984:MCA

[Woo84] William H. Woodall. On the Markov chain approach to the twosided CUSUM procedure. Technometrics, 26(1):41–46, February 1984. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1268414.

Woodall:1986:BRB

[Woo86] William H. Woodall. Book review: Understanding Statistical Process Control by D. J. Wheeler; D. S. Chambers. Technometrics, 28(4):402, November 1986. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268993.

Worsley:1983:TTP

[Wor83] K. J. Worsley. Testing for a two-phase multiple regression. Technometrics, 25(1):35–42, February 1983. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1267724.

Walter:1988:EEO

[WP88] S. D. Walter and M. M. Panjabi. Experimental errors in the observation of body joint kinematics. Technometrics, 30(1):71–78, February 1988. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1270322.

Wolde-Tsadik:1980:CSP

[WTA80] Girma Wolde-Tsadik and A. A. Afifi. A comparison of the “sometimes pool”, “sometimes switch” and “never pool” procedures in the two-way ANOVA random effects model. Technometrics, 22 (3):367–373, August 1980. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268321.

Wu:1981:ICN

[Wu81] Chien-Fu Wu. Iterative construction of nearly balanced assignments. I. Categorical covariates. Technometrics, 23(1):37–44, February 1981. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1267974.

Wu:1987:PMI

[Wu87] C. F. J. Wu. [Performance Measures Independent of Adjustment]: Discussion. Technometrics, 29(3):279–281, August 1987. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1269337. See [LSK87a].

Wasserman:1989:MBP

[WW89] Gary S. Wasserman and Harrison M. Wadsworth. A modified Beattie procedure for process monitoring. Technometrics, 31(4):415–421, November 1989. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1269992.

Yang:1983:RPA

[Yan83] Grace L. Yang. A renewal-process approach to continuous sampling plans. Technometrics, 25(1):59–67, February 1983. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1267727.

Yang:1985:MES

[Yan85] Shie-Shien Yang. [Multiresponse Estimation with Special Application to Linear Systems of Differential Equations]: Discussion. Technometrics, 27(4):349–351, November 1985. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1270200. See [BW85a].

Yandell:1989:BRB

[Yan89] Brain S. Yandell. Book review: Spline Smoothing and Nonparametric Regression by Randall L. Eubank. Technometrics, 31(3):380, August 1989. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/3556148.

Yashchin:1989:WCS

[Yas89] Emmanuel Yashchin. Weighted cumulative sum technique. Technometrics, 31(3):321–338, August 1989. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/3556142.

Young:1987:BCG

[YB87] David H. Young and Saad T. Bakir. Bias correction for a generalized log-gamma regression model. Technometrics, 29(2):183–191, May 1987. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1269773.

Yfantis:1987:BRB

[Yfa87] Evangelos Yfantis. Book review: Analytical Measurement and Information Advances in the Information Theoretic Approach to Chemical Analyses by K. Eckschlager; V. Stepanek. Technometrics, 29(3):391, August 1987. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1269365.

Young:1980:BRB

[You80] Walter R. Young. Book review: Outliers in Statistical Data by Vic Barnett; Toby Lewis. Technometrics, 22(4):631, November 1980. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1268205.

Young:1982:BCS

[You82] A. S. Young. The bivar criterion for selecting regressors. Technometrics, 24(3):181–189, August 1982. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268677.

Yang:1984:ETC

[YSOE84] S. S. Yang, B. O. Solomon, M. D. Oner, and L. E. Erickson. Estimating and testing common parameters for some multiresponse models associated with microbial growth and bioenergetics. Technometrics, 26(4):355–361, November 1984. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1269498.

Yanagimoto:1987:UML

[YY87] Takemi Yanagimoto and Masakatsu Yanagimoto. The use of marginal likelihood for a diagnostic test for the goodness of fit of the simple linear regression model. Technometrics, 29(1):95–101, February 1987. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1269887.

Zahn:1989:BRB

[Zah89] Douglas A. Zahn. Book review: The Teaching of Practical Statistics by C. W. Anderson; R. M. Loynes. Technometrics, 31(3):382–383, August 1989. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/3556150.

Zaino:1980:BRB

[Zai80] Nicholas Zaino, Jr. Book review: Reliability: Management, Methods, and Mathematics by D. K. Lloyd; M. Lipow. Technometrics, 22(1):131, February 1980. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268398.

Zaino:1987:BRB

[Zai87] Nicholas A. Zaino, Jr. Book review: Potential Pattern Recognition in Chemical and Medical Decision Making by D. Coomans; I. Broechaert. Technometrics, 29(4):499–500, November 1987. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1269481.

Zemroch:1986:CAE

[Zem86] Peter J. Zemroch. Cluster analysis as an experimental design generator, with application to gasoline blending experiments. Technometrics, 28(1):39–49, February 1986. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1269602.

Ziegel:1980:KMM

[ZG80] Eric R. Ziegel and John W. Gorman. Kinetic modelling with multiresponse data. Technometrics, 22(2):139–151, May 1980. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268453.

Ziegel:1984:EDR

[Zie84] Eric R. Ziegel. [Experimental Design: Review and Comment]: Discussion. Technometrics, 26(2):98–104, May 1984. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268098. See [Hah84a].

Ziegel:1985:BRB

[Zie85a] Eric Ziegel. Book review: Applied Factorial and Fractional Designs by Robert A. McLean; Virgil L. Anderson. Technometrics, 27(3):315–316, August 1985. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1269714.

Ziegel:1985:MES

[Zie85b] Eric R. Ziegel. [Multiresponse Estimation with Special Application to Linear Systems of Differential Equations]: Discussion. Technometrics, 27(4):352–357, November 1985. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1270201. See [BW85a].

Ziegel:1986:BRBb

[Zie86a] Eric Ziegel. Book review: A Celebration of Statistics by Anthony C. Atkinson; Stephen E. Fienberg. Technometrics, 28(4):404, November 1986. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1268997.

Ziegel:1986:BRBc

[Zie86b] Eric Ziegel. Book review: Multivariate Analysis VI by P. R. Krishnaiah. Technometrics, 28(4):404–405, November 1986. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268998.

Ziegel:1986:BRBa

[Zie86c] Eric R. Ziegel. Book review: The Analysis of Linear Models by Ronald R. Hocking. Technometrics, 28(1):87–89, February 1986. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1269609.

Ziegel:1987:BRBh

[Zie87a] Eric Ziegel. Book review: Data: A Collection of Problems from Many Fields for the Student and Research Worker by D. F. Andrews; A. M. Herzberg. Technometrics, 29(4):502–503, November 1987. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1269485.

Ziegel:1987:BRBb

[Zie87b] Eric Ziegel. Book review: Encyclopedia of Statistical Sciences, Volume 6: Multivariate Analysis-Plackett Burman Designs by S. Kotz; N. Johnson. Technometrics, 29(3):384, August 1987. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1269355.

Ziegel:1987:BRBa

[Zie87c] Eric Ziegel. Book review: Encyclopedia of Statistical Sciences, Volume 7: Plackett Family of Distribution-Regression: Wrong by S. Kotz; N. Johnson. Technometrics, 29(3):383, August 1987. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1269354.

Ziegel:1987:BRBg

[Zie87d] Eric Ziegel. Book review: Numerical Recipes: The Art of Scientific Computing by William Press; Brian Flannery; Saul Teukolsky; William Vetterling. Technometrics, 29(4):501–502, November 1987. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1269484.

Ziegel:1987:BRBc

[Zie87e] Eric Ziegel. Book review: Statistical Methods in Discrimination Litigation by D. H. Kaye; Mikel Aikin. Technometrics, 29(3):385–386, August 1987. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1269358.

Ziegel:1987:BRBf

[Zie87f] Eric Ziegel. Book review: Statistics and the Law by M. DeGroot; S. Fienberg; J. Kadane. Technometrics, 29(4):500–501, November 1987. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1269482.

Ziegel:1987:BRBe

[Zie87g] Eric Ziegel. Book review: Testing Statistical Hypotheses by E. L. Lehmann. Technometrics, 29(4):494, November 1987. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1269474.

Ziegel:1987:BRBd

[Zie87h] Eric Ziegel. Book review: Volume 5: How to Run Mixture Experiments for Product Quality by John Cornell. Technometrics, 29(4):489, November 1987. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1269463.

Ziegel:1988:BRBv

[Zie88a] Eric Ziegel. Book review: A Folio of Distributions by Edward B. Fowlkes. Technometrics, 30(4):462, November 1988. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1269822.

Ziegel:1988:BRBaa

[Zie88b] Eric Ziegel. Book review: Applied Statistics: A Handbook of BMDP Analyses by E. J. Snell. Technometrics, 30(4):468, November 1988. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1269830.

Ziegel:1988:BRBx

[Zie88c] Eric Ziegel. Book review: Applied Statistics: Analysis of Variance and Regression by O. J. Dunn; V. A. Clark. Technometrics, 30(4):462–463, November 1988. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1269824.

Ziegel:1988:BRBu

[Zie88d] Eric Ziegel. Book review: Nonparametric Statistical Inference by Jean Gibbons. Technometrics, 30(4):457, November 1988. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1269817.

Ziegel:1988:BRBa

[Zie88e] Eric Ziegel. Book review: Practical Experimental Designs and Optimization Methods for Chemists by C. K. Bayne; I. B. Rubin. Technometrics, 30(1):119–120, February 1988. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1270334.

Ziegel:1988:BRBt

[Zie88f] Eric Ziegel. Book review: Reliability and Quality Control by Asit Basu. Technometrics, 30(4):455, November 1988. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1269813.

Ziegel:1988:BRBy

[Zie88g] Eric Ziegel. Book review: Statistical Inference in Linear Models by H. Bunke; O. Bunke. Technometrics, 30(4):463, November 1988. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1269825.

Ziegel:1988:BRBz

[Zie88h] Eric Ziegel. Book review: Statistical Packages for the IBM PC Family by Linda O’Keeffe; J. Klagge. Technometrics, 30(4):468, November 1988. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1269831.

Ziegel:1988:BRBb

[Zie88i] Eric Ziegel. Book review: Statistics in Research and Development by Roland Caulcott. Technometrics, 30(1):122–123, February 1988. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1270338.

Ziegel:1988:BRBw

[Zie88j] Eric Ziegel. Book review: The History of Statistics by Stephen Stigler. Technometrics, 30(4):462, November 1988. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1269823.

Ziegel:1988:BRBc

[Zie88k] Eric Ziegel. Book reviews: Regression: A Second Course in Statistics by T. H. Wonnacott; R. J. Wonnacott; Multivariate Data Analysis by W. Cooley; P. Lohnes. Technometrics, 30(1):130–131, February 1988. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1270350.

Ziegel:1988:BRBe

[Zie88l] Eric R. Ziegel. Book review: A Guide to Statistical Methods and to the Pertinent Literature by Lothar Sachs. Technometrics, 30(2):241, May 1988. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1270185.

Ziegel:1988:BRBg

[Zie88m] Eric R. Ziegel. Book review: Biostatistics: A Foundation for Analysis in the Health Sciences by Wayne Daniel. Technometrics, 30(2): 243, May 1988. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1270189.

Ziegel:1988:BRBk

[Zie88n] Eric R. Ziegel. Book review: Compstat: Proceedings in Computational Statistics by F. DeAntoni; N. Lauro; A. Rezzi. Technometrics, 30(2):245–246, May 1988. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1270194.

Ziegel:1988:BRBd

[Zie88o] Eric R. Ziegel. Book review: Counterexamples in Probability and Statistics by J. Romano; A. Siegel. Technometrics, 30(2):240, May 1988. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1270183.

Ziegel:1988:BRBh

[Zie88p] Eric R. Ziegel. Book review: Filtering in the Time and Frequency Domains by H. Blinchikoff; A. Zverev. Technometrics, 30(2):245, May 1988. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1270191.

Ziegel:1988:BRBp

[Zie88q] Eric R. Ziegel. Book review: Human Resources Management by J. Kern; J. Riley; L. Jones. Technometrics, 30(3):359, August 1988. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1270103.

Ziegel:1988:BRBo

[Zie88r] Eric R. Ziegel. Book review: Integrating Productivity and Quality Management by J. A. Edosomwan. Technometrics, 30(3):358–359, August 1988. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1270102.

Ziegel:1988:BRBl

[Zie88s] Eric R. Ziegel. Book review: Introduction to Probability and Mathematical Statistics by Lee J. Bain; Max Engelhardt. Technometrics, 30(3):353, August 1988. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1270094.

Ziegel:1988:BRBi

[Zie88t] Eric R. Ziegel. Book review: Mathematical Methods by M. Potter; J. Goldberg. Technometrics, 30(2):245, May 1988. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1270192.

Ziegel:1988:BRBm

[Zie88u] Eric R. Ziegel. Book review: New Perspectives in Theoretical and Applied Statistics by M. Puri; J. Vilaplana; W. Wertz. Technometrics, 30(3):356, August 1988. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1270098.

Ziegel:1988:BRBj

[Zie88v] Eric R. Ziegel. Book review: Numerical Methods for Computer Science, Engineering, and Mathematics by John H. Mathews. Technometrics, 30(2):245, May 1988. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1270193.

Ziegel:1988:BRBr

[Zie88w] Eric R. Ziegel. Book review: Operations Research: Applications and Algorithms by Wayne Winston. Technometrics, 30(3):361–362, August 1988. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1270107.

Ziegel:1988:BRBn

[Zie88x] Eric R. Ziegel. Book review: Quality Control and Industrial Statistics by A. J. Duncan. Technometrics, 30(3):357, August 1988. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1270100.

Ziegel:1988:BRBf

[Zie88y] Eric R. Ziegel. Book review: Statistical Design: Theory and Practice by C. McCulloch; S. Schwager; G. Casella; S. Searle. Technometrics, 30(2):242, May 1988. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1270187.

Ziegel:1988:BRBq

[Zie88z] Eric R. Ziegel. Book review: Stochastic Parameter Regression Models by P. Newbold; T. Bos. Technometrics, 30(3):361, August 1988. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1270105.

Ziegel:1988:BRBs

[Zie88-27] Eric R. Ziegel. Book review: The Collected Works of John W. Tukey [Vol. 3 (1949–1964) and Vol. 4 (1965–1986): Philosophy and Principles of Data Analysis] by Lyle Jones; John W. Tukey. Technometrics, 30(3):363, August 1988. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1270109.

Ziegel:1988:RBC

[Zie88-28] Eric R. Ziegel. Review: The Collected Works of John W. Tukey, Volumes 3 and 4. Technometrics, 30(3):363, August 1988. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1270109.

Ziegel:1989:BRBl

[Zie89a] Eric Ziegel. Book review: A Dictionary of Scientific Units by H. Jerrard; D. McNeill. Technometrics, 31(2):275, May 1989. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268849.

Ziegel:1989:BRBd

[Zie89b] Eric Ziegel. Book review: Artificial Intelligence and Statistics by William Gale. Technometrics, 31(1):130, February 1989. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1270393.

Ziegel:1989:BRBe

[Zie89c] Eric Ziegel. Book review: Computer Science and Statistics: Proceedings of the 17th Symposium on the Interface by David Allen. Technometrics, 31(1):130, February 1989. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1270394.

Ziegel:1989:BRBa

[Zie89d] Eric Ziegel. Book review: Design, Data, and Analysis by Colin L. Mallows. Technometrics, 31(1):118–119, February 1989. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1270377.

Ziegel:1989:BRBk

[Zie89e] Eric Ziegel. Book review: Directory of Statistical Microcomputer Software by W. Woodward; A. Elliott; H. Gray; D. Matlock. Technometrics, 31(2):274, May 1989. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268848.

Ziegel:1989:BRBm

[Zie89f] Eric Ziegel. Book review: Handbook of Mathematical Sciences by William H. Beyer. Technometrics, 31(2):275, May 1989. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268850.

Ziegel:1989:BRBw

[Zie89g] Eric Ziegel. Book review: Introduction to the Theory and Practice of Econometrics by G. Judge; P. Hill; W. Griffiths; H. L. Lütkepohl; T. Lee. Technometrics, 31(4):501, November 1989. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1270023.

Ziegel:1989:BRBb

[Zie89h] Eric Ziegel. Book review: Kendall’s Advanced Theory of Statistics, Vol. 1: Distribution Theory by Alan Stuart; J. Keith Ord. Technometrics, 31(1):128, February 1989. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1270388.

Ziegel:1989:BRBc

[Zie89i] Eric Ziegel. Book review: Matrices for Statistics by M. J. R. Healy. Technometrics, 31(1):128, February 1989. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1270390.

Ziegel:1989:BRBx

[Zie89j] Eric Ziegel. Book review: Matrix Differential Calculus with Applications in Statistics and Econometrics by J. Magnus; H. Neudecker. Technometrics, 31(4):501–502, November 1989. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1270024.

Ziegel:1989:BRBp

[Zie89k] Eric Ziegel. Book review: Modern Mathematical Statistics by E. Dudewicz; S. Mishra. Technometrics, 31(3):391, August 1989. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/3556161.

Ziegel:1989:BRBq

[Zie89l] Eric Ziegel. Book review: Multivariate Statistical Modelling and Data Analysis by H. Bozdogan; A. Gupta. Technometrics, 31(3): 393, August 1989. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
3556166.

Ziegel:1989:BRBo

[Zie89m] Eric Ziegel. Book review: Practical Statistics for Analytical Chemists by Robert L. Anderson. Technometrics, 31(3):387, August 1989. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/3556156.

Ziegel:1989:BRBt

[Zie89n] Eric Ziegel. Book review: Proceedings of the Second International Tampere Conference in Statistics by Tarmo Pukkila; Simo Puntanen. Technometrics, 31(3):398, August 1989. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/3556174.

Ziegel:1989:BRBu

[Zie89o] Eric Ziegel. Book review: Quality Assurance of Chemical Measurements by John K. Taylor. Technometrics, 31(4):491–492, November 1989. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1270011.

Ziegel:1989:BRBh

[Zie89p] Eric Ziegel. Book review: Robustness of Statistical Methods and Nonparametric Statistics by D. Rasch; M. Tiku. Technometrics, 31(2):267–268, May 1989. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1268837.

Ziegel:1989:BRBf

[Zie89q] Eric Ziegel. Book review: Scientific and Engineering Applications with Personal Computers by R. Annino; R. Driver. Technometrics, 31(1):130–131, February 1989. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1270395.

Ziegel:1989:BRBi

[Zie89r] Eric Ziegel. Book review: Statistical Data Analysis by Y. Dodge. Technometrics, 31(2):272, May 1989. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1268844.

Ziegel:1989:BRBj

[Zie89s] Eric Ziegel. Book review: Statistical Methods for the Analysis of Biomedical Data by Robert F. Woolson. Technometrics, 31(2):274, May 1989. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1268847.

Ziegel:1989:BRBv

[Zie89t] Eric Ziegel. Book review: Statistics by J. McClave; F. Dietrich. Technometrics, 31(4):499, November 1989. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1270019.

Ziegel:1989:BRBg

[Zie89u] Eric Ziegel. Book review: Statistics Today: A Comprehensive Introduction by Donald Byrkit. Technometrics, 31(2):263–264, May 1989. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1268832.

Ziegel:1989:BRBs

[Zie89v] Eric Ziegel. Book review: The Collected Works of John W. Tukey (Vol. 5, Graphics, 1965–1985) by W. S. Cleveland. Technometrics, 31(3):398, August 1989. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/3556173.

Ziegel:1989:BRBn

[Zie89w] Eric Ziegel. Book review: The Statistical Consultant in Action by D. J. Hand; B. S. Everitt. Technometrics, 31(3):383, August 1989. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/3556151.

Ziegel:1989:BRBr

[Zie89x] Eric Ziegel. Book reviews: Annotated Computer Output for Analysis of Variance of Unbalanced Data: SAS GLM by S. R. Searle; R. P. Yerex; Annotated Computer Output for Analysis of Variance of Unbalanced Data: SPSSX ANOVA by S. R. Searle; R. P. Yerex. Technometrics, 31(3):397–398, August 1989. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/3556172.